A Complete Bibliography of Scandinavian Journal of Statistics. Theory and Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 1.08

Title word cross-reference

3 [Huc11]. α [BW19]. β [DQYZ23]. c [LYW22]. c|ĵ|^{1/2}\overline{L} [BN85a]. D [DW95, YC22, Höp99]. D_{A} [LYW22]. Δ [Jac01]. ℓ_{1} [ST10, SBV11]. F [But86, Hög79, ZL14]. f^{2}(x) [SHD94]. h [CMN08]. I [Chr89]. J [CV15]. K [CCH98, Ter14, Jan91]. L [HJS90, PC99]. L^{p} [GGS20]. L_{1} [Kuh04, Wan90]. L_{2} [HS98, KST95]. L_{p} [Arc98]. M [Adi97, BNK74, BBG97, Che91, Che15b, DT05, EHR88, GR10, Joh77, Ped75a, Ped75b, Stu83, Wan95, Wan99, Zha00, ZKP^{+}24, vZ03, BH84, GHH95, Hol80b, Hol81b, Sjö00, TGM17, Zet88]. n [BG98, PWN22]. P [Ano24, Gas23, Gre23a, Gre23b, Lav23, Ric23, Vos23, ZXL^{+}18, GHD20, GMMT06, SS09]. P (X_{2} ≤ X_{1}) [GK91]. Q [FHSZ19, Nai82]. R^{2} [DM80]. r^{∗} [Jen97]. R^{k} [Blæ78]. S [CDY11]. S^{2} [KT95]. T [KT19, And23, BCCAVMO21, Rom04, SW87]. T^{2} [But86]. × [SG04]. U [Ahm81, Ahm17, DBS10, Fre89, HS06, HLP23, KS22, SW93b, Web81]. U^{2} [Per79]. W [Cur80b]. Z [BW08].

* [Toc01].

-Ancillarity [BNK74, Joh77, Ped75a, Ped75b]. -consistent [BG98]. -Dependent [CMN08, GHH95, Hol80b, Hol81b, Sjö00, TGM17, Zet88]. -Dimensional [Höp99, PWN22]. -Estimation [GR10, Che15b, Zha00, ZKP^{+}24]. -Estimator [Che91]. -Estimators [Arc98, CDY11, EHR88, Stu83, Wan95, Wan99, BBG97, vZ03]. -Exponential [Cur80b]. -function [CV15]. -functionals [Adi97]. -Functionals* [DT05]. -learning [FHSZ19]. -means [Ter14]. -mixing [BW19]. -model [DQYZ23]. -Optimal [DW95]. -optimality [Jac01, LYW22]. -optimization [GGS20]. -Overlapping [BH84]. -Penalization [SBV11]. -Projections [Chr89]. -quantile [GW24]. -Sample [Jan91, CCH98]. -Statistics [Ahm81, Fre89, HJS90, HS06, SW93b, Web81, DBS10, HLP23, KS22, PC99]. -Test [SW87, Rom04, Hög79, ZL14]. -tests [Ahm17]. -theorem [BW08]. -value [GHD20]. -values [Ano24, Gas23, GMMT06, Gre23a, Ric23, SS09, Vos23, ZXL^{+}18, Gre23a, Gre23b, Lav23]. -vine [YC22]. -year [MS23].

1 [Blo74, Hok75, Hok76].

2011 [DGGM16]. 2019 [NDH^{+}21].

33 [AVA22]. 35 [Ano24].

4th [NDH^{+}21].

50 [Ano24]. 574 [AVA22].

Aalen [BC15, KHSS12b, BDP13, GJ05, Grø97, GNPM07, Kle91, MP14, SZS02]. Abelian [Seg02]. Absence [HH16]. Absolute [BNS05, Dal77, DR00, Ess75, HL99, TB22]. absorption [ABB24]. Abstract [Hol93]. Abundance [HH16]. Accelerated [Bor99, Doo18, HP09, QQZ16, RFK22]. Accelerating [VR08]. Acceptance [Sto11]. Accessibility [Ano75a]. Accidents [GS76]. Accounting [GK21]. Accuracy [KH16, Mam92, ZXL^{+}18, RMG19].

Accurate [DHH24, DSD^{+}14, LA16, SMB14]. Acknowledgement [YL14a]. Acknowledgements [Ano05a, Ano07a, Ano10a, Ano11a]. Acknowledgments [Ano97a]. across [CSW79, CH22]. Action [Lai83]. actions [TPH21]. Acyclic [AMP97, CL12, Gås16, Ric03, XG09]. adapted [AFV14]. Adaptive [AGR13, Cha15, CM82, CV02, CMMR12, Des23, ES00, EB08, JSDT11, Kle16, LPW21, LCZ09, LRT23, Mab17, Mar98, SS02, Sar09, SG15, SBB05, XBQF15, CFR19, CM20c, CGC23, Gås03, OHIS24, ZSJT24, ZLZZ21]. adaptive-to-model [ZLZZ21]. Adaptively [SK19]. Adaptivity [GM98]. Addendum [Nie99]. addition [Hua24]. Additional [MS98a, Sen88]. Additive [Bor99, CDZ11, CPWZ13, FMS15, GJ05, KO03, LG09, LDY16, LCZ09, LC11b, LS15, MMS16, MW12, MS01, MS09, MP14, MU91, SZS02, SMZ11, XL10, YL04, ZL18, ZHL15, HYZ22, LJZ^{+}18, MS24, SM24, VW19, YM22, YLZ^{+}19]. Adjusted [CV14, CLP17, DS94, HTK15, JM93, MS01, ŞM05, CMY24, HESZ16, KXZA20, Thu14, VD18]. Adjustment [DC00, JK04]. Adjustments [Møl86]. Admissible [FM90]. Advantages [MS12]. adversarial [Mei24]. Affects [GOV15]. Affine [Erl81, Oja99]. after [AHP^{+}18, KK19, Sve86]. Against [LW12, SBR98, AL79, AL81, Ber81, CFS95, Kle83, THF18, Xie89]. Age [Ber79a, HJO15, OKK^{+}00, SSD15, SS79]. Ageing [DS90, SD85]. Aggregate [Gil86, Tho83]. Aggregation [LM16, TC05, BLG20, LZL^{+}24, VMG22]. Aging [Kle82]. AIC [IYW14, LC11a]. Aid [Sib80]. Akaike [Sak19, Yu16]. al [PC99]. al. [KHSS12b]. Algebraic [Rap03, Rap12]. Algebras [GH89]. Algorithm [ANO96k, Bie07, CV02, FM89, GH14a, GWT00, GJW08, HT08, Hol80b, Hol81b, LWY97, SR03, SLB06, VR08, FHSZ19, Gås03, Mar99]. Algorithms [ACR16, Bjö10, CC12, CW16, Doo18, HJKQ18, HS17, JR07, JN16a, JN16b, MP84, NX17, Oja16, Ron16, SV10, Zwa16]. Allergy [BMG82]. Allocation [Laa75]. Allowing [SG78]. Almost [Fer91]. Alternative [AMP01, DE06, SBR98]. Alternatives [Ban05, Ber81, GS19a, HK99, JM93, LW12, Wył16]. amateur [GG23, VV24]. ambiguous [Gre23a]. Among [CO07, NM24]. Analogues [IS99]. Analyses [BQ09, Gar82]. Analysing [Bjö10]. Analysis [ABKT80, Aal87b, Aal95, And82, AK07, AL98, AL99, ABN12, Asm00, BF02, BO11, Bøl83b, øBFHB07, BGL13, Bro87b, BW04, Car07, Cav16, CHWY05, CLSZ16, CM17a, Chr74, CC12, CGL^{+}81, Cro91, DL89, DH07, DEL92, DSWH09, FHTT16, FMS15, GN95, GK00, HVV14, HVA00, HJO15, Hel98, HT14, HM02, HJKQ18, HKD02, Hor85, HCS15, HL00, JVA11, JLP09, Jen87a, JLY06, Joh97, JR76, JAL^{+}81, KKC17, KSM87, Kou79, Kou84, Kre87, KC11, Laa88, Lan07, LC00a, LB88, LHML16, LZZ14, LMH14, LB80, LQ17, Mad76, MS01, Mol94, MS94, MT14, MW97, MNS07, Mur95, NGMS94, NM87, OKK^{+}00, OR94, PT92, PS89, PBBM12, PS10, QL15, Ros89, Rov15, ŠBD05, SÅS07, Sch79, See93, ŞM05, Sib80, SMS12, Spj74, SLCR14, SS98, Tju82, TSH91, TCC^{+}95, Tra11]. Analysis [Van07a, VHK11, Van07b, Von96, WD98, WLS15, WW01, fWZY16, Xue10, YL96, ZHF03, APM19, BG16, BMXT20, Cav23, CFR19, DZ21, FGLT23, HFP24, HLP23, ICM19, JV24, LGL19, LCZW22, MP22, NR23, SLCN19, WCJ18, ZLK21, ZCL22, ZXLL23, CM17b]. Analytic [Asm00, GHH95, Hoe76, Ber23]. Ancestor [DW16]. anchoring [CZT20]. Ancillarity [BNK74, FG96, Joh77, Ped75a, Ped75b]. Ancillary [BNC91]. angular [Hel23]. Anisotropic [MT14, CDO24, MMO23]. Annihilation [SV10]. ANOVA [LCZ14, PFJGE15, ZL14]. ANOVA-Type [LCZ14, PFJGE15]. anterior [AHP^{+}18]. Antibody [EGM^{+}03]. Any [VR08]. aoristic [vLM23]. Applicability [Var79]. Applicable [WL18]. Application [Ano83i, BM16, øBFHB07, BW07, BW08, BKO11, CSW79, CB84, CYM93, DBS10, FS10, FMHB16, FM89, GM08a, GM83, HS87, Höp87, JN19, Kol97, KHT14, LYZ15, LB88, LLY18, MG95, OBL18, Ohl86, Rov02, San14, Sas92, See93, Van13, CCV23, CL21, GRS22, KHSJ19, KL22, LM23, RMG19, YA20, ZXLL23]. Applications [AK07, AHJ15, ABN12, BH99, BS00, Ber74, Ber75, BG80, BAR^{+}85, BB11, BO11, BF03, Bor84a, Bor84b, CM01, CP98, DPV06, Doo18, DGGM16, ES91, FMS11, GCL87, Gui80, Gui86, GJ83, HM99, Hor85, HYWC18, ICM19, Ist96, KP77b, LRT^{+}87, Mur95, NH15, NX17, Nor86, Ped75a, Ped75b, SV10, SMV05, Sjö00, SLCR14, STZ01, WHF98, WR93, ZLY14, BKB23, DHH24, HFP24, HYLT23, JTT21, LYW22, SZ20, Ano98e]. Applied [GS76, Hok75, HS17, MT03, Sch02, AHP^{+}18]. Approach [Aal12, AK07, AH84, AFV14, Ber79b, BN13, Cer17, Dem17, FV06, GH00, GH08, Gri80, Gua07, Hel98, Hou12, HS04, KHSS12a, KHSS12b, KF07, Kni98, KKMP18, KV98, Kou79, Kur18, Lai79, Lai80, LR06, LBNE^{+}78, Lin77, MSP01, MS98a, MW97, NGAS92, OSG08, Par01, Ped95, PS13, Rov05, SN13, SJ93, Sun83, SJS08, TGM17, Toc01, WC12, dCJV82, AHWKP19, Ber23, CLR19, DEV20a, DQR21, FHTT18, GMvdM20, LKT^{+}23, LW23, LCZW22, XT20, ZHW19, ZLZZ21]. Approaches [DY17, OB16, WL18, WC20]. Approaching [BS00]. Appropriate [Häg07]. Approximate [AL98, AL99, Bac11, BNS05, DH23, Dem17, Dia23, EMR09, Kol97, LPPW22, MAR11, Uch04, VMG22, Wan00, WW01, VHF20]. Approximately [DS94]. Approximating [RS94]. Approximation [Che09, CD01, CB84, Eri78, GH00, GH08, HJR06, JP06, KV98, KRV07, MR14, MB91, MZ11, Sør01, HNRT22, PL23, ZD24]. Approximations [BJ89, BNK99, BH84, EGPS98, HM99, IKL94, JKR02, Kün83, RLOS18, Sør03, WW11, BS21, WC21]. AR-ARCH [HNNS19]. AR-Processes [AOH00]. Arbitrarily [Jen87a]. Arbitrary [LQ17, BDS22]. ARCH [BSC24, HNNS19, Mil85]. Archimedean [HS12]. Area [ADL15, DSD^{+}14, GM08a, MRS14, MSZ16, PBHMC09, STMC16, STK17, SKO17, TDR09, ZZLZ16, CK23, DR18, ELLV^{+}22, JN19, SKR19]. Area-interaction [PBHMC09]. Area-level [SKO17]. Areas [PBB06]. Arguments [Edw78]. ARIMA [Nie84]. Arising [SB85]. Arjas [Cor23, KPS23]. ARMA [CZL24, KS88, Nie83, Ter77b, Wu13]. Armitage [SW87]. Arrangements [Ber79b]. Article [ML75, BG14b]. Artificial [Wan06]. ask [LL20]. Aspects [Eri04, GN95, KP02]. Assessing [Zha95]. Assessment [Cer17, La 08, SA15]. Assessments [Pap08]. Assessments* [GIA02]. Assigned [Efr16]. assignments [LYW22]. Assisted [DGGM16, XMW15, MT19, WZ22]. Associated [Höp87, Ste91]. Association [LAE^{+}89, Mei06, QZP12, SWS06, VOG11, ZLY14, ZXL^{+}18, HBD^{+}20, ZXLL23]. assumption [DLS^{+}24, OH21]. Assumptions [GPM04, ZV21]. asymmetric [ADMP19]. Asymmetries [BPW14]. Asymmetry [CJ08, Dok75]. Asymptotic [Aab83, AB85, Ahm81, AR80, AALM17, AH87, AOH00, Awa81, BL83, BIP14, BDW16, BP05, CZL24, CGL14, CM84a, CY17a, Cha84a, CP98, CDG16, CYM93, CM15, CV22, DGSL02, Eng80, Gar82, GA86, Gui80, Ham88, Hjo86a, Hjo86b, Hol75a, Hol81a, Höp90, HL00, Irl90, JTT21, JSG86, Jen79, Jen87b, Jen89, Jen93a, Joh82, JN16a, JMT94, KR01, Kle99, KR20, Laa88, Lin00, LQ17, Lus94, MSR16, MG95, McG88, Miu81, Næs82, Nie97b, Nor80, Ohl86, Oja16, Pal09, PL23, PC99, Ran75, Ron16, Ros74a, Ros74b, Sai83, Sam89, Sch75, Sch81, SB90, SW76, Sve90a, Tak23, Tho77, TZ95, VU05, Vie99, Wan99, Wre78, WW11, YK20, Zet88, Zwa16, vP92, BKT20, CM20a, CCWZ19, GM23, ZGZ22, JN16b, ACR16, CW16]. Asymptotically [Ber82, FGD12, GG13, KKP08, Pfa93, Ryd95, PRV21]. Asymptotics [BE10, HB06, SMB14, Sko01, TT17, VBJ97]. Asynchronous [Bib11, MV20]. asynchronously [Koi14]. Attraction [Mar98, Sei24]. attributes [LZL^{+}24]. AUC [BB15]. Augmentation [EGM^{+}03, Far09]. Autocorrelated [LB80]. Autocovariance [TGM17]. Automated [MT19]. Automatic [BRH83, FK98, Wył16, Kle24]. Automatized [LN95]. Autopsy [GN98]. Autoregression [SN13]. Autoregressions [McK87, LM23]. Autoregressive [BIP14, BCS00, Cav16, DW97, KL14, Law82, LC00a, LG09, Lus94, PCW02, Ris80, Ris81, Rob78, SJ94, Wal00, CM20b, Kar20, KP21]. Auxiliary [AHJ15, ADL15, FMS15, HW98, LDW06, Sto11, Sei24]. Availabilities [Nat93]. Availability [BL94]. Average [Awa81, BDY85, HP00, LP01, OB16, SW04, Vel12, WLS15, KP21, YZ23]. Averaged [KWA16, Fan19, MS24]. averages [GRS22, LP22, PPS21]. averaging [GZZM23, GH21, LW23, XWH14].

B [ACMLM03, KSR13]. B-Splines

[ACMLM03, KSR13]. BA [HVV14]. Back [Ano74a, Ano74b, Ano74c, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b]. Back [Ano96c, Ano96d, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05b, Ano05c, Ano05d, Ano05e, Ano06a, Ano06b, Ano06c, Ano06d, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d]. backcross [LLLP20]. Backfitting [WL04]. Backward [Gup76, OS24]. Bahadur [Kni98, Tor88, Xia94]. Balance [PT92, Ano23, JJCYG21]. Balanced [GM18]. balancing [LRT23]. Balayages [Lyn88]. Bands [BBL87, BL90, CGL14, FZ00, HST74, LV02, LAKZ12, Nai82, SU92, SR11, SN24]. Bandwidth [CMN08, CY17a, Dab92, DH05, EL96, GM98, GM08b, Gua07, Haz96, JK92, Pre03, SHD94, WG96]. Barrier [ML86]. Bartlett [MT03, Møl86]. Based [AJ78, ACFS83a, AK07, BBQ18, BZF08, BM03, BB11, BN15, BDY85, BL90, CL05, CSW79, Cey10, CQ02, Che09, CJ08, DLS96, DS03b, DSS13, Eks13, Far15, FZ06, GN98, Gås16, Gär03, GQR06, Guo11, GS02, HKJ11, Hoe78, Imo15, JM16, Jen87a, JLY06, Joh08, KR01, KB04, Kle16, Kou85, LO16, LQZR09, LC11a, LA16, MS98a, MM93, Nor80, OFFL12, PTF09, Pap00, PFJGE15, Ped95, PM03, PdT91, Qin98, RV04, RD10, SSD15, STH^{+}78, SHD94, Sti82, SR01, Sve90a, SV05, TGM17, Toc01, Uch04, VHK11, Waa06, Wan87, Wan99, WR02, WZ10, ZHH10, vE92, BQ22, BNS03, BSO22, BKB23, CDQ20, CXW23, DQR21, EK22, FGY23, FMS11, GH16, HNRT22, JTT21, ýJRNMJ13, Kar20, KS22, KK23, KZ17, LZC23, LR06, LLS^{+}22, LZL^{+}24, Lu21, MM24]. based [MPV19, OHN21, ORL20, PP16, PD22, PG13, RAQ21, Sak19, dRSS22, TKU23b, VHF20, WC20, WWW15, WC21, XZ09, XWH14, YK16, ZHS22, ZLZZ21, vdWBM19, Hoe78]. Baseline [CC98, LN13b, LM18]. Bases [FM89, Rap03]. Basic [LY08, SN88]. basis [TB22]. Basu [KP77b]. Bathtub [Aar85, Xie89]. Bathtub-Shaped [Xie89]. Bayes [Lai80, CL12, CLP17, CG99, DGSL02, Edw78, EG02, GM08a, Hjo86a, Hjo86b, Joh08, KSSR21, Lai79, Lai83, OKK^{+}00, SS18, Sti82, SKO17, SKR19, TDR09, vH80, vdWBM19]. Bayesian [AAA04, AO11, AGR13, ADL15, AT15, Ave86, BBBS19, BR97, BM03, BR23, BLM20, BGL13, BO02, CLP18, CHWY05, CO07, DGN07, DLS96, DW02, DB03, DH07, DCIK14, DO05, DL06, EGM^{+}03, EMR09, Erh08, EPM15, Eva16, FT16, FKA04, GN09, Gås16, GK00, GR05, GK03, GSK06, GJ16, GWH11, GH21, GMvdM20, HVA00, HBH17, HA98, HJKQ18, JKM19, JvdMP22, KKC17, KKMP18, Kol81, KK23, KK09, KA06, KMG21, La 08, LC00a, LV13, LBNE^{+}78, Lo81, MSP01, MAR11, MRM09, MR12, MBR03, MW97, NE87, NG23, NBY08, NJG18, NM87, OHIS24, PEK22, PKH17, Rov02, SA11, SA15, SGR11, Scr07, SG15, SS09, STK17, SO97, Thy75, Uta17, VMG22, VW15, WD98]. Be [Pfa93]. Before [Sve86]. Before-and-after [Sve86]. Behavior [BIP14, Bér94, Gui80, Lus94]. Behaviour [BKS76, Bie07, FR00, GJ83, JMT94, Pal09, Rom04, Sch79, Væt79]. Bell [RVG15]. Bemoulli [Ste88]. Benchmark [LPB15]. Benchmarked [CK23, KHT14]. Benefits [PS13]. Bent [DH16, Oja16, Zwa16, Doo16]. Bernoulli [Fra77b, Lin78b]. Bernstein [BC99, CHWY05, Pet99]. Berry [BBG97, HJS90]. Besov [Koo99]. Bessel [Eie83]. Best [AM84, Bon79, Cac77, Nor75, ELLV^{+}22, Min81]. Beta [BC99, DLS96, NBW02, APM19, HMP22]. Beta-Bernstein [BC99]. beta-Stacy [APM19]. Better [BEK83, MS86]. Betti [KH22]. Between [DDL14, PW06, VB99, ABKT80, BMG82, CLR19, CJGPL07, DNCZ21, DDK04, HKD02, Jen79, LL06, LZL^{+}24, ML74, Pon86, Tju82, vL18]. between-series [CLR19]. between-subject [LZL^{+}24]. Beyond [GGS20, CLP18]. Bi [LMB09]. Bi-directed [LMB09]. Bias [AG85, And79, AOH00, BS10, DNL10, ES91, IYW14, JK04, KT95, Lun00, MR10, Nie98, NT01, NGZ18, SS02, Sak19, Seg02, SBH03, Stu94, WC21, YF12, DHH24]. Bias-corrected [IYW14]. Bias-reduced [Sak19]. Biased [GK86, Tre83, BCCH19, CLSZ16, HCS15, QQZ16, RFK22]. Biases [BKW10]. Bidimensional [FS10]. big [Kop23]. Bilateral [CM84a]. Bilinear [GP89a, GP89b]. Binary [Amu74, Amu76, AT15, BBG06, Got94, Kab78, Nor81, Pan02, Pap08, QZP12, SV10, SW93a, SBB05, TS91, XMW15, DM19, LT21]. Binning [PS99]. Binomial [AL79, AL81, DM83, HH16, Lai79, Lai80, McK87, Thu14, Vai91, BSO22, GME24]. Bioassays [SMB14]. Biological [FS10]. biomarkers [MP22]. Birch [Lan13]. Birnbaum [BNHH95]. Birth [BKS76, Höp87, Höp90, HL99, MS94, Ros77, Ros78, SM12]. Birth-and-Death [Höp87, Höp90, MS94]. Birth-Death [Ros77]. Bivariate [BNP79, Gho06, GL15, KY12, KS99, Llo88, MH10a, Mes22, PdT91, Que12, SWS06, WCXS15, ELLV^{+}22, HV22, WCJ18]. Blackwell [Tor88]. Block [AFV14, BDL^{+}17, De 06, KHR02, Mej85, RL06, HFS23]. block-missing [HFS23]. Block-threshold-adapted [AFV14]. Blockwise [BDL^{+}17, BK95, Efr05]. BLUE [Bon76]. BMT [SMZ11]. Bole [Huc11]. Bone [AK07]. Bonferroni [Bøl88]. Boolean [HS17, Mol94]. Boosting [Vid21]. Bootstrap [BDP13, BK95, CMN08, Che15b, CL01b, FVV10, FSGMM16, GK13, HJS90, HL08, Hol93, LY03, Mam92, Neu09, PTF09, Præ95, PW10, SB90, BHLP19]. Bootstrap- [HL08]. bootstrapped [Kop23]. Bootstrapping [FKA04, HW98, JN19, LB94]. Both [MRS14, STK17, YZ07, HKŠ22, LLLP20]. Bound [BBG97, Lin94]. Boundaries [BDL^{+}17, GJ03, JP06]. Boundary [Mül93, NT01, Yao96, STM22]. Bounded [Wal97, PPS21]. Bounding [GK91]. Bounds [BDY85, Bøl88, CL01b, Dal77, Efr05, Ess75, KT95, Nat93, Van11, QB23]. Box [Lau76, LL12, LT77]. Brain [ýJRNMJ13]. Branching [Bro87a, CM84a, HHL02, KL78, Lin76, Ove98]. Brazil [øBFHB07]. Break [JFO23]. Breakdown [DT05]. Bregman [Ano23, JJCYG21, Zha08]. Bridge [See93, See96, YL96]. Brief [Ano74d, Ano74e, Ano74f, Ano75e, Ano75f, Ano75g, Ano76e, Ano76f, Ano76g, Ano76h, Ano77e, Ano77f, Ano77g, Ano77h, Ano78e, Ano78f, Ano78g, Ano78h, Ano79e, Ano79f, Ano79g, Ano79h, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano82e, Ano82f, Ano82g, Ano82h, Ano83e, Ano83f, Ano83g, Ano83h, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h]. Brown [KL22]. Brownian [LMT14, LM24]. buckets [GHD20]. Buckley [Yu11]. Building [KS08]. Bumpy [ZL10]. bundle [LGL19]. bundles [LGL19]. Burr [CK97]. Busy [Hok76]. butterfly [HWC20]. Byrne [Dry14, DSH14, Per14].

Calanoida [Sch79]. Calculating [IYW14, Sas92]. calibrated [DZ21]. Calibration [Bel03, GMMT06, LdM80, OS96, Sun96, SBM^{+}99]. callbacks [GLQ18]. can [AF07]. Cancer [LMH14]. Canonical [PS92]. Cantelli [SS80]. Capability [VK95]. Capture [BL08, Ber74, Hol80a]. care [LYW22]. Caries [HVA00]. Carlo [BG14b, Dry14, Ken14, SLS14, Sim14, BG13, BG14c, CDMR02, GH14a, GH16, GHD20, JR07, JSDT11, Kle24, LRSS23, LET22, NH15, PWN22, Sak19, SPR^{+}13, SW75, SW76, VKY^{+}14, VHF20]. CARMA [Fas16]. Case [BN85a, BNM^{+}06, BO99, Bro87b, Cer17, Guo11, KKP08, KKMP18, KA06, Lan07, Lau76, LT77, MNS07, SÅS07, SM04a, SY00, Sko81a, WL18, dMR88, ABY22, BKN23, KXZA20, MV20, NG22, ZZLC21, ZMH^{+}24]. Case-Cohort [SÅS07, BKN23, KXZA20, ZZLC21]. Case-Control [BO99, Lan07, BKN23, NG22, ZMH^{+}24]. Case-Study [Lau76]. cases [ORL20]. Catchability [NC92]. Categorical [LMB09]. Category [SC06]. Cauchy [DF90, Gro21]. Causal [AF07, AGR^{+}18, AP04, GL02, Kar15, KH16, MP14, Pal04, RVG15, Rub04a, DM19, KP21, WHR22]. Causality [Aal04, Lau04]. Causation [KC11]. Cause [DDK04, MW93]. Cause-of-Death [MW93]. CC [CH22]. Cell [And77a]. Cells [ADZ15, BBP21]. Censored [AJ78, AHK91, AG90, AH84, Bor84a, Bor84b, BJMP14, BCG08, Dab87, DD88, DBS10, DLP08, DH12, EV08, GSYB05, Gho06, Hua13, HC10, Jan91, JM01, Joh17, Jon01a, Kim03, LWY97, LV02, LO16, LHWS18, LdUád15, LS96, MW08, Ols96, PFV06, Pon86, PdT91, QJ01, Ren03, Sam89, SY00, SV04, SW93b, SZZ05, SWS06, SLB06, TEV15, VBJ97, VJ01, Wan87, Wan95, Wan99, WLT15, YLW00, YWK06, Yu11, ZHH10, BCCH19, BC15, CLSZ16, CZT20, CAVGM21, DEV20a, DR22, HCS15, Par20, WCY22, ZVD22, ZCL22]. Censoring [ABKT80, AGM00, BB11, DBS10, DSWH09, GL15, GG01, Gui04, HESZ16, HW95, JFKC05, KKP99, LZ08, MS98b, DT20, KXZA20, OH21, SLCN19, XLY20]. Censorship [CH82, Stu96, Zha96, ZYT02]. Census [Ber76, DGGM16]. Central [AM84, BMP19, BW04, FL11, Hel82, LLY18, LdUád15, Mur95, SW18, SZ95, BW19, KH22]. centred [NX17]. Certain [Aab83, Ber77a, Ber81, Bøl83a, Gil86, GJ83, Höp90, Sve86]. Certainly [Jag77]. CG [DE04]. CG-Regression [DE04]. Chain [Aal87a, AMP97, AMP01, DE06, Fry90, Got94, HV08, HJR06, JR07, Lin77, Lin78c, NH15, Ran75, Ran78, Rov05, SPR^{+}13, Tho81, LGL19, LRSS23, LLS^{+}22, PWN22, VHF20, Wol24]. Chain* [Häg07]. chain-of-bundles [LGL19]. Chains [AJ78, Edw80, FW03, Höp87, JXCK14, Jen89, Nic14, PKH17, RR01, Ste91, SPK23]. Challenges [Sch02]. Change [AGM00, CM20c, DRT13, GS02, HJ04, KL14, Lee97, LHNN03, Neu97, NV09, NDH^{+}21, SN13, Swe88, Var76, VW09, Yao96, ADMP19, BS21, CWZ21, DEV20b, ELY22, KS22, Mes22, MN21, Tak23, TKU23a, WHZ20, ZLK21]. change-plane [Tak23]. Change-Point [AGM00, DRT13, GS02, Neu97, NV09, NDH^{+}21, Yao96, CM20c, CWZ21, WHZ20]. Changed [FL11]. Changepoint [HM02, Uta17]. Changes [HS95, HHM17, LP01, BBP21, CN16, HMR21, Mes22]. Channel [CYM93]. Chao [Böh10]. Chaos [TCC^{+}95]. Chaotic [LS98, PPS21]. chaotic-driven [PPS21]. Character [Hol75a]. Characteristic [BBQ18, HC10, Kou85, PFJGE15, PG13, SB00, DEH21, HJG21]. Characteristics [AP07, Gui77, SS00]. Characterization [Mac93, Rov05, Sei24]. Characterizations [Cro00, Gup76]. Chart [EPM15]. Charts [GK13]. Checking [LS96, PR07, Ris81, YZ12]. Checks [BDP12, DH12, SZ02]. Chernoff [Bly93]. Chi [ADZ15, BR81, Hoe76, LL96, Min79, Min81]. Chi-Square [ADZ15, BR81, Hoe76, LL96, Min79, Min81]. Choice [CSW79, Cer17, CJGPL07, LC11a, Rud82, SV76, LL20]. Choices [SGR11]. Choosing [DS03a, VHK11]. Chronic [SA15]. Chronical [dCJV82]. Circular [DPT13]. claim [YC22]. Class [Azz85, BNLSV14, CCH98, DC00, EVP15, FPW11, FS08, HL00, JK92, LHWS18, LB98, MG04, MR14, NC15, NYR18, PKH17, PS89, RD10, TF12, WF79, Yu16, vE92, vP92, CN16, CWZ21]. Classes [Cha84b, Dam75, GS76, Hol75a, Jac01, LQ17, NC92, Rov05]. Classical [Deg96, OS96, Wil77]. Classification [BKM18, BCCA11, BJFG15, Bro80, Mül05b, Zha08, vH80, DDM20, DDBEMT24, FGY23]. Classifications [Nor77]. Classifier [Swe88, DDBEMT24, WCSftADNI23]. Classifiers [GC05]. Clayton [GS99, PM03]. Climate [BKW10, OBL18]. Climatic [BCS00]. Clinical [CV14, SBB05]. clipped [SH21]. Closed [BL08, NHS^{+}19]. CLT [BW04, Ohl86]. Cluster [QMP15, CFR19, MS24]. cluster-correlated [MS24]. Cluster-Specific [QMP15]. Clustered [EMS15, HHVA03, HZZ07, LDY16, MGSFB08, Van07b, Xue10, CHI23, LLXH19, NJG18]. Clustering [ACMLM03, HS87, SM12, Ter14, VS21, AKP22, GH21, LPR23, MM24]. Co [KR01, Kur18, KC11, Koi14]. Co-integrated [KR01]. Co-Jumps [Kur18]. Co-variate [KC11]. co-volatility [Koi14]. Coarsened [CJGPL07, CGP07]. Cochran [Ber16, Ohl86]. Coefficient [AC99, CY17b, FZ00, GCLP92, GCJ94, GG13, HS04, HYWC18, Jac00, Joh82, Man09, Mur93, Nor77, Rob78, ŞM05, WL04, XL10, MP21, ZKP^{+}24]. Coefficients [AH78, AALM17, CS03, KKC17, SW05, SSZ09, Vel12, ZHL15, CFL24]. COGARCH [BN15, dRSS22]. Coherence [Bro80]. Cohort [GL07, Lan07, OKK^{+}00, SÅS07, SM04a, BKN23, KXZA20, KA06, ZZLC21]. Coiflets [Ant96]. Coincidences [BP89]. Cointegrated [Doo18]. Collapsibility [DE04, GL02, KK06, LG13, Vel12, XG09]. Collapsing [NX17]. Collective [LHML16]. Column [See96]. Combination [DP04, PT92, Vid21]. Combinations [Ano83i, BNB93, BJ78, GM83, Jon78]. Combinatorial [Eri04]. Combined [CSW79, Par20, VKY^{+}14]. Combining [CH22, Han16, SMB14, Swe88]. Comment [Dry14]. Commentary [Edw78, VV24]. Comments [BNHH95, HOF^{+}94, KHSS12b, Lav23, Vos23, Gre23a]. Committee [Hol80a]. common [MV20]. Communication [GW24]. Communication-efficient [GW24]. Community [CO07]. Compact [BC99]. Comparative [Böh10, Eri04]. Compared [FWW77]. Comparing [And83, DWV11, LB98, Lon12, Nai82]. Comparison [Aab83, DW02, Hjo88, IKL94, Kor82, Kou79, MC97, NS06, PFV06, SG78, SG04, Sør98, SA80, ST81, SR01, Sve90b, TJL^{+}76, HD22]. Comparisons [BM15, Kou76, Kou84, OS96]. Compatible [AP07]. Compensator [Ave85, Nor86].

Competing [BDP13, CHW^{+}07, Cro91, Cro00, DS09, DL89, DSWH09, Gar82, GK00, HESZ16, KS01, LB98, WCXS15, APM19, JH17, OPP18, YY15]. Competing-Risks [CHW^{+}07]. Competition [Sko86]. Competitions [See93, See96]. Complementary [JSW91]. Complete [MS91]. Completely [GL02]. Completeness [LR76]. Complex [KM95a, DR18, DQR21, GK21, Gre23a]. Complexity [Bro80]. Complicated [AH84]. Component [And90, BDV06, CFJP07, Chr74, CPWZ13, GN98, HT14, HT17, MT03, QL15, FB20, FGLT23, JLRT19, YA20]. Components [Car07, Fra78, HST12, LZ10, LCZ14, Lin88, MW12, NH93, Nat85, NS06, Hua24, LT21, vL18]. Composable [Did07]. Composite [Bon82, DB03, Gua07, HC10, LMH14, Par01, QQZ16, TWL18, BSO22, KL22, XWH14]. composites [LGL19]. Compositional [BH14, FHTT16, FHTT18]. Compositions [BLBEO92]. Compound [CDGCK15, Var79]. Computational [KP02]. Computationally [HBD^{+}20]. Computations [LDW06]. Computed [CR13]. computer [BPR22]. Computing [Agr93, GJW08, VKY^{+}14]. Concave [FM89, THSS09, Che15a, RSTU21]. Concentration [BB15]. Concept [Ber79a]. concerning [Kri95, Lai83]. Concordance [MP14, WC20, FGY23]. Concordance-based [WC20, FGY23]. Condition [DL01, KJH16]. Conditional [Agr93, ADGP14, BL83, BIP14, Bar76, Bér94, BCC17, BM16, BR17, Bon82, BTL06, Bon10, BT13, Car07, CD03, Dab92, DB03, DGSL02, DWV11, DDK04, EGG14, EBGG18, FT16, FM22, GK13, GCJL03, JM01, Jen78, KN12, KH99, Kol97, Kou79, Kre87, LGP11, LET22, LG13, NL16, OBL18, OS97, PS92, SK20, Swe83, VOG11, WC12, Yao96, YY15, Yu16, GGS20, GSUC22, KHBK22, LLYC22, LPYZ24, PS20, PPS21, ZGZ22, ZLZZ21].

Conditionality [BN84, BN85b]. Conditionally [CV01, PF08]. Conditioning [Ner98, SO97]. Conditions [DH08, GL02, Ran78]. Confidence [Ano83i, BW05, BN85a, BL94, Bér94, BBL87, BDY85, BL90, BBdW20, CGL14, CQ02, CFJP07, CL01b, CK06, DGSL02, DSD^{+}14, EW94, FZ00, FB20, GM08a, GM83, GH18, Gui86, GG01, Gui04, HL08, Hol93, HST74, JS12, KWA16, LL90, LV02, LAKZ12, LHW^{+}16, LA16, MW12, MV87, Nai82, PWY97, RGS03, SN24, SH02, SU92, Thu14, Tri03, VM15, Xue09, GAC23, GM23, GME24, KK19, PLL23]. Conflict [DGN07, GN09, Gås16]. Confounding [GL02, Lue15, PLHS17, Van07b]. Conjecture [Kri95]. Conjugacy [JLP06]. Conjugate [CV01, Dam75, GMS93, GPM04, Mac93, Pic00]. Connected [Fra78]. Connecting [Gre23a]. Connection [Aal76, Tju82]. Connections [But98, DSH14]. Consistency [CD96, Che15a, CL01a, Che15b, Eks01, Kim03, Ran78, SY00, SV04, Tan09, Ter14, Wan95, YLW00, YWK06, TKLM23, YA20]. Consistent [DDM20, FSGMM16, JFKC05, KM00, KK23, KSN95, KM95b, LWY97, Pen95, Ryd95, SW04, TvdM96, Van01, Wij95, YLW00, ZSJT24, BG98]. Constancy [CDMGR06]. Constant [Aar85, GNPM07, Xie89]. Constrained [LAKZ12, LN13b, VHK11, Wan00, ZHL17, CFL24]. Constraints [BBG06, FM89, LC11b]. constrictivity [NHS^{+}19]. Construct [HJR06]. Constructed [GL02]. Constructing [DS94, PCW02, ZBS20]. Construction [CL01b, EB08, vHV85]. Contact [GWH11, HGB96]. contaminated [SW19]. content [LKT^{+}23]. Contents [Ano97a, Ano05j, Ano07k, Ano11j, Ano10k]. Context [GGG13, Høj04]. Contextual [PNC17]. Contingency [And74, Cey10, Høj04, Jen78, KK06, Kre87, Kuh04, Mad76, Ped75a, Rap03, Rap12, Sun75, LET20]. Continuity [DR00, HL99]. Continuous [AR80, BB10, BG16, BJ93, CW99, DPV06, GM18, GJW12, HM99, Hel82, Höp87, Jon91, KS94, LPB15, LM23, MW08, MG98, Ove98, SH96, Slu92, Sør01, Sti82, TC05, Vet12, Win13, JvdMP22, KP21, NHMW22, dRSHK19]. Continuous-Time [BB10, DPV06, Höp87, Slu92, Sør01, TC05, LM23, KP21, dRSHK19]. Continuum [BS99]. Contour [RS94]. Contrast [AG20, BL17, Lud04]. Contribute [AF07]. Contribution [Ken14, SLS14, Sim14]. contributions [MR24]. Contributors [FH04]. Control [BRH83, BO99, CM01, CFS95, Far07, Far09, FGD12, GK13, Guo11, HT08, Lan07, Mei06, BNM^{+}06, BKN23, DBNR20, HBD^{+}20, NG22, VS21, ZMH^{+}24]. Controlled [Bel03, Efr08, Van11]. Controlling [BS16]. Controls [Nor90]. controversies [GG23, VV24]. Convergence [BDP13, CDMR02, CV02, Dab96, Fer91, GR05, GR10, HV05, HJKQ18, Hol80b, HS98, JR07, KST95, MT02, SJKS22, Scr07, SJ94, Stu96, Swe83, VR08, Wan90, Yuk92, vH80, vZ03, vdV94, HMP22, LM24, Hol81b]. Convergent [VR08]. conversation [BG24a, Cor23]. Converse [Tor88]. Convex [BD07, Blæ78, JP06, PW06, vEvZ96, GH23]. Convex-Hull [JP06]. Convexity [Lyn88, PS83]. Convolution [BDP12, RD10, ST12]. Convolution-Based [RD10]. Convolution-Type [BDP12]. Cooking [Häg07]. coordinate [FHTT18]. Coordinates [FHTT16]. Copepoda [Sch79]. Copula [BDS22, DSS13, GQR06, GOV15, GL15, GH14b, JFKC05, KSR13, Son00, VOG11, BQ22, KG18, LLYC22, OHN21, YC22]. Copula-Based [DSS13]. Copulas [BBQ18, BNL07, BQ09, HS12, KY12, KHH19]. Core [BF03]. Corners [HQR08]. correct [GME24]. Corrected [ABN12, Aug04, CY17a, HWH15, IYW14]. Correcting [MR10]. Correction [AL81, AL99, Ano83i, BR03, BN85b, Bor84a, BW08, CM17b, Dok82, DGGM16, GH08, GHU03, GP89a, Guo11, Hal01, Hjo86b, Hoe78, Hol81b, Lai80, Nie98, NT01, NGZ18, PC99, Sch81, YF12, SW19]. Corrections [LHHF13]. Correlated [Boe10, JM16, Kos99, Pan02, Pap08, Puk82, Ris80, Ris81, YZ07, ADN21, MS24, VW19]. Correlation [Abt99, Bly93, Gua07, LHHF13, PLKP06, RGS03, Sai83, STMC16, ŞM05, Vet12, XY15, CRS24, CLR19, DNCZ21, GBB^{+}24, MWW24]. Correlations [BNP92, Bøl82, HV22]. Corresponding [Eie83, Oja99]. Corrigenda [Ano96e]. Corrigendum [Ano10f, Ano24, AVA22]. corruption [DBJ^{+}22]. Cosine [Eub00]. Cosmetic [SW84]. Cost [MC97]. Cost-Efficiencies [MC97]. Count [Boe10, PK18, SJS08, THSS09, WLS15, CWZ21, Lu21, YLZ^{+}19]. Counter [BO99]. Counter-Matched [BO99]. Counterexample [Kri95]. Counting [ABH^{+}85, Ave85, Ave86, Bor84a, Bor84b, CCH98, CYL11, CC12, CGP07, Gré93, GS02, Hjo86a, Hjo86b, NGAS92, Sch94, SJ93, STZ01, Sve90a]. Counts [DM83, Eri78, Fok01, HSW03, HHM17]. Coupling [HN99, Tho95]. Covariables [Stu96]. Covariance [Ahm17, Bib11, BZ82, CR13, Eri96, Erl81, FSHK13, Hou86, JAL^{+}81, Jun11, LC00a, LLY18, MM24, MWY15, MNS07, Nor75, PSS10, PT92, RD10, SG04, ZLL^{+}16, BMP19, BS21, CLP^{+}19, JFO23, Kle24, SK20, SP22, ZL22, ZXLL23]. Covariance-based [MM24]. Covariances [Møl86, RR95, ST81, ZSJT24]. Covariate [CV14, CLP17, GOV15, Grø97, HL02, JW10, Jon91, KHL98, LR06, Mar99, MSSM02, MS09, ŞM05, THF18, Ano23, JJCYG21, KXZA20, LJZ^{+}18, VD18, WZ22, ZSJT24]. covariate-adaptive [ZSJT24].

Covariate-Adjusted [CV14, CLP17, KXZA20, VD18]. covariate-assisted [WZ22]. Covariate-Varying [THF18]. Covariates [AH84, BIPV13, BHC88, DBD18, DFG00, FMS15, HESZ16, HCS15, HS98, LAKZ12, MSP01, NYR18, SMS12, TDR09, WL04, YZ07, Yu11, ZIS09, DBJ^{+}22, LLXH19, SW19, XNL23]. Covariates-adjusted [HESZ16]. Cover [KSN95]. Coverage [DGSL02, MW12, Thu14, KK19]. Coverage-adjusted [Thu14]. Covolatility [BR14]. Cox [BW08, GH87, AFL10, Aug04, Bed93, BKT20, BC15, BW07, BHC88, BM01b, CS03, CC98, CMW17, DP16, HL02, JGW13, KYZC21, LL12, LO16, LLXH19, LN13b, LM18, LJZ^{+}18, MMO23, Mar99, MM93, MS98b, MSW98, MDA10, MT14, MR24, Mur93, Næs82, NC18, Nie97a, Nie99, PR07, PdT87, PV00, Sas92, SZS02, SM04a, SSZ09, Tak23, Vai91, Wan08, fWZY16, XNL23, ZHH10, ZHS22]. Cox-Aalen [BC15]. Cramér [BB11, ELY22]. Creation [SV10]. Creation/Annihilation [SV10]. Credible [SR11]. Cressie [OT09]. Criminology [ABN12]. Crisis [Cro91]. Criteria [GH14b, LC11a, SW93a, TM86, Yu16, GME24, KHH19, Sak19, YLGL20]. Criterion [BL08, GC18, Imo15, LPPS82, LVV09, NW06, ST10, Tra11, VW15, YF12, KK23, XWH14]. Critical [FR00, Ner77, VKY^{+}14]. Cross [DRM96, DH05, Gho06, Gré93, Gua07, Jun11, Sai83, Van07b, XZ09, YF12, ZV21, ZHS22]. Cross-Covariance [Jun11]. Cross-Ratio [Gho06]. Cross-sectional [Van07b]. cross-validated [ZHS22]. Cross-Validation [DRM96, DH05, Gré93, YF12, Gua07, XZ09]. Crossing [Yao96]. Crossings [Ber77b]. Crossover [HVV14, SG04]. cruciate [AHP^{+}18]. Cumulant [GHH95, JTT21]. Cumulants [BNB93, PS92]. Cumulative [BDP13, BBL87, CH82, PW06, SBR98, Koi14, Wal24]. Cure [NBY08, MPV24]. Current [Aal12, Ano74d, Ano74e, Ano74f, Ano75e, Ano75f, Ano75g, Ano76e, Ano76f, Ano76g, Ano76h, Ano77e, Ano77f, Ano77g, Ano77h, Ano78e, Ano78f, Ano78g, Ano78h, Ano79e, Ano79f, Ano79g, Ano79h, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano82e, Ano82f, Ano82g, Ano82h, Ano83e, Ano83f, Ano83g, Ano83h, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h, BW05, FMS15, Gro12, GJW12, GH18, HKK^{+}76, Hou12, KHSS12a, KHSS12b, LS15, VBJ97, VJ01, WC12, JvdMP22]. curvature [YH20]. curvatures [DD22]. Curve [ACMLM03, BB15, FGD12, HV06, KRV07, LC00a, Mül93, VM00, WWP14]. Curved [BN84, BN85b, Jen97, KR15a, LS98, Sun10]. Curves [Bly93, BC99, DWV11, GMPFV11, HC10, KHL98, LZ08, PFV06, PFJGE15, Wan87, SN24, TPH21]. cusp [Kut19]. Customers [Nat75]. Cusum [LP01, LHNN03]. Cuts [CK94, EGPS98]. Cyclic [BSV13, Gad85, AAFO20]. cylindrical [KH22].

D [SW93a, Huc11]. DAG [KK06]. DAGs [Rov05]. Daisee [LRT23]. Danish [BMG82]. Dantzig [AFL10]. Data [ABKT80, ADZ15, ATV17, ABH^{+}85, AG90, Ant96, AH84, AJ00, BZF08, BW05, BB11, Bib11, BM16, Bie07, BRM14, Boe10, BCH16, Bor84b, øBFHB07, BC15, BJMP14, BHC88, Bro87b, CGL14, CHW^{+}07, Che09, CLSZ16, CY17b, CWH05, Dab87, DD88, DP18, DBS10, DCIK14, DRT13, DLP08, DPFV09, DSWH09, EGM^{+}03, Eks08, EV08, EMS15, FMS15, FRZ16, GN98, GSYB05, Gär03, GK91, GWT00, Gho06, Gil86, Got94, GWH11, Gro12, GHC92, Gro96, HBH17, HJO15, HESZ16, HT17, HCS15, HZZ07, HW17, HC10, JXCK14, JWL00, Jan91, JT07, JH17, JLY06, JW10, Jon91, Jon01a, Kim03, Kou79, KA06, LYZ15, LWY97, LV02, LR06, LZ10, LO16, LDY16, LLY17, LHWS18, LdUád15, LS96, LZZ14, LPB15, LFL16, LC11b, LS15, LMB09, MW10, ML74, MS01, MGSFB08, MW93, Mur95, Mus81, MZ11]. Data [NGMS94, Nie97b, OBL18, OKK^{+}00, Ols96, Pan02, Par01, PR07, PdT91, PW10, QZP12, QST08, QQZ16, Ren03, RV04, RR95, SSD15, Sam89, San14, SM12, SC06, Sch94, SY00, SFW16, SJ93, SHD94, Sib80, SMS12, SA80, SLCR14, SW93b, SR01, SZZ05, SW05, SWS06, Sun74, SBB05, SJS08, SLB06, SV05, Taq02, TW04, THSS09, TZ95, VBJ97, VJ01, VHK11, Van07b, Wan87, Wan95, Wan99, WR02, WLS15, WL18, WC12, fWZY16, Xue09, Xue10, YZ12, YK16, YL96, YLW00, YWK06, Yu11, YY15, ZHH10, ZL14, ZLS14, ZYX14, AHP^{+}18, ABY22, AH19, BCCH19, Ber23, BMXT20, BBP21, CW19, CHI23, CCWZ19, CDQ20, CH23, CAVGM21, CXW23, DZ21, DR22, DEV20b, DM19, DQR21, HFP24, HFS23, HYZ22, HLP23, ICM19, JKM19, JN19, KV23, KHSJ19, KL22, Kop23, LZC23, LCZW22, LZL^{+}24, LMH22]. data [MP22, MM24, MS24, MWW24, NR23, NHMW22, NJG18, SLCN19, SZ20, SH21, TKU23b, WCJ18, WGT19, WCSftADNI23, XLY20, YC22, YLZ^{+}19, ZWS19, ZLK21, ZCL22, Bor84a]. Data* [AP04, Mül05b]. Data-Based [SHD94]. Data-Driven [BB11]. Datasets [LM16]. De-initializing [RR01]. Death [AG85, BKS76, Bro87b, Höp87, Höp90, HL99, HHL02, MS94, MW93, Ros77, Ros78, MBMG23]. Debugging [Slu97]. Decent [Jag77]. Decision [FR00, Gas23, Lav23, Lin77, Ric23, Var79, Vos23, vHV85, Ano24, Gre23a, Gre23b]. Decomposable [Sun75, LET20, Rov02, Van14]. Decomposition [Bla99, DW16, Hok76, JGW13]. Decompounding [GMvdM20]. Deconvolution [AGJ07, BV09, CSS14, DP06, HT10, HB06, Jon01b, Koo99, Mab17, TW04, VU05, vEvZ96]. Decreasing [KY12, Pal09]. Decrement [Aal76]. Deep [WCSftADNI23]. Deficient [Huc11]. Definition [Van11, HYLT23]. Degeneracy [Bie07]. Degenerate [Web81]. Degree [ORL20]. Degree-based [ORL20]. Degrees [ZHL17]. Delay [KK00]. Delays [Gad85]. Delta [Kni98, Sve90b]. Demographic [HKK^{+}76]. denoising [TB22]. Dense [CY17b]. Densities [GHU03, KV98, LZ97, Pal09, Pic00, Sat96, TvdM96, APCG24, RSTU21]. Density [ATV17, Arc98, BB10, CL05, CGC06, CDGCK15, CJ08, DS03b, DL01, DH05, Efr16, EL96, FK98, GS80, GH00, GH08, GM08b, GHU03, Haz96, HT10, HL08, HW95, JV24, JKN12, JK92, KSR13, KKP99, Lou98, Mac82, MBN17, NW06, Pap00, PS99, Rud82, SS02, ST10, SW04, Scr07, SR03, ST12, Tri03, Žur79, vE92, BV09, HD22, Kle24, RW13, TKLM23]. Density-Based [CJ08]. Dental [LDY16]. Departure [MS78]. Dependence [BPS17, BJ12, CR98, Dab96, DSS13, DDK04, DL01, Ege92, EBGG18, Far07, Fas16, GG13, KKP08, KT95, Lin78b, Mei06, Que12, SS06, VB99, Vel12, GBB^{+}24, MdCCD19, TCK^{+}23, WZ22, vL18]. Dependent [ADZ15, BDW16, CS03, CMN08, DRT13, DPFV09, DSWH09, EV08, EHR88, FNR09, GHH95, Hol80b, Hol81b, HC10, KKP08, LLY18, LdUád15, LFL16, Lin88, LL99, Lue15, ML86, Mur93, Sha12, Sjö00, SMS12, SLCR14, Ste88, TGM17, Wan86, WCXS15, WL04, Zet88, AALM17, BLG20, BBD^{+}21, DEV20b, HS98, LJZ^{+}18, QB23, SN24, SLCN19, SSZ09, Wol24, YLZ^{+}19, ZCL22]. Depends [LPPS82]. Depletion [SB85]. Depth [GC05]. Derivation [Blo74, Jen97, Wre78]. Derivative [Wal97].

Derivatives [CYL11, ES00, GM84, HG85]. Derived [BPS17]. descent [Fan19]. Design [AJN02, Ber16, BNM^{+}06, CSW79, DL06, Efr08, Hua24, Jen87a, Kar15, KM91, Lan07, LT08, OB16, PW06, STH^{+}78, See93, TSH91, Yu11, NG22, Hoe78]. Design- [Jen87a]. Design-Based [CSW79, STH^{+}78, Hoe78]. Designs [AB85, BM01a, BB15, BCC17, BT13, CGL14, DW95, DR96, Det04, DM80, FMHB16, GH12, HVV14, KM95a, Mej85, SÅS07, See96, SW93a, SW76, ARP23, DMK24, LKT^{+}23, LS23, PEK22]. desparsified [KYZC21]. Detect [Bri97]. Detecting [Ber79b, HMR21, OSG08]. Detection [ACR16, CW16, GS02, HJ04, JN16a, JN16b, Oja16, Ron16, TWL18, VW09, Zwa16, BKB23, BBP21, CM20c, DEV20b, Hei19, JFO23, LET20, Mes22, MN21, RW13]. detectors [AHWKP19]. Determinantal [BL17, LPW21, PL23]. determinants [GH21]. Determine [VW15]. Determining [Ave85]. Deterministic [ADGP14, BM15]. Developed [SW87]. Development [Sch80, ARP23]. Developments [BHR^{+}76, CGL^{+}81, HKK^{+}76]. deviatile [CMY24]. Deviation [MWY15]. Deviations [EG02, Lou98]. Diagnosis [LMH14]. Diagnostic [SGR11, ZIS09]. Diagnostics [BCS13, Mül92]. Diagonal [BDL^{+}17, GP89a, GP89b]. Diagonals [Agr93]. Diagram [BCG08, Lun00, PW06]. Diagrams [GL02, Kos99]. Diarrhoea [øBFHB07]. Dichotomizations [Rov15]. Difference [Deg96, Jen79, PWY97, TGM17, Wil77]. Difference-Based [TGM17]. Different [HJ04, LC11a, NS06]. Differentiability [vdV91]. Differential [Bac11, DGCS13, Ped95, PDD10, EU21, HNRT22, JKM19]. Diffusion [BS01, Cle97, DS04, FS08, GCL87, GCLP92, GCJ94, Glo06, Huc11, Jac00, Kes97, Kes00, KP02, KK00, Man09, Ped00, SW18, Van01, AG20, DDM20, LP20, MR23, TKU23a]. Diffusion-Type [KK00]. Diffusions [BDW16, BD13, FS08, FMS11, HHL02, Jac01, KR01, Lud04, SPR^{+}13, Sør01, SJ94, Uch04, vZ03, LPPW22, NU19, SJKS22]. Digital [HS17]. Digraphs [AMP97]. Dimension [Ahm17, Haa08, LRSS23, Lue15, NGZ18, PS20, PS10, WWW15, CXW23, LPYZ24, RAQ21, WC20, ZLZZ21]. Dimension-independent [LRSS23]. Dimensional [BS01, BW04, Glo14, Höp99, JQ15, LLY18, MH97, Ped75b, PW10, SBV11, Wij95, BC23, BM16, BBS23, BS21, CM20b, CLP^{+}19, CL19, CDO24, FHY24, FGLT23, GW24, GPST23, GC18, HHHL23, HFS23, HT17, JB20, KYZC21, KK23, LZL^{+}24, LPPW22, LJZ^{+}18, PWN22, RMG19, YM22, YA20, ZL22, ZHS22, ZHL15, ZLK21, vdWBM19]. Dimensionality [BF02, BLM20]. Dimensions [HS12, MvdG15, BDS22, SM24, TKU23b]. Direct [Kur16, Rub04a, SZ07, SBM^{+}99, Van11, ZV21]. Directed [CL12, Gås16, Ric03, XG09, DQYZ23, LMB09]. Direction [JM93, PS10]. Directional [ATV17, BRM14, FRZ16, LL06, GPVCGM16, HKŠ22, KHSJ19]. Directions [Arj11, Gre11]. Dirichlet [CCV23, GR01, GH21, JLP06, Kim03, LPR23, Sib80, WWP14, ZD24]. Disagreement [SV10]. Disc [BF03]. Discontinuity [OB16]. Discontinuous [TGM17]. Discouraged [Nat75]. Discovery [Far07, Far09, FGD12, Mei06, XBQF15]. Discrepancy [EGB13, ML74]. Discrete [CW99, Hel82, Jac89, Kes97, Lau75, LZ97, Ped75b, Ped95, Ran75, Rov15, SJ93, Ter77b, Van13, ACF^{+}21, BKT20, DDM20, GMvdM20, LCF24]. discrete-time [ACF^{+}21]. Discretely [BD13, Glo06, Jac01, Kes00, KP02, Sør01, Uch04, KV23, TKU23a]. Discriminant [BO11, LQ17, ŠBD05]. Discrimination [Sri97, RAQ21]. Discs [MH10b]. Discussion [Aal04, Aal12, ABH^{+}85, Ano07f, AKB^{+}89, Arj02, Arj04, ACR16, Azz05a, BAR^{+}85, BHR^{+}76, BRH83, BG14b, CSJ^{+}77, CGL^{+}81, CW16, DSH14, Doo16, DH16, Eri84, Gas23, GI02, Gen05, GWP89, Gus02, Heu05, HOF^{+}94, HKK^{+}76, Hoe78, Høs02, Hou12, Ize05, Jan02, JNS^{+}83, Joh02, JAL^{+}81, Ken14, LBND^{+}84, LAE^{+}89, Lau04, LRT^{+}87, LBNE^{+}78, Mak05, ML74, Mül05a, Oja16, Per14, Ram05, Ric23, Ron16, Rub04b, STH^{+}78, SN88, SLS14, Sim14, SKBBN79, SBM^{+}99, Sze05, TSH91, TCC^{+}95, TJL^{+}76, Zwa16]. Disease [BM01a, KHR02, KHT14, Lin14, RD17a, SA15, PD22]. Diseases [dCJV82]. Dispersion [AJ00, Jør86, PS83, Sch82, Son00, Vid01]. Distance [AM84, CD96, HV05, HK97, JS12, LZL^{+}24, PW06, DEV20a, RAQ21]. Distance-based [LZL^{+}24]. distance-weighted [RAQ21]. Distances [BB11, Ano23, JJCYG21, PEK22]. Distant [DE82]. Distinction [Gas23, Lav23, Ric23, Ano24, Gre23b]. Distinctive [JLP06]. Distinguish [DDL14]. Distributed [HLP23, Law82, Nor80, Wal00, AH19]. Distribution [Aar85, AV01, AW79, AM84, Awa81, ADGP14, Azz05b, BN82, BS00, Blæ78, Blo74, BCH16, BJD82, Bøl83a, Bon75, Bon82, CRCV12, CR98, Cha84a, Chr74, Cur80a, Cur80b, DP04, Deg96, FWW77, Fer91, GN98, GM16, Gar82, GS80, Gui79, Gui80, Hen86, HV05, HJR06, Hok76, HST74, HC17, ICG12, JGØ79, Jen81b, Jen86, JP06, Joh17, JSW91, KN12, KM00, KRV07, KS99, Kou85, Law82, LM04, LN13b, MW08, Miu78, Næs82, NV09, OH16, RS83, Rei81, Roj98, Rov02, RS94, SSD15, SN13, SH96, Sun75, SV05, Ter77a, Væt79, Vel12, Vid09, Wal00, Wre78, YY15, Zet88, vR88, CCV23, FM22, Gås03, HV22, HNRT22, HYLT23, Kar20, PWN22]. Distribution-Free [Chr74, Kou85, SH96, GM16]. Distributional [BNR00, Stu96, VK95]. Distributions [Aal87a, Aal95, And83, AL98, AL99, AVA06, AVCRG13, Asm89, ANO96k, Azz85, BL83, BN78, BN97, BNS05, BEK83, Ber74, Ber77a, BR17, BO11, BP05, Bøl83a, BGL13, But98, BV14, Cac77, Cha77, CMW17, CK97, Dam75, DLR18, DM80, Dok80, EGG14, GMS93, GM08a, GOV15, GM82, Gup76, HGB96, HMG06, Huz99, Jac89, JQ15, Kri95, KR15b, Laa88, Lee97, Lin78a, MG04, Mac93, MS86, McK87, MRM09, MH10a, Nor86, Oja81, Ols96, Ped75b, Præ95, PK18, SB85, SY93, Sko81a, SLB06, Swe83, Tan09, Vai91, VM15, Wan86, Whi86, WW11, Xie89, vEvZ96, vHV85, ABB24, ABY22, AVA22, BMP19, BKKR23, GH23, GAC23, GMvdM20, HPS24, KHSJ19, LCF24, Lu21, Sei24, Wal24, Dok82]. Divergence [Ano24, Gas23, Gre23a, Gre23b, Lav23, LA16, OT09, Ric23, Zha08, XT20, Lav23, Vos23]. Diverging [NC18, XNL23]. diverse [CH22]. Diversity [BNHH95]. Divisibility [SKBBN79]. Divisible [BLBEO92, BNLSV14, Kri95, GRS22, HOT21]. DNA [FH04]. Do [Lav23, Ric23, Gre23b]. Does [Lin77]. Domain [Mar98, MS91, PP16, SP09]. Domain-based [PP16]. domains [Sei24]. dominance [AKP22]. Dominant [ZL10]. Dose [LPB15]. Dose-Response [LPB15]. Double [CM17a, CM17b, DP04, MZ11, PKH17, YY15, BKKR23, PPS21]. double-bounded [PPS21]. Double-sampled [YY15]. Doubly [BCH16, LO16, LHWS18, Sam89, VBJ97, VJ01, Des23]. Down [FGD12, SS79]. Downsampling [FMS11]. Downscaling [OBL18]. Drift [FMS11, AG20, LP20, OS24]. Driven [BB11, Bol14, JWL00, JSDT11, ST10, JKM19, MR24, PPS21, dRSHK19, Uta17]. dropout [YLZ^{+}19]. Drs [KHSS12b]. Dual [FM89]. Dualization [BR03, Kau96]. Dune [MS94]. Duplicate [YL96]. Duration [Aal12, Hou12, KHSS12a, KHSS12b]. Dutch [GG23, VV24]. Dynamic [øBFHB07, DSS14b, DSS14a, DT20, KH99, LT77, MS01, MR10, SY93, Van07a, XBQF15, DQYZ23, FHSZ19, NR23, PPS21]. dynamical [Kut19, SP22]. Dynamics [AKB^{+}89, BNL07].

Early [Sch80, HMR21]. earthquake [ICM19]. Ecology [ABN12]. Eczema [BMG82]. Edge [HKD02]. Edges [HQR08]. Edgeworth [Sko81a, Sko81b, ZXL^{+}18]. Editorial [AS10, Ano74g, Ano79i, Arj92, Arj94, Bon01, BL07, DSS14b, GS19b, HD16, KPS22, KPS24, Lau98, RR13, Sch04, Tjø95, Tjø96]. Effect [BNP92, Eub00, NC92, OB16, SMZ11, Ter77a, MWW24, NM24, XZ09, ZGZ22]. Effective [EL96]. Effects [BHC88, CLSZ16, DGCS13, Grø97, HS87, Kou79, KH16, Kur16, LHML16, LFL16, MSZ16, MSSM02, Nie83, PDD10, QST08, Rub04a, SM04b, SZ07, SBV11, Sve86, Van11, Waa06, Yu16, ZLL^{+}16, CHI23, EB23, LLLP20, LYW22, Sak19, WGT19, WHR22, YZ23, ZV21]. Efficiencies [MC97]. Efficiency [Aab83, And77b, BO99, DP04, FRS99, GNPM07, Hjo86a, Hjo86b, LN95, MT02, RFK22, Van07b, ZLY14, Žur79, vdL96, vdV91, VD18, ZGZ22]. Efficient [BM01a, Bib11, CR98, CW19, CH96, CCH98, Che09, CDZ11, Che13, Det04, FGH20, HKJ11, HZZ07, HYZ22, Jon78, LV13, LFL16, LLS^{+}22, LS15, Mar98, MSSM02, MS23, NM14, OS24, PEK22, Pre03, TTZZ18, Von96, WCY22, XY15, XLS16, BBBS19, FHY24, GW24, HBD^{+}20, LLCW21, WLX19, CCWZ19]. efficient-GMM [CCWZ19]. EGARCH [Win13]. Eigenanalysis [WR93]. Eigenvalue [LLY18]. Eigenvector [Gui77]. EIV [Wil77]. Elasticities [LT77]. Electrical [BB14]. electricity [LM23]. Elja [KPS23, Cor23]. Elliptical [AVCRG13, HMG06]. EM-algorithm [Mar99]. Embedded [BG13, BG14b, BG14c, Dry14, Ken14, SLS14, Sim14]. embedding [HFP24]. Emigration [AG85]. Emission [Ped00]. Emphasis [FGD12]. Empirical [AJ78, Adi97, Ber16, BCC17, BN13, CDG16, CK06, DGSL02, Deg96, ELLV^{+}22, Fer91, GSK06, GM08a, Gui80, HKJ11, Kou85, Lai79, Lai80, Lai83, LGP11, LXZ16, LQZR09, LZ99, LVV09, MWY15, Neu09, OKK^{+}00, Pol95, PdT91, Por16, QW96, QJ01, Ran75, Rud82, SN13, Sti82, SKO17, SSZ09, Tho83, TDR09, WR02, Wol24, WLT15, Xue09, Xue10, Yuk92, ZG03, ZWS19, ZKP^{+}24, vH80, vZ03, vdWBM19, vdV94, Ber23, EPM15, HJG21, KSSR21, NG23, Par20, ZHW19]. Employment [Laa78]. Emulation [PD22]. Emulation-based [PD22]. End [KA06, Pal09]. End-Point [Pal09]. End-points [KA06]. Endogenous [BVV17]. endpoint [NM24]. Endpoints [KR15b, Mül93]. Enhancements [NGZ18]. Enriched [CV01]. Enriching [LYZ15]. Ensemble [DY17, LT21, SO13, LZC23]. Entropy [AVCRG13, CGL14, JM83, ML86, TvdM96, WW11]. envelope [GZZM23, LLS^{+}22]. Envelopes [CFL24, ZVD22]. Environment [Guo11]. Environments [SY93]. Epidemic [Bri97, GWH11, LY08, NH15, NX17]. Epidemics [BO02, BKO11, CO07, DO05]. Epidemiological [Det04]. Epistemic [PLL23]. Equality [FSHK13, Præ95, CLP^{+}19]. Equation [And91, Bac11, Dem17, DF90, DW16, PM03, QQZ16, EU21, GRS22, HNRT22]. Equations [AJ00, CYM93, DGCS13, Imo15, Jun08, Kos99, Kün83, LL96, Li01, NYR18, PTF09, Ped95, Wan99, Wil77, JKM19]. Equivalence [AMP97, Bac11, LYW22, Rov05, CCWZ19]. Equivariant [Bon79, HK24]. Ergodic [Arf98, Gui77, Höp87, Jen87b, Kes97, OS97, Van01, vZ03, AG20, EU21, NU19, OS24, TKU23a]. Ergodicity [LDM15]. Errata [Ano23]. Error [Abt99, ADL15, Aug04, BJFG15, BS16, GSK06, Guo11, HT14, HB06, HL02, HWH15, HW17, KN12, LR06, LB94, MRS14, NV09, OS96, PLKP06, SFW16, TL03, TDR09, TRL15, Wan08, WWW15, XLS16, YD07, Zha95, Zha08, HBD^{+}20, JV24, SW19, SAS24, VS21]. error-contaminated [SW19]. Error-prone [HW17]. Errors [BG98, CK06, DBD18, DR97, FWW77, HJKQ18, KHL98, Kos99, LLY17, RV04, Sha12, TGM17, TDR09, VM00, WC12, Wil77, XMW15, YZ07, ZC03, DR18, HMR21, JN19, VW19]. Errors-in-Variable [WC12]. Errors-in-Variables [CK06, Wil77]. Esséen [BBG97, HJS90]. establishing [KPS23]. Estimable [Tre83]. Estimate [AL98, AL99, BL83, Ber76, CD96, Dab92, Hög78, JK92, KHT14, Laa88]. Estimated [AC99, BW08, Sun96, OHN21, ZGZ22]. Estimates [Agr93, And83, Arc98, Efr05, Eks01, Eng80, GM98, GJ03, GJW08, Jen93a, KL78, Kol97, Kuh04, Laa78, LdM80, Mac82, ML86, MH97, NC92, Oja99, Ryd95, Žur79, CZT20]. Estimating [Aal12, AGJ07, Abt99, ABB24, ABK96, AJ00, BSV13, BB15, BS01, BD07, BDL^{+}17, CTYF13, Dem17, DK80, DN15, DY17, FRS99, GM84, Gil86, Gui77, HKJ11, HS12, Hou12, HH16, Imo15, Ist96, Jac01, JK04, Jun08, KHSS12a, KHSS12b, Kes00, KP02, KN12, Kol81, KHL98, KM95b, LY08, Lee97, LL96, LXZ16, Lin00, LZ97, LJZ^{+}18, MS98a, MR10, NYR18, OPP18, PTF09, Ped00, PM03, QQZ16, Sun95, Sve86, Swe88, THSS09, Uch04, VW19, Wan87, WG96, Wan99, Wan06, ZNJ15, vE92, FHSZ19, LPW21]. Estimation [AV01, AH78, AAFO20, ATV17, AG85, AHK91, AG90, AGM00, AGR13, Arf98, AGGM06, AOH00, ADGP14, Bar03, Bed93, BG98, BD13, BN15, Bib11, BR14, Bie07, BG01, BVV17, BL17, BDV06, Bor84a, Bor84b, BCG08, Cac77, CR98, CS03, CM20a, CGL14, CM82, CM84b, CHW^{+}07, Cha84b, Che09, Che15a, CMY24, CDZ11, Che13, CWH05, Cle97, CR13, CSS14, CDGCK15, Cuc08, CWZ21, DE82, DSJP14, DGCS13, DLH14, DS03b, Det04, DPV06, DRM96, DSD^{+}14, Doo18, DL01, DE06, DH05, DP16, EVP15, ES00, Efr16, EL96, Eks08, EGG14, EBGG18, FT16, FHT94, FK98, FLS05, Fas16, FS10, FW03, FL11, FZ06, Fra77a, Fra77b, Fra78, FSGMM16, GGG13, Gär03, GS80, GR05, GCLP92, GCJ94, GR10, GSK06, GM08a, GOV15, GM08b, GS99, Glo06, GG13, GSG96]. Estimation [GL15, GH87, Gro96, Gua07, GP89a, GP89b, Gui79, HS10, Ham88, Han16, HG85, HA98, HK97, HHL02, HK15, HSW03, HW95, HP00, HZZ07, HC17, Jac00, JFKC05, JKN12, Jon01b, JV06, KO03, KSR13, KL78, Kes97, KB04, KM00, KP21, KD84, KKP99, KK00, KR15b, KR15c, KH16, KS01, La 08, Lan13, Lan74a, LT77, LT08, LCZ09, LPB15, Lin88, LS15, Lud04, MSZ16, Man09, MSSM02, MP14, MF97, MBN17, Mic09, Miu81, MS78, MW93, Mül93, NE87, Neu97, NHS^{+}19, NHMW22, NGAS92, Nie98, NT01, NM24, OB16, OKW88, Ols96, OFFL12, OS97, Ove98, PLHS17, Pal09, PSS10, PS99, Ped95, Pen95, PBB06, PdT87, PS13, PK18, QZP12, Qin98, Rah86, Ran84, Rei81, Roj98, RR95, Sae15, SSD15, San14, Sar09, ST10, SM04b, SM04a, SMZ11]. Estimation [SBV11, STMC16, SS06, Sch75, Sch81, Scr07, Shi17, SR03, ST81, SWS06, SG12, Sve90a, SLB06, Tan94, TWL18, TTZZ18, TGM17, Ter81, TKU23a, TDR09, Uch04, Uta17, VBJ97, Van01, VOG11, Vet12, VM00, Von96, Wan08, WLS15, WGT19, Wij95, WF79, Win13, WZ10, XY15, XL10, YZZ11, YY15, Zha08, ZHH10, ZX96, ZZLZ16, ADN21, ACF^{+}21, APCG24, AG20, AV21, BGH19, BBBS19, CW19, CM20b, CFR19, CL21, CDQ20, Che15b, CGC23, CDO24, DHH24, Des23, DR18, FGH20, FHY24, GW24, GGS20, GSUC22, GKL21, HNRT22, JV24, JvdMP22, KV23, KL22, Kut19, LZC23, LP20, LLCW21, LM23, LLS^{+}22, LC22, LLYC22, LP22, LPPW22, LMH22, MBMG23, MPV24, NG22, OS24, OHN21, PRV21, RSTU21, SAS24, dRSS22, SKR19, TKU23b, VVI^{+}22, WLX19, WC20, WCY22, Wol24, Wu13, YH20]. estimation [Zha00, ZHW19, ZL22, ZHS22, ZSJT24, ZLK21, ZKP^{+}24, vLM23]. Estimator [BB10, BKS76, BDP13, CL05, CGL14, CY17a, Cha84a, CH96, Che91, CD01, CRI03, CGC06, GM16, GL07, GNPM07, HL00, Joh78, KD84, KSN95, KSM87, LL09, LN13a, LB88, LWY97, LFL16, Lou98, LL99, MGSFB08, Næs82, Ohl86, PLKP06, PdT87, Ros74a, Ros74b, Seg02, SV04, SW18, SW76, Sti82, ST12, Sun96, SV05, Tan09, TZ95, Wei93, YLW00, Yu11, BD20, BW19, CL19, GME24, Koi14, LM24, MR23, MT19, Tak23]. Estimators [AB85, AAA04, Ala77, AALM17, Arc98, ADL15, AFV14, BIP14, BB11, Böh10, Bon79, BZ82, Buh93, Cha15, CC98, CYL11, CP98, CL01a, CJGPL07, CDY11, DP13, DNL10, EHR88, FM90, GCJL03, GWP89, GV93, GHU03, GDS88, GJW12, HGB96, Haz96, Hjo86a, Hjo86b, HW17, JM16, JP06, Kle91, LC11a, LN13b, LM18, MS98b, Miu81, Mül85, NM14, Nor80, OS96, Pfa93, PS83, RS83, Rud82, RS94, SS02, Sam89, SW84, SW04, Sch75, Sch81, SHD94, Sko81a, SA80, SS00, Stu83, STK17, Ter83, Tre83, TvdM96, VU05, Wan95, WG96, Wan99, Wan00, Xia94, XLS16, Zha96, ZHF03, vEvZ96, BBG97, CCWZ19, HMP22, JN19, KR20, MPV19, RFK22, SJKS22, VD18, VHF20, WHR22, WC21, YZ23, vZ03]. Ethernet [Taq02]. Euchaeta [Sch79]. EV [YZ07]. Evaluating [ACFS83a, CTYF13, HTK15, ZXL^{+}18]. Evaluation [AG85, Laa78, Min79, ST76, dCJV82]. Even [Lav23, Ric23, Gre23b]. Event [øBFHB07, CWH05, DM80, HBH17, HS98, SC06, Van07a, fWZY16, LLXH19, MP22, NJG18, NM24, PD22]. Events [ABKT80, ADZ15, ABN12, DS09, ADN21, CZT20]. Evidence [BNHH95, Lav23, Ric23, Bic23, Gre23b]. Evolution [BM15]. Exact [AL98, AL99, Aug04, BNR00, BKS76, BLM20, CO07, Dem17, GAC23, GG01, Gui04, HN99, Kim97, Kre87, Laa88, MG98, Nat93, NM87, OS96, SPR^{+}13, TF12, VKY^{+}14, Wan90, dCCU17]. Exactness [BNK99]. Example [Doo16, MS91, PS10, Sen88]. Examples [Aal87b, Mad76]. Exceedance [Far09]. Excess [Zah96]. Exchangeability [KY12, BQ22, Dia23]. exchangeable [ORL20]. Exchangeably [Che15b]. exciting [DK06]. exclude [YLGL20]. excursion [CDO24]. exhibiting [VMG22]. Existence [Buh93, GDS88, Jac89, MH97]. Expansion [Sko81b]. Expansions [CP98, Jen87b, Jen89, Sko81a, ZXL^{+}18]. Expectation [BEK83, GGS20, Mes22]. Expectations [ST76, HK24]. Expected [FT16, MC03, Nie97a, Nie99, CM20a, LT21]. Expectile [EK22, KZ17, GSUC22]. Expectile-based [EK22, KZ17]. Experiment [Kou84]. Experimental [GPM04, KM91, DM19, LS23]. Experimentation [Lai79, Lai80]. Experiments [AR94, BM01a, FOS^{+}14, GT98, GH87, Kou79, Laa75, SW93a, Sko86, SA80, TS91, TSH91, TJL^{+}76, Tor88, WY03, BPR22, Hua24, LGL19]. Expert [DMPV02]. Explanatory [Kou84, Nor81]. Explicit [Kes00]. exploitation [LRT23]. Exploiting [Guo11]. exploration [LRT23]. Exploring [HBH17]. Exponential [Abt99, AOH00, Asm89, AJRN16, BL83, Bar03, BNK74, BN84, BN85b, Blæ78, BJ85, Bøl83a, CP07, Chr89, CV01, Cur80b, DDL14, EGPS98, Eri84, FM90, GM08a, GT98, Hol75a, Huz99, Jac89, Jen79, Jen97, Jør86, Kim97, KR15a, KS99, KL89, KS94, KM95b, Lau75, Lee97, LZ97, Mac93, MH97, NC15, Nor80, PS92, Pic00, RS83, Sør98, Ste91, Sti82, Sun74, Sun10, Væt79, VM15, WF79]. Exponential-Type [FM90]. Exponentiality [Kle83]. Exposure [Væt79]. Exposures [Gil86]. Expressed [Gui82]. expressing [GH23]. Expression [SM12]. Extended [Bon79, CFJP07, HV06, WZH16]. Extendibility [But86]. Extending [MR14]. Extension [BG11, Wil79, Tor88]. Extensions [Aug04, DSH14, GN09, ZG03]. Extensively [HWH15]. Extinction [Bro87a]. Extraction [Van13]. Extrapolation [SAS24]. Extrema [Mül85]. Extremal [BPS17, MdCCD19]. Extreme [BR17, EGG14, GGS20, GJ03, GA86, KY12, KL89, Lau74, LBND^{+}84, LRT^{+}87, Que12, BKB23, BBdW20, CMY24, CV22, GM23, GSUC22, HPS24, LPYZ24, Sei24]. extreme-type [LPYZ24]. Extreme-Value [KY12]. Extremes [BG14a, FNR09, HPR21, HOT21, KL22]. Extremogram [CDG16].

Factor [AL98, AL99, Laa88, WW01, BS21, CLR19, FGH20, KG18, LC22]. Factorial [Ber74, Ber77a, KM95a, TS91]. Factorizations [CW99]. Factorized [ZL22]. Factors [CL12, Joh08]. Failure [Aar85, AH84, BIPV13, Bor84a, Bor84b, CLSZ16, CL20, Cro98, DDK04, GS99, HHVA03, HCS15, JLY06, LHWS18, NGMS94, QQZ16, SZZ05, SWS06, Xie89, GH23, MP22, RFK22, ZZLC21, ZCL22]. Failures [HS87, LHML16, Thy75]. Fallible [Swe88]. False [Far07, Far09, FGD12, Mei06, XBQF15, DBNR20]. Familial [SJS08]. Familial-Longitudinal [SJS08]. Families [AVA06, Asm89, Bar03, BNK74, BPS17, CP07, Chr89, CV01, EGPS98, Erl81, Jen79, Jen97, KR15a, KL89, KS94, Lee97, Mac93, MH97, PS92, Pic00, Sør98, Ste91, Sun10, VM15, AVA22]. Families* [Azz05b].

Family [AJRN16, BBQ18, BCCA11, BL83, Bri97, GM08a, GM82, Guo11, Jac89, Kim97, LZ97, LA16, Nor80, Sti82, Sun74]. Family-Based [Guo11]. Family-Quadratic [GM08a]. Familywise [BS16, HBD^{+}20]. Fast [CR13, Hua13, Lin77, RD17a, VLIN21]. Father [SS79]. Fatigue [Ryc96]. Favorable [AR94]. Fay [CK23, DSD^{+}14]. Feature [JLP06, WL18, ZZLC21, LZL^{+}24]. Feedback [AGR^{+}18, Ter77a]. feedbacks [MP22]. Female [BMG82]. FF [CH22]. Fibre [KR15b, SLB06]. Fibres [JH05]. fibrous [LGL19]. Fiducial [VM15]. Field [AT15, KHR02, Kou84, SA80]. Fields [Bol14, De 06, GHH95, HN99, Kün83, LL99, MS94, PSS10, RLOS18, RT02, SV10, TB98, VS07, WB15, DD22, CDO24, DH23, KP21, LCF24]. Fieller [SW93a]. figure [KPS23]. Fiksel [CDDL12]. Filter [MS98a, SO13]. filters [HWC20]. Financial [DPV06]. Finding [Jør92]. fine [LPPW22]. Finetti [BJ93, DF90, DEL92, Dia23]. Finite [ABC11, CSW79, CL01a, CM15, DEL92, Did07, Edw80, Erh08, FT16, GSK06, HW98, Hög78, HMG06, JR76, Kol81, NG23, Ste91, SW75, Sun83, Tan09, Tan94, Tho81, JB20, SK19]. finite-population [SK19]. Finland [KPS23]. Fires [MDA10]. First [DW97, GCL87, ML86, PCW02, SS98]. First-Order [DW97, SS98]. Fish [HJO15]. Fisher [SJKS22]. Fisherie [GIA02]. Fit [BBQ18, BR81, CL05, CH04, CS90, DPV06, DPFV09, DF03, DR10, FMS11, GJ05, GQR06, IKL94, JM93, KP77a, Mac82, MM93, MU91, MRM09, MH10a, Min79, Mül92, Mun02, NDH^{+}21, NEV13, Pan02, Pap00, Qin98, Ren03, Rit04, SMSD92, SH96, STZ01, Sun10, Waa06, dCCU17, BQ09, BRM14, CCH01, CAVGM21, GPÁLÁPGM21, ORL20, Min81]. fitness [MP21]. Fitting [ABY22, ANO96k, KS08, LO16, LMB09, RT02, WWP14, MWW24]. Fixed [Ano83i, Ban05, GM83, MSSM02, Nat93, KT19]. fixed- [KT19]. Fixed-Width [Ano83i, GM83]. Flags [Rov05]. Flat [Sun10]. Flexibility [Sto11]. Flexible [ICG12, KSR13, KKW24, LPR23, MG04]. Floating [Lai79, Lai80, Lai83]. Flocculated [BDH03]. Flow [HL99, dMR88]. Flowgraph [Huz99]. fluorescence [BBP21]. Focused [BL08, GC18, KHH19, WLS15, XWH14, DHH24]. follow [CL20]. follow-up [CL20]. Forecasting [BCS00, Kop23, Lau76, Sjö00, Tho83]. Forecasting* [GH02]. Forensic [DMPV02, EM02, CCV23]. Forest [DK80, MDA10]. Forests [LN95]. Forget [Lin77]. Forgetting [AOH00]. Form [PC99, SW84]. Forms [Mun02]. Formula [IYW14, Gro21]. Formulas [BJ85, DW95]. Forward [But86, Gup76, HH82, NH15]. Foundation [Lok07]. Fractional [Lud04, RLOS18, BHLP19]. fragmentation [HNRT22]. Frailties [CH96, CCH98]. Frailty [DNL10, HHVA03, JH17, NGAS92, WCXS15, WCJ18]. Frame [DGGM16]. Framework [GH16, GH12, NEV13, SO97, Ano23, CLP18, JJCYG21, KSSR21, LZL^{+}24, YK20]. Free [Chr74, Kou85, SH96, DM19, GM16, WL18]. Freedom [ZHL17]. Frequencies [GS76, Ran75]. Frequency [Bib11, DE82, HJO15, MZ11, PP16, SP09, BM16, Fas16, JKM19, TKU23b]. Frequentist [Cer17, DB03, GZZM23]. Front [Ano74h, Ano74i, Ano74j, Ano75h, Ano75i, Ano75j, Ano76i, Ano76j, Ano76k, Ano76l, Ano77i, Ano77j, Ano77k, Ano77l, Ano78i, Ano78j, Ano78k, Ano78l, Ano79j, Ano79k, Ano79l, Ano79m, Ano80i, Ano80j, Ano80k, Ano80l, Ano81i, Ano81j, Ano81k, Ano81l, Ano82i, Ano82j, Ano82k, Ano82l, Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano84k, Ano84l, Ano85i, Ano85j, Ano85k, Ano85l, Ano86i, Ano86j, Ano86k, Ano86l, Ano87h, Ano87i, Ano87j, Ano87k, Ano88h, Ano88i, Ano88j, Ano88k, Ano89h, Ano89i, Ano89j, Ano89k, Ano90i, Ano90j, Ano90k, Ano90l, Ano91i, Ano91j, Ano91k, Ano91l, Ano92e, Ano92f, Ano92g, Ano92h, Ano93e, Ano93f, Ano93g, Ano93h, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g, Ano95h, Ano96f, Ano96g]. Front [Ano96h, Ano96i, Ano97f, Ano97g, Ano97h, Ano97i, Ano98f, Ano98g, Ano98h, Ano98i, Ano99e, Ano99f, Ano99g, Ano99h, Ano00e, Ano00f, Ano00g, Ano00h, Ano01e, Ano01f, Ano01g, Ano01h, Ano02e, Ano02f, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano04g, Ano04h, Ano05f, Ano05g, Ano05h, Ano05i, Ano06e, Ano06f, Ano06g, Ano06h, Ano07g, Ano07h, Ano07i, Ano07j, Ano08e, Ano08f, Ano08g, Ano08h, Ano09e, Ano09f, Ano09g, Ano09h, Ano10g, Ano10h, Ano10i, Ano10j, Ano11f, Ano11g, Ano11h, Ano11i, Ano12e, Ano12f, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14e, Ano14f, Ano14g, Ano14h]. frontier [BBBS19]. Full [WD98, WL18]. Fully [AT15]. Function [Aly90, ABK96, AGM00, Aug04, Ban05, BNHJP76, BL94, BH99, BDP13, BBL87, BG01, BD07, BCH16, BL90, CR98, CC98, CYL11, CWH05, Deg96, Efr05, EBGG18, FHT94, GM08a, GM08b, GHH95, GJW08, Gua07, Gui80, HST74, Ist96, Jag77, JMT94, KM00, KS08, KKP08, Kou85, KS01, LT08, Lin00, Lin94, MS98a, OS97, PdT91, PG13, Puk82, Rah86, Roj98, Rue97, Sør01, SV05, THSS09, AG20, BD20, CDQ20, CV15, DEH21, LPW21, QB23, SKR19]. Function-based [PG13]. Functional [BCCA11, BIPV13, BCS00, CFMS03, Car07, CGL14, DL01, FVV10, FV06, FSHK13, Har02, HS04, JT07, KH22, LB94, MSZ16, MS78, Mül05a, Mül05b, OBL18, SLCR14, TPH21, WFC16, ZL14, AHP^{+}18, CZL24, CZT20, FM22, GPÁLÁPGM21, HFP24, HJG21, HYZ22, JFO23, KV23, MM24, SZ20, WCSftADNI23, YH20, ZKP^{+}24, ZXLL23, Heu05, Ize05, Ram05]. functional-on-scalar [AHP^{+}18]. Functionally [AP07]. Functionals [Dre98, GM08b, Mam92, vE92, Adi97, BMP19, Gro21, LAO23]. Functionals* [DT05]. Functions [Adi97, BBQ18, BZF08, BLBEO92, BS01, BN13, Bøl83a, CTYF13, CM04, CJ08, DP13, DR10, Eks13, FRS99, Fer91, GM84, Gui82, HG85, HL08, Jac01, JS12, Jen79, JK04, Kes00, KP02, KSM87, KL89, KS94, LDW06, LWY97, LXZ16, LVV09, MF97, Mat79, Mül85, PSS10, PFJGE15, PWY97, Pre05, PS10, Rei81, SH96, Sko81b, Tre83, Uch04, Vid09, WG96, Wan06, WW11, YLW00, Zha96, ZL18, DQR21, HD22, HK24, PS20, Sei24, SP22, Wal24, ZBS20]. fused [CM20c]. Future [Arj11, Gre11].

G [Blo74, DMK24, Hok75, Hok76, Hok76]. G-optimal [DMK24]. gait [TPH21]. Galton [Ner77]. Gamma [Bon75, Dam75, Jen86, Law82, LPS03, McK87, NBW02, Wal00, Whi86]. GARCH [Cav23, LL09, LL12, LN13a, LKN15, dRSHK19]. Gauss [Bon79, OKW88]. Gaussian [Abt99, BCCA11, BR03, BB10, BN97, BNS03, BNS05, BPW14, DD22, BCCAVMO21, Bie07, BKKR23, Bol14, Bøl88, BM01b, BDH03, Buh93, But98, CMW17, CL12, De 06, DH23, DE06, GBB^{+}24, HJKQ18, Ist96, JSW91, Kau96, KR01, KS08, Kos99, Kün83, LL99, MR14, Miu78, MSW98, MR24, MZ11, OT09, RZM16, Rov02, RT02, Son00, SR11, SJ94, TKLM23, WB15, Whi86, ZHF03]. Gaussianity [HJG21]. GCV [LDA12]. GEE [Pan02]. Gene [Guo11, SM12]. General [And77b, Böh10, BK95, CH23, DH78, FHTT18, FPW11, FW03, Gär03, GH12, Gui79, Irl90, Kol81, KSM87, LY08, Lau75, LC00a, Mei06, OKW88, PT92, RV04, SÅS07, Sko81a, SS80, ZX96, ABB24, BCCH19, BW19, KK19, QB23, SK19, Wu13]. Generalization [AM84, Blæ78, Lan13]. Generalizations [Höp99]. Generalized [Agr93, AC99, BIPV13, BS10, BNS05, BH99, BB11, BS99, But98, CP07, CG99, DW97, EMR09, Eie83, Eks01, Far09, Far15, FOS^{+}14, GA86, HJS90, Huz99, IYW14, Imo15, ICG12, JM16, Jun08, KT19, KO03, KR15c, LL96, LDY16, LQZR09, Lin14, LC11b, LA16, Lyn88, MW12, Mic09, Mse22, NGZ18, PKH17, PS99, PC99, See96, SZ02, SZZ05, Sun96, SJS08, Waa06, WHF98, XZ09, XL10, ZHL17, ZG03, ZIS09, DLS^{+}24, GPST23, HBD^{+}20, HFS23, Kar20, KR20, LLCW21, VVI^{+}22, XT20, YM22, ZX19]. Generalizing [LKT^{+}23]. Generally [CGP07]. Generated [Asm89, KM91, Son00]. Generating [GHH95, KS08]. Generation [Ner77, Sto11]. generative [Mei24]. Generic [DW16]. Genetic [DMPV02, ZLY14, ZXL^{+}18, LLLP20]. Genetical [EM02]. Genetics [EM02, MG95]. genome [HBD^{+}20, ZXLL23]. genome-wide [HBD^{+}20, ZXLL23]. Genomic [San14, RMG19]. Genomics* [Eri04]. geo [Kop23]. geo-locations [Kop23]. Geoadditive [KF07]. Geodesic [BG13, BG14c, Sim14, NHS^{+}19, Dry14, Ken14, BG14b, SLS14]. Geometric [MT14, YA20, Hei19, MMO23]. Geometrically [ABN12]. Geometry [Har02, Sko84, ZHL17, BKKR23]. Geostatistical [DL06, WB15, BCCAVMO21]. Germ [DLH14]. Germ-Grain [DLH14]. German [DGGM16]. Germs [DLH14]. GI [Blo74]. GI/G/1 [Blo74]. Gibb [Pic00]. Gibbs [BC23, BG01, CLP18, CDDL12, CR13, GSG96, LDM15]. Gibbs-type [CLP18]. Gini [DNCZ21]. Girolami [BG14b, DSH14, Dry14, Per14]. Given [GWH11]. Glivenko [SS80]. Global [Gho06, Glo14]. Globalizing [ZL14]. Globally [VR08]. Globe [Jun11]. GMLE [SY00]. GMM [CCWZ19, SW19]. Golfer [BM15]. Good [HT17]. Goodness [BBQ18, BQ09, BR81, BRM14, CL05, CCH01, CH04, DPV06, DPFV09, DF03, DR10, FMS11, GJ05, GQR06, IKL94, JM93, KP77a, Mac82, MM93, MU91, MRM09, MH10a, Min79, Min81, Mül92, Mun02, NDH^{+}21, Pan02, Pap00, Qin98, Ren03, Rit04, STZ01, Waa06, dCCU17, GPÁLÁPGM21, ORL20]. Goodness-of-Fit [BBQ18, CL05, DPV06, DPFV09, DF03, DR10, FMS11, GJ05, GQR06, IKL94, Mac82, MU91, MH10a, Min79, Mül92, NDH^{+}21, Pan02, Pap00, Rit04, Waa06, dCCU17, BQ09, BRM14, CCH01, GPÁLÁPGM21, ORL20]. Gradient [CJ08, Fan19, Kle24]. Graduation [Hoe76]. Grain [DLH14]. Grains [Mol94]. Graph [DE06, Fra77a, Fra78, Fry90, LMB09, ORL20]. Graphical [AP07, BR03, Buh93, CAS03, CL12, CLP17, Cor03, DE04, Did07, GWT00, GMA11, Kau96, LAE^{+}89, LG13, Rov02, RL06, Rov15, SGR11, CRS24, LET20, NHMW22]. Graphs [AMP97, AMP01, Eva16, Fra77b, Gås16, Ric03, Rov02, Rov05, XG09, LCF24]. Greatest [BM15, PW06]. Greenland [Gas23, Ano24, Lav23, Ric23]. Greenwood [AKP22]. Grenander [Gro21]. Grey [Rue97]. grid [DMK24, LPPW22]. Group [CV14, HW17, KM91, MC97, BKM18, LYW22, TPH21]. Group-Sequential [CV14, MC97]. Grouped [Bie07, GK91, PR07, RV04]. Grouping [AKC80, BR81, Hal01, Min79, Min81]. Groups [CS82, HJO15, GH21]. groupwise [ZLZZ21]. Growth [HV06, Huc11, KRV07, LC00a]. Guaranteed [GK13]. guide [APCG24]. Guided [Gla98, TEV15]. Gumbel [Mar98].

Haar [GJ03]. Haavelmo [And91]. Hadamard [vdV91]. Haenszel [GL07]. half [OHIS24]. Hand [BMG82]. handling [CH23]. Haplotype [Cer17, SMZ11]. Hard [BF03, SJZI19]. Hard-Core [BF03]. Harmonic [BEK83, DEH21].

Harmonizable [MS91]. Harris [Jen89]. Hartley [Ber16, Ohl86]. Harvesting [LN95]. Hastings [CV02, Gås03, HT08, SR03, Sto11]. Hausdorff [PEK22]. Having [Huz99]. Hazard [And83, AGM00, AH87, BBL87, CC98, CH82, DH07, Grø97, HO93, HW95, KF07, KST95, La 08, LM18, MW93, Nie98, NT01, NBW02, PdT91, SA15, SBR98, YZ12, Zha96, ZHL15, Wal24]. Hazards [ACFS83a, ACFS83b, ABK96, AFL10, Aug04, CH96, FLS05, FMS15, HESZ16, KKC17, KHL98, LDY16, LN13b, LS15, MU91, MV87, NC18, PM03, SM04b, SBR98, SG12, Von96, XLS16, KXZA20, NG22, XNL23]. health [LYW22]. Heart [BIPV13]. Heavy [Blo74, EGG14, FWW77, JR07, Taq02, AKP22, LP22]. Heavy-tailed [EGG14, JR07, Taq02, LP22]. Helmers [PC99]. Helsinki [Cor23]. Hermite [BNP79]. Herring [GG23, VV24]. Herriot [CK23, DSD^{+}14]. Hessian [Kle16, PS10]. Hessian-Based [Kle16]. Heterogeneity [Aal87b, Boe10, Far15, GR01, THF18]. heterogeneous [BR23, LMH22, ORL20, WGT19, XLY20]. Heteroscedastic [CM01, LGP11, MNS07, HMR21]. Heteroscedasticity [NL16]. Heteroskedasticity [Bac11]. Hidden [DE82, DH08, DSJP14, GCJL03, HJ16, AGGM06, LPR23]. Hierarchical [CGGI19, DGN07, GN09, Gås16, GSK06, GDS88, KSR13, SGR11, SS09, SKR19, LPR23, ZD24]. High [BS01, Bib11, BM16, CGL14, CY17a, Fas16, Glo14, GC18, HT17, HS12, JQ15, KYZC21, MvdG15, MZ11, SBV11, WWW15, ZHL15, BC23, BBS23, BS21, CM20b, CLP^{+}19, CL19, FHY24, FGLT23, GW24, HHHL23, HFS23, JKM19, KK23, LPYZ24, LZL^{+}24, LJZ^{+}18, RMG19, SM24, TKU23b, YA20, ZL22, ZHS22, ZLK21, vdWBM19]. High-Dimensional [BS01, Glo14, JQ15, SBV11, GC18, HT17, KYZC21, ZHL15, BC23, BBS23, BS21, CM20b, CLP^{+}19, CL19, FHY24, FGLT23, GW24, HHHL23, HFS23, KK23, LZL^{+}24, LJZ^{+}18, RMG19, YA20, ZL22, ZHS22, ZLK21, vdWBM19]. High-Frequency [Bib11, BM16, Fas16, JKM19]. High-order [CY17a]. Higher [BN15, Mam92, TB98]. Higher-order [TB98]. Hilbert [FR21, HJG21, QL15]. Hilbertian [HK15]. Hill [CD01, Seg02]. Hirsch [BE10, PBBM12]. Histogram [PdT87]. Histogram-Type [PdT87]. Histograms [JKN12, Rud82]. Historical [Jen93b, Agr23]. History [ABKT80, ABH^{+}85, Hal01, HS98, SJ93, Van07a]. Hitting [GCL87]. Hjort [BG24a, BG24b]. HNBUE [Kle81, Kle83]. hoc [DBNR20]. Hoeffding [Lan74b]. Homogeneity [LPS03, NC15, THF18]. Homogeneous [AJ78, AHK91, BSV13, Fwu85, KR01, OBL18, Ter83, ZX96]. honoring [BG24b]. hop [AHP^{+}18]. Hopf [Hok76]. Horvitz [Fra77b]. Hougaard [KHSS12b]. Households [CO07]. Howell [See96]. Hull [JP06]. human [TPH21]. Hybrid [HM09]. Hyper [Rov02]. Hyperbolae [BN78]. Hyperbolic [BN78, BN82, BNS05, Kar20]. Hyperboloid [Jen81b]. hypersphere [APCG24]. hyperspheres [JV24]. Hypotheses [BS16, BH97, CTYF13, CFMS03, DK06, DB03, DBS10, Erl81, HTK15, Jen81a, Mej85, NM14, PVD13, SP09, DBNR20]. Hypothesis [Bon82, FZ00, HS06, LLLP20, ML74, PW10, SN88, TS91, KK19, MV20, SZ20].

I.I.D. [MR23, Sko81a]. Ideal [Var76]. Identifiability [Cro91, DW16, GKL21, HMG06, JLRT19, WCXS15, Wil77]. Identifiable [CTYF13]. Identification [Kuh04, Kur16, LC22, Puk82, Rah86, Van11, TPH21, ZV21]. Identifying [GH21, SGR11].

IFRA [Roj98, Wan87]. Ignorance [Wal97, GH23]. II [CH22, GV93, GG01, Gui04, Ros74b]. II-CC-FF [CH22]. IID [Gui82]. Ill [FS12, HM09]. Ill-Posed [FS12, HM09]. Illness [AG85, MBMG23]. illness-death [MBMG23]. Illness-Death-Emigration [AG85]. Illustrated [AGR^{+}18]. Image [HQR08, HKJ11, OR94, SV10, BBP21]. Image-Based [HKJ11]. Images [Har02, RS94, Rue97]. Imaging [JT07, ýJRNMJ13, KHSJ19, LLS^{+}22]. imaging-based [LLS^{+}22]. imbalances [ZSJT24]. Impact [Cer17, HT14]. Imperfect [KKMP18, Slu97]. Implementation [DLS96]. Implementing [GH14a, GHD20]. Implications [VK95]. implicit [Fan19]. Importance [CMMR12, HKJ11, Nat85, SBH03, VHF20, LRT23]. Imposed [Tan09]. Improper [TTL22]. Improve [Han16, LN95]. Improved [Jen93b, PKR^{+}97, SBH03, Vid09]. Improving [CV02, GNPM07, PRS^{+}22, SS00, Ter81, VD18]. Imputation [GWT00, Han16, LR06, Zha98, BKN23, DZ21, GK21, YK20]. Imputation-based [LR06]. In-Slide [YZZ11]. Incidence [BDP13, OKK^{+}00]. Includes [Azz85, Gre23a]. Including [BNM^{+}06]. Inclusion [AJN02, Bon12, BT13]. Incomplete [Hol81a, Mar99, Mej85, SR01, Sun74, Web81]. Inconsistencies [DGN07, GN09]. Inconsistency [MW08]. Incorporating [Kou79, VKY^{+}14, PD22]. Increasing [MW93, SBR98, GH23]. Increments [JLP09, KL89, SFW16, WF79]. Incurred [Lin78c]. Independence [BJFG15, BM16, CD03, DBD18, Guo11, Pic00, Pon86, PdT91, Rov15, CM20a, GPST23, ZLZZ21]. Independent [Cro00, JLP09, KL89, Nor80, ORL20, WF79, Gås03, LRSS23, OH21]. Independent-Risks [Cro00]. Index [Ano98e, Ano00i, Ano01i, BE10, CM17a, CM17b, CP98, HC17, LG09, PBBM12, TWL18, Var76, BGH19, BVV17, CW19, HHHL23, WZH16, WFC16, YH20, ZYX14]. Indices [CM04, VB99, VK95, vL18]. Indirect [Kur16, Rub04a, dRSHK19, SBM^{+}99]. Individual [LHML16, See93, ZV21]. Individuals [LY08, Vai91]. Induced [JV06, SY93, WCXS15, SP22]. Industry [dMR88]. Inefficient [Bar76]. Inequalities [Bøl82, Gui86, IS99, Mol98]. Inequality [Bly93, FM89, Lan74b, RVG15]. Infant [øBFHB07, Bro87b]. Infections [EGM^{+}03]. Infectious [RD17a, PD22]. Infectivity [Bri97]. Inferactive [BMXT20]. Inference [Aal76, AGR^{+}18, ABC11, AR80, Ave86, BC23, BZF08, BS10, Bar76, BDW16, BG16, Ber16, BM03, BM16, BR17, BCH16, BO02, BKO11, CV14, CL21, Che09, CY17b, CK94, CO07, CD18, CM15, DLS96, DFI14, DMPV02, DMV16, Dem17, DS04, DFG00, DQR21, EMR09, Erh08, FMHB16, GM94, Glo14, GA86, Gri09, GWH11, HM02, Hoe78, Høj04, Höp87, HC17, HYWC18, Huc11, JSDT11, Jen78, Jen86, Joh82, KR01, Kol97, Kou85, KZ17, KA06, LL12, LR06, LZ08, Lo81, MAR11, MM93, MH10b, MR12, MBR03, NU19, NV17, Nie97b, NV04, Nor77, PF08, PTF09, PKH17, Pal04, PG13, Rob78, RD17a, Rov02, STH^{+}78, SJ93, SPR^{+}13, Sha12, SMB14, dRSHK19, SJ94, SSZ09, Sun83, Sve76, TT17, Toc01, TRL15, WR02, WW01, YK16, ZYX14, AHP^{+}18]. inference [BSO22, BT21, CZL24, CLP18, Cav23, CV22, DR22, DO05, DM19, DT20, Fan19, GW24, GLQ18, HHHL23, HYZ22, HLP23, HOT21, JKM19, KYZC21, LPW21, MS24, Mei24, NG23, PD22, SH21, TPH21, TCK^{+}23, ULK23, VMG22, WHZ20, WZ22, XNL23, YK20]. Inferences [LQZR09]. Infinite [Bac11, Fre89, Kol97, LP01, SKBBN79, Taq02, Wu13]. Infinitely [BLBEO92, BNLSV14, Kri95, GRS22, HOT21]. Inflated [HS10, LC11b].

Influence [BCS13, Bro87b, HVV14, HO93, Jun08, Pre05, PS10, ZHF03]. Influential [SGR11, vR95]. Information [Ano74d, Ano74e, Ano74f, Ano74k, Ano75e, Ano75f, Ano75g, Ano75k, Ano76e, Ano76f, Ano76g, Ano76h, Ano76m, Ano77e, Ano77f, Ano77g, Ano77h, Ano77m, Ano78e, Ano78f, Ano78g, Ano78h, Ano78m, Ano79e, Ano79f, Ano79g, Ano79h, Ano79n, Ano80e, Ano80f, Ano80g, Ano80h, Ano80m, Ano81e, Ano81f, Ano81g, Ano81h, Ano81m, Ano82e, Ano82f, Ano82g, Ano82h, Ano82m, Ano83e, Ano83f, Ano83g, Ano83h, Ano83n, Ano84e, Ano84f, Ano84g, Ano84h, Ano84m, Ano85e, Ano85f, Ano85g, Ano85h, Ano85m, Ano86e, Ano86f, Ano86g, Ano86h, Ano86m, Ano87e, Ano87f, Ano87g, Ano87l, Ano88e, Ano88f, Ano88g, Ano88l, Ano89e, Ano89f, Ano89g, Ano89l, Ano90e, Ano90f, Ano90g, Ano90h, Ano90m, Ano91e, Ano91f, Ano91g, Ano91h, Ano91m]. Information [Ano92i, Ano93i, Ano94i, Ano95i, Ano96j, Ano98j, Ano99i, Ano00j, Ano01j, Ano02i, Ano03i, Ano04i, Ano05k, Ano06i, AVCRG13, ADL15, BL08, GH00, GH08, GH14b, GC18, HW98, KC11, Lan07, Lin78c, MG95, MS98a, MS98b, MBR03, PT92, ST10, Sas92, Sen88, Yu16, CH22, KHH19, KK23, LKT^{+}23, Sak19, XWH14]. Information-Theoretic [GH00, GH08]. Informative [DP18, DLR18, JFKC05, SS18, SMS12, SLCN19]. Ingersoll [BKT20]. Inhomogeneous [CV15, DP16, FSGMM16, JVA11, MNS07, NV04, ABY22, BC23, ICM19, vL18]. Initial [LY08, Lin77]. initializing [RR01]. Initially [DE82]. injury [AHP^{+}18]. Innovation [Law82, McK87, SN13, Wal00]. Innovations [Che15a, CR13]. Inputs [FV06, Puk82]. Instability [Doo16]. instance [Wol24]. instance-dependent [Wol24]. Instrument [XMW15]. instrumental [DLS^{+}24, SW19, VD18]. insurance [YC22]. Integer [BIP14, BNLSV14, BSC24]. Integer-valued [BIP14, BNLSV14, BSC24].

Integral [Bar03, Dab96, ES00, GQR06, GM08b, SHD94, GRS22]. Integrals [Erh08, SMV05, Stu94]. Integrated [BNS03, BG16, DPV06, DS04, Glo06, HB06, MR14, PSS10, SZ95, KR01]. Integration [MP80, MP84, CH23]. Integrative [LMH14]. Intensities [Gil86, Gré93, HKJ11, Sve90a]. Intensity [BDW16, CYL11, Cuc08, DLH14, FSGMM16, HBH17, HA98, MSSM02, PdT87, Zah96, Des23, RW13, WCY22]. intensity-reduction [WCY22]. Inter [RGS03]. Inter-Rater [RGS03]. Interacting [MH10b, CGC23, VS07]. Interaction [ABKT80, BMG82, DLH14, DFG00, Høj04, TS91, PBHMC09]. Interactions [BPR22, San14, TB98, HWC20, RVG15]. Interactive [Mus81]. Intercept [GM16, JGØ79]. intercepts [NR23]. Interjumping [ADGP14]. Intermediate [CCH01, Kur16]. intermittent [CL20]. Interpolating [Dam80]. Interpretation [Kos99, HKŠ22]. Interpreting [FH04]. Intersecting [GGG13]. Interval [AM84, BC15, DSD^{+}14, Gui79, Lan74a, LWY97, LdM80, MW08, Nat93, PR07, Ren03, SY00, SZZ05, SWS06, TZ95, Wij95, YLW00, YWK06, ZHH10, CZT20, GME24, PLL23, Van98, ZBS20, ZCL22]. Interval-Censored [LWY97, MW08, SY00, SZZ05, SWS06, YLW00, YWK06, ZHH10, BC15, CZT20, ZCL22]. Interval-grouped [PR07]. Interval-Truncated [TZ95]. interval-valued [ZBS20]. Intervals [Ano83i, BW05, BL94, Bér94, BBL87, BL90, CQ02, CFJP07, DGSL02, GM08a, GM83, GH18, GG01, Gui04, HL08, Hol93, KWA16, LL90, LGP11, LAKZ12, LY03, MW12, MC03, PWY97, RGS03, Thu14, Tri03, Vid09, Xue09, BBdW20, FB20, GM23, GME24, KK19]. Intraclass [RGS03]. Intractable [DSJP14, HJR06]. Introduction [KM91, Pal04]. Invalid [LHHF13]. Invariance [Lan13]. Invariant [BNC91, DF03, Jen81a, KL89, LM04, Oja99, OKW88, vR94, Ahm17]. Inverse [BN97, BNS05, BDP12, BO11, But98, DBS10, DRM96, FV06, FS12, GJW12, Han16, HM09, JSW91, Lue15, Miu78, Neu97, OS96, Por16, Pre05, PS13, Rov02, Whi86, AV21, BR23, HK24]. Inverse-regression [Lue15]. Inversion [Hol75b]. Inverted [vR88]. Invertibility [Ter77b, Win13]. Invertible [HP00]. invited [Gas23]. Involving [BBG06, Jon91, WD98]. Ion [CYM93]. Irregular [LR08, MV20]. Ising [dCCU17]. Isotonic [CC98, CDMGR06, DP13, Dha16, HM02, LM18, BD20, OHIS24]. Issues [GIA02, GH02, JJ02]. Item [Chr74, Tju82]. Iterated [Bjö10, HL08, Zha96]. Iterations [Che13]. Iterative [KR15a, RV04, YF12].

J [AVA22]. Jack [Che91, SL90]. Jack-Knife [Che91, SL90]. Jackknife [LXZ16, Tho77]. JADE [VLIN21]. James [LB88, Yu11]. Jenkins [Lau76, LT77]. Johansen [ACR16, DH16, Oja16, Ron16, Zwa16, BDP13, Doo16]. Joint [DP18, EGM^{+}03, GGG13, LLY18, MW08, Mes22, PdT91, CLR19]. Jonckheere [Jon01a]. Jonckheere-Type [Jon01a]. Journal [Ano98e, Ano24, Lav23]. Jump [BDW16, CDGCK15, EB08, ABB24, AG20, RW13]. Jumping [TH01]. Jumps [Kur18, Man09, SW18, Koi14, LM23, MV20]. Juxtaposition [KM95a].

Kalman [SO13]. Kaplan

[Dab92, HGB96, Kle91, Stu94, Wei93].

Keeping

[Lav23, Ric23, Ano24, Gas23, Gre23b]. Kendall [HV22, Que12]. Kernel [CL05, Cha15, CWH05, DH05, EL96, FSGMM16, GM84, GM98, Gri09, HG85, HFP24, Jac00, JK92, Lou98, MBN17, MGSFB08, Mül85, Mül93, Nie98, NT01, PS99, Rud82, SHD94, SV05, VU05, Vie99, Wan90, WG96, Xia94, XZ09, YD07, Zha96, CL19, CXW23, FR21, HMP22, HNRT22]. Kernel-based [XZ09, CXW23]. Kernel-Type [VU05]. Khintchine [IS99]. Kiefer [Kni98]. Killed [BD13]. Killing [DD22]. Kind [Ber75, Lin00]. kinematic [AHP^{+}18]. knee [AHP^{+}18]. Knife [Che91, SL90]. Knowledge [GL02]. Known [BDV06, LB94, MS78, ST81, BR17, JLRT19]. Kolmogorov [AM84, Gui86, Præ95]. Kriging [LY03, DMK24]. Krylov [Bjö10, LCZW22]. Kurtosis [CJ08, Oja81].

Labelled [Cor03, PW10]. Lack [Sun10, CAVGM21]. lack-of-fit [CAVGM21]. Lagrange [SL88]. Lambda [BW04]. Landmark [PW10]. Landmarking [Van07a]. Langevin [Kle16]. Lanke [KD84]. Laplace [EGPS98, MR14, WLX19, WWW15]. Large [Ahm17, BS21, Edw80, Hjo88, JM16, JP06, LM16, LL90, LLY18, LCZW22, Lou98, RZM16, Sch02, TCK^{+}23, WZ22, BMP19, KK19]. Large-Sample [Hjo88, LL90, BS21]. Large-scale [RZM16, TCK^{+}23, WZ22]. Lasso [KYZC21, PRS^{+}22, ZHL17, CRS24, DBJ^{+}22]. Lasso-Zero [DBJ^{+}22]. Last [Stu83]. late [HMR21]. Latent [And82, BG14a, CDMR02, Far15, ICG12, KS08, MR14, PCW02, SRH07, SR11, ZLSL14, CGGI19, CV22]. Latin [BM01a]. Lattice [And90, Eks08]. Law [Zha96]. Layout [DFI14]. LCV [LC11a]. leading [PVFF24]. Learning [CCV23, PNC17, vdWBM19, EB23, FHSZ19, LZC23]. Least [AR94, AC99, AOH00, BIP14, Gré93, GP89a, GP89b, Hel90, LC00b, MSR16, Nor75, SS98, Sun96, Ter81, Ter83, ZG03, ZZLZ16, BD20, TB22]. Left [AJRN16, EMS15, Gär03, KY12, WCJ18]. Left-Tail [KY12]. Left-Truncated [Gär03]. Left-truncation [EMS15]. Lemma [KP77b]. Length [CLSZ16, FW03, GK86, HJO15, HCS15, KR15b, Nat75, Nor86, QQZ16, SLB06, VS07, BCCH19, RFK22]. Length-biased [CLSZ16, HCS15, QQZ16, BCCH19, RFK22]. Length-Frequency [HJO15]. Length-interacting [VS07]. Lengths [JGØ79, MC03]. Level [Gås16, GR05, Rue97, SV76, DR18, Hua24, Pap08, SKO17, ZV21]. Levels [Bøl88, BKO11, DO05]. Leverage [BCS13, Jør92, WHF98]. Lévy [BQ22, BNS05, BNL07, FL11, JSDT11, JKM19, ýJRNMJ13, Mse22, dRSHK19]. Lévy-based [ýJRNMJ13]. Lévy-Driven [JSDT11, JKM19, dRSHK19]. Lexis [BCG08, Lun00]. Lexis-Diagram [BCG08]. Lid [BG24a, BG24b]. Life [ABKT80, Aly90, ABH^{+}85, Bor99, BJMP14, Dok80, Dok82, GK03, MMS16, Nor86, SJ93, Thy75, BCCH19]. Life-Testing [Thy75]. Lifetimes [Asm89, BCG08, GN98]. ligament [AHP^{+}18]. Light [Nat85]. Likelihood [Adi97, Agr93, AL98, AL99, AG90, AH92, Aug04, Ban05, BL83, BNHJP76, BN84, BN85b, BN90, BNC91, BNB93, BH99, Bel03, BG16, Ber16, Ber79b, BM03, BKS76, BR17, Bie07, Bor84a, Bor84b, BW07, BW08, Buh93, CDMR02, CM84b, CHW^{+}07, Che15a, CM04, CFJP07, CRI03, CH04, CYM93, CGP07, CW99, CK06, DD88, DH08, DL89, DGCS13, DS94, DNL10, DE06, DC00, FLS05, GCJL03, GMS93, GM94, GWP89, Gil92, GV93, GS99, GDS88, GSG96, Gri09, Gro12, GH87, Gua07, HST74, HL99, HP00, HW17, Imo15, Jag77, Jen79, Jen93b, JQ15, Joh78, Joh97, Jon01b, KO03, KL78, KR01, KS94, Kuh04, Kün83, Laa88, Lan13, LL09, Li01, LV02, LGP11, LO16, LXZ16, LQZR09, Lok07, LVV09, MG95, MH97, MS78, MH10b, MW93]. Likelihood [NGAS92, NGZ18, Nor80, Pal09, Par01, Ped95, PM03, QW96, Qin98, QJ01, Ran84, RS94, ST10, SM04a, Sko81a, Sko01, Slu92, Sør03, SLCR14, SSZ09, Sun74, Tan09, Toc01, TZ95, WR02, Wan06, Wan08, WLT15, Xue09, Xue10, YK16, Ytt91, Zha98, ZG03, ZHH10, ZX96, ZHF03, vHV85, BSO22, Ber23, CDQ20, GLQ18, KL22, LP20, LM24, LPPW22, OHN21, Par20, PNC17, PL23, RW13, RSTU21, SJKS22, SJS08, Tak23, WC21, XT20, ZHW19, ZBS20, ZWS19, ZKP^{+}24]. Likelihood* [SH02]. Likelihood-Based [LQZR09, WR02, YK16, BSO22, CDQ20, WC21]. Likelihoods [DSJP14, Sun10]. Likely [HMB18]. Limit [BJ89, Deg96, FL11, GCL87, Hel82, Joh78, KM94, LLY18, LdUád15, Lou98, Mol98, Mur95, Ner98, Pol95, SMV05, SW18, SZ95, BW19, BMP19, KH22, MS98b]. Limiting [GJ83, Jon78]. Limits [BN85a, Höp90, TWL18]. Lindisfarne [HS95]. Line [CM01, Mab17, PS89, Van98, Wij95, vdL96]. Line-Segment [vdL96, Van98]. Linear [And90, ABN12, Ano83i, ACR16, BZF08, BNP92, BIPV13, BS10, BG98, BBM06, Bla01, CS03, CFMS03, CD03, CM82, CQ02, CTGS14, CLSZ16, CW16, CK06, DP04, DEV20a, DEL92, DW16, EMR09, FT16, Far15, FRS99, FS12, FMS11, FZ06, Gao98, GMMT06, GMA11, GM83, Grø97, Gro96, HGB96, HO93, HJKQ18, Hög79, Hol93, HK15, Hor85, Hou86, HS04, HZZ07, Hua13, HC17, IYW14, ICG12, JM16, Jen81a, JLY06, JN16a, JN16b, KM95a, KSM87, KS88, KG18, LQZR09, LB94, LAKZ12, LHW^{+}16, LS15, LQ17, MSR16, MBN17, Mej85, MS78, MS94, MR12, MvdG15, Nie84, Nor75, Oja16, OS96, PLHS17, PF08, PS83, QJ01, Rah86, RV04, Ron16, Rov15, ŠBD05, Sas92, SBV11, SFW16, SU92, SZ20, SV76, ST81, SZ02, TTZZ18, Ter81, Ter83, Tjø94, TCC^{+}95]. Linear [TDR09, Tra11, Tre83, Waa06, WR02, WFC16, WZ10, WLT15, YL14b, YZZ11, YL04, YWK06, Zah96, ZLL^{+}16, ZL18, ZHF03, ZIS09, Zwa16, vR94, vR95, AHP^{+}18, BS21, CW19, CHI23, CK23, CM20c, CFL24, DLS^{+}24, EB23, FB20, FHY24, GPVCGM16, GPÁLÁPGM21, GPST23, GKL21, GRS22, GJ16, HBD^{+}20, HFS23, HMR21, KK23, LW23, LYW22, Lue15, MR24, Mse22, RMG19, Sak19, Toc01, TKU23b, VD18]. linear-directional [GPVCGM16]. Linear-Index [HC17]. Linear-Representation [FZ06]. Linearities [SL88]. Linearity [LST88]. Linearly [OKW88, ZHL17, CFL24]. Lines [Sch80]. Link [CM04, ZL18, SKR19]. Lipschitz [DD22]. List [BT08]. living [BBP21]. Loading [WW01]. Loads [Ryc96]. Local [AP07, Ban05, BJ89, BKT20, BQ09, BR14, CS03, CQ02, CYL11, CK94, CYM93, Die92, EGB13, EBGG18, FK98, GM98, GNPM07, HVV14, Haz96, Höp90, HL99, HS17, Jun08, KO03, MG95, MF97, Mur95, OB16, PP16, SW18, SSZ09, WG96, Xue10, XT20, YH20, YL04, ZHF03, BHLP19, GBB^{+}24]. Locally [Che09, GME24, OHIS24, PVD13, SS02, SP09, VM00, XLS16, PRV21]. Location [Ahm17, AO11, Ano83i, AGM00, Arc98, Dok75, EHR88, GMPFV11, GM83, GM82, KRV07, Kou85, Miu81, Mül85, NL16, Oja81, Sch75, Tan09, DD22, BMP19, KV23, Kut19, LLLP20, Sch81]. Location-invariant [Ahm17]. Location-Scale [GMPFV11, NL16]. Locations [LR08, Nor90, Kop23]. locus [LLLP20]. Loess [BH14]. Log [Aug04, BN90, BNC91, BV09, BM01b, Che15a, CMW17, De 06, FT16, GMA11, Jen93b, Kle24, MSW98, Rov15, SZZ05, Var76, FR21, RSTU21]. Log-Change [Var76]. Log-concave [Che15a, RSTU21]. Log-density [BV09, Kle24]. Log-Gaussian [De 06]. Log-Likelihood [Aug04]. Log-Linear [GMA11]. Log-mean [Rov15]. Log-Normal [FT16]. Log-Rank [SZZ05, FR21]. Logarithm [Zha96]. Logistic [AH87, DH07, FLS05, GM16, Kol97, TL03, fWZY16, MS23, SW19]. Logistic-Cox [fWZY16]. Logistic/Proportional [FLS05].

Logistic/Proportional-Hazards [FLS05]. Logit [Nor81]. Loglinear [Agr93, GDS88]. Lognormal [AJRN16]. Logrank [BJMP14]. Logspline [Koo99, KKP99]. Long [BG98, DRT13, TRL15, YZ12, AAFO20, BT21, BBD^{+}21]. Long-memory [BG98, TRL15, AAFO20]. Long-Range [DRT13, BBD^{+}21]. Long-Term [YZ12]. Longitudinal [AP04, AJ00, BZF08, CY17b, DP18, HZZ07, LZ10, LLY17, LZZ14, MS01, Mül05b, Mur95, QST08, Sch94, SPK23, SMS12, SLCR14, SW05, SBB05, SJS08, Van07b, AV21, CW19, CZT20, DQR21, HYZ22, MS24, MWW24, NR23, SLCN19, YC22, ZLK21, ZXLL23]. Longitudinal/Clustered [HZZ07]. longitudinal/functional [HYZ22]. Look [BHR^{+}76]. Loop [SV10]. Loss [FT16, FRS99, Lin78c, Rue97, Sen88, Tra11]. losses [CL20]. lot [vdWBM19]. Low [Hög79, BC23, GW24]. low- [BC23]. low-dimensional [GW24]. Lower [Efr05]. LR [Cer17]. Lumping [Lin78c].

M [Wu13, Hok75]. M-estimation [Wu13]. MA [Ter77b]. machine [EB23]. Macroeconomic [JJ02]. Magnetic [Har02, JT07, LLS^{+}22]. Majorized [Tor88]. Make [WL18]. Making [Lav23, Ric23, Ano24, Gas23, Gre23b]. Malignant [DE82]. Manifold [BG14b, Kle16, Kle24]. Manifolds [BG13, BG14c, Dry14, Sim14, SLS14, Ken14]. Mann [FOS^{+}14, Zet88]. Mantel [GL07]. Many [Che13, Ros89]. Map [LS98]. Mapping [KHR02]. Maps [HH16]. March [Cor23]. Marcinkiewicz [IS99]. Marginal [DSWH09, GOV15, Gil92, HZZ07, JLY06, Joh17, KL78, LO16, MS24, PNC17, PM03, Rov15, SBB05, YY15, CM20a, CGGI19, LKT^{+}23, Sak19, VHF20]. Marginally [QZP12]. Marginals [KS99, Sun75]. Margins [Eva16]. Mark [GJW12, Joh17, SG12, JvdMP22].

Mark-Specific [SG12]. Marked [AH84, CDDL12, GSG96, PBHMC09, SO97, ICM19, vL18]. Markers [DMPV02, HC10]. markets [LM23]. Markoff [OKW88]. Markov [AJ78, Aal87a, ABB24, AR80, AHK91, AMP97, AMP01, AT15, AGGM06, Asm89, ADGP14, Bon79, Cav23, CCH01, CJGPL07, DH08, DSJP14, Did07, DT20, DE06, Edw80, Far15, FW03, Fry90, GCJL03, Gil86, Got94, HN99, Häg07, HV08, HJR06, Höp87, HK97, HPS24, JXCK14, JR07, Jen89, KHR02, Kos99, LCF24, LRSS23, Lin78a, Lin77, Lin78c, Lin88, MBMG23, NH15, Nic14, NBW02, PWN22, PKH17, PNC17, Ran78, Rap03, Ric03, RR01, RT02, Ryd95, SV10, SPK23, SPR^{+}13, Ste91, Sun75, Tho81, TB98, VS07, VHF20, Wol24]. Markov-Type [Sun75]. Markovian [BKKR23, Die92]. Marks [Joh17, MW08, SG12]. Marrow [AK07]. Martingale [AKB^{+}89, Gri80, KP02, Ohl86, Uch04, Wal24]. Martingales [HM99, Hel82, Höp90, Mur95, SZ95]. massive [HLP23]. Match [Cer17, SMZ11]. Matched [BO99, Bro87b, Dok80, Dok82, GH87, Ros89, WGT19]. matched-pairs [WGT19]. Matching [AKC80, CRCV12, YK20, YZ23]. Matérn [Bol14, RLOS18, WB15]. Mathematical [BRH83, Lok07]. Mating [Sch79]. Matrices [Ahm17, BDL^{+}17, DH05, LLY18, LHHF13, Nor75, PS83, SG04, ST76, Tho83, BS21, CLP^{+}19, HK24]. Matrix [AJ78, Asm00, BZ82, Gui77, Hol80a, MG95, WR93, BMP19, HK24, LC22]. Matrix-Analytic [Asm00]. matrix-valued [HK24]. matrix-variate [BMP19, LC22]. Matter [Ano74a, Ano74b, Ano74c, Ano74h, Ano74i, Ano74j, Ano75b, Ano75c, Ano75d, Ano75h, Ano75i, Ano75j, Ano76a, Ano76b, Ano76c, Ano76d, Ano76i, Ano76j, Ano76k, Ano76l, Ano77a, Ano77b, Ano77c, Ano77d, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79j, Ano79k, Ano79l, Ano79m, Ano80a, Ano80b, Ano80c, Ano80d, Ano80i, Ano80j, Ano80k, Ano80l, Ano81a, Ano81b, Ano81c, Ano81d, Ano81i, Ano81j, Ano81k, Ano81l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82i, Ano82j, Ano82k, Ano82l, Ano83a, Ano83b, Ano83c, Ano83d, Ano83j, Ano83k, Ano83l, Ano83m, Ano84a, Ano84b, Ano84c, Ano84d, Ano84i, Ano84j, Ano84k, Ano84l, Ano85a, Ano85b, Ano85c, Ano85d]. Matter [Ano85i, Ano85j, Ano85k, Ano85l, Ano86a, Ano86b, Ano86c, Ano86d, Ano86i, Ano86j, Ano86k, Ano86l, Ano87a, Ano87b, Ano87c, Ano87d, Ano87h, Ano87i, Ano87j, Ano87k, Ano88a, Ano88b, Ano88c, Ano88d, Ano88h, Ano88i, Ano88j, Ano88k, Ano89a, Ano89b, Ano89c, Ano89d, Ano89h, Ano89i, Ano89j, Ano89k, Ano90a, Ano90b, Ano90c, Ano90d, Ano90i, Ano90j, Ano90k, Ano90l, Ano91a, Ano91b, Ano91c, Ano91d, Ano91i, Ano91j, Ano91k, Ano91l, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d]. Matter [Ano96f, Ano96g, Ano96h, Ano96i, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano98d, Ano98f, Ano98g, Ano98h, Ano98i, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano07b, Ano07c, Ano07d, Ano07e]. Matter [Ano07g, Ano07h, Ano07i, Ano07j, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano10b, Ano10c, Ano10d, Ano10e, Ano10g, Ano10h, Ano10i, Ano10j, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Häg07]. Max [HOT21, OBL18, GKL21, Sei24]. max-domains [Sei24]. Max-infinitely [HOT21]. max-linear [GKL21]. Max-Stable [OBL18]. Maximal [Ner77]. maximin [LMH22]. maximizing [LT21]. Maximum [Agr93, AL98, AL99, BL83, BKS76, Bie07, Bor84b, Buh93, CDMR02, CM84b, CHW^{+}07, Che15a, CRI03, DGCS13, DE06, Eks01, FLS05, GC05, GWP89, GV93, GDS88, GSG96, GH87, HW17, Joh78, Jon01b, JM83, KL78, KL22, Kuh04, KR15c, Kün83, Laa88, Lan13, LP20, LM24, Lin76, MG95, MH97, MS78, MW93, NGAS92, Nor80, OHN21, Ped95, Ran84, RSTU21, RS94, SM04a, Sko81a, Sun74, Tan09, TZ95, Wan08, WW11, ZHH10, ZX96, GLQ18, KR20, LL09, LPPW22, Tak23, Bor84a]. Maximum-Entropy [JM83]. Maximum-Likelihood [CHW^{+}07]. Maxiset [AFV14]. MCAR [ZHW19]. MCMC [BM03, EB08, HV05, HJKQ18, NX17, Rap03, TH01]. Mean [Abt99, Ala77, Aly90, Bon76, BF03, CGL14, DP04, Dha16, DN15, DY17, Erl81, FT16, Ham88, Jen86, KL78, Kor00, MMS16, Miu78, OS96, Pen95, RV04, SC06, SW76, Sun95, Xue09, BMP19, HFP24, JN19, LPYZ24, PS20, Rov15, YK20, YLZ^{+}19]. Mean-Based [RV04]. mean-squared [JN19]. Means [Bac11, Chr74, CS82, GS80, GM08a, Hin79, Huc11, JS12, JKR02, MRS14, Ped00, RR95, STK17, TDR09, WW11, Ter14]. Measles [BKO11]. Measure [Dab96, DGN07, DSS13, Erh08, GN09, Lav23, ML74, Ric23, Xie88, Gre23b]. Measured [ADL15]. Measurement [Aug04, GSK06, Guo11, HT14, HL02, HWH15, KHL98, LR06, Sch94, SFW16, TL03, TDR09, TRL15, Wan08, XLS16, JV24, SAS24]. Measurements [Cro98, Mar99, Swe88, EB23]. Measures [Dok75, EGG14, FPW11, Gås16, GH89, GK86, JLP09, KZ17, Nat85, NEV13, TF12, VOG11, ZIS09, BQ22, BDS22, EK22, HFP24, vL18]. Measuring [ACFS83b, EGB13, KHT14]. Mechanics [RVG15]. mechanism [CDQ20]. mechanisms [PVFF24]. Median [AM84, DP04, GK03, Gui79, Lan74a, GGS20]. Mediated [Van11]. Mediation [AGR^{+}18]. medical [KHSJ19]. Meier [Dab92, HGB96, Kle91, Stu94, Wei93]. MEL [CCWZ19]. Melanoma [DE82]. Members [LA16]. Membrane [JGØ79]. memory [AAFO20, BG98, BT21, TRL15]. Menarcheal [SSD15]. Merge [SM12]. Meta [BG16]. Meta-analysis [BG16]. Metastases [DE82]. Meter [TW04]. Method [BKM18, BR81, BG01, BT08, BG11, CDMR02, CB84, CDDL12, GM84, GWP89, GV93, GSG96, GP89a, GP89b, HM02, Kni98, KM95b, LDY16, LCZ09, Lue15, Mam92, Min79, Min81, Mül93, Qin98, QQZ16, Ran84, SA11, Sve90b, Tri03, ZHH10, CM20c, CXW23, DZ21, JN19, Sak19, ZSJT24, ZMH^{+}24]. method-of-payments [JN19]. Methodologies [Arj11]. Methodology [BKO11, SBM^{+}99]. Methods [Bel03, Bon10, CTGS14, Cle97, DDL14, Eub00, GK13, GCLP92, Hjo88, LT77, LZ99, PS10, See93, SKO17, Tho95, VHK11, VR08, ZYT02, Kle24, NG23, SW19, YM22]. metric [Kle24, MM24]. Metropolis [CV02, Gås03, HT08, SR03, Sto11]. Microarray [FOS^{+}14]. microeconometrics [Kop23]. Microscopy [AHJ15, BBP21]. microstructure [WLX19]. Mind [HK99]. Mineo [BR81]. Minimal [BNHJP76, Jag77, KL89, Lin94, Sat96]. Minimax [Ala77, BG80, Bla01, Hol75a, Kle99, ZXLL23, Mei24]. Minimaxity [Sun95]. minimization [ZSJT24]. Minimum [CD96, DMV16, Glo14, Hoe76, HK97, Lud04, DEV20a]. minimum-distance [DEV20a]. Minorant [PW06, vEvZ96]. Mises [BB11, ELY22, GWP89, GV93]. Mismeasured [YZ07]. Missing [BCH16, øBFHB07, Che09, Dam80, DR96, DY17, GWT00, HCS15, JW10, MS98b, Nie97b, NYR18, RR95, SC06, SFW16, SG12, TTZZ18, WR02, WZH16, WL18, Xue09, YK16, Yu11, ZIS09, CCWZ19, CDQ20, DBJ^{+}22, DLS^{+}24, HFS23, LZC23, LW23, LLXH19, ZWS19]. Missing-at-Random [YK16]. missingness [CDQ20, ZMH^{+}24]. Misspecified [LL96]. Mixed [CDGCK15, CW99, EMR09, Far15, FM89, GWT00, Höp90, KF07, LAE^{+}89, LZ10, LCZ14, Lin78a, MSZ16, Mej85, PDD10, QST08, Ric03, Rit04, SBV11, SY00, Ter81, Waa06, WZ10, XZ09, ZLL^{+}16, ZLSL14, CHI23, EB23, FB20, LYW22, SK20, Sak19, SK19]. Mixed-effect [XZ09]. Mixed-Effects [PDD10, QST08, SBV11, ZLL^{+}16, CHI23, EB23, Sak19]. mixed-model [SK19]. Mixed-outcome [ZLSL14]. Mixing [Aal87a, BKO11, Ege92, Jen89, LZ97, MF97, BW19, DO05, Wol24]. Mixture [BF02, BB11, BDV06, Cav16, CHW^{+}07, CL01a, CK97, FLS05, FRZ16, GJW08, HJO15, HYWC18, JSW91, Kor00, LL09, LPS03, NC15, Qin98, WWP14, fWZY16, Yu16, BMP19, KK23, LLYC22, MPV19, MPV24, YA20]. Mixtures [Bie07, BP05, BV14, Cha77, CP07, CRI03, FNR09, FH04, HMG06, Kim03, Laa75, Lyn88, MRM09, Sch82, Tan09, Whi86, JLRT19]. ML [HV08]. MLE [MW08, MB91, Ran78, YWK06]. MLEs [Jac89]. MMCTest [GH14a]. Modal [YL14b]. Mode [CSW79, JKN12, OS97, Sar09, TH01, ZL10]. Model [ACFS83a, ACFS83b, AG85, ABK96, And77b, AL98, AL99, AFL10, AH87, ACR16, Aug04, Ave86, BBG06, Bed93, BH14, BN15, BVV17, BDP12, Bon79, BDV06, BG14a, BC15, BHC88, BDH03, BCG08, CFMS03, CSW79, Cer17, CC98, CHW^{+}07, CDZ11, Chr74, CO07, CYM93, CLP17, Cro98, CPWZ13, DS03a, DLH14, DSD^{+}14, DM19, DH16, Doo18, EGB13, FLS05, FMS15, FRZ16, GN98, GJ05, GM16, Gär03, GS99, GA86, GJW12, GH18, Grø97, GP89a, GP89b, GS76, HV06, HK99, HHVA03, HBH17, Har02, HJO15, HESZ16, Hel00, HH82, Hoe78, HH16, Imo15, Jen87a, JT07, JFKC05, JH17, JW10, Joh97, JH05, KWA16, KKC17, Kle99, KF07, KM95a, KRV07, KHL98, Kor00, Kou84, Laa88, LY08, LC00a, LKN15, LDY16, LLY17, LS96, LR08, LB98, LN13b, LM18, LS15]. Model [MSZ16, MMS16, MW12, MS09, MP14, MM93, Mej85, MS94, MR12, MV87, MvdG15, Mur93, Næs82, Nat75, NC18, Nic14, Nie97a, Nie99, NBY08, Nor81, OH16, OKW88, OR94, OFFL12, Ped00, PBHMC09, PM03, PR07, PdT87, PV00, QW96, QQZ16, Ris81, Rob78, Rue97, San14, Sar09, STH^{+}78, Sas92, SGR11, SZS02, SM04b, SM04a, SMZ11, SMSD92, SG78, Sko84, Sko86, SV04, SZ02, SW05, SSZ09, Sun10, TM86, Tju82, TDR09, Tri03, Uta17, VW15, Von96, Wal97, WY03, WLS15, WZH16, WFC16, WL18, WC12, Win13, WW01, fWZY16, XL10, YZ12, YL14b, Ytt91, YD07, Zah96, Zha00, ZHH10, ZYT02, ZC03, dCCU17, BCCH19, BGH19, BSO22, CZL24, CN16, CK23, CM20c, CLR19, CV22, DHH24, DBJ^{+}22, DQYZ23, EU21, FGH20, GZZM23, GPÁLÁPGM21, GK21, GH21]. model [HFS23, JvdMP22, Kar20, KXZA20, LGL19, LW23, LLYC22, LJZ^{+}18, MS23, MT19, Mse22, MBMG23, NJG18, PRS^{+}22, PPS21, SK19, Tak23, Vid21, VMG22, XWH14, YK20, YM22, YH20, YLZ^{+}19, ZHS22, ZKP^{+}24, ZLZZ21, vdWBM19]. model-assisted [MT19]. Model-Averaged [KWA16]. Model-Based [CSW79, Hoe78, Jen87a, OFFL12, STH^{+}78, vdWBM19]. Model-free [DM19, WL18]. Modeling [HPS24, ZX19, BCCAVMO21, BEP20, LKT^{+}23, LCZW22]. Modelling [Aal87b, AO11, BN97, BM01b, DP18, DGGM16, EGM^{+}03, GR01, HS87, HKD02, Hou87, ýJRNMJ13, KK09, LM16, MS94, MDA10, Mül05a, Mül05b, NBW02, Pal04, RL06, SZ07, SRH07, Taq02, TL03, TW04, WB15, Heu05, Ize05, Ram05]. Modelling* [JJ02]. Models [Aal76, AGR^{+}18, Agr93, AALM17, And77a, ABH^{+}85, AK07, And91, And90, AC99, ADL15, AKB^{+}89, AT15, AGGM06, AJ00, AP07, Asm00, BZF08, BCCA11, BM15, BR03, BIPV13, BIP14, BS10, BN84, BNS03, BL08, BG98, BPS17, BB11, BBM06, BS01, BDP12, BJ85, Boe10, Bøl88, Bor99, Bor84a, Bor84b, BW07, BW08, Buh93, BB14, CS03, CAS03, CDMR02, Car82, CD03, Cav16, CM01, CH96, CCH01, CTGS14, Che15a, CM17a, CY17b, CL01a, CO07, CJGPL07, CGP07, CL12, Cor03, Cro00, CK06, DD88, DGN07, DH08, DH07, DCIK14, DSJP14, DS03b, DPV06, DPFV09, DRS09, DEL92, DE04, Did07, DW97, DSS14b, DSS14a, DFG00, DNL10, DE06, DW16, DR97, DC00, EMR09, Eri84, Eri96, FT16, FZ00, Far15, Fas16, FL11, FNR09, FZ06, GN95, GN09]. Models [Gås16, GGG13, GQR06, GWT00, GCJL03, GSK06, Gho06, GMMT06, GDS88, GMPFV11, GMA11, Got94, GJW08, GHC92, Hel90, Hel98, Hel00, HO93, Hjo86a, HV08, Hög79, Høj04, Hol93, HL99, Höp99, HK15, HS17, HS04, HZZ07, HP09, HWH15, HC17, HW17, HYWC18, Huc11, Huz99, IYW14, ICG12, JM16, JM01, JSDT11, Jen87b, JLY06, Joh82, JN16a, Jør86, Jun11, KL14, Kar15, Kau96, KO03, KK06, Kle81, KS08, KKP08, KH99, KHR02, KKMP18, Kol81, KSM87, Kou76, KS88, KL89, Lan13, LDW06, Lau74, Lau75, LBND^{+}84, LAE^{+}89, LS98, LHNN03, LL09, LL12, LN13a, LG09, LZ10, LGP11, LCZ14, LO16, LHWS18, LQZR09, LCZ09, Lin78a, Lin88, LPS03, LC11b, LG13, Llo88, Lo81, LZ97, Lok07, LMB09, LST88, MT03, Man09, MAR11, MR14, MS01, MSSM02]. Models [MU91, Mic09, Mol94, MDA10, Mül92, Mun02, NH15, NX17, NGAS92, Nie83, Nie84, NC15, NYR18, NL16, Nor77, Nor80, PLHS17, PF08, Pap00, PLKP06, PDD10, PCW02, PG13, PS83, Puk82, Qin98, QST08, QMP15, Rah86, Ran78, Rit04, RV04, RD10, RD17a, Ron16, RR95, Rov02, Rov15, SL88, SGR11, SZ07, SBV11, Sch02, ŞM05, Sha12, SB85, Son00, SR11, Sør03, SS09, ST81, SH21, SZ02, STK17, SKO17, STZ01, SG12, Sun75, Sun83, TWL18, TTZZ18, Ter81, TM86, Tho81, Toc01, Van01, Vid01, Waa06, WWP14, WR02, WCXS15, Wil77, Wil79, WL04, WZ10, XZ09, XMW15, XLS16, YZ07, YZZ11, YWK06, Yu16, ZLL^{+}16, ZL18, ZHL15, ZLSL14, ZHF03, ZIS09, ZYX14, Zwa16, vP92, vR94, vR95, AHP^{+}18, ACF^{+}21, ADMP19, BS21, CW19, CM20b, Cav23]. models [CHI23, CCWZ19, CZT20, CGGI19, CWZ21, DMK24, DR18, DLS^{+}24, EU21, EB23, FB20, FHY24, GZZM23, GPVCGM16, GPST23, GKL21, GJ16, HBD^{+}20, HHHL23, HMR21, HYZ22, HOT21, HNNS19, KK19, KL22, KHBK22, KYZC21, KK23, KG18, KMG21, KKW24, LW23, LET20, LYW22, LC22, LLYC22, MS24, MdCCD19, MWW24, MPV24, NR23, NHMW22, OHN21, ORL20, PRV21, PD22, RMG19, RFK22, SK20, Sak19, SM24, SPK23, STM22, dRSHK19, ULK23, VW19, WCJ18, WGT19, WHZ20, WCY22, XNL23, YC22, YA20, ZL22, ZD24, vLM23, CM17b, CW16, BN85b, Hjo86b, JN16b]. Moderate [EG02, MWY15]. Modern [MW07, Ano07f]. modes [BT21]. Modification [BNC91]. Modified [CFJP07, Hoe76, LB88, LC11a, LDA12]. Modulated [Ryd95]. Moment [Che15b, Cle97, Dal77, DW95, Ess75, KS08, Mil85, dRSS22]. Moment-based [dRSS22]. Moment-Generating [KS08]. Moments [BNS05, BN15, CS90, Kle91, LM04, LDY16, Lin94, Swe83, WZ10, vR88]. Monitor [EPM15]. monitoring [BBS23]. Monotone [Ban05, DR10, HW95, LM18, Nat93, Sti82, vHV85, BGH19, SFW16]. Monotonic [DS90, SA11]. Monotonicity [Aly90, BBG06, BN13, BJD82, Ner98, PS83, LAO23]. Monte [BG14b, Dry14, Ken14, SLS14, Sim14, BG13, BG14c, CDMR02, GH14a, GH16, GHD20, JR07, JSDT11, Kle24, LRSS23, LET22, NH15, PWN22, Sak19, SPR^{+}13, SW75, SW76, VKY^{+}14, VHF20]. Morphisms [KM91]. Mortality [Bro87b, Gar82]. Mosaics [Van13]. Most [DF03, HMB18, GAC23, GJ16]. Mother [Bro87b]. Motion [LMT14, LM24]. Motor [dMR88]. Moving [BDY85, HP00, LP01, SW04, GRS22, KP21, LP22]. Multi [Ahm17, BM16, BM01b, CJGPL07, MH97, OS96, Pap08, PBHMC09, SW93b]. Multi-Dimensional [MH97, BM16]. Multi-level [Pap08]. Multi-Sample [SW93b, Ahm17]. Multi-scale [PBHMC09]. Multi-state [CJGPL07]. Multi-Univariate [OS96]. multiclass [DDM20, DDBEMT24]. Multicolour [Van13]. Multidimensional [And74, Kre87, LP22, Sun75, WCSftADNI23]. Multilayer [FV06, AHWKP19]. Multimodal [Sun10]. Multinomial [Ber81, Hol81a, HS95, Lan13, Wan86]. Multinomial-Poisson [Lan13]. multioutcome [KMG21]. multiparameter [BEP20]. Multiple [Aal76, Amu74, BM15, BS16, BH97, Ber76, CPS20, CM04, CYM93, CMMR12, Far09, FR00, GH14a, GH16, GPP96, Han16, Hjo88, Hol79, Kab78, Kor82, KA06, LB88, LHML16, Mad76, Mei06, NM14, SA11, SC06, Spj74, Sve76, Sve90b, Vie99, ZLSL14, BKN23, GK21, LL20, Par20, WHZ20]. multiple-choice [LL20]. Multiple-output [CPS20]. Multiple-Recapture [Ber76]. Multiple-Sequence [SC06]. Multiplicative [And77a, DPFV09, MMS16, MSSM02, Nie98, SZS02, Tju82, Vid21, YLZ^{+}19]. Multiplier [SL88]. Multiply [Cac77, WHR22, YZ23, CH23]. Multiscale [DEV20b]. Multistage [GH12]. Multistate [AK07, BB14, CGP07, MP22, Nat93, SZ07, dMR88]. multitreatment [LYW22]. multitype [MP22]. Multivariate [Ala77, AVCRG13, Azz05b, BNK99, BEK83, Ber77a, BR17, BJD82, Bøl82, BZ82, BW04, CM84b, Cha84a, CCH98, CDY11, CK97, Cro98, DH78, DP13, DCIK14, DR96, DEL92, DT05, DH05, Erl81, Fas16, GA86, Got94, HHVA03, HPR21, HKŠ22, HJKQ18, Hou87, Imo15, JLY06, JKN12, Jun11, Kim97, KHBK22, KKP08, KV98, KR15c, LV13, LHW^{+}16, MS86, MMO23, MH10a, Mic09, Nic14, Nor86, Oja99, Par01, QZP12, ŠBD05, STM22, SMB14, SY93, SG04, Sko84, Son00, Sun95, STZ01, Sun96, SBM^{+}99, VB99, Van01, Wil79, ZLK21, vR95, ABY22, BD20, BQ22, BS21, CN16, CFL24, DEH21, HD22, HPS24, KK23, KR20, Par20, dRSS22, vL18]. Munch [McG88]. Mutual [AVCRG13].

n [SW04, TvdM96]. Nadaraya [CL19, MR23]. Naive [GNPM07, RFK22]. Natural [Bar03, CV01, DLR18, EGPS98, GM08a, Pic00]. Natvig [Xie88]. NBU [Wan87]. Near [HST12, MR14, BJ89, Mül93]. Near-Gaussian [MR14]. Nearest [Cey10, DM80]. Nearly [BSC24]. necessarily [LCF24]. Necessary [KJH16, Ran78]. Need [Pfa93]. Negative [Bøl82, BJ12, DM83, HH16, McK87, Ros77, Vai91, BSO22, Mab17]. Negative-Binomial [HH16]. Neighbour [Cey10, DM80]. Neighbourhood [MWY15]. Nelson [GNPM07, Kle91]. Nested [BO99, Lok07, MR14, Nor77, See96, TDR09, BKN23, NG22, ZMH^{+}24]. Network [ABN12, MBN17, RD17a, Van13, SPK23, WCSftADNI23]. Networks [BB14, Eva16, GWH11, JT07, PNC17, DQYZ23, KH22, Mei24, MR24]. Neumann [Che09]. neural [WCSftADNI23]. Neutral [WD98]. neutron [AHWKP19]. Neyman [JVA11, Lus94, SN88]. Nickel [BMG82]. Nielsen [ACR16, Doo16, DH16, Oja16, Ron16, Zwa16]. Nils [BG24a, BG24b]. Nitrous [Ped00]. No [Eub00, Gil86, Rov05, TS91]. Node [Gås16]. Nodular [DE82]. Noise [Bol14, Cuc08, DP16, JH05, Kle99, Kur18, MDA10, MNS07, Nie83, Puk82, Shi17, NU19, WLX19]. noises [CZL24]. Noisy [Ant96, Bib11, BR14]. Non [AJ78, AV01, AHK91, AGM00, Arf98, ADGP14, BB10, BNS03, BPW14, BL94, BSV13, BM03, BCCAVMO21, Bol14, Bøl82, BDH03, BW04, CCH98, CHW^{+}07, CH04, CJGPL07, Dab87, DLS96, DW02, DH08, DS09, DRT13, DS90, DSS13, DLR18, DBD18, DPT13, DW97, Die92, Efr08, Eks08, EV08, EGG14, Erh08, EW94, FHT94, FVV10, FL11, GSYB05, Gao98, GL02, GCLP92, GWP89, GV93, Gla98, GMPFV11, GL15, GJW08, Gui82, GG01, Gui04, Haa08, HVA00, HHVA03, HBH17, HA98, HJKQ18, HHL02, Hor85, Hou86, HTK15, HP00, HS04, Huz99, Jac00, Jen87b, JH05, Jun11, KB04, KN12, KS08, KKMP18, KY12, KS88, KS01, La 08, Li01, LV02, LT08, LLY17, LdUád15, LB98, LVV09, LN13b, Lue15, Mab17, Man09, Mol94, MW97, Mül93, MNS07, NX17]. Non [Ner77, NV09, Neu09, NGZ18, OT09, OSG08, OFFL12, PFJGE15, PKR^{+}97, PLKP06, PV00, Pre03, PK18, RD10, Rom04, Ros77, Rov02, SA15, SL88, Sam89, Sas92, SS06, SFW16, Sko81a, SR01, SJS08, SV05, Tjø94, Toc01, TCC^{+}95, THSS09, TC05, VM00, Von96, WD98, WB15, Wan90, Wan08, Wij95, XBQF15, XZ09, XY15, Zah96, Zha95, Zha08, ZX96, dCJV82, vEvZ96, CCWZ19, DT20, MBMG23, TKLM23]. Non- [DLR18, GWP89, GV93]. Non-and [AHK91]. Non-Bayesian [KKMP18]. Non-Central [BW04]. Non-centred [NX17]. Non-Concave [THSS09]. Non-Confounding [GL02]. Non-Critical [Ner77]. Non-decomposable [Rov02]. Non-discovery [XBQF15]. non-equivalence [CCWZ19]. Non-Ergodic [Jen87b]. Non-Exponential [Huz99]. Non-Gaussian [BNS03, BPW14, BCCAVMO21, Bol14, BDH03, HJKQ18, KS08, OT09, WB15, TKLM23]. Non-Homogeneous [AJ78, AHK91, BSV13, ZX96]. Non-IID [Gui82]. Non-Invertible [HP00]. Non-Linear [Gao98, Hor85, Hou86, HS04, Tjø94, TCC^{+}95, Lue15, Toc01]. Non-Linearities [SL88]. Non-Markov [CJGPL07, DT20, MBMG23]. Non-Markovian [Die92]. Non-monotone [SFW16]. Non-Monotonic [DS90]. Non-Negative [Bøl82, Ros77, Mab17]. Non-Orthogonal [Sas92]. Non-Parametric [AGM00, Arf98, ADGP14, BB10, BL94, CH04, Dab87, DLS96, DRT13, DW97, EW94, FHT94, FVV10, FL11, GSYB05, GCLP92, Gla98, GMPFV11, GJW08, Gui04, HHVA03, HHL02, Jac00, KB04, KS01, La 08, Li01, LB98, Mül93, PKR^{+}97, PLKP06, Pre03, Sam89, SS06, SR01, SV05, Wan90, Wij95, Zah96, Zha95, Zha08, dCJV82, vEvZ96, AV01, BM03, CCH98, CHW^{+}07, DW02, DS09, DSS13, DBD18, DPT13, Efr08, EV08, EGG14, Erh08, GL15, GG01, Haa08, HVA00, HBH17, HA98, KN12, KY12, LT08, LdUád15, LN13b, Man09, MW97, MNS07, NV09, Neu09, NGZ18, OFFL12, PFJGE15, PK18, Rom04, SA15, VM00, WD98, Wan08, XZ09]. Non-Proportional [Von96]. Non-Random [Mol94]. Non-Separable [RD10]. Non-Smooth [LVV09]. Non-Standard [HTK15, DH08]. Non-stationarities [OSG08]. Non-Stationary [KS88, SJS08, TC05, Eks08, Jun11, LLY17, PV00, XY15]. Non-Uniform [JH05]. Noncurved [Ste91]. nonignorable [CDQ20, GLQ18, ULK23, ZWS19]. Nonlinear [DLP08, Dem17, PF08, ST12, CXW23, GBB^{+}24, VW19]. nonneutral [SJKS22]. Nonparametric [Aal76, ABKT80, AHP^{+}18, APCG24, BD07, BBP21, Cha84b, CGC23, DDBEMT24, ES91, GM23, GSUC22, Hei19, HNRT22, Joh17, LPB15, LPYZ24, Lo81, Mac82, MN21, MW93, QB23, Rei81, SN13, TZ95, vL18, CLP18, CL19, GMvdM20, HMP22, HNNS19, JvdMP22, MP21, SW19, SAS24, ZL22]. nonpolynomial [GPST23]. nonprobability [CH23]. nonproportional [NG22]. Nonresponse [And79, FMHB16, GLQ18, ULK23]. Nonsmooth [LXZ16]. nonstationarities [VW19]. Nonstationary [Nie83, SP22, LPW21]. Normal [Ala77, AVA06, Azz85, Azz05a, BN97, BNS05, BF02, BO11, BJD82, Bøl82, Bon75, CRCV12, CAS03, CRI03, Cur80a, DP04, DS94, DC00, Erl81, FT16, FWW77, Ham88, Hen86, HST74, JQ15, Kor00, LL09, Llo88, Lon12, MH10a, Nor90, Ryd95, Sko84, Sun95, Whi86, AVA22, Azz05b, BMP19, DR18, Gen05, HV22, JLRT19, OH16, PRV21]. Normal-Gamma [Whi86]. Normalised [KM94]. Normality [Awa81, BP05, Eng80, Höp90, Jen93a, McG88, Nor80, Ohl86, SW76, Tho77, VU05, DEH21, KR20]. normalization [ZLS14]. Normalized [JLP09, MSR16, TF12]. norvegica [Sch79]. Note [AL81, AL99, Ano83i, BR03, BNK74, BN85b, BN90, Ber76, BR81, Bly93, Bor84a, BL90, BW08, Bro87a, CM17b, DS03b, Dok82, DR00, Fra77b, GH08, GP89a, GP89b, Hin79, Hjo86b, Hoe78, Hol81b, HK15, JM16, JM83, KK06, Lai80, LR76, Lau76, LZ99, Lin78b, Min81, OS97, PS92, Sch81, Swe83, Væt79, Van01, Var79, Wal00, Xie88]. Notion [ML74]. Novel [YLGL20, DEV20a]. November [Cor23]. NPML [SV04]. NPMLE [VJ01, vdL96]. Nuisance [BW08, LL12]. Null [Aar85, Höp90, MR10, NM14, MV20]. Number [BDL^{+}17, Cha84b, DLH14, Fra78, LY08, Lee97, LQ17, NC92, NC18, Nor90, PK18, See93, Thy75, HPR21, LT21, XNL23]. Numbers [Ber75, McG88, Var76, KH22]. Numerical [Lan13]. Nurminen [Ano96e].

Oakes [GS99, PM03]. Objective [BGL13, CL12, CLP17, GMMT06, KSSR21, VW15]. Objects [Ber79b]. Observation [AHK91, AH92, DP18, HJ04, SMS12, CL20, OPP18, YLZ^{+}19, ZCL22]. Observational [ML74, Ros89]. Observations [AJ78, BR14, øBFHB07, CJGPL07, CGP07, CSS14, DS04, Hol80b, HS95, Kes97, KR15c, Laa75, Lau75, LC11a, LB80, Nor80, Ped95, Ste91, ST12, SS80, Zet88, vR95, BKT20, HPR21, JM16, KR20, MV20, Hol81b]. Observed [BB10, BD13, Bri97, Cuc08, DLH14, FS10, Glo06, Jac01, Kes00, KP02, Sør01, Uch04, Wij95, Koi14, LCF24, LPPW22, PLL23, SW19, TKU23a]. Obtain [Per79]. Occupancy [Hol80a]. occupation [MBMG23]. Occurrence [Væt79]. Occurrence/Exposure [Væt79]. Occurrences [Gil86]. Odds [MP14, WC12]. Off [CM01, SH21]. Off-Line [CM01]. Offspring [KL78]. Often [Häg07]. Old [PVFF24]. Omitting [BHC88]. Omnibus [LL06]. One [Bon82, BDV06, CCH01, Hol75a, LW12, MS78, Nor77, Sti82, Sve90a, Wij95, ZL14, JLRT19, LPPW22]. One-Dimensional [Wij95, LPPW22].

One-Parameter [Hol75a, Sti82]. One-Sided [LW12]. One-Way [Nor77, ZL14]. Ones [Azz85]. only [GOV15]. Onset [CSS14]. Opacity [BDH03]. Operating [HC10]. operator [TB22]. Operators [BDP12, FSHK13]. Optimal [AHJ15, AGR13, ARP23, BKM18, BB15, BJ78, Ber82, CY17a, CL01b, De 06, DW95, DR96, Efr08, FGD12, GM16, HST12, HJ04, HC10, Jan91, Laa75, LPPS82, Lai79, Lai80, Lai83, MP84, MR10, Neu97, Pfa93, RW13, SW04, SW93a, VW09, WG96, CPS20, DMK24, FHSZ19, KV23, WC20]. optimalities [YLGL20]. Optimality [AAA04, But86, CDY11, Hoe76, Irl90, LPPS82, LQ17, SB90, Wei93, Jac01, LYW22]. optimization [GGS20, PEK22]. Oracle [KJH16, CL19]. Order [AJN02, AW79, ABN12, Ano83i, Bac11, BIP14, BNP79, Bøl83a, CC98, DFI14, DW97, Eng80, Fre89, GM83, Kou79, LP01, LR08, Mam92, PCW02, SB00, SS98, SS00, Wan86, Wei93, Bon12, CY17a, HJ16, Hua24, ICM19, LA16, TB98]. Order-Dependent [Wan86]. order-of-addition [Hua24]. Ordered [AL79, AL81, Ber81, Ber79b, BT13, HJO15, RD17b, Ros89]. Ordering [GT98, RL06]. orders [LAO23]. Ordinal [BBD^{+}21, FGY23]. Ordinary [Ter83]. Öresund [CSW79]. Orientation [JB20, ZNJ15]. Orientations [JH05]. Oriented [JS12]. Ornstein [Die92, Eie83, FS10, NV17, SH21, VVI^{+}22]. Orthogonal [And90, Bla99, JGØ79, Sas92]. Orthogonality [JK04, WZ10]. Orthogonality-Based [WZ10]. Orthogonalized [QZP12]. oscillator [DEH21]. Other [Asm89, CK97, Doo18, Mei24]. Otherwise [AGM00]. OU-based [BNS03]. out-of-sample [Kop23]. Outbreak [BKO11]. Outcome [BIPV13, SN24, SLCN19, ZZLC21, ZLSL14, ZCL22]. outcome-dependent [SN24, SLCN19, ZCL22]. Outcomes [BG16, WHR22]. Outcomes* [Rub04a]. Outlier [ACR16, BKB23, CW16, JN16a, JN16b, Kuh04, LET20, Oja16, Ron16, STMC16, Zwa16, Hei19]. Outliers [BNP92, DH16, Rap12, And23]. output [CPS20]. overall [HKŠ22]. overdispersion [ZX19]. Overlapping [BH84]. Overview [TCC^{+}95, Agr23]. Oxide [Ped00].

p [GS19a, Toc01, Sim14]. Pair [AGGM06, BG01, Dok80, Dok82, GH87, Gua07, Kou79]. Pair-hidden [AGGM06]. Paired [HH16, Kou76, Kou79, Kou84, SG78]. pairs [WGT19]. Palm [CMW17]. Panel [Got94, HSW03, THSS09, WLS15, ZYX14, KT19, YLZ^{+}19]. Paper [BDH03, DSH14, Doo16, DH16, JH05, Ken14, Per14, Ron16, SLS14, Sim14, Zwa16, Gas23]. Papers [DSS14b, DSS14a]. Parabolic [HL00, TKU23b]. paradigm [CH22, LZL^{+}24]. Paradox [FG96, YLGL20, DRS09]. Parallelizing [HWC20]. Parameter [Arc98, AGGM06, AOH00, Bar03, BN85a, BS01, CRCV12, DSJP14, DLH14, DRM96, DP16, FS10, Glo06, GSG96, Hol75a, HSW03, Jen86, JK04, KL14, Kol97, KRV07, KSM87, LHNN03, Næs82, NC18, Ryd95, Sti82, TKU23b, Uta17, Wal97, AG20, GW24, LP22, OS24, VVI^{+}22, ZD24]. Parameter-driven [Uta17]. Parameterization [Rov15, Wan86]. Parameterizations [LMB09]. Parameterized [NX17]. Parameters [Agr93, AO11, Awa81, BW08, BK95, Car82, CK06, DSD^{+}14, GN98, GM83, GM82, GP89a, GP89b, LL12, MV87, NC18, QZP12, Tan09, Tan94, Van01, AAFO20, CCV23, DR18, ELLV^{+}22, SK19, Ano83i]. Parametric [AHK91, AGM00, Arf98, Ave86, ADGP14, BB10, BL94, BL17, Bor84a, Bor84b, CH04, Dab87, DLS96, DRT13, DW97, Eks13, EGB13, EW94, FHT94, FVV10, FL11, FKA04, GSYB05, GCLP92, GWP89, GV93, Gla98, GMPFV11, GJW08, Gui04, HHVA03, Hjo86a, Hjo86b, HHL02, Jac00, JV06, KB04, KKP08, KS01, La 08, Li01, LS96, LB98, MP21, Mic09, Mül93, Mun02, PKR^{+}97, PLKP06, Pre03, QW96, RR95, SSD15, Sam89, Sch94, SS06, Sha12, SR01, Sve90a, SV05, Wan90, Wij95, YWK06, Zah96, Zha95, Zha08, ZC03, dCJV82, vEvZ96, AV01, APCG24, BM03, BVV17, CCH98, CHW^{+}07, CJGPL07, DW02, DS09, DSS13, DBD18, DPT13, Efr08, EV08, EGG14, Erh08, GPVCGM16, GL15, GG01, Haa08, HVA00, HBH17, HA98, JH17, KN12, KY12, LV02, LT08]. parametric [LdUád15, LN13b, Man09, MW97, MNS07, NV09, Neu09, NGZ18, OFFL12, PFJGE15, PK18, Rom04, SA15, VM00, WD98, Wan08, XZ09]. Parametrically [Gla98]. Parametrization [Fwu85]. Pareto [BTL06, Bon10, BBdW20, Mic09, PKH17]. Pareto-type [BBdW20]. Part [GWP89, YC22]. Partial [AR94, Bac11, BZF08, BG98, Bøl82, CRS24, DD88, DF74, GWT00, Gil92, Hel90, JKR02, LC00b, LC11b, Lok07, MBR03, MvdG15, Slu92, SS98, SSZ09, WFC16, FHY24, HNRT22, Tak23]. Partially [CLSZ16, FS10, HJO15, HZZ07, LQZR09, ST81, YZZ11, ZL18, ZHF03, CW19, EB23, LCF24, LW23, NJG18, SW19]. Particle [LDM15, ZNJ15, CGC23, HWC20]. Particles [Jen87a]. Partition [QMP15]. Partitioning [CS82, Nor90, Sun75]. Partly [BBM06, LS15, Sas92]. Passage [ML86, Ros77, Stu83]. Past [HN99]. Path [BPW14, BM03, GR10, Kos99, Lin94, SV10]. paths [DDM20, DDBEMT24, MR23]. Patients [DE82]. Pattern [LM16, MT14, MB91, PBHMC09]. Patterns [DF74, MR12, Rap12, VB99, BBD^{+}21]. payments [JN19]. PCA [BLM20]. Pearson [FS08, Lus94, SN88]. Penalization [BDL^{+}17, LMH14, SBV11]. Penalized [AG90, CRI03, CGC06, DFG00, GR10, Hel23, KSR13, Pal09, ST10, THSS09, ZHF03, ZZLZ16]. penalties [SK20, Van14]. Penalty [Tan09, WWW15]. Penalty-based [WWW15]. Penultimate [CD01]. Percentile [CL01b]. Percentiles [EPM15]. Perceptron [FV06]. Perfect [BM03, VS07]. Performance [GK13, Kor82, LDA12, Sve90b]. perimeter [CDO24]. Perimetry [OR94]. Period [BSV13, Hok76, OKK^{+}00]. Periodic [DP06, PdT87]. Periodogram [Bøl83b, KM94]. Permanent [HVA00]. Permutation [BJMP14, Præ95, SPK23]. permutation-uniform [SPK23]. persistent [KH22]. Personal [TCC^{+}95]. personalized [ARP23, Hel23]. Perspective [TCC^{+}95]. Perspective* [EM02]. Perturbation [Huc11]. Perturbed [Deg96, Kut19]. PH [NJG18]. Phase [Aal95, Asm89, ANO96k, Huz99, Ols96, Sae15, ABY22, BW07, BW08, FMHB16]. Phase-Type [Asm89, ANO96k, Ols96, ABY22]. phenomenon [ZWS19]. Physics [BN82]. Piecewise [ADGP14]. Piecewise-Deterministic [ADGP14]. Pilot [CMN08]. Pitfalls [GG23, VV24]. Pitman [Jon78, ZD24]. pivotal [CL21]. Pivots [DS94]. pixelated [CDO24]. PL [HV08]. Planar [SV10]. Plane [JGØ79, VS07, Tak23]. Plans [FM90, Thy75]. Plant [BM01a]. Plausibility [Jen78, Jen79]. Players [See93]. Playground* [Nor05]. pleiotropic [ZV21]. Plot [Ber77b, PS13]. Plots [Nai82]. PLS [vR94]. Plug [EB23, GM98, GJW12, DDBEMT24]. Plug-in [EB23, GM98, GJW12, DDBEMT24]. plus [NU19]. Point [AGJ07, Ano07f, AGM00, AH84, BCS13, BL94, BM03, BG01, BL17, CDDL12, CR13, CV15, Cuc08, De 06, DRT13, DM80, DM83, DP16, Eri78, FSGMM16, GM94, GJ03, GSG96, Gri80, Gui80, GS02, HJ16, HS87, HJ04, Jen93a, KL89, LBND^{+}84, LM16, LZ99, MW07, MDA10, MR12, MT14, MB91, Neu97, NV09, NDH^{+}21, NV04, Pal09, PBHMC09, SB00, SS00, SR01, SO97, VB99, VW09, Yao96, BC23, BW19, CM20c, CWZ21, DEV20b, ICM19, JFO23, KS22, LPW21, Mes22, PL23, TKU23a, VMG22, WHZ20, ZX19]. Point-Event [DM80]. pointed [Arc98]. Points [Böh10, Lee97, PW06, ELY22, KA06, ZLK21]. Pointwise [HD22, LL90, ZL14]. Poisson [And77a, BJ89, BSV13, BCC17, BTL06, CCV23, CDGCK15, DK06, Des23, DR97, Fok01, FSGMM16, GS80, GS76, HS10, HA98, HJ04, Hol81a, Joh90, KL14, Lan13, LZ99, RW13, Ryd95, Tju82, Yu16, ZD24, ZX96]. Poissonian [San14]. Polemical [ML75]. Policy [VW09]. Poll [BH14]. Polls [BH14]. Polya [DR00, MW97, NBM12]. Polychotomous [ABC11]. Polygonal [VS07]. Polygons [PBB06]. Polynomial [Bon75, CYL11, DDL14, EGB13, HS04, JKN12, MF97, SW18, Xue10, XL10]. Polynomials [BNP79, CHWY05, Pet99]. Polytomeous [Amu76]. Population [Böh10, BMG82, Cha84b, CM15, DY17, GSK06, Hel98, Hög78, Kol81, Lun00, MS98a, NC92, SW75, Sun83, Tan94, MS24, SK19]. population-averaged [MS24]. Populations [ABC11, BL08, BO02, Bro87a, CSW79, DF74, FH04, GM18, HW98, JR76, Nai82, Nor90, Tho81]. Portfolio [BO11]. Portfolios [Glo14]. Portmanteau [Ter77a]. Posed [FS12, HM09]. Positive [Bar03, DBNR20, LCF24, RSTU21]. Positively [Lin88]. possibility [Bic23]. Possible [DN15]. Possibly [LL96, KK23]. Post [DBNR20, GS19a, MT19]. post-processing [GS19a]. post-strata [MT19]. Posterior [AT15, Awa81, GN98, JLP09, Kim03, NW06, SM12, Sar09, SS09, TKLM23, GS19a, Wal24]. Posteriors [DR00, EG02, FS12, TTL22]. postselection [GAC23]. Potential [BG01, Nat75, Rub04a]. Power [BQ09, BP05, Cac77, Car82, CB84, IKL94, JSG86, Jen81a, Kur18, LA16, MG95, OT09, PP16, Sen88, Sve75, Sve77, ML75]. Power-Divergence [OT09]. Powerful [DF03, Eks13, GAC23, GJ16, ZXLL23]. Powers [Bar03]. pp [AVA22]. PPS [AHJ15]. Practice [Lav23, Ric23, SKBBN79, Ano24, Gre23b, Gas23]. Pre [GPM04]. Pre-Experimental [GPM04]. precipitation [KL22]. precise [Gre23a]. Precision [Sun96]. Prediction [Abt99, BIPV13, BH99, BR97, BN15, CSW79, CG99, De 06, DR97, FKA04, Gui04, Hel00, HST12, Joh90, Lau74, LN95, LY03, MRS14, MSP01, Mat79, OS97, PLHS17, Van07a, Van07b, Vid09, Zha95, Zha08, CK23, ELLV^{+}22, PRS^{+}22, SK19, vdWBM19]. Prediction-Based [BN15]. predictions [Hel23]. Predictive [FPW11, MC03, NW06, SS09, Sun83, TF12, WWP14, Ytt91, GS19a, Lu21, YK20]. Predictors [KS88, KHT14, MRS14, Nie84, Ytt91]. Preface [Ano74g]. Pregnancy [CSS14, KHSS12a, KHSS12b, Aal12, Hou12]. Preliminary [PRV21, Rah86, KK19]. Prequential [SMSD92]. Prescribed [AJN02, Bon12]. Presence [ABKT80, BCH16, HH16, PLHS17, RR95, DLS^{+}24, JV24, XLY20]. Present [Stu96]. presentations [Agr23]. Presented [PC99]. Price [BBK07, LT77]. Primal [FM89]. Primal-Dual [FM89]. Primary [DH78]. Principal [Car07, HT14, HT17, JM93, PS10, QL15, FGLT23, YA20, ZV21]. Principle [BS16, JM83]. Principles [CSW79, MWY15]. Prior [AGR13, AT15, CRCV12, GH23, KH99, Mac93, MBR03, VW15, Wal97, APM19, KK23]. Priority [YL14a]. Priors [AP07, CV01, DLR18, DR00, GPM04, Kim03, Lon12, PKH17, SG15, VHK11, DH23, Dia23, KSSR21, LPR23, OHIS24, TTL22]. Probabilistic [CC12, DMPV02, Hen86]. Probabilities [AJN02, AL79, AHK91, And79, Bon12, BT13, DGSL02, DDK04, KC11, Ros78, SZ07, Sto11, VW15, Yao96, DT20, MBMG23, AL81]. Probability [BJD82, Bøl82, Bøl88, Bon10, BG11, Bro87a, CGL14, Cha77, DBS10, FPW11, GH89, GQR06, Gui77, Han16, HJR06, Irl90, MC97, Pap08, Tho95, AV21, HFP24, HKŠ22, KK19]. probit [WGT19]. Problem [BF02, Dok80, Dok82, Häg07, Hin79, HS95, JWL00, Jon01b, LdM80, Mac82, Van98, vdL96, FR21]. Problems [Ban05, BH99, Ber74, BG80, Bro80, Che09, DRM96, FS12, Gri80, HM09, Hol80a, Jon01b, Kou85, Neu97, NDH^{+}21, Oja16, Rom04, Wan00, WF79, vEvZ96, BR23, CDQ20, Mei24]. Procedure [Hol79, JW10, SM12, SL90, Vie99, XBQF15, MPV24]. Procedures [CM82, FR00, GQR06, GM94, GMMT06, HC17, Kor82, Kou85, Kuh04, Mic09, RV04, SG15, Sør98, CH23, DDM20]. Process [AGR^{+}18, ABH^{+}85, AHK91, AH84, Ave85, Ave86, BB10, BSV13, BKS76, BN13, Bor84a, Bor84b, BG14a, Bri97, BDH03, CYL11, CR13, CGP07, CDGCK15, Die92, Eie83, Eri78, FS10, GCL87, Gil86, GJ03, Glo06, GR01, Gré93, Gup76, HJ04, Hjo86a, Hjo86b, Ist96, JLP06, KL78, Kes00, Kim03, Law82, LZ99, Lin76, MS91, Møl76, MS94, MDA10, MR12, Ner77, NGAS92, OBL18, PBHMC09, Por16, Que12, Ros77, Ros78, Sch94, SB00, SN13, SJ93, SS00, STZ01, WWP14, WD98, Wal00, Wij95, ZX96, AG20, And23, AHWKP19, APM19, BSO22, BSC24, BKT20, Des23, GH21, RW13, dRSS22, VMG22, ZD24, ZX19]. Processes [Abt99, AR80, AOH00, ADGP14, BCS13, BNS03, BNS05, BNLSV14, BPW14, BH84, BS00, BCC17, BM03, BG01, BL17, BJ93, BP89, BF03, BM01b, CM84a, CCH98, Cle97, CDDL12, CMW17, CD18, CV15, CH82, Cuc08, DLS96, Did07, DM83, DS04, DP16, EVP15, FW03, FS08, FM90, FSGMM16, GCJ94, GM94, GSG96, Gri80, GS02, HJ16, HS87, Höp87, Höp90, HK97, HL99, Höp99, HP00, JVA11, JGW13, Jen93a, Joh90, JV06, Jun11, KM94, KL89, KS94, KK00, LP01, LKN15, LZ99, Lo81, Lus94, MF97, Mil85, MSW98, MT14, Neu09, NV17, NV04, NBW02, OT09, OS97, Ove98, PS89, PCW02, Pol95, PVD13, Ris80, Ris81, RR01, RD10, Ryd95, SW04, SP09, Shi17, Slu92, SW18, Sør98, SJ94, SR01, Sve90a]. Processes [SO97, Ter77b, TC05, Vai91, WF79, Wu13, Yuk92, vZ03, vdV94, ABB24, ADMP19, BC23, BCCAVMO21, BW19, ICM19, LPW21, Lu21, LJZ^{+}18, MR23, MMO23, MR24, Mse22, PL23, SH21, TKU23a, VVI^{+}22, YLZ^{+}19, Ano07f]. Processes* [MW07]. Processing [Mus81, GS19a]. Procrustes [Huc11, MM24]. Product [BO11, Dab96, DR96, Joh78, MS98b, QMP15, SMV05]. Product-limit [MS98b]. Production [dMR88]. Products [BLBEO92, ST76, WR93]. Professor [KPS23]. Profile [DC00]. Prognoses [ACFS83a]. Prognosis [ACFS83b]. Progressive [GG01, Gui04]. Projection [CGC06, DP13, ZHS22]. Projection-based [ZHS22]. Projection-type [DP13]. Projections [BKW10, Chr89, Sas92]. prominent [KPS23]. prone [HW17]. Proof [RVG15]. propensity [ZGZ22]. Proper [Vid01]. Properties [Ahm81, AALM17, Aly90, ABK96, AMP01, BL83, BH99, Ber75, Ber77a, BBL87, CL21, CDG16, Chr89, Die92, DT05, GA86, GM82, Gup76, JM16, Joh08, Jør86, Kle82, LM23, MW12, MT02, Nie84, PP16, PS83, Ran75, Ric03, Sai83, SD85, TZ95, VK95, VVI^{+}22, Wan99, ZLY14, vP92, BKT20, HYLT23, KG18, Tak23]. Property [DE06, Fry90, KJH16, CL19]. Proportion [NM14, Swe88, Thu14, GME24]. Proportional [ACFS83a, ACFS83b, ABK96, AFL10, Aug04, CH96, DH07, HESZ16, KKC17, KHL98, LN13b, MU91, MV87, NC18, PM03, QST08, SM04b, SG12, Von96, WC12, XLS16, KXZA20, XNL23]. Proportional-Hazards [FLS05]. Proportionality [SBR98]. Proportions [Aab83, NM87, XBQF15]. Proposal [CV02, EB08, Sto11, Gås03]. Proposals [KH99, TH01, PVFF24]. Prospective [ABKT80]. Protection [And77b]. Proxy [Cro00]. Pseudo [AK07, CDQ20, GSG96, JM16, JFKC05, PNC17, OHN21, OPP18]. Pseudo-Likelihood [GSG96, PNC17, OHN21]. pseudo-observation [OPP18]. Pseudo-observations [JM16]. Pseudo-value [AK07]. Pseudolikelihood [DFG00, NC15]. Pseudovalues [Che91]. public [Ber23]. published [AVA22]. Pump [NGMS94]. Pump-Failure [NGMS94]. Pure [BKS76, ZV21]. Purpose [DH78, CH23]. pursuit [TB22]. Putting [BBK07].

QL [HV08]. QML [AALM17, Win13]. Quadrant [CR98]. Quadrat [DM83]. Quadratic [BZF08, Bon75, FT16, GM08a, PTF09, Wan06, DQR21, YH20]. qualitative [DNCZ21]. Quality [CM01, NEV13]. Quantal [OR94]. Quantification [KKMP18, SO13]. Quantile [CLSZ16, CDZ11, DWV11, GSYB05, KK09, LN13a, NEV13, NL16, Roj98, Vel12, Wre78, Xia94, YL04, BCCH19, CPS20, CM20c, CAVGM21, DEV20a, GW24, KKW24, XWH14, YZ23, ZVD22, ZGZ22]. Quantiles [CM15, GG01, BBdW20, GM23, HKŠ22]. Quantitative [ML74, DNCZ21, LLLP20]. Quantities [Xie89]. quantization [CPS20]. Quantum [RVG15]. Quasi [CM04, DF03, GMA11, HP00, Imo15, LL09, Li01, Lin00, SJS08, THF18, LZC23]. Quasi-Likelihood [CM04, HP00, Imo15, SJS08]. Quasi-maximum [LL09]. quasi-randomization-based [LZC23]. Quasi-Score [Lin00, THF18]. Quasi-Symmetric [GMA11]. Quermass [DLH14]. Quermass-Interaction [DLH14]. questions [LL20]. Queue [Hok75, Hok76, Nat75]. Queueing [Gad85]. Queuing [Nat75].

R [Toc01]. Raj [Ros74a, Ros74b]. Random [AC99, AGM00, AT15, BB10, BNR00, BPW14, BO02, CH82, De 06, DGCS13, FPW11, GM16, GCJ94, GHH95, Gui80, Gui82, HN99, HW95, JLP09, JS12, JW10, Joh82, JH05, KHR02, Kün83, LDW06, LR08, Lon12, ML86, Mol94, Mun02, Nor77, PSS10, Pet99, Rob78, RT02, SG15, Sjö00, SW75, SS80, Stu96, TTZZ18, TB98, TF12, Van13, Waa06, WZH16, WL04, Xue09, YK16, Yu16, Zha96, ZLL^{+}16, ZYT02, DD22, CDO24, DH23, DT20, DLS^{+}24, HPR21, KP21, LCF24, LW23, LLXH19, NR23, NHS^{+}19, ORL20, VMG22, ZBS20, vL18]. random-intercepts [NR23]. Randomization [AB85, LZC23, ZSJT24]. Randomized [And77b, CV14, Sun83]. Randomly [BJMP14, Jan91, JM01]. Range [Cur80a, DRT13, BBD^{+}21]. Rank [Dok80, HJO15, Huc11, JWL00, JLY06, Jør92, KSM87, Miu81, Oja99, Sch75, Sch81, Sri97, SZZ05, Tra11, FR21, HV22, Dok82]. Rank-Ordered [HJO15]. Rao [Ber16, Ohl86, Tor88]. Raoblackwellization [Blo75]. Rare [Cer17]. Rasch [Agr93, Chr74, Møl76, Tju82]. Rate [Aar85, BBL87, CWH05, Far07, FGD12, HHL02, HS98, KST95, MW93, NBY08, Ped00, SC06, WG96, WLX19, Xie89, HBD^{+}20, LM24]. Rate/Mean [SC06]. Rater [RGS03]. Rates [And77a, AG90, BS16, Gar82, GR05, HKK^{+}76, La 08, NBW02, Scr07, Væt79, Wan90, vH80, vZ03, GH23, HMP22]. Ratio [Adi97, And83, Ban05, BN84, BN85b, BN90, BNC91, Ber75, CFJP07, DH08, DS94, Gho06, Gro12, HST74, HL99, Irl90, Jen93b, JQ15, LV02, LB94, MG95, MC97, NC15, NGZ18, Pal09, Pen95, Pol95, SBR98, SLCR14, SS00, Sun95, YZ12, vHV85, RW13]. Rational [Bøl83a]. Ratios [CH04, Tan09, FB20, SJKS22]. Rayner’s [Min81]. Re [NGMS94]. Re-Analysis [NGMS94]. Reaction [NGMS94]. Read [OT09]. Reader [NGMS94]. Real [BT08, Mab17, VM15, Wal97]. Real-Time [BT08]. Real-Valued [Wal97]. realities [Gre23a]. Realization [Sve90a]. realized [WLX19]. Reasoning [AP04]. Recall [SSD15]. Recapture [BL08, Ber74, Ber76, Hol80a, WY03]. Receiver [HC10]. Reciprocal [JSW91]. Reconstruction [JGØ79]. Recorded [MS98a]. recovery [LCF24, TB22]. Rectangles [BM01a]. Rectangular [BJD82]. Recurrence [DW95, Gup76, HP09]. Recurrent [ADZ15, øBFHB07, CWH05, DS09, EGM^{+}03, Höp90, HK97, Jen89, SC06, ADN21, MP22]. Recursive [AOH00, EHR88, Hol80b, Hol81b, GKL21]. Reduce [Bon82]. Reduced [SBH03, Sri97, Ter14, Sak19]. Reducible [CV01]. Reducing [Gad85]. Reduction [BS10, DNL10, ES91, GJW08, Haa08, Hel00, Lue15, NGZ18, PSW09, PS10, CXW23, PS20, RAQ21, WC20, WCY22, ZLZZ21]. reduction-based [ZLZZ21]. Redundancy [ML74]. reference [DH23]. Reflections [Gre11]. Regime [GBB^{+}24, Lin78a]. Regime-switching [GBB^{+}24]. Regimes [MR10, FHSZ19]. Region [BJD82, CK06, Gui86, RS94]. Regions [EW94, JS12, LA16, MV87, SU92]. Register [DGGM16]. Register-Assisted [DGGM16]. Regressand [Amu76]. Regression [AH78, ABK96, AK07, AC99, AH87, ACR16, BNP92, BBG06, Bed93, BD07, BN13, BDP12, BDY85, BS99, Bjö10, BW07, BZ82, BC99, CS03, Car82, CD03, CM01, Cha84a, CC98, CTGS14, CLSZ16, CDZ11, CM04, CW16, CPWZ13, Dab87, DLP08, Dem17, DR96, DH12, DSS13, Dha16, DBD18, DPT13, DE04, DSWH09, DC00, ES00, Efr08, EV08, ES91, EW94, FT16, FMS15, FWW77, FVV10, FV06, Fok01, FM89, GSYB05, Gao98, GM84, GK03, Gla98, GMPFV11, GHC92, Gro96, GS76, Haa08, HS10, HG85, HK99, HESZ16, Hel90, Hel00, HJKQ18, Hou86, HL02, HSW03, HCS15, HS98, HS04, Hua13, HW17, JLY06, Joh82, JNS^{+}83, JN16a, JN16b, JR76, Jør92, Kab78, KB04, KN12, KF07, Kol97, KST95, KK09, LN13a, LYZ15, LB88, LV02, LGP11, LV13, LLY17, LdUád15]. Regression [LS96, LZZ14, LAKZ12, LHW^{+}16, MSR16, Mar99, MS01, MM93, MF97, MS98b, MdCCD19, Mül85, Mül92, Mun02, Mur93, Næs82, NH93, NV09, Neu09, NEV13, NL16, Nor77, OB16, Oja16, PF08, PFV06, PFJGE15, PKR^{+}97, PLKP06, Por16, Pre05, Pre03, QJ01, RS83, Ron16, SA11, SA15, SC06, Sch94, SZS02, SM04a, SZ07, SU92, SLCN19, SMS12, SV76, ST12, SW05, SBM^{+}99, SBB05, TEV15, TGM17, TM86, Toc01, TDR09, Tra11, VBJ97, Vel12, Wan90, WG96, Wan00, Wan08, WLS15, WLT15, XMW15, Xue10, YL14b, YZ07, YZZ11, YL04, YWK06, YD07, Zah96, Zha95, Zha08, ZC03, Zwa16, BD20, BEP20, CPS20, CN16, CAVGM21, CL19, CFL24, DEV20a, FM22, FHY24, GG23, GW24, GPVCGM16, Hel23, JV24, KK19, KYZC21, KK23, KKW24, LLCW21, LLXH19, Lue15, MS23, MT19]. regression [MPV19, OHIS24, STM22, SW19, SAS24, SJZI19, VV24, XWH14, ZVD22, BW08]. Regression-type [MdCCD19]. Regressions [Amu74, Amu76, CDMGR06, Lin78a, RD17b, ŠBD05]. Regular [KM95a]. Regularity [LM16, VMG22]. Regularization [GR10, NR23, Van14]. Regularized [FS12, HYLT23]. Regulating [EVP15]. Reiersøl [Wil79]. Reinforced [APM19]. Reinforcement [SB85]. Reinforcement-Depletion [SB85]. Rejection [BS16, FGD12]. Rejective [Hol79]. Rejoinder [Azz05a, BG14c, HOF^{+}94, JN16b, Min81, Mül05a, Sve77, TSH91, TCC^{+}95, Gre23a]. Related [Azz05b, FH04, GC05, Gui77, Jon01b, KP02, NDH^{+}21, Ran84, Rom04, vE92]. Relation [HKD02, Wil77]. Relations [GK86]. Relationship [CM82, LL06, LB94, MS78, JB20]. Relationships [CM84b]. Relative [DH07, Die92]. Relatively [BJ78]. Relatives [BNM^{+}06]. Relevant [Bøl88, HST12, NH93]. Reliability [BAR^{+}85, BR97, Chr74, Ege92, GN95, GK86, Lin94, NE87, RGS03, Slu97, dMR88, vP92]. Remainder [Eng80, Hög78]. Remark [Höp87]. Remarks [Ham88, Joh77]. Remove [LS23]. Renewal [BL94, BP89, CD18, Gup76, GS02, Hor85]. Rényi [JM93]. Repair [Gär03]. Repairable [Lin88, WCY22]. Repeated [Cro98, EB23]. Replacement [Ber79a]. Replicate [CSS14]. Replicates [TL03]. Reply [ABH^{+}85, AKB^{+}89, Arj04, BNHH95, BAR^{+}85, BHR^{+}76, BRH83, CSJ^{+}77, CGL^{+}81, Eri84, GI02, GWP89, Gus02, HKK^{+}76, Hoe78, Jan02, JNS^{+}83, JAL^{+}81, KHSS12b, LBND^{+}84, LAE^{+}89, LRT^{+}87, LBNE^{+}78, ML75, Rub04b, STH^{+}78, SN88, SKBBN79, SBM^{+}99, TJL^{+}76]. Represent [GPM04]. Representation [Ano83i, FZ06, GM83, Hen86, Nor86, SMV05, Wan00, Xia94, FHTT18]. Representations [Sat96, Sei24]. Representative [GS14]. reproducing [CXW23, FR21]. Reproductive [BJ85, LY08]. Resampling [BS00, CM15, Sjö00, SBH03, ZYT02]. Research [Ano74d, Ano74e, Ano74f, Ano75e, Ano75f, Ano75g, Ano76e, Ano76f, Ano76g, Ano76h, Ano77e, Ano77f, Ano77g, Ano77h, Ano78e, Ano78f, Ano78g, Ano78h, Ano79e, Ano79f, Ano79g, Ano79h, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano82e, Ano82f, Ano82g, Ano82h, Ano83e, Ano83f, Ano83g, Ano83h, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h]. Residual [AV01, Aly90, GK03, JVA11, MMS16, Neu09, BCCH19]. Residuals [HV06, KB04, Neu09, QZP12, OHN21]. Resistance [BNR00]. Resnick [KL22]. Resolution [BN84, BN85b]. Resonance [Har02, JT07, LLS^{+}22]. Respect [Nor90, ST81, Erh08]. responders [NM24]. Response [And77b, And79, DY17, EGM^{+}03, LPB15, NGMS94, OR94, PS13, SW93a, SG04, TTZZ18, TS91, WZH16, WL18, XMW15, Xue09, GPÁLÁPGM21, ZKP^{+}24]. Responses [ABC11, DPT13, PFV06, TWL18, FM22, LW23]. Resting [JT07]. Restoration [Rue97]. Restricted [AB85, Bon10, GJ16, MMS16, SU92, Ter83]. Restriction [CC98]. Restrictions [CD03, DFI14, FRS99, MTA99]. Result [WR93]. Results [AB85, AH87, BNR00, CM15, Far07, HJS90, Hol81a, Jen93b, Lai83, Lan13, Nie97b, Sun75, Ter81, Wil79]. retrieves [LS23]. Retrospective [ABKT80, OKK^{+}00, VW09]. Reversals [LS98]. Reversibility [Edw80]. Reversible [EB08]. Review [ABH^{+}85, BHR^{+}76, HKK^{+}76, Oja99, Tjø94, MR24]. Reviews [BHR^{+}76]. revisited [LET22]. Riemann [Kle24]. Riemannian [Sko84]. Right [BJMP14, CLSZ16, DBS10, HCS15, Jon01a, Min79, SV04, WD98, BCCH19, DT20, OH21]. Right-Censored [DBS10, Jon01a, SV04, CLSZ16, HCS15, BCCH19]. Risk [BDP13, DL89, DH07, Det04, Efr16, EGG14, GJ05, Kle99, KZ17, KHT14, LGP11, MS09, SA15, SMZ11, Sun95, vH80, GGS20, MS23]. Risks [CHW^{+}07, Cro91, Cro00, DS09, DSWH09, Gar82, GK00, HESZ16, JH17, KS01, LB98, WCXS15, YY15, APM19, CMY24, OPP18]. Road [DK80, GS76, ZL10]. Robust [ACF^{+}21, BS10, Bed93, BB15, BBM06, BCH16, CTGS14, DGN07, DR22, Det04, FHT94, FMHB16, GPST23, HS10, HHHL23, HG85, Jon91, KV23, LZC23, LDA12, PF08, PS13, STMC16, SA80, TTZZ18, ZLY14, CFR19, CH23, JN19, KYZC21, LW23, WHR22, YZ23, DBJ^{+}22]. Robustness [AO11, BHR^{+}76, Han16, LL90, And23, VD18]. ROC [GMPFV11, LZ08]. Role [CSJ^{+}77, ZGZ22]. Root [BG98, MWY15, SW04, TvdM96, BSC24, KT19]. Root- [BG98]. Root-n [TvdM96]. Roots [DS94]. Rosen [LGL19, Lan74b]. Rosenbrock [PWN22]. Rosenthal [IS99]. Ross [BKT20]. rotational [TPH21]. Row [See96]. Rubbery [NBM12]. Rule [BJFG15, CG99, DMV16]. Rules [BG80, Lai83, vHV85, ARP23, CN16]. Runs [JSG86].

S [Ano24]. S.D.E. [DDBEMT24].

Saddlepoint

[BNK99, BJ85, JKR02, PTF09]. Saddlepoint-Based [PTF09]. Safe [GH14a]. Sampford [BTL06, BG11]. Sample [AL79, AL81, ABK96, BPW14, BG16, BG80, BBL87, Bon76, CL01b, Cur80a, CS90, DP04, Edw80, Gro12, Hjo88, JM16, JWL00, Jan91, JP06, Joh17, Kle91, LPPS82, Lai79, Lai80, Lai83, LW12, LL90, LLY18, MW10, MC97, NW06, OH16, PW10, SMB14, SW93b, Wre78, WW11, Ytt91, Zha00, Ahm17, BMP19, BS21, CCH98, FR21, HLP23, KK19, Kop23, NG23, TPH21, WZ22]. sample-specific [TPH21]. Sampled [CGL14, Fas16, Fra78, GT98, KV23, YY15].

Sampler [LDM15, Pic00]. Samplers [HM09, Kle16, MT02]. Samples [Arf98, BW07, BW08, FSHK13, GS14, HJO15, Hol81a, JR76, LB98, Lon12, NS06, SBR98, Sun83, CH23, SN24]. Sampling [AJN02, AHJ15, Ber16, BCC17, BM03, Bon82, BTL06, BT08, Bon10, BG11, Bon12, BT13, BO99, BJ12, CGL14, CMMR12, DH78, DF74, DM80, FMHB16, Fra77b, GCJ94, GSK06, GL07, GM18, GK86, HKJ11, Hög78, Lun00, MS01, MG98, OBL18, OFFL12, QQZ16, Sae15, SÅS07, STH^{+}78, SW84, SS18, SM04a, SBH03, SW75, SW76, Tho81, Thy75, TF12, CL21, DQR21, GK21, LRT23, RFK22, SLCN19, VHF20, Hoe78]. Sampling-Importance [SBH03]. Sander [Ric23, Gas23]. Scalable [Fan19, SS18]. scalar [AHP^{+}18]. Scale [AO11, Ano83i, DH12, EHR88, GMPFV11, GM83, KSM87, LM16, LP01, NL16, Oja81, OSG08, Tan09, DD22, LLLP20, PBHMC09, RZM16, TCK^{+}23, WZ22]. Scale-space [OSG08]. Scales [HBH17]. Scaling [BNS05, KR15a]. Scan [Lin14, AH19]. Scand. [AVA22]. Scandinavia [Ano74d, Ano74e, Ano74f, Ano75e, Ano75f, Ano75g, Ano76e, Ano76f, Ano76g, Ano76h, Ano77e, Ano77f, Ano77g, Ano77h, Ano78e, Ano78f, Ano78g, Ano78h, Ano79e, Ano79f, Ano79g, Ano79h, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano82e, Ano82f, Ano82g, Ano82h, Ano83e, Ano83f, Ano83g, Ano83h, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h]. Scandinavian [Ano24, Sch80, Ano98e, Lav23]. Scheme [AH92, TF12, OPP18]. Schemes [GL07, HJ04, Var79, ZCL22]. Schwartz [EU21]. Schwartz-type [EU21]. Science [BJMP14, Agr23]. Sciences [Sch02].

Scientific [SN88]. Score

[BGH19, Che09, CK97, HTK15, HWH15, KSM87, Lin00, Sør01, THF18, WC12, HBD^{+}20, WC20, ZGZ22]. Score-Type [HTK15]. Scores [HT17, dCJV82, YA20]. Scoring [DMV16]. Scott [JVA11]. Screening [WL18, GPST23, XLY20, ZZLC21]. Scribes [HS95]. SDEs [OS24]. Seasonal [Lau76, TRL15, Zet88]. Second [ABN12, Ber75, Bon12, HJ16, ICM19, LR08, LA16, SS00, Wei93]. Second-Order [LR08, Bon12, HJ16, ICM19, LA16]. secondary [DZ21]. Section [JGØ79, BG24b]. sectional [Van07b]. Seeger [Ano96e]. Segment [PS89, Wij95, vdL96, Van98]. Segmentation [SV10, CLR19]. Segregation [Cey10]. Seismic [La 08]. Select [GS14]. Selecting [Gua07, MSP01]. Selection [BG80, But86, CMN08, CHI23, CTGS14, CO07, CLP17, CPWZ13, DS03a, DRM96, DH16, Eri96, Haz96, HH82, Imo15, Kle16, MSR16, MS09, NC18, Nor81, OH16, Pre03, QMP15, Sar09, SHD94, TM86, THSS09, Tra11, WWW15, WLT15, ZLL^{+}16, ZLSL14, BLM20, CM20b, DHH24, DBJ^{+}22, EU21, HFS23, KMG21, LLYC22, MT19, PRS^{+}22, RMG19, SJKS22, TB22]. Selective [TT17, Tjø94]. Selector [AFL10, EL96]. Selectors [GM98, JK92]. Self [DK06, EVP15, JFKC05, KM94, LWY97, YLW00, ZLS14]. Self-Consistent [JFKC05, LWY97, YLW00]. Self-exciting [DK06]. Self-Normalised [KM94]. Self-normalization [ZLS14]. Self-Regulating [EVP15]. Semi [AR80, AHK91, BM01a, BVV17, CCH01, CJGPL07, CK06, DSWH09, EPM15, GWP89, GV93, JH17, KKP08, QW96, RR95, YWK06, YY15, ZC03]. Semi-Competing [DSWH09, JH17, YY15]. Semi-empirical [EPM15]. Semi-Latin [BM01a]. Semi-Linear [CK06]. Semi-Markov [AR80, CCH01]. Semi-Parametric [AHK91, GWP89, GV93, KKP08, QW96, RR95, YWK06, ZC03, BVV17, CJGPL07, JH17]. Semimartingales [Vet12, Koi14]. Semiparametric [ADN21, BKM18, BBG06, BBM06, Bor99, BDV06, BW07, BW08, BEP20, BV14, CHW^{+}07, CP07, Che15a, CLSZ16, Che13, Che15b, DFI14, DNL10, FLS05, FHY24, Gao98, GK03, Gho06, GS99, GLQ18, HT10, HC17, Kor00, KK09, LZ10, LLY17, LHWS18, LZ08, LZZ14, LLYC22, MS09, NBY08, NYR18, PG13, PVD13, Qin98, SP09, SW05, WY03, XLS16, YZ07, YD07, Zha00, ZHH10, ZYT02, ZCL22, LLCW21, LZL^{+}24, MPV24, NJG18, OHN21, ULK23, WCY22, ZMH^{+}24]. Semiparametrically [LFL16]. Semivariogram [KB04]. sensitive [LL20]. Sensitivity [PS13, Ros89]. Separable [NS06, RD10]. Separate [CK94]. Sequence [JXCK14, Sar09, SC06, Sko81b, Var79]. Sequences [Bjö10, EHR88, Pfa93, SS80]. Sequential [BS16, Ber82, BBS23, BT08, CV14, Efr08, FM90, GS02, Hol75b, Hor85, Irl90, JSDT11, KS22, MR12, MC97, SN13, SW87, Sør98, WF79]. Sequentially [GT98, Hol79, MSR16]. Serially [YZ07]. Series [Ant96, ACR16, Ber74, Ber77a, BP05, BK95, Cac77, Che09, Che15a, CGL^{+}81, CW16, Dam80, DS03b, Eub00, Fok01, Gao98, GJ03, Gri09, HS04, HHM17, JN16a, JN16b, Lau76, LS98, LHNN03, LG09, LST88, NL16, Oja16, OSG08, Pap00, Ron16, SL88, Sha12, SG15, Tho83, Tjø94, TCC^{+}95, TRL15, YZZ11, Zwa16, BLG20, BBD^{+}21, BBS23, BS21, CLR19, CWZ21, FGLT23, PS20, PPS21, TKLM23]. Services [Gad85]. Set [AL79, AL81, GR05, HJO15, ML74, Nor90]. Sets [Blæ78, JS12, CDO24, NHS^{+}19, vL18]. Setting [DP06]. Setup [NM14]. Several [LLY18, Nai82, Puk82]. Shannon [AVCRG13]. Shape [CRCV12, HKD02, Jen86, LN13b, MTA99, ZNJ15, LGL19].

Shaped [Jen87a, Xie89]. shared [KMG21]. Sharp [Arc98]. Shock [GN95, GN98, Kle81, Lo81]. Short [YZ12]. Short-Term [YZ12]. shortfall [CM20a]. Shot [DP16, JH05, MDA10]. Shot-Noise [DP16, JH05, MDA10]. Shrinkage [Bla99, BZ82, LC00b, PKH17, CK23, KMG21, OHIS24, TB22]. Shrinking [MRS14, STK17]. Sided [LW12]. Sieve [AGR13, HL00, LO16, ZG03]. Sieved [Jon01b]. Sigma [GH89]. Sigma-Algebras [GH89]. Sign [Oja99, CLP^{+}19, TB22]. Signal [TGM17]. Signals [KM95b, Taq02]. Signed [BN90, BNC91, DS94]. Significance [Bøl83b, Bøl88, CSJ^{+}77, SN88, SV76, Sve76]. Similar [Gui86]. Simple [Agr93, AL98, AL99, BDY85, Blo74, Bon82, BO99, EL96, FG96, Hol79, IYW14, Jen97, Kes00, Laa88, PKR^{+}97, PLKP06, Ris80, Ris81, SW75, VR08, WW01, Wre78, dCJV82, Gre23a]. Simplex [QST08]. Simplified [BS01]. Simpson [DRS09]. Simulated [Cle97, Sør03]. Simulation [AT15, BM03, FWW77, GM94, HN99, Lai83, MV87, NH15, NV17, RS83, SW87, VS07, Waa06, Ytt91]. Simulation-Based [Waa06]. Simulations [VKY^{+}14]. Simultaneous [And91, DP06, FZ00, LAKZ12, LHW^{+}16, PWY97, SR11, TCK^{+}23]. Single [BN85a, BVV17, CM17a, CM17b, CYM93, Cur80a, Jon91, LG09, TWL18, Thy75, Uta17, WZH16, WFC16, ZYX14, BGH19, CW19, HHHL23, YH20]. Single-Index [LG09, TWL18, BVV17, WZH16, WFC16, ZYX14, BGH19, CW19]. Single-Parameter [BN85a]. Singly [VBJ97]. Singular [MG95]. Singularity [DR00, Ist96]. Six [BNP79, DSS14b, DSS14a]. Size [ABK96, BG80, Böh10, CL01b, CDGCK15, Kle16, Lai79, Lai80, Lai83, MC97, Mül85, Ner77, NW06]. SJS [Gas23]. Sketch [Jen93b]. ’Skew [Hen86, AVA06, AVCRG13, Azz05a, Azz05b, BPS17, BGL13, CRCV12, CAS03, DLR18, Gen05, LMT14, MG04, MH10a, OH16, PG13, AVA22, BCCAVMO21, DR18, HV22, JLRT19, LM24, NHMW22]. Skew- [BGL13]. Skew-Elliptical [AVCRG13]. ’Skew-Normal [Hen86, AVA06, Azz05a, CRCV12, CAS03, MH10a, Azz05b, Gen05, OH16, AVA22, DR18]. Skew-Symmetric [MG04, BPS17, DLR18, PG13]. Skewed [GM82, LZZ14]. Skewness [CJ08, Oja81, EK22]. Sklar [BDS22]. Slice [MT02]. Sliced [Pre05]. Slide [YZZ11]. Slopes [Kim97]. Small [AL79, ABK96, ADL15, BG16, BBL87, DR18, DSD^{+}14, GM08a, GS76, Jac01, KHSJ19, Kle91, Kur18, LM16, MRS14, MSZ16, STMC16, See93, SMB14, Shi17, STK17, TDR09, Uch04, Ytt91, ZXL^{+}18, ZZLZ16, CK23, ELLV^{+}22, JN19, SKR19, AL81]. Small-Area [STMC16, JN19, SKR19]. Small-sphere [KHSJ19]. Smirnov [Præ95]. Smooth [Adi97, AGM00, CM04, Dre98, FSGMM16, GJW12, HK99, IKL94, La 08, LVV09, Mam92, Neu09, Rei81, Sko81b, WW11]. Smoothed [CD96, Fer91, LM18, Wei93, Yuk92, vdV94, FHSZ19, HMP22, Tak23]. Smoother [CQ02]. Smoothers [Bla01]. Smoothing [Ant96, BC99, EGB13, FK98, Gao98, GK00, LWY97, LDA12, MW10, MTA99, PSW09, WL04]. Social [BO02]. Soft [LMH22, MM24]. Software [Lok07, Slu97, vP92]. Soil [Ped00]. Solution [Bac11, GRS22]. Solutions [CYM93, Mol98, OS24]. Solving [Häg07]. Some [Ahm81, Arj11, Aug04, Ber75, BCS00, BHR^{+}76, Böh10, Bøl88, CGL^{+}81, Die92, DT05, DR97, Far07, GN95, Gui80, Gup76, GJ83, Hol81a, Höp99, Jen93b, Joh77, JNS^{+}83, Jør86, Kle81, Kle91, KM94, KT95, Kor82, KK00, Kün83, LL90, Mac93, Mac82, PVFF24, Sai83, Sch80, SD85, Ste88, Sun75, Ter81, Wil79, Xie89, ZLY14]. Søren [Oja16, Zwa16, Doo16, DH16]. Sought [Jon01a]. Sound [Kou84]. sources [CH22].

Space [BM01b, DP16, Koo99, KL89, MG98, QL15, Vid01, ACF^{+}21, AH19, FR21, JB20, OSG08, SP22, TKU23b]. Space-time [BM01b]. Spaced [PW06]. Spaces [GH89, vR94, DEH21, HJG21]. Spacing [Eks01, KR15c, Miu81, Ran84, KR20]. Spacings [BH84, Eks13, Sti82, vE92]. Sparse [CY17b, CLP17, FGY23, FGLT23, JXCK14, LQ17, QL15, RLOS18, Sar09, SM24, DBJ^{+}22]. Sparsity [ST10, Van14]. Spatial [Ano07f, BCS13, BM03, Bol14, BDH03, Cey10, CDG16, CR13, Cuc08, DFG00, EMR09, Eks08, FSGMM16, GM94, HJ16, HA98, HOF^{+}94, JGW13, Jen93a, KB04, Kop23, LM16, MS94, MW07, MR12, MT14, MB91, PLHS17, PBHMC09, ŠBD05, STMC16, ZLS14, CM20b, CL21, CLP^{+}19, HOT21, KL22, LCZW22, PEK22, PD22, VMG22, ZX19]. Spatially [GM18, ADN21]. Spatio [CV15, JT07, MDA10, NV17, RD10, Vai91]. Spatio-Temporal [JT07, MDA10, RD10, Vai91, CV15, NV17]. spatiotemporal [ICM19]. SPC [HS87]. SPC-Systems [HS87]. SPDEs [HL00, TKU23b]. Spearman [HV22]. Special [FGD12, BG24b]. Specific [Høj04, QMP15, SG12, Wan08, TPH21]. Specification [AT15, AP07, DS03b, HNNS19, LKN15, KKW24]. Specified [QZP12]. Spectral [BB10, BR14, BR17, Cav23, DS03b, Efr16, FK98, MS91, Pap00, Žur79, TKLM23]. Spectroscopy [LdM80]. Speculations [Arj11]. Sphere [FRZ16, Kle99, KHSJ19, LRSS23]. Spheres [BG01]. Spherical [Jen81a]. Spike [Efr05]. Spline [HS04, LO16, LDA12, XL10, ZHH10]. Spline-Based [LO16, ZHH10]. Splines [ACMLM03, HZZ07, MTA99, PSW09, KSR13]. Split [SM12]. SPR [Hol75a]. Spread [RD17a]. Square [ADZ15, BR81, Hoe76, HB06, LL96, Min79, Min81, Pen95]. Squared [Abt99, ES00, OS96, JN19].

Squares [AC99, AOH00, BIP14, Gré93, GP89b, Hel90, LP01, LC00b, MSR16, Nor75, SS98, Sun96, Ter81, Ter83, ZG03, ZZLZ16, BD20, GP89a]. Stabilization [MS12]. Stabilizing [Hou86]. Stable [AV21, BNB93, EBGG18, FNR09, Höp99, KM94, OBL18, Win13, JKM19]. Stacy [APM19]. Stage [Ham88, HS06, Sun83, Yao96, LT08, LCZ09]. Standard [DS94, HTK15, Lok07, CMY24, DH08]. standard-deviatile [CMY24]. Standardized [BN90]. Stat. [AVA22]. State [BDW16, CCH01, JT07, Lin77, MG98, Ste91, Vid01, vLM23, ACF^{+}21, CJGPL07, LT21, MBMG23]. State-Dependent [BDW16]. States [Lin78c]. stationarities [OSG08]. Stationarity [Edw80, PP16]. Stationary [BNLSV14, BL17, BK95, CDDL12, Eri78, FW03, HJ16, KS88, KL89, KK00, LL99, OT09, PSS10, PCW02, PVD13, Ran75, SP09, Sha12, SS80, SJS08, TC05, VM00, Eks08, FGLT23, HNRT22, Jun11, LLY17, NHS^{+}19, PL23, PV00, TKLM23, XY15]. Statistic [CFJP07, Jen93b, Sat96, Ter77a, THF18, AH19, AKP22, LPYZ24]. Statistical [Arj11, BN82, Bic23, CTYF13, Cav16, Chr89, CC12, CD18, CGL^{+}81, Dem17, Die92, DSS14b, DSS14a, Dre98, EVP15, EM02, Eri04, GIA02, GCL87, Glo14, Gri80, GH12, GH02, Hel90, HKK^{+}76, Høj04, Höp87, HL99, Höp99, HYWC18, JJ02, KZ17, KC11, LL06, Lo81, LHHF13, Mad76, ML74, Mei24, Mol94, MS94, Mus81, NV17, NV04, PS89, PBBM12, Rob78, Sch79, See93, Sko86, SJ94, Sve76, TW04, TJL^{+}76, ULK23, VKY^{+}14, WHZ20, WR93, XNL23, ZLY14, ZYX14, vP92, Agr23, Cav23, Fan19, LAO23, VS21]. Statistically [FS08]. Statistics [AF07, Ahm81, AW79, Ano83i, Ano98e, Ano07f, Ano24, BJ78, DD22, BRH83, Bøl83a, Bon75, CB84, DS94, EM02, Eng80, Fre89, GH00, GH08, GA86, GM83, Gui86, HJS90, HOF^{+}94, HS06, HTK15, IS99, Joh08, Kim97, LBND^{+}84, LLY18, Lin14, MW07, OT09, Rap12, Sch80, Sch02, SW93b, Web81, vE92, DBS10, GRS22, HLP23, JTT21, KS22, KPS23, PC99, RW13, Lav23]. Statistiscs [Jon78]. Status [BW05, FMS15, Gro12, GJW12, GH18, LS15, VBJ97, VJ01, WC12, JvdMP22]. Stein [CN16, LB88]. Stein-rules [CN16]. Stems [LN95]. Step [FGD12, HK97, Kle16, MPV24]. Step-Up-Down [FGD12]. stepped [LKT^{+}23]. Stepwise [Nor81]. Stereological [Jen87a]. Stirling [Ber75]. Stochastic [AGR^{+}18, Abt99, BN97, BNS03, BS00, BJ93, BO02, CZT20, CO07, DGCS13, Eri78, FWW77, FZ06, GCJL03, Gui86, Har02, Hol80b, Hol81b, JSDT11, KL89, KS94, LS98, Man09, MB91, NH15, Ner98, Nor05, Ped95, PDD10, Ryc96, Shi17, Slu92, Sør98, Sør03, ST76, Vid01, Wil77, AHWKP19, AKP22, Des23, EU21, Fan19, JKM19, LAO23]. Stock [GIA02]. Stopped [GJ83]. Stopping [Höp87, Sør98, Ste88]. Straits [CSW79]. strata [MT19]. Strategies [BCG08, Kre87]. Strategy [DS03a, ZLL^{+}16]. Stratification [AKC80, DH78, MM93]. Stratified [BW07, BW08, KXZA20, SÅS07, SG12, SN24]. stratum [ZV21, ZSJT24]. Stream [ATV17]. Strength [BM15]. Strict [Nat93]. Strip [DK80]. Strip-Road [DK80]. Strong [Ano83i, Deg96, GM83, GPM04, Hor85, KJH16, Ped75a, Ped75b, Tan09, Ter14, Wan95, XY15]. Strongly [Jen89, LL99, BLG20]. Structural [Bon82, CM82, CM84b, DW16, GSK06, HHM17, KM00, Kos99, Lok07, CN16]. Structure [Abt99, And82, And90, BO02, FPW11, LC00a, LCF24, LC00b, Mil85, PNC17, Pic00, RD10, ZLL^{+}16, HYZ22, ZGZ22]. Structured [Møl86, MDA10, Sch02, DBNR20, Van14].

Structures [Far15, JAL^{+}81, MR12, GBB^{+}24]. Student [And23]. Student- [And23]. Studentized [Cur80a, JM01]. Studies [AK07, Det04, Guo11, Kur16, Lan07, LT08, LMH14, Lun00, Ros89, SMZ11, Sve86, ZLY14, ZXL^{+}18, AV21, BKN23, CL20, HBD^{+}20, KXZA20, LLLP20, LLS^{+}22, ZZLC21, ZXLL23]. Study [ABK96, Bro87b, CSW79, DM80, Efr05, Hok76, Kar15, Lau76, LT77, MV87, PS10, RS83, SW87, SW75, SW76, WWP14, dMR88, BD20, ZMH^{+}24]. Studying [LYZ15]. Subdistribution [HESZ16, APM19, KXZA20]. Subdivided [FH04]. Subgraph [Fra78]. Subgroup [fWZY16]. Subject [SG04, TWL18, Wan08, LZL^{+}24]. Subject-Specific [Wan08]. subordinated [Mse22]. Subordinators [JV06]. Subsample [HV05]. Subsampling [Eks08, BW19]. subsequent [ZHW19]. Subset [BG80, MSR16]. Subsets [Jør92]. subspace [LCZW22]. Sufficiency [AR94, BNHJP76, KP77a, Lau74]. Sufficient [Che13, Jag77, OKW88, RAQ21, Ran78, Sat96, WC20]. Suitable [BT08]. Sum [AW79, Eks13, Hol75b, PW06]. Sum-Functions [Eks13]. Summary [BH99, GH00, GH08]. Sums [GJ83, Ter77b]. Superimposed [KM95b]. Superiority [Ter83]. superpopulation [YK20]. supersampled [BKN23]. Supersmooth [HB06]. Superstructure [VK95]. Supervised [BCCA11]. Supplementary [Hok75, KH16, LYZ15]. Support [AGJ07, BC99, GJW08, TvdM96, Vie99]. Supremum [HTK15]. Sure [Fer91, GPST23]. Surface [LV13, Ped00]. surprise [Bic23]. surrogate [YLGL20]. Surveillance [Lin14]. Survey [And82, BCC17, BCH16, DH78, Hoe78, MW10, MS86, STH^{+}78, SW84, Ber23, DZ21]. Survey* [SRH07]. Surveys [LYZ15, MP80, MP84, DQR21, JN19].

Survival [Aal87b, Aal95, And83, ABK96, AG90, AKB^{+}89, BL90, BHC88, Bro87b, BB14, CCH01, CHWY05, Dab87, DP18, DL89, DH07, DCIK14, DRS09, DN15, EMS15, GHC92, Gro96, Hou87, HC10, JM01, JH17, Jon91, Kle81, KHL98, KS01, LDW06, LWY97, LR06, LT08, LHML16, LDY16, LFL16, MW08, MAR11, MS98a, MW97, Nie97a, Nie99, Par01, PWY97, PR07, Pon86, Væt79, Von96, Wan87, YZ12, YLW00, BEP20, CXW23, DR22, Par20, SN24, WHR22, XLY20]. survivor [QB23]. Susceptible [LY08]. Sverdrup [ML75]. Switching [Doo18, Lin78a, Cav23, GBB^{+}24]. Symmetric [Arc98, BV14, GMA11, IS99, MG04, vHV85, BPS17, DLR18, PG13]. Symmetrizable [WR93]. Symmetry [Jen81a, Kou85]. Syndrome [SS79]. Synthetic [DLP08, Laa78]. System [BR97, Ege92, Nat85, NE87]. Systematic [OFFL12]. Systems [Chr74, DMPV02, Gad85, HS87, Kos99, Lin88, Nat93, CGC23, Kut19, SP22, WCY22].

t [BGL13, HL08]. Tables

[And74, Cey10, FHTT16, Høj04, Jen78, KK06, Kre87, Kuh04, Mad76, Ped75a, Rap03, Rap12, Sun75, VKY^{+}14, FHTT18, LET20]. Tail [AJRN16, BJ89, BNK99, CP98, Cur80b, Dre98, EBGG18, GG13, KKP08, KY12, SS06, GGS20]. Tailed [FWW77, EGG14, JR07, LP22, Taq02]. Tails [DDL14, AKP22]. Takacs [CDDL12]. Tapered [Ant96]. Targeted [CV14, MBMG23]. Tauberian [Seg02]. Technical [LRT^{+}87]. Technique [Hok75]. Techniques [AJRN16, Haa08, Kre87, LS96, PKR^{+}97]. Teeth [HVA00]. Telemonitoring [BIPV13]. Teletraffic [Nor05, Mak05, Sze05]. Temperature* [BBK07]. Temporal [BG14a, JT07, MDA10, RD10, ŠBD05, TC05, Vai91, CV15, NV17]. Temporally [HHVA03]. tensorial [VLIN21]. Tensors [Huc11, ZNJ15, Kle24]. Term [Eng80, Hög78, YZ12]. Terms [Cro00, DR96]. tesselation [CL21]. Tessellation [Sib80]. Test [Aar85, ADZ15, Ber77b, Ber79a, BJ78, Ber82, BR81, BRM14, Bri97, CK97, DS03b, DW97, DR10, GA86, Hög79, Hol79, JWL00, Joh08, Jon78, KL14, Kim97, Kle82, KY12, LW12, LP01, LHNN03, LKN15, MT03, Mar98, Min79, Min81, NGZ18, PFJGE15, Pon86, SH96, SW87, Sti82, Ter77a, Waa06, Wył16, Xie89, Zet88, CAVGM21, FR21, GPÁLÁPGM21, HJG21, KK19, PRV21, Rom04, ZL14]. Testing [Bac11, BNP92, BH97, BM16, BN13, BP89, CFMS03, CLP^{+}19, CFS95, CDMGR06, DK06, DH08, DB03, DBS10, DPV06, DDK04, DBD18, DEH21, DLS^{+}24, Eub00, FZ00, Far09, FSHK13, GH16, GPVCGM16, GMMT06, Grø97, GBB^{+}24, HK99, HS06, HS95, HW17, Kle83, Kur18, LCZ14, Lin88, LPS03, LST88, MRM09, Mej85, Mun02, Mur93, NM14, OKW88, PVD13, Que12, Rom04, SL88, SN88, SN13, SBR98, Slu97, Spj74, Sve76, Thy75, THF18, Xie89, ZYT02, ZLZZ21, dCCU17, BSC24, BS21, CDQ20, HNNS19, KKW24, LLLP20, LPYZ24, PWN22, Par20, SZ20, ZHW19]. Tests [Aab83, Aly90, AL79, AL81, BBQ18, BQ22, Ban05, BS16, BQ09, Ber81, BBM06, BNM^{+}06, Bøl83b, BJMP14, CL05, Car82, CM84a, Cey10, CCH98, CCH01, CFJP07, CSJ^{+}77, CS90, DS09, DRT13, DPFV09, Dok80, Dok82, DF03, Edw80, Eks13, Eri96, FWW77, FGD12, FOS^{+}14, GJ05, GH14a, GPP96, Gro12, Haa08, Hol75a, Hol75b, HHM17, IKL94, Irl90, Jan91, JM93, JM01, JSG86, Jen81a, JQ15, Joh17, Jon91, Jon01a, Kre87, KP77a, LL90, LL96, LL06, LB98, Lus94, MG95, MU91, Mei06, MH10a, Miu78, MC97, NV09, NC15, Oja99, Pan02, Pap00, PP16, PdT91, Præ95, PW10, Qin98, Ren03, Rit04, Sai83, SL88, SM04b, SMSD92, SP09, SLCR14, STZ01, SR01, SZZ05, Sun75, Sve75, Sve77, TM86, VKY^{+}14, Yao96, ZLY14, ZHL15, Ahm17, ELY22]. tests [GHD20, GJ16, HBD^{+}20, KT19, KS22, ORL20, ZXLL23, ML75]. textbook [Agr23]. Their [Asm00, BBL87, GM84, Gup76, LS98, Nor90, Sas92, But98, KG18, Sei24, YLGL20]. Theorem [Ave85, DF90, ES91, Hor85, LLY18, LdUád15, Mol98, Mur95, SZ95, SS80, BDS22, BW19, BW08, BNHH95, Tor88]. Theorems [BJ93, Deg96, DEL92, FL11, GCL87, Hel82, Kni98, Lou98, Pol95, Seg02, SW18, BMP19, KH22, LYW22]. Theoretic [GH00, GH08]. Theoretical [Lin77, Nie84]. Theory [Ano98e, ACR16, BAR^{+}85, BO11, CSW79, CYM93, CW16, DGSL02, Ham88, HKK^{+}76, JN16a, JN16b, KM94, Kol81, Lav23, LZ99, LRT^{+}87, MSR16, NV17, Oja16, Ric23, Ron16, Ros74a, Ros74b, Sam89, SN88, SKBBN79, Sun74, Tho95, Zwa16, dMR88, Ano24, BKB23, Bic23, CV22, Gas23, Gre23b, JTT21, YK20, YM22]. Thickness [JGØ79]. thinning [Lu21]. thinning-based [Lu21]. Third [Dal77, Ess75, SB00]. Thomas [Edw78]. Thompson [Fra77b]. Three [BKO11, FWW77, MP84, PW10, SW76]. Three-Dimensional [PW10]. Threshold [BD13, LL12, Man09, NBY08, Shi17, SW18, AFV14, LP20, LM23, LC22]. Thresholded [BJFG15, TB22]. Thresholding [BV09, San14, VM00, SJZI19]. Ties [Kou76, SG78]. Tighter [Dal77]. Tilt [NC15]. Time [Aal12, ADMP19, AR80, AALM17, ACR16, BB10, Ber77b, Ber79a, BJ93, Blo74, BT08, Bor84a, Bor84b, BK95, CS03, Che15a, CLSZ16, CY17b, CGL^{+}81, CW16, Cro98, Dab87, Dam80, DS03b, DPV06, DDK04, DN15, DQYZ23, DP16, FL11, Fok01, Gad85, Gao98, GS99, Gri09, HM99, HBH17, Hel82, Hok76, Höp87, Hou12, HCS15, HS98, HS04, HP09, HHM17, HC10, JLY06, JN16a, JN16b, KHSS12a, KHSS12b, KKC17, Kle82, KL89, KS94, Lau76, LS98, LHNN03, LG09, LHWS18, LFL16, LST88, ML86, MS98a, MSSM02, Mur93, Nat93, NL16, Oja16, OSG08, Pap00, QQZ16, Ran75, Ron16, SL88, SM04b, Sha12, Slu92, SMS12, Sør01, SZZ05, SW05, SWS06, SSZ09, Tho83, Tjø94, TCC^{+}95, TC05, TRL15, WL04, fWZY16, Xie89, Zwa16, ACF^{+}21, ABB24, AH19, BLG20, BBD^{+}21, BBP21, BBS23]. time [BS21, BM01b, CZL24, CL20, CWZ21, FGLT23, GBB^{+}24, HYZ22, KP21, LM23, LJZ^{+}18, MP22, NJG18, NM24, PS20, PD22, PPS21, RFK22, SP22, dRSHK19, TKLM23, WCY22, Wol24, ZZLC21, ZCL22]. time-censored [WCY22]. Time-change [ADMP19]. Time-Changed [FL11]. Time-Continuous [KS94]. Time-Dependent [CS03, HC10, LFL16, SMS12, WL04, AALM17, HS98, SSZ09, LJZ^{+}18]. time-functional [CZL24]. Time-to-Event [HBH17, fWZY16, NJG18, NM24]. Time-Varying [KKC17, LFL16, SM04b, SW05, CY17b, DQYZ23, MSSM02, GBB^{+}24, HYZ22]. Times [ADGP14, BB10, DP18, GCL87, Gup76, HHVA03, Höp87, Huz99, MW08, Pon86, Ros77, Sch94, SMS12, Sør98, Ste88, Stu83, LLXH19, Par20]. Toeplitz [Mun02]. Tolerance [Gui04]. Tool [Bjö10, HT17]. Topics [HOF^{+}94, JNS^{+}83]. tortuosity [NHS^{+}19]. Total [Ber77b, Ber79a, Kle82, Kol81, ST10, Xie89]. totally [RSTU21]. Totals [Fra77a]. Tournament [YL96]. Trace [PKH17]. Tracking [HQR08]. Tractable [FS08, HJR06, BSO22]. Traffic [Blo74]. Training [VHK11]. trait [LLLP20]. Transfer [Puk82, Rah86]. Transform [LAO23, WLX19]. Transformation [Car82, DD88, DCIK14, Eri84, GQR06, JH17, LHWS18, Llo88, Lue15, MW10, NV04, NYR18, PKR^{+}97, Sko81b, SV04, SV05, KHBK22]. Transformations [BL90, HMB18, Hou86, OKW88]. Transformed [JKR02, LL12, MR24, SK19]. Transforms [BNL07, Kle82]. Transient [BB14]. Transition [AJ78, AHK91, Gil86, Gui77, Ran75, Ros78, SZ07, Tho83, DT20]. transmission [PD22]. Transplantation [AK07]. Transportation [CSW79]. Trawl [BNLSV14, CV22]. Treatment [LFL16, OB16, dCJV82, ARP23, FHSZ19, NM24, WGT19, YZ23, ZGZ22]. Treatments [CFS95, DW02, Ros89]. Tree [DR00, GGG13, Huc11, LN95, NBM12, MT19]. Trees [BNR00, MW97, PBHMC09, HPS24]. Trend [Jon01a, XY15, Zet88]. Trends [Ber79b]. Trento [NDH^{+}21]. Trials [CV14, Lin78b, SG04, Ste88, SBB05]. Triangular [MRM09]. Trimmed [Dha16, Miu81]. trimming [CFR19]. Triple [Ham88]. True [NM14, Swe88]. Truncated [Cac77, DP04, Fok01, Gär03, GHC92, Gro96, GS02, LdUád15, PK18, Ros78, TZ95, ZD24, WCJ18]. Truncation [KKP99, Thy75, EMS15, QB23]. Trygve [And91]. Tubes [LHW^{+}16]. Tuning [NC18]. Two [Aal87b, BKM18, Bon10, BDV06, BW07, BW08, CM82, Chr74, DW02, DFI14, FMHB16, GA86, Gro12, HBH17, HS06, JWL00, Joh17, KP77b, Lai83, LW12, LT08, LCZ09, LB98, Mad76, MP84, NM87, PWY97, Ped75b, Pon86, Præ95, PW10, Rap03, RS94, Sae15, SBR98, Sun83, Van98, YC22, Yao96, Zha00, CCV23, CLP^{+}19, CDO24, DO05, FR21, FGH20, HD22, HLP23, Hua24, JLRT19, MPV24, TKU23b, WZ22, ZD24]. Two-Action [Lai83]. Two-Component [BDV06, Chr74, JLRT19]. Two-Dimensional [Ped75b, CDO24]. two-factor [FGH20]. Two-group [BKM18]. Two-interval [Van98]. two-level [Hua24]. two-parameter [ZD24]. Two-part [YC22]. Two-Phase [Sae15, BW07, BW08, FMHB16]. Two-Region [RS94]. Two-Sample [Gro12, JWL00, Joh17, LW12, PW10, Zha00, FR21, HLP23, WZ22]. Two-Stage [HS06, Sun83, Yao96, LT08, LCZ09]. two-step [MPV24]. Two-Way [Rap03, DFI14]. Type [Aal95, Asm89, ANO96k, BNL07, BDP12, BJ93, BM01b, BCG08, FOS^{+}14, FM90, Gil86, Gro12, HGB96, HTK15, Huz99, JM93, Jon01a, KK00, LCZ14, Ols96, PdT87, Sun75, Sun83, VU05, Yu11, ABY22, BBdW20, CLP18, DP13, EU21, KL22, LPYZ24, MdCCD19, PFJGE15, VHF20, GG01, Gui04]. Type-II [GG01, Gui04]. Types [VB99].

Uhlenbeck [Die92, Eie83, FS10, NV17, SH21, VVI^{+}22]. Ultrahigh [YM22]. Ultrahigh-dimensional [YM22]. Umbrella [Wył16]. UMVU [Ber76]. unbalanced [FB20]. Unbiased [Bar03, Bon79, CFS95, GG13, Joh90, KD84, Nor75, Pfa93, ST81, Tan94]. Unbiasedness [GPP96]. Unbounded [TvdM96]. Uncertain [MW93, SKO17]. Uncertainty [BKW10, KKMP18, KHT14, Lon12, SO13, CGGI19, GK21, PD22]. unchanged [LT21]. Unconditional [Ber23, DGSL02, AH19]. underdispersion [PVFF24]. Underlying [PdT87]. Understanding [AF07]. Undirected [AMP97]. Unequal [And77a, BNP92, BG80, Bon10, BG11, CGL14, Kor82, NC92]. Unequally [Nor80, PW06]. Unicity [Jac89]. Unification [AVA06, AVA22]. Unified [CY17b, DB03, KV98, Rov05, SJS08, YM22, Bic23, ZHW19]. Uniform [BDY85, GM08b, HMP22, JH05, LL90, LDM15, Pol95, Rom04, SMV05, SPK23]. Uniformity [Gås16]. Uniformly [XY15, GAC23, PRV21]. Unimodality [Ped75a, Ped75b]. Unions [MH10b]. Unit [FMHB16, MWY15, BSC24, DR18, KT19]. unit-level [DR18]. Units [DH78]. Univariate [Bie07, LST88, Oja81, OS96, ABY22]. Unknown [BZ82, Erh08, Jen86, LY08, NBY08, Sun95, ZL18, DD22, LS23, SKR19]. Unmasking [Jon01a]. Unmeasured [Kur16]. Unobserved [Boe10]. Unpublished [Ano74d, Ano74e, Ano74f, Ano75e, Ano75f, Ano75g, Ano76e, Ano76f, Ano76g, Ano76h, Ano77e, Ano77f, Ano77g, Ano77h, Ano78e, Ano78f, Ano78g, Ano78h, Ano79e, Ano79f, Ano79g, Ano79h, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano82e, Ano82f, Ano82g, Ano82h, Ano83e, Ano83f, Ano83g, Ano83h, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h]. Unstable [BIP14, BSC24]. Unsupervised [ACMLM03]. unwanted [LS23]. Updates [RZM16]. Updating [KHR02, LT21]. Upgraded [Qin98]. Upper [AGJ07, Cur80b, Lin94]. Upsilon [BNL07]. Upwards [Ros77]. Urn [SB85]. urns [APM19]. Use [BKN23, Hok76, Lan07, Lun00, ML74, Tho81, TL03, Ber23]. Used [BEK83, MS86, See96]. Useful [Xie89]. Usefulness [Sve90b]. Using [Agr93, ACFS83b, AGR13, AOH00, BR14, CTYF13, CMN08, CHWY05, Che91, CFJP07, CS90, DLP08, FNR09, Fra78, GK00, Haa08, HN99, HJR06, HZZ07, JS12, KB04, KC11, LO16, MS98a, MW97, Que12, RVG15, SBV11, SMB14, SG15, Tho83, Toc01, WB15, Wan06, ZXL^{+}18, ZNJ15, dCJV82, ACMLM03, APCG24, Ano23, DD22, BCCAVMO21, GGS20, GH21, HBD^{+}20, JJCYG21, KYZC21, Kuh04, KH16, LGL19, LET20, MSR16, Mar99, MP14, MS23, MT19, OHIS24, PEK22, PNC17, SM12, WCJ18, YK20]. Utilities [Lon12].

Vacancy [BF03]. valid [Sei24]. validated [ZHS22]. Validation [DRM96, DH05, Gré93, HH82, JW10, YF12, Gua07, XZ09]. Validity [FVV10, Kos99, SL90]. Value [BR17, FR00, GA86, KY12, LGP11, AK07, BKB23, GHD20, HPS24, Sei24]. Value-at-Risk [LGP11]. Valued [Wal97, BIP14, BNLSV14, BSC24, HK24, VVI^{+}22, ZBS20]. Values [Dam80, GJ03, Hög79, LRT^{+}87, VKY^{+}14, ZXL^{+}18, Ano24, CV22, GS19a, Gas23, GMMT06, Gre23a, Gre23b, Lav23, Ric23, SS09, ZWS19, Vos23]. Vanishing [AHJ15]. VAR [Doo18]. Variable [Amu74, But86, CM20b, CDMR02, CTGS14, Dab92, FW03, HFS23, Hok75, ICG12, KS08, Kur16, Lue15, MC97, QMP15, SRH07, Sto11, THSS09, Tra11, WC12, WWW15, WLT15, XLY20, KMG21, VD18]. Variable-Sample-Size-Sequential [MC97]. Variables [AHJ15, BVV17, CW99, CK06, DR97, ICG12, Kab78, Kou84, KH16, LZ97, Nor81, Pap08, RL06, Sjö00, Wil77, XMW15, ZLSL14, ZC03, DNCZ21, DLS^{+}24, SW19, ZBS20]. Variance [And90, Arf98, BLBEO92, BG80, BF03, CGL14, CY17a, CM04, CFJP07, Eks08, Eri78, GM08a, Glo14, Hel98, JGW13, JMT94, KB04, Kle91, KT95, KD84, Kor00, LZ10, LCZ14, LB94, LB80, MT03, Miu81, MS12, OFFL12, PSW09, PC99, Sae15, Sch75, Sch81, Spj74, Sun95, TSH91, VHK11, Wu13, BW19, BR23, CZL24, CL21, FB20, Mes22, OPP18, PS20, ZLK21]. Variance* [Taq02]. Variance-Mean [Kor00]. Variances [BNP92, Kor82, MRS14, MS78, STK17]. Variate [DF74, BMP19, KC11, LC22]. Variates [CAS03, HT08]. Variation [ST10, LS23]. Variations [BCS00, Kur18]. Variogram [MGSFB08]. VARMA [AALM17]. Varying [CLSZ16, FZ00, HC17, HYWC18, KKC17, LFL16, SM04b, ŞM05, SW05, THF18, WL04, CY17b, DQYZ23, GBB^{+}24, HYZ22, MSSM02, ZKP^{+}24]. Varying-Coefficient [FZ00, WL04, ZKP^{+}24]. Vector [Cav16, LG09, OT09, VVI^{+}22, BMP19]. Vector-valued [VVI^{+}22]. Vectors [BR17, Gui82, JTT21, LT21]. Version [PR07, SN88, Gås03]. Versus [DK06, Aar85, And77b, BTL06, CM15, Gre23b, LL90, LHML16, MP21]. Vertex [HKD02]. Vertices [HQR08]. Via [BIPV13, Efr05, And83, ANO96k, AFV14, CH04, DDK04, EGB13, GK13, HJG21, HPS24, HH16, JSDT11, KS08, LDW06, LXZ16, LPR23, LLS^{+}22, LPYZ24, LZL^{+}24, OS97, PCW02, Rub04a, ŞM05, SH21, SSZ09, THSS09, ZL14, ZX19]. View [BS99, BM01b, JT07, MS94, MDA10, Por16]. vine [YC22]. Visualization [RZM16]. Visualizing [HT17]. vol [AVA22]. Volatility [BN97, BNS03, CGC06, DPV06, FZ06, GCJL03, JSDT11, Sør03, Vid01, XY15, FGH20, Koi14, MN21, WLX19]. Volume [Ano74k, Ano75k, Ano76m, Ano77m, Ano78m, Ano79n, Ano80m, Ano81m, Ano82m, Ano83n, Ano84m, Ano85m, Ano86m, Ano87l, Ano88l, Ano89l, Ano90m, Ano91m, Ano92i, Ano93i, Ano94i, Ano95i, Ano96j, Ano97a, Ano98e, Ano98j, Ano99i, Ano00i, Ano00j, Ano01i, Ano01j, Ano02i, Ano03i, Ano04i, Ano05j, Ano05k, Ano06i, Ano07k, Ano10k, Ano11j, ZNJ15]. vs [Ano24, Gas23, Gre23a, Lav23, Ric23, Vos23].

Wage [LYZ15]. Waiting

[Blo74, Gad85, Hok76, Huz99]. Walk [LR08, ML86]. Waring [ZX19]. Warped [Cha15]. Wasserstein [MM24]. Watson [CL19, MR23, Ner77, Per79]. Wavelet [AAA04, BV09, DP06, Efr05, MPV19, San14, Tri03, VM00]. Wavelet-based [MPV19]. Wavelets [GCLP92, SM24]. Way [HKJ11, Nor77, Per79, Rap03, ZL14, DFI14]. Weak [BDP13, Dab96, DL01, GR10, SJ94, XY15, Yuk92, vdV94]. Weak/Strong [XY15]. Weakly [CTYF13, Van14]. Weakly-Identifiable [CTYF13]. Wear [Lo81]. Weather [GH02]. wedge [LKT^{+}23]. Weeds [BM01b]. Weibull [CK97, EPM15, Møl76]. Weight [HESZ16, KXZA20]. Weighted [BB11, BJMP14, BW07, Che15b, DBS10, Gro96, HM99, Hin79, NYR18, SZ95, ZZLZ16, RAQ21, SZ20, XWH14, BW08]. Weighting [Han16, AV21]. Weights [BH97, BR14, LL96, Li01]. Where [BDV06, Nat75]. Which [Azz85, Nor75]. White [Kle99]. Whitney [FOS^{+}14]. Whittle [LL99, XT20]. Who [BM15]. Whole [Bri97]. Whose [BR17, MS91]. Wicksell [Jon01b]. wide [HBD^{+}20, ZXLL23]. Width [Ano83i, DK80, GM83]. Wiener [Hok76]. Wilcoxon [CB84]. Wild [BDP13, CFR19, HW98]. Wilks [BW04, ZWS19]. Window [CC98, Sch75, Sch81]. Wishart [BO11, But98, HK24, LM04, Rov02, vR88]. Within [Bri97, Kou79, SG04, Chr89, SO97]. Within-Family [Bri97]. Within-Pair [Kou79]. Within-Subject [SG04]. Without [BDL^{+}17, GR01, GL02, ML75, Sve75, Sve77, ZV21]. Wood [SLB06]. working [MWW24]. Workshop [NDH^{+}21]. world [ZV21]. Worth [Lav23, Ric23, Ano24, Gas23, Gre23b]. Wright [SJKS22].

year [MS23]. Yield [TW04]. Yokes [BNB93]. Yor [ZD24]. Yule [Joh90].

Zelterman [Böh10]. Zero [HS10, LC11b, SS02, DBJ^{+}22]. Zero-Bias [SS02]. Zero-Inflated [HS10, LC11b]. Zeros [Mül85, PK18, DEH21]. ZIP [AH19]. ZIPdistributed [AH19]. Zygmund [IS99].

References

Abramovich:2004:OBW

[AAA04] Felix Abramovich, Umberto Amato, and Claudia Angelini. On optimality of Bayesian wavelet estimators. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 217–234, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                                       /www.jstor.org/stable/
                                                       4616825.

Aaberge:1983:AEC

[Aab83]  Rolf Aaberge.  Asymp-

totic efficiency of certain tests for comparison of proportions. Scandinavian Journal of Statistics. Theory and Applications, 10 (2):125–133, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                                       https://www.jstor.org/
                                                       stable/4615911.

Alomari:2020:ECL

[AAFO20] Huda Mohammed Alomari, Antoine Ayache, Myriam Fradon, and Andriy Olenko. Estimation of cyclic long-memory parameters. Scandinavian Journal of Statistics. Theory and Applications, 47(1): 104–133, March 2020. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Aalen:1976:NIC

[Aal76] Odd Aalen. Nonparametric inference in connection with multiple decrement models. Scandinavian Journal of Statistics. Theory and Applications, 3 (1):15–27, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615603.

Aalen:1987:MDM

[Aal87a] Odd O. Aalen. Mixing distributions on a Markov chain. Scandinavian Journal of Statistics. Theory and Applications, 14(4): 281–289, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616071.

Aalen:1987:TEM

[Aal87b] Odd O. Aalen. Two examples of modelling heterogeneity in survival analysis. Scandinavian Journal of Statistics. Theory and Applications, 14(1):19– 25, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616045.

Aalen:1995:PTD

[Aal95] Odd O. Aalen. Phase type distributions in survival analysis. Scandinavian Journal of Statistics. Theory and Applications, 22(4):447–463, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616373.

Aalen:2004:DC

[Aal04] Odd O. Aalen. Discussion of causality. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 193–195, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616823. See [AP04].

Aalen:2012:DCD

[Aal12] Odd O. Aalen. Discussion of ‘The Current Duration Approach to Estimating Time to Pregnancy’. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 205–206, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

41679788. See [KHSS12a].

Alj:2017:APQ

[AALM17] Abdelkamel Alj, Rajae Azrak, Christophe Ley, and Guy Mélard. Asymptotic properties of QML estimators for VARMA models with time-dependent coefficients. Scandinavian Journal of Statistics. Theory and Applications, 44(3):

617–635, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Aarset:1985:NDT

[Aar85] Magne Vollan Aarset. The null distribution for a test of constant versus “bathtub” failure rate. Scandinavian Journal of Statistics. Theory and Applications, 12(1):55–61, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615972.

Aalen:1985:ARE

[AB85] Odd O. Aalen and Edvin Bredrup. Asymptotic results for estimators in restricted randomization designs. Scandinavian Journal of Statistics. Theory and Applications, 12(3): 203–211, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615989.

Ahmad:2024:EAT

[ABB24] Jamaal Ahmad, Martin Bladt, and Mogens Bladt. Estimating absorption time distributions of general Markov jump processes. Scandinavian Journal of Statistics. Theory and Applications, 51(1):171–200, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Adhya:2011:IPR

[ABC11] Sumanta Adhya, Tathagata Banerjee, and Gaurangadeb Chattopadhyay. Inference on polychotomous responses in finite populations. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 788–800, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41411151.

Andersen:1985:CPM

[ABH^{+}85] Per Kragh Andersen, Ørnulf Borgan, Nils Lid Hjort, Elja Arjas, Jon Stene, and Odd Aalen. Counting process models for life history data: a review [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 12(2): 97–158, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615980.

Andersen:1996:ESF

[ABK96] Per Kragh Andersen, Michael Weis Bentzon, and John P. Klein. Estimating the survival function in the proportional hazards regression model: a study of the small sample size properties. Scandinavian Journal of Statistics. Theory and Applications, 23(1):1– 12, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616381.

Aalen:1980:IBL

[ABKT80] Odd O. Aalen, Ørnulf Borgan, Niels Keiding, and Jens Thormann. Interaction between life history events. nonparametric analysis for prospective and retrospective data in the presence of censoring. Scandinavian Journal of Statistics. Theory and Applications, 7(4):161–171, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615796.

Ang:2012:GCS

[ABN12] Qi Wei Ang, Adrian Baddeley, and Gopalan Nair. Geometrically corrected second order analysis of events on a linear network, with applications to ecology and criminology. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 591–617, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357213.

Abt:1999:EPM

[Abt99]  Markus Abt.  Estimat-

ing the prediction mean squared error in Gaussian stochastic processes with exponential correlation structure. Scandinavian Journal of Statistics. Theory and Applications, 26(4):563–578, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616579.

Albrecher:2022:FIP

[ABY22] Hansjörg Albrecher, Mogens Bladt, and Jorge Yslas. Fitting inhomogeneous phase-type distributions to data: the univariate and the multivariate case. Scandinavian Journal of Statistics. Theory and Applications, 49(1):44– 77, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anh:1999:EGL

[AC99] V. V. Anh and T. Chelliah. Estimated generalized least squares for random coefficient regression models. Scandinavian Journal of Statistics. Theory and Applications, 26(1):31– 46, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616539.

Aeberhard:2021:RED

[ACF^{+}21] William H. Aeberhard, Eva Cantoni, Chris Field, Hans R. Künsch, Joanna Mills Flemming, and Ximing Xu. Robust estimation for discrete-time state space models. Scandinavian Journal of Statistics. Theory and Applications, 48(4): 1127–1147, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Andersen:1983:EPB

[ACFS83a] Per Kragh Andersen, Erik Christensen, Lis Fauerholdt, and Poul Schlichting. Evaluating prognoses based on the proportional hazards model. Scandinavian Journal of Statistics. Theory and Applications, 10(2):141–144, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615913.

Andersen:1983:MPU

[ACFS83b] Per Kragh Andersen, Erik Christensen, Lis Fauerholdt, and Poul Schlichting. Measuring prognosis using the proportional hazards model. Scandinavian Journal of Statistics. Theory and Applications, 10(1):49–52, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615901.

Abraham:2003:UCC

[ACMLM03] C. Abraham, P. A. Cornillon, E. Matzner-Løber, and N. Molinari. Unsupervised curve clustering using B-splines. Scandinavian Journal of Statistics. Theory and Applications, 30(3): 581–595, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616785.

Atkinson:2016:DAT

[ACR16] Anthony C. Atkinson, Andrea Cerioli, and Marco Riani. Discussion of ‘Asymptotic Theory of Outlier Detection Algorithms for Linear Time Series Regression Models’ by Johansen and Nielsen. Scandinavian Journal of Statistics. Theory and Applications, 43(2):349–352, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JN16a].

Azais:2014:NPE

[ADGP14] Romain Azaïs, François Dufour, and Anne Gégout-Petit. Non-parametric estimation of the conditional distribution of the interjumping times for piecewise-deterministic Markov processes. Scandinavian Journal of Statistics. Theory and Applications, 41(4):950–969, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586815.

Adimari:1997:ELR

[Adi97] Gianfranco Adimari. On the empirical likelihood ratio for smooth functions of M-functionals. Scandinavian Journal of Statistics. Theory and Applications, 24(1):47–59, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616436.

Arima:2015:BES

[ADL15] Serena Arima, Gauri S.

Datta, and Brunero Liseo. Bayesian estimators for small area models when auxiliary information is measured with error. Scandinavian Journal of Statistics. Theory and Applications, 42(2):518–529, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ailliot:2019:TCM

[ADMP19] Pierre Ailliot, Bernard Delyon, Valérie Monbet, and Marc Prevosto. Timechange models for asymmetric processes. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1072–1097, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Adekpedjou:2021:SES

[ADN21] Akim Adekpedjou and Sophie Dabo-Niang. Semiparametric estimation with spatially correlated recurrent events. Scandinavian Journal of Statistics. Theory and Applications, 48(4): 1097–1126, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Adekpedjou:2015:DDC

[ADZ15] Akim Adekpedjou, Withanage A. De Mel, and Gideon Kd Zamba. Data dependent cells chi-square test with recurrent events. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1045–1064, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Aalen:2007:WCS

[AF07] Odd O. Aalen and Arnoldo Frigessi. What can statistics contribute to a causal understanding? Scandinavian Journal of Statistics. Theory and Applications, 34(1):155–168, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548544.

Antoniadis:2010:DSC

[AFL10] Anestis Antoniadis, Piotr Fryzlewicz, and Frédérique LetuÉ. The Dantzig selector in Cox’s proportional hazards model. Scandinavian Journal of Statistics. Theory and Applications, 37(4):531–552, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000407.

Autin:2014:BTA

[AFV14] Florent Autin, Jean-Marc Freyermuth, and Rainer Von Sachs. Block-threshold-adapted estimators via a maxiset approach. Scandinavian Journal of Statistics. Theory and Applications, 41(1):240–258, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586740.

Andersen:1985:EEB

[AG85] Per Kragh Andersen and Anders Green. Evaluation of estimation bias in an illness-death-emigration model. Scandinavian Journal of Statistics. Theory and Applications, 12(1):63– 68, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615973.

Antoniadis:1990:PLE

[AG90] Anestis Antoniadis and Gérard Grégoire. Penalized likelihood estimation for rates with censored survival data. Scandinavian Journal of Statistics. Theory and Applications, 17 (1):43–63, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616154.

Amorino:2020:CFE

[AG20] Chiara Amorino and Arnaud Gloter. Contrast function estimation for the drift parameter of ergodic jump diffusion process. Scandinavian Journal of Statistics. Theory and Applications, 47(2):279–346, June 2020.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Arribas-Gil:2006:PEP

[AGGM06] Ana Arribas-Gil, Elisabeth Gassiat, and Catherine Matias. Parameter estimation in pair-hidden Markov models. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 651–671, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616951.

Aarts:2007:EUS

[AGJ07] Lucie Aarts, Piet Groeneboom, and Geurt Jongbloed. Estimating the upper support point in deconvolution. Scandinavian Journal of Statistics. Theory and Applications, 34(3): 552–568, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548569.

Antoniadis:2000:NPE

[AGM00] Anestis Antoniadis, Irène Gijbels, and Brenda Macgibbon. Non-parametric estimation for the location of a change-point in an otherwise smooth hazard function under random censoring. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 501–519, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616619.

Agresti:1993:CCM

[Agr93] Alan Agresti. Computing conditional maximum likelihood estimates for generalized Rasch models using simple loglinear models with diagonals parameters. Scandinavian Journal of Statistics. Theory and Applications, 20(1):63– 71, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616260.

Arbel:2013:BOA

[AGR13] Julyan Arbel, Ghislaine Gayraud, and Judith Rousseau. Bayesian optimal adaptive estimation using a sieve prior. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 549–570, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586686.

Aalen:2018:FMC

[AGR^{+}18] Odd O. Aalen, Jon Michael Gran, Kjetil Røysland, Mats Julius Stensrud, and Susanne Strohmaier. Feedback and mediation in causal inference illustrated by stochastic process models. Scandinavian Journal of Statistics. Theory and Applications, 45(1):62– 86, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Agresti:2023:HOT

[Agr23]  Alan Agresti.  A his-

torical overview of textbook presentations of statistical science. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1641–1666, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Alam:1978:ERC

[AH78] Khursheed Alam and James S.

Hawkes. Estimation of regression coefficients. Scandinavian Journal of Statistics. Theory and Applications, 5(3):169–172, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615708.

Arjas:1984:MPP

[AH84] Elja Arjas and Pentti Haara. A marked point process approach to censored failure data with complicated covariates. Scandinavian Journal of Statistics. Theory and Applications, 11(4):193–209, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615959.

Arjas:1987:LRM

[AH87] Elja Arjas and Pentti Haara. A logistic regression model for hazard: Asymptotic results. Scandinavian Journal of Statistics. Theory and Applications, 14(1): 1–18, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616044.

Arjas:1992:OSL

[AH92] Elja Arjas and Pentti Haara. Observation scheme and likelihood. Scandinavian Journal of Statistics. Theory and Applications, 19(2):111–132, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616232.

Allevius:2019:UST

[AH19] Benjamin Allévius and Michael Höhle. An unconditional space–time scan statistic for ZIP-distributed data. Scandinavian Journal of Statistics. Theory and Applications, 46(1): 142–159, March 2019. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Andersen:2015:OPS

[AHJ15] Ina Trolle Andersen, Ute Hahn, and Eva B. Vedel Jensen. Optimal PPS sampling with vanishing auxiliary variables — with applications in microscopy. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1136–1148, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Andersen:1991:NSP

[AHK91] Per Kragh Andersen, Lars Sommer Hansen, and Niels Keiding. Non-and semiparametric estimation of transition probabilities from censored observation of a non-homogeneous Markov process. Scandinavian Journal of Statistics. Theory and Applications, 18(2): 153–167, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616198.

Ahmad:1981:SAP

[Ahm81] Ibrahim A. Ahmad. On some asymptotic properties of U-statistics. Scandinavian Journal of Statistics. Theory and Applications, 8(3):175–182, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615830.

Ahmad:2017:LIM

[Ahm17] M. Rauf Ahmad. Locationinvariant multi-sample Utests for covariance matrices with large dimension. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 500–523, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Abramowicz:2018:NIF

[AHP^{+}18] Konrad Abramowicz, Charlotte K. Häger, Alessia Pini, Lina Schelin, Sara Sjöstedt de Luna, and Simone Vantini. Nonparametric inference for functional-on-scalar linear models applied to knee kinematic hop data after injury of the anterior cruciate ligament. Scandinavian Journal of Statistics. Theory and Applications, 45(4):1036–1061, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anevski:2019:SPA

[AHWKP19] Dragi Anevski, Richard Hall-Wilton, Kalliopi Kanaki, and Vladimir Pastukhov. A stochastic process approach to multilayer neutron detectors. Scandinavian Journal of Statistics. Theory and Applications, 46(2):621–635, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Aalen:1978:ETM

[AJ78] Odd O. Aalen and Søren Johansen. An empirical transition matrix for non-homogeneous Markov chains based on censored observations. Scandinavian Journal of Statistics. Theory and Applications, 5(3): 141–150, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615704.

Artes:2000:LDE

[AJ00] Rinaldo Artes and Bent Jørgensen. Longitudinal data estimating equations for dispersion models. Scandinavian Journal of Statistics. Theory and Applications, 27(2): 321–334, June 2000. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616606.

Aires:2002:OSD

[AJN02] Nibia Aires, Johan Jonasson, and Olle Nerman. Order sampling design with prescribed inclusion probabilities. Scandinavian Journal of Statistics. Theory and Applications, 29 (1):183–187, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616708.

Asmussen:2016:EFT

[AJRN16] Søren Asmussen, Jens Ledet Jensen, and Leonardo Rojas-Nandayapa. Exponential family techniques for the lognormal left tail. Scandinavian Journal of Statistics. Theory and Applications, 43(3):774–787, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Andersen:2007:RAM

[AK07] Per K. Andersen and John P. Klein. Regression analysis for multistate models based on a pseudo-value approach, with applications to bone marrow transplantation studies. Scandinavian Journal of Statistics. Theory and Applications, 34(1):3–16, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548535.

Arjas:1989:SMM

[AKB^{+}89] Elja Arjas, Niels Keiding, Ørnulf Borgan, Per Kragh Andersen, and Bent Natvig. Survival models and martingale dynamics [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 16(3):177–225, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616135.

Anderson:1980:SGM

[AKC80] Dallas W. Anderson, Leslie Kish, and Richard G. Cornell. On stratification, grouping and matching. Scandinavian Journal of Statistics. Theory and Applications, 7(2):61– 66, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615774.

Arendarczyk:2022:GSS

[AKP22] Marek Arendarczyk, Tomasz J.

Kozubowski, and Anna K. Panorska. The Greenwood statistic, stochastic dominance, clustering and heavy tails. Scandinavian Journal of Statistics. Theory and Applications, 49 (1):331–352, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Amundsen:1979:SST

[AL79] H. T. Amundsen and H. Ljøgodt. Small sample tests against an ordered set of binomial probabilities. Scandinavian Journal of Statistics. Theory and Applications, 6(2):81– 85, December 1979. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615736. See correction [AL81].

Amundsen:1981:CNS

[AL81] H. T. Amundsen and H. Ljøgodt. Correction note: Small Sample Tests against an Ordered Set of Binomial Probabilities. Scandinavian Journal of Statistics. Theory and Applications, 8(1):56, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615814. See [AL79].

Anderson:1998:EAD

[AL98] T. W. Anderson and Petter Laake. Exact and approximate distributions of the maximum likelihood estimate in a simple factor analysis model. Scandinavian Journal of Statistics. Theory and Applications, 25(1):39–51, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616483. See cor-

rection [AL99].

Anderson:1999:CNE

[AL99] T. W. Anderson and Petter Laake. Correction note to “Exact and Approximate Distributions of the Maximum Likelihood Estimate in a Simple Factor Analysis Model”. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 477, September 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616570. See [AL98].

Alam:1977:MEM

[Ala77] Khursheed Alam. Minimax estimators of a multivariate normal mean. Scandinavian Journal of Statistics. Theory and Applications, 4(3): 125–130, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615663.

Aly:1990:TMP

[Aly90] Emad-Eldin A. A. Aly. Tests for monotonicity properties of the mean residual life function. Scandinavian Journal of Statistics. Theory and Applications, 17(3):189–200, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616168.

Averous:1984:BCD

[AM84] Jean Averous and Michel Meste. Best central distribution and interval for the Kolmogorov distance: a generalization of the median. Scandinavian Journal of Statistics. Theory and Applications, 11(1):17– 27, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615937.

Andersson:1997:MEC

[AMP97] Steen A. Andersson, David Madigan, and Michael D. Perlman. On the Markov equivalence of chain graphs, undirected graphs, and acyclic digraphs. Scandinavian Journal of Statistics. Theory and Applications, 24(1):81–102, March 1997.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616438.

Andersson:2001:AMP

[AMP01] Steen A. Andersson, David Madigan, and Michael D. Perlman. Alternative Markov properties for chain graphs. Scandinavian Journal of Statistics. Theory and Applications, 28(1):33– 85, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616644.

Amundsen:1974:BVM

[Amu74] H. T. Amundsen. Binary variable multiple regressions. Scandinavian Journal of Statistics. Theory and Applications, 1(2):59– 70, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615554.

Amundsen:1976:BRP

[Amu76]  H. T. Amundsen.  Bi-

nary regressions for a polytomeous regressand. Scandinavian Journal of Statistics. Theory and Applications, 3(1):39–41, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615607.

Andersen:1974:MCT

[And74] A. Holst Andersen. Multidimensional contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 1(3):115– 127, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615563.

Andersen:1977:MPM

[And77a] Erling B. Andersen. Multiplicative Poisson models with unequal cell rates. Scandinavian Journal of Statistics. Theory and Applications, 4(4):153–158, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615671.

Anderson:1977:EVP

[And77b] Harald Anderson. Efficiency versus protection in a general randomized response model. Scandinavian Journal of Statistics. Theory and Applications, 4(1):11–19, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615642.

Anderson:1979:NBR

[And79] Harald Anderson. On nonresponse bias and response probabilities. Scandinavian Journal of Statistics. Theory and Applications, 6(3): 107–112, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615743.

Andersen:1982:LSA

[And82] Erling B. Andersen. Latent structure analysis: a survey. Scandinavian Journal of Statistics. Theory and Applications, 9(1):1– 12, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615848.

Andersen:1983:CSD

[And83] Per Kragh Andersen. Comparing survival distributions via hazard ratio estimates. Scandinavian Journal of Statistics. Theory and Applications, 10(2):77– 85, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615907.

Andersson:1990:LSO

[And90] Steen Arne Andersson. The lattice structure of orthogonal linear models and orthogonal variance component models. Scandinavian Journal of Statistics. Theory and Applications, 17 (4):287–319, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616179.

Anderson:1991:THS

[And91] T. W. Anderson. Trygve haavelmo and simultaneous equation models. Scandinavian Journal of Statistics. Theory and Applications, 18(1):1–19, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616185.

Andrade:2023:ROS

[And23] J. Ailton A. Andrade. On the robustness to outliers of the Student-t process. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 725–749, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 1(1):??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615551.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 1(2):??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615559.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 1(3):??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615568.

Anonymous:1974:BICa

[Ano74d] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 1(1):41– 48, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615550.

Anonymous:1974:BICb

[Ano74e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 1(2):93– 96, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615558.

Anonymous:1974:BICc

[Ano74f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 1(3):153– 156, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615567.

Anonymous:1974:EP

[Ano74g] Anonymous. Editorial preface. Scandinavian Journal of Statistics. Theory and Applications, 1(1):1, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615543.

Anonymous:1974:FMa

[Ano74h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 1(1):??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615542.

Anonymous:1974:FMb

[Ano74i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 1(2):??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615552.

Anonymous:1974:FMc

[Ano74j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 1(3):??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615561.

Anonymous:1974:VI

[Ano74k] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 1(3): ??, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615560.

Anonymous:1975:A

[Ano75a] Anonymous. Accessibility. Scandinavian Journal of Statistics. Theory and Applications, 2(2):??, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:1975:BMa

[Ano75b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 2(1):??, December 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615575.

Anonymous:1975:BMb

[Ano75c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 2(2):??, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615588.

Anonymous:1975:BMc

[Ano75d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 2(3):??, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615600.

Anonymous:1975:BICa

[Ano75e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 2(1):43– 48, December 1975. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615574.

Anonymous:1975:BICb

[Ano75f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 2(2):109– 112, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615587.

Anonymous:1975:BICc

[Ano75g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 2(3):166– 171, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615599.

Anonymous:1975:FMa

[Ano75h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 2(1):??, December 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615569.

Anonymous:1975:FMb

[Ano75i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 2(2):??, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615576.

Anonymous:1975:FMc

[Ano75j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 2(3):??, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615590.

Anonymous:1975:VI

[Ano75k] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 2(3): ??, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615589.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 3(1):??, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615609.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 3(2):??, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615619.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 3(3):??, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615630.

Anonymous:1976:BMd

[Ano76d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 3(4):??, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615639.

Anonymous:1976:BICa

[Ano76e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 3(1):42– 48, September 1976. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615608.

Anonymous:1976:BICb

[Ano76f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 3(2):95– 96, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615618.

Anonymous:1976:BICc lished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 3(3):140– 144, September 1976. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615629.

Anonymous:1976:BICd

[Ano76h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 3(4):218– 220, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615638.

Anonymous:1976:FMa

[Ano76i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 3(1):??, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615601.

Anonymous:1976:FMb

[Ano76j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 3(2):??, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615610.

Anonymous:1976:FMc

[Ano76k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 3(3):??, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615620.

Anonymous:1976:FMd

[Ano76l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 3(4):217, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615632.

Anonymous:1976:VI

[Ano76m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 3(4): ??, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615631.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 4(1):??, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615650.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 4(2): ??, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615657.

Anonymous:1977:BMc

[Ano77c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 4(3):??, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615667.

Anonymous:1977:BMd

[Ano77d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 4(4):??, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615679.

Anonymous:1977:BICa

[Ano77e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 4(1):44– 48, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615649.

Anonymous:1977:BICb

[Ano77f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 4(2):93– 96, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615656.

Anonymous:1977:BICc

[Ano77g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 4(3):139– 144, December 1977. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615666.

Anonymous:1977:BICd

[Ano77h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 4(4):186– 188, September 1977. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615678.

Anonymous:1977:FMa

[Ano77i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 4(1):??, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615640.

Anonymous:1977:FMb

[Ano77j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 4(2): ??, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615651.

Anonymous:1977:FMc

[Ano77k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 4(3):??, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615658.

Anonymous:1977:FMd

[Ano77l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 4(4):185, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615669.

Anonymous:1977:VI

[Ano77m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 4(4): ??, September 1977. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615668.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 5(1):??, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615690.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 5(2):??, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615701.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 5(3):??, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615710.

Anonymous:1978:BMd

[Ano78d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 5(4): ??, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615721.

Anonymous:1978:BICa

[Ano78e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 5(1):75– 80, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615689.

Anonymous:1978:BICb

[Ano78f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 5(2):125– 128, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615700.

Anonymous:1978:BICc

[Ano78g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 5(3):173– 176, December 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615709.

Anonymous:1978:BICd

[Ano78h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 5(4):215– 220, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615720.

Anonymous:1978:FMa

[Ano78i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 5(1):??, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615680.

Anonymous:1978:FMb

[Ano78j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 5(2):??, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615691.

Anonymous:1978:FMc

[Ano78k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 5(3):??, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615702.

Anonymous:1978:FMd

[Ano78l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 5(4): 213, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615712.

Anonymous:1978:VI

[Ano78m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 5(4): ??, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615711.

Anonymous:1979:BMa

[Ano79a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 6(1): ??, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615729.

Anonymous:1979:BMb

[Ano79b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 6(2):??, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615740.

Anonymous:1979:BMc

[Ano79c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 6(3):??, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615750.

Anonymous:1979:BMd

[Ano79d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 6(4):??, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615762.

Anonymous:1979:BICa

[Ano79e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 6(1):45– 48, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615728.

Anonymous:1979:BICb

[Ano79f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 6(2):92– 96, December 1979. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615739.

Anonymous:1979:BICc

[Ano79g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 6(3):140– 144, December 1979. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615749.

Anonymous:1979:BICd

[Ano79h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 6(4):184– 188, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615761.

Anonymous:1979:E

[Ano79i]  Anonymous.  Editorial.

Scandinavian Journal of Statistics. Theory and Applications, 6(4):179, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615758.

Anonymous:1979:FMa

[Ano79j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 6(1): ??, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615722.

Anonymous:1979:FMb

[Ano79k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 6(2):??, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615730.

Anonymous:1979:FMc

[Ano79l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 6(3):??, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615741.

Anonymous:1979:FMd

[Ano79m] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 6(4):??, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615752.

Anonymous:1979:VI

[Ano79n] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 6(4): ??, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615751.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 7(1):??, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615771.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 7(2):??, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615785.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 7(3):??, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615793.

Anonymous:1980:BMd

[Ano80d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 7(4):??, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615804.

Anonymous:1980:BICa

[Ano80e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 7(1):44– 48, December 1980. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615770.

Anonymous:1980:BICb

[Ano80f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 7(2):106– 111, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615784.

Anonymous:1980:BICc

[Ano80g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 7(3):155– 160, December 1980. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615792.

Anonymous:1980:BICd

[Ano80h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 7(4):216– 220, September 1980. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615803.

Anonymous:1980:FMa

[Ano80i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 7(1):43, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615763.

Anonymous:1980:FMb

[Ano80j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 7(2):??, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615772.

Anonymous:1980:FMc

[Ano80k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 7(3):??, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615786.

Anonymous:1980:FMd

[Ano80l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 7(4):??, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615795.

Anonymous:1980:VI

[Ano80m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 7(4): ??, September 1980. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615794.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 8(1):64, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615816.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 8(2): ??, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615824.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 8(3):??, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615835.

Anonymous:1981:BMd

[Ano81d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 8(4): ??, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615846.

Anonymous:1981:BICa

[Ano81e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 8(1):57– 63, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615815.

Anonymous:1981:BICb

[Ano81f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 8(2):125– 128, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615823.

Anonymous:1981:BICc

[Ano81g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 8(3):189– 192, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615834.

Anonymous:1981:BICd

[Ano81h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 8(4):250– 252, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615845.

Anonymous:1981:FMa

[Ano81i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 8(1):??, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615805.

Anonymous:1981:FMb

[Ano81j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 8(2): ??, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615817.

Anonymous:1981:FMc

[Ano81k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 8(3):??, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615825.

Anonymous:1981:FMd

[Ano81l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 8(4): ??, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615837.

Anonymous:1981:VI

[Ano81m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 8(4): ??, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615836.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 9(1): ??, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615857.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 9(2): ??, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615866.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 9(3):??, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615879.

Anonymous:1982:BMd

[Ano82d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 9(4):??, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615892.

Anonymous:1982:BICa

[Ano82e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 9(1):59– 64, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615856.

Anonymous:1982:BICb

[Ano82f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 9(2):122– 128, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615865.

Anonymous:1982:BICc

[Ano82g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 9(3):187– 192, September 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615878.

Anonymous:1982:BICd

[Ano82h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 9(4):248– 252, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615891.

Anonymous:1982:FMa

[Ano82i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 9(1): ??, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615847.

Anonymous:1982:FMb

[Ano82j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 9(2): ??, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615858.

Anonymous:1982:FMc

[Ano82k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 9(3):??, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615867.

Anonymous:1982:FMd

[Ano82l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 9(4):??, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615881.

Anonymous:1982:VI

[Ano82m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 9(4): ??, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615880.

Anonymous:1983:BMa

[Ano83a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 10(1): 64, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615904.

Anonymous:1983:BMb

[Ano83b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 10(2): ??, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615916.

Anonymous:1983:BMc

[Ano83c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 10(3): 256, September 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615924.

Anonymous:1983:BMd

[Ano83d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 10(4): 316, September 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615934.

Anonymous:1983:BICa

[Ano83e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 10(1):57– 63, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615903.

Anonymous:1983:BICb

[Ano83f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 10(2): 150–160, March 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615915.

Anonymous:1983:BICc

[Ano83g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 10(3): 251–255, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615923.

Anonymous:1983:BICd

[Ano83h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 10(4): 307–315, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615933.

Anonymous:1983:CNS

[Ano83i] Anonymous. Correction note: A Strong Representation for Linear Combinations of Order Statistics with Application to Fixed-Width Confidence Intervals for Location and Scale Parameters. Scandinavian Journal of Statistics. Theory and Applications, 10(4): ??, September 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615927. See [GM83].

Anonymous:1983:FMa

[Ano83j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 10(1): ??, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615893.

Anonymous:1983:FMb

[Ano83k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 10(2): ??, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615905.

Anonymous:1983:FMc

[Ano83l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 10(3):??, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615917.

Anonymous:1983:FMd

[Ano83m] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 10(4):??, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615926.

Anonymous:1983:VI

[Ano83n] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 10(4): ??, September 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615925.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 11(1):

??, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615943.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 11(2): ??, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615949.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 11(3): ??, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615956.

Anonymous:1984:BMd

[Ano84d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 11(4):??, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615968.

Anonymous:1984:BICa

[Ano84e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 11(1):59– 63, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615942.

Anonymous:1984:BICb

[Ano84f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 11(2): 123–128, March 1984. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615948.

Anonymous:1984:BICc

[Ano84g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 11(3): 187–192, March 1984. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615955.

Anonymous:1984:BICd

[Ano84h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 277–282, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615967.

Anonymous:1984:FMa

[Ano84i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 11(1): ??, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615935.

Anonymous:1984:FMb

[Ano84j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 11(2): ??, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615944.

Anonymous:1984:FMc

[Ano84k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 11(3):

??, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615950.

Anonymous:1984:FMd

[Ano84l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 11(4):??, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615958.

Anonymous:1984:VI

[Ano84m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 11(4): ??, September 1984. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615957.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 12(1): ??, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615978.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 12(2):??, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615986.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 12(3):??, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615995.

Anonymous:1985:BMd

[Ano85d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 12(4): ??, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616004.

Anonymous:1985:BICa

[Ano85e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 12(1):89– 96, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615977.

Anonymous:1985:BICb

[Ano85f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 12(2): 183–190, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615984.

Anonymous:1985:BICc

[Ano85g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 12(3): 249–256, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615994.

Anonymous:1985:BICd

[Ano85h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 12(4): 309–316, June 1985. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616003.

Anonymous:1985:FMa

[Ano85i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 12(1): ??, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615969.

Anonymous:1985:FMb

[Ano85j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 12(2):??, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615979.

Anonymous:1985:FMc

[Ano85k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 12(3):??, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615987.

Anonymous:1985:FMd

[Ano85l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 12(4): 308, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615997.

Anonymous:1985:VI

[Ano85m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 12(4): ??, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615996.

Anonymous:1986:BMa

[Ano86a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 13(1):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616012.

Anonymous:1986:BMb

[Ano86b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 13(2):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616020.

Anonymous:1986:BMc

[Ano86c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 13(3):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616030.

Anonymous:1986:BMd

[Ano86d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 13(4):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616042.

Anonymous:1986:BICa

[Ano86e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 13(1):57– 62, September 1986. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616011.

Anonymous:1986:BICb

[Ano86f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 13(2): 153–158, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616019.

Anonymous:1986:BICc

[Ano86g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 13(3): 230–238, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616029.

Anonymous:1986:BICd

[Ano86h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 13(4): 307–312, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616040.

Anonymous:1986:FMa

[Ano86i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 13(1):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616005.

Anonymous:1986:FMb

[Ano86j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 13(2):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616013.

Anonymous:1986:FMc

[Ano86k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 13(3):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616021.

Anonymous:1986:FMd

[Ano86l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 13(4):??, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616032.

Anonymous:1986:VI Journal of Statistics. Theory and Applications, 13(4): ??, September 1986. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616031.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 14(1): ??, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616052.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 14(2):??, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616059.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 14(3):??, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616067.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 14(4):??, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616078.

Anonymous:1987:BICa

[Ano87e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 14(1):87– 96, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616051.

Anonymous:1987:BICb

[Ano87f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 14(3): 234–240, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616066.

Anonymous:1987:BICc lished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 14(4): 353–363, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616077.

Anonymous:1987:FMa

[Ano87h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 14(1): ??, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616043.

Anonymous:1987:FMb

[Ano87i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 14(2):??, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616053.

Anonymous:1987:FMc

[Ano87j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 14(3):??, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616060.

Anonymous:1987:FMd

[Ano87k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 14(4):??, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616069.

Anonymous:1987:VI

[Ano87l] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 14(4): ??, September 1987. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616068.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 15(1): ??, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616087.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 15(2):??, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616096.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 15(3):??, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616105.

Anonymous:1988:BMd

[Ano88d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 15(4):??, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616116.

Anonymous:1988:BICa

[Ano88e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 15(1):69– 80, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616086.

Anonymous:1988:BICb

[Ano88f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 15(3): 215–224, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616104.

Anonymous:1988:BICc

[Ano88g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 15(4): 311–316, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616115.

Anonymous:1988:FMa

[Ano88h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 15(1): ??, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616079.

Anonymous:1988:FMb

[Ano88i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 15(2):??, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616088.

Anonymous:1988:FMc

[Ano88j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 15(3):??, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616097.

Anonymous:1988:FMd

[Ano88k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 15(4):??, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616107.

Anonymous:1988:VI

[Ano88l] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 15(4): ??, December 1988. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616106.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 16(1):??, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616125.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 16(2): ??, June 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616133.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 16(3):??, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616139.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 16(4): ??, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616148.

Anonymous:1989:BICa

[Ano89e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 16(1):89– 96, September 1989. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616124.

Anonymous:1989:BICb

[Ano89f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 16(3): 263–272, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616138.

Anonymous:1989:BICc

[Ano89g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 16(4): 351–361, March 1989. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616146.

Anonymous:1989:FMa

[Ano89h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 16(1):??, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616117.

Anonymous:1989:FMb

[Ano89i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 16(2): ??, June 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616126.

Anonymous:1989:FMc

[Ano89j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 16(3):??, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616134.

Anonymous:1989:FMd

[Ano89k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 16(4): ??, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616141.

Anonymous:1989:VI

[Ano89l] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 16(4): ??, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616140.

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 17(1):??, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616157.

Anonymous:1990:BMb

[Ano90b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 17(2):??, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616165.

Anonymous:1990:BMc

[Ano90c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 17(3): ??, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616175.

Anonymous:1990:BMd

[Ano90d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 17(4): ??, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616183.

Anonymous:1990:BICa

[Ano90e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 17(1):79– 94, December 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616156.

Anonymous:1990:BICb

[Ano90f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 17(2): 161–172, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616164.

Anonymous:1990:BICc

[Ano90g] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 17(3): 263–270, March 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616174.

Anonymous:1990:BICd

[Ano90h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 17(4): 355–366, March 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616182.

Anonymous:1990:FMa

[Ano90i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 17(1):??, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616149.

Anonymous:1990:FMb

[Ano90j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 17(2):??, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616158.

Anonymous:1990:FMc

[Ano90k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 17(3): ??, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616166.

Anonymous:1990:FMd

[Ano90l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 17(4): ??, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616177.

Anonymous:1990:VI Journal of Statistics. Theory and Applications, 17(4): ??, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616176.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 18(1):??, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616192.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 18(2):??, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616200.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 18(3):??, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616210.

Anonymous:1991:BMd

[Ano91d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 18(4): ??, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616219.

Anonymous:1991:BICa

[Ano91e] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 18(1):91– 96, September 1991. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616191.

Anonymous:1991:BICb

[Ano91f] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 18(2): 169–174, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616199.

Anonymous:1991:BICc lished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 18(3): 263–272, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616209.

Anonymous:1991:BICd

[Ano91h] Anonymous. Brief information on current unpublished research in Scandinavia. Scandinavian Journal of Statistics. Theory and Applications, 18(4): 357–363, March 1991. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616218.

Anonymous:1991:FMa

[Ano91i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 18(1):??, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616184.

Anonymous:1991:FMb

[Ano91j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 18(2):??, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616193.

Anonymous:1991:FMc

[Ano91k] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 18(3):??, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616201.

Anonymous:1991:FMd

[Ano91l] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 18(4): ??, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616212.

Anonymous:1991:VI

[Ano91m] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 18 (4):365–366, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616211.

Anonymous:1992:BMa

[Ano92a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 19(1):96, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616229.

Anonymous:1992:BMb

[Ano92b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 19(2):??, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616238.

Anonymous:1992:BMc

[Ano92c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 19(3): 288, September 1992. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616246.

Anonymous:1992:BMd

[Ano92d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 19(4): ??, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616254.

Anonymous:1992:FMa

[Ano92e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 19(1):??, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616220.

Anonymous:1992:FMb

[Ano92f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 19(2):??, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616230.

Anonymous:1992:FMc

[Ano92g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 19(3):??, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616239.

Anonymous:1992:FMd

[Ano92h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 19(4): ??, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616248.

Anonymous:1992:VI

[Ano92i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 19(4): 381, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616247.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 20(1): 96, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616265.

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 20(2): 176, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616273.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 20(3):??, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616282.

Anonymous:1993:BMd

[Ano93d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 20(4):??, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616293.

Anonymous:1993:FMa

[Ano93e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 20(1): ??, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616255.

Anonymous:1993:FMb

[Ano93f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 20(2): ??, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616266.

Anonymous:1993:FMc

[Ano93g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 20(3):??, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616274.

Anonymous:1993:FMd

[Ano93h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 20(4):??, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616284.

Anonymous:1993:VI

[Ano93i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 20 (4):379, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616283.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 21(1): 96, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616302.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 21(2): 192, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616311.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 21(3): 288, September 1994. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616318.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 21(4):??, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616335.

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 21(1): ??, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616294.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 21(2): ??, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616303.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 21(3):??, September 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616312.

Anonymous:1994:FMd

[Ano94h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 21(4):??, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616320.

Anonymous:1994:VI

[Ano94i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 21 (4):505, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616319.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 22(1): 142, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616348.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 22(2): 270, June 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616357.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 398, September 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616369.

Anonymous:1995:BMd

[Ano95d] Anonymous. Back matter.

Scandinavian Journal of Statistics. Theory and Applications, 22(4):524, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616379.

Anonymous:1995:FMa

[Ano95e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 22(1): ??, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616336.

Anonymous:1995:FMb

[Ano95f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 22(2): ??, June 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616349.

Anonymous:1995:FMc

[Ano95g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 22(3):??, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616358.

Anonymous:1995:FMd

[Ano95h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 22(4):??, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616371.

Anonymous:1995:VI

[Ano95i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 22(4): 523–526, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616370.

Anonymous:1996:BMa

[Ano96a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 23(1): 128, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616391.

Anonymous:1996:BMb

[Ano96b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 23(2): 256, June 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616401.

Anonymous:1996:BMc

[Ano96c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 23(3):??, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616412.

Anonymous:1996:BMd

[Ano96d] Anonymous. Back matter.

Scandinavian Journal of Statistics. Theory and Applications, 23(4):638, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616431.

Anonymous:1996:CNS

[Ano96e] Anonymous. Corrigenda to Nurminen and Seeger. Scandinavian Journal of Statistics. Theory and Applications, 23(4):635, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616430. See [NM87, See93].

Anonymous:1996:FMa

[Ano96f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 23(1): ??, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616380.

Anonymous:1996:FMb

[Ano96g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 23(2): ??, June 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616392.

Anonymous:1996:FMc

[Ano96h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 23(3):??, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616402.

Anonymous:1996:FMd

[Ano96i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 23(4):??, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616414.

Anonymous:1996:VI

[Ano96j] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 23(4): 637–640, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616413.

Asmussen:1996:FPT

[ANO96k] Søren Asmussen, Olle Nerman, and Marita Olsson. Fitting phase-type distributions via the EM algorithm. Scandinavian Journal of Statistics. Theory and Applications, 23(4): 419–441, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616418.

Anonymous:1997:ACV

[Ano97a] Anonymous. Acknowledgments and contents of volume 24. Scandinavian Journal of Statistics. Theory and Applications, 24(4): 573–576, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616467.

Anonymous:1997:BMa

[Ano97b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 24(1): 144, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616443.

Anonymous:1997:BMb

[Ano97c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 24(2): 288, June 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616454.

Anonymous:1997:BMc

[Ano97d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 24(3):??, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616466.

Anonymous:1997:BMd

[Ano97e] Anonymous. Back matter.

Scandinavian Journal of Statistics. Theory and Applications, 24(4):574, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616477.

Anonymous:1997:FMa

[Ano97f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 24(1): ??, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616432.

Anonymous:1997:FMb

[Ano97g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 24(2): ??, June 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616444.

Anonymous:1997:FMc

[Ano97h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 24(3):??, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616455.

Anonymous:1997:FMd

[Ano97i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 24(4):??, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616468.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 25(1): 254, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616498.

Anonymous:1998:BMb

[Ano98b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 25(2): ??, June 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616511.

Anonymous:1998:BMc

[Ano98c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 25(3):??, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616522.

Anonymous:1998:BMd

[Ano98d] Anonymous. Back matter.

Scandinavian Journal of Statistics. Theory and Applications, 25(4):730, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616535.

Anonymous:1998:SJS

[Ano98e] Anonymous. Scandinavian Journal of Statistics, Theory and Applications: Index to volume 25, 1998. Scandinavian Journal of Statistics. Theory and Applications, 25(4):729–732, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:1998:FMa

[Ano98f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 25(1): ??, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616478.

Anonymous:1998:FMb

[Ano98g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 25(2): ??, June 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616499.

Anonymous:1998:FMc

[Ano98h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 25(3):??, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616512.

Anonymous:1998:FMd

[Ano98i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 25(4):??, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616524.

Anonymous:1998:VI

[Ano98j] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 25 (4):729, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616523.

Anonymous:1999:BMa

[Ano99a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 26(1): 160, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616548.

Anonymous:1999:BMb

[Ano99b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 26(2): ??, June 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616558.

Anonymous:1999:BMc

[Ano99c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 478, September 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616571.

Anonymous:1999:BMd

[Ano99d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 26(4): 635–636, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616584.

Anonymous:1999:FMa

[Ano99e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 26(1): ??, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616536.

Anonymous:1999:FMb

[Ano99f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 26(2): ??, June 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616549.

Anonymous:1999:FMc

[Ano99g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 26(3):??, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616559.

Anonymous:1999:FMd

[Ano99h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 26(4):??, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616573.

Anonymous:1999:VI

[Ano99i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 26 (4):634, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616572.

Anonymous:2000:BMa

[Ano00a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 27(1): ??, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616598.

Anonymous:2000:BMb

[Ano00b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 27(2): ??, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616610.

Anonymous:2000:BMc

[Ano00c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 27(3):??, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616625.

Anonymous:2000:BMd

[Ano00d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 27(4):766, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616640.

Anonymous:2000:FMa

[Ano00e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 27(1): ??, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616585.

Anonymous:2000:FMb

[Ano00f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 27(2): ??, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616599.

Anonymous:2000:FMc

[Ano00g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 27(3):??, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616611.

Anonymous:2000:FMd

[Ano00h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 27(4):??, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616627.

Anonymous:2000:IV

[Ano00i] Anonymous. Index to volume 27, 2000. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 765–768, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2000:VI

[Ano00j] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 27 (4):765, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616626.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 28(1): ??, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616655.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 28(2): ??, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616667.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 28(3):??, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616680.

Anonymous:2001:BMd

[Ano01d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 28(4):??, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616693.

Anonymous:2001:FMa

[Ano01e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 28(1): ??, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616641.

Anonymous:2001:FMb

[Ano01f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 28(2): ??, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616656.

Anonymous:2001:FMc

[Ano01g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 28(3):??, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616668.

Anonymous:2001:FMd

[Ano01h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 28(4):??, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616682.

Anonymous:2001:IV

[Ano01i] Anonymous. Index to volume 28, 2001. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 733–736, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2001:VI

[Ano01j] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 28(4): ??, December 2001. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616681.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 29(1): 188, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616709.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 29(2): ??, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616718.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 29(3):??, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616733.

Anonymous:2002:BMd

[Ano02d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 29(4):??, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616744.

Anonymous:2002:FMa

[Ano02e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 29(1): ??, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616694.

Anonymous:2002:FMb

[Ano02f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 29(2): ??, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616710.

Anonymous:2002:FMc

[Ano02g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 29(3):??, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616719.

Anonymous:2002:FMd

[Ano02h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 29(4):??, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616735.

Anonymous:2002:VI

[Ano02i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 29(4): ??, December 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616734.

Anonymous:2003:BMa

[Ano03a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 30(1): ??, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616761.

Anonymous:2003:BMb

[Ano03b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 30(2): ??, June 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616774.

Anonymous:2003:BMc

[Ano03c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 30(3):??, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616791.

Anonymous:2003:BMd

[Ano03d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 30(4):??, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616805.

Anonymous:2003:FMa

[Ano03e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 30(1): ??, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616745.

Anonymous:2003:FMb

[Ano03f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 30(2): ??, June 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616762.

Anonymous:2003:FMc

[Ano03g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 30(3):??, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616775.

Anonymous:2003:FMd

[Ano03h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 30(4):??, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616793.

Anonymous:2003:VI

[Ano03i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 30(4): ??, December 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616792.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 31(1): ??, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616818.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 31(2): ??, June 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616831.

Anonymous:2004:BMc

[Ano04c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 31(3):??, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616844.

Anonymous:2004:BMd

[Ano04d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 31(4):??, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616856.

Anonymous:2004:FMa

[Ano04e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 31(1): ??, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616806.

Anonymous:2004:FMb

[Ano04f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 31(2): ??, June 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616819.

Anonymous:2004:FMc

[Ano04g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 31(3):??, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616832.

Anonymous:2004:FMd

[Ano04h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 31(4):??, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616846.

Anonymous:2004:VI

[Ano04i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 31(4): ??, December 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616845.

Anonymous:2005:A

[Ano05a] Anonymous. Acknowledgements. Scandinavian Journal of Statistics. Theory and Applications, 32(4):i, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2005:BMa

[Ano05b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 32(1): ??, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616866.

Anonymous:2005:BMb

[Ano05c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 32(2): ??, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616882.

Anonymous:2005:BMc

[Ano05d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 32(3):??, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616892.

Anonymous:2005:BMd

[Ano05e] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 32(4):??, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616903.

Anonymous:2005:FMa

[Ano05f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 32(1): ??, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616857.

Anonymous:2005:FMb

[Ano05g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 32(2): ??, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616867.

Anonymous:2005:FMc

[Ano05h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 32(3):??, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616883.

Anonymous:2005:FMd

[Ano05i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 32(4):??, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616894.

Anonymous:2005:VC

[Ano05j] Anonymous. Volume 32 contents. Scandinavian Journal of Statistics. Theory and Applications, 32 (4):ii–iii, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2005:VI

[Ano05k] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 32(4): ??, December 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616893.

Anonymous:2006:BMa

[Ano06a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 33(1): ??, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616915.

Anonymous:2006:BMb

[Ano06b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 33(2): ??, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616932.

Anonymous:2006:BMc

[Ano06c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 33(3):??, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616945.

Anonymous:2006:BMd

[Ano06d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 33(4):894, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616964.

Anonymous:2006:FMa

[Ano06e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 33(1): ??, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616904.

Anonymous:2006:FMb

[Ano06f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 33(2): ??, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616916.

Anonymous:2006:FMc

[Ano06g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 33(3):??, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616933.

Anonymous:2006:FMd

[Ano06h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 33(4):??, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616947.

Anonymous:2006:VI

[Ano06i] Anonymous. Volume information. Scandinavian Journal of Statistics. Theory and Applications, 33(4): ??, December 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616946.

Anonymous:2007:A

[Ano07a] Anonymous. Acknowledgements. Scandinavian Journal of Statistics. Theory and Applications, 34(4):v– vi, December 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2007:BMa

[Ano07b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 34(1): ??, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548549.

Anonymous:2007:BMb

[Ano07c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 34(2): ??, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548562.

Anonymous:2007:BMc

[Ano07d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 34(3):??, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548574.

Anonymous:2007:BMd

[Ano07e] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 34(4):??, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548586.

Anonymous:2007:DMS

[Ano07f] Anonymous. Discussion of ‘Modern Statistics for Spatial Point Processes’. Scandinavian Journal of Statistics. Theory and Applications, 34(4):685–711, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

41548577. See [MW07].

Anonymous:2007:FMa

[Ano07g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 34(1): ??, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548533.

Anonymous:2007:FMb

[Ano07h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 34(2):

??, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548550.

Anonymous:2007:FMc

[Ano07i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 34(3):??, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548563.

Anonymous:2007:FMd

[Ano07j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 34(4):??, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548575.

Anonymous:2007:VC

[Ano07k] Anonymous. Volume 34 contents. Scandinavian Journal of Statistics. Theory and Applications, 34(4): vii, December 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2008:BMa

[Ano08a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 35(1): ??, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000393.

Anonymous:2008:BMb

[Ano08b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 35(2): ??, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548600.

Anonymous:2008:BMc

[Ano08c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 35(3):??, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000288.

Anonymous:2008:BMd

[Ano08d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 35(4):??, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000302.

Anonymous:2008:FMa

[Ano08e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 35(1): ??, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000381.

Anonymous:2008:FMb

[Ano08f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 35(2): ??, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548587.

Anonymous:2008:FMc

[Ano08g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 35(3):??, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000277.

Anonymous:2008:FMd

[Ano08h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 35(4):??, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000289.

Anonymous:2009:BMa

[Ano09a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 36(1): ??, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000315.

Anonymous:2009:BMb

[Ano09b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 36(2): ??, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000326.

Anonymous:2009:BMc

[Ano09c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 36(3):??, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000339.

Anonymous:2009:BMd

[Ano09d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 36(4):??, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000356.

Anonymous:2009:FMa

[Ano09e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 36(1): ??, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000303.

Anonymous:2009:FMb

[Ano09f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 36(2): ??, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000316.

Anonymous:2009:FMc

[Ano09g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 36(3):??, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000327.

Anonymous:2009:FMd

[Ano09h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 36(4):??, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000340.

Anonymous:2010:A

[Ano10a] Anonymous. Acknowledgements. Scandinavian Journal of Statistics. Theory and Applications, 37(4):iii– iv, December 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2010:BMa

[Ano10b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 37(1): ??, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000922.

Anonymous:2010:BMb

[Ano10c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 37(2): ??, June 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000884.

Anonymous:2010:BMc

[Ano10d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 37(3):??, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000405.

Anonymous:2010:BMd

[Ano10e] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 37(4):??, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000417.

Anonymous:2010:C

[Ano10f]  Anonymous.  Corrigen-

dum. Scandinavian Journal of Statistics. Theory and Applications, 37(1): 176, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [PDD10].

Anonymous:2010:FMa

[Ano10g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 37(1): ??, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000911.

Anonymous:2010:FMb

[Ano10h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 37(2): ??, June 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000873.

Anonymous:2010:FMc

[Ano10i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 37(3):??, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000394.

Anonymous:2010:FMd

[Ano10j] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 37(4):??, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000406.

Anonymous:2010:VC

[Ano10k] Anonymous. Volume contents. Scandinavian Journal of Statistics. Theory and Applications, 37(4):i–ii, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2011:A

[Ano11a] Anonymous. Acknowledgements. Scandinavian Journal of Statistics. Theory and Applications, 38(4):i–ii, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2011:BMa

[Ano11b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 38(1): ??, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015390.

Anonymous:2011:BMb

[Ano11c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 38(2): ??, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015501.

Anonymous:2011:BMc

[Ano11d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 38(3):??, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015584.

Anonymous:2011:BMd

[Ano11e] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 38(4):??, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411152.

Anonymous:2011:FMa

[Ano11f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 38(1): ??, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015380.

Anonymous:2011:FMb

[Ano11g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 38(2): ??, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015487.

Anonymous:2011:FMc

[Ano11h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 38(3):??, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015570.

Anonymous:2011:FMd

[Ano11i] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 38(4):??, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41411140.

Anonymous:2011:VC

[Ano11j] Anonymous. Volume 38 contents. Scandinavian Journal of Statistics. Theory and Applications, 38 (4):iii–iv, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Anonymous:2012:BMa

[Ano12a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 39(1): ??, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411164.

Anonymous:2012:BMb

[Ano12b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 39(2): ??, June 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679801.

Anonymous:2012:BMc

[Ano12c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 39(3):??, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679814.

Anonymous:2012:BMd

[Ano12d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 39(4):??, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357229.

Anonymous:2012:FMa

[Ano12e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 39(1):

??, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411153.

Anonymous:2012:FMb

[Ano12f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 39(2): ??, June 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679786.

Anonymous:2012:FMc

[Ano12g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 39(3):??, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679802.

Anonymous:2012:FMd

[Ano12h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 39(4):??, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357212.

Anonymous:2013:BMa

[Ano13a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 40(1): ??, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357260.

Anonymous:2013:BMb

[Ano13b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 40(2): ??, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586718.

Anonymous:2013:BMc

[Ano13c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 40(3):??, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586690.

Anonymous:2013:BMd

[Ano13d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 40(4):??, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586706.

Anonymous:2013:FMa

[Ano13e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 40(1): ??, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23357248.

Anonymous:2013:FMb

[Ano13f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 40(2): ??, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586707.

Anonymous:2013:FMc

[Ano13g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 40(3):??, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586676.

Anonymous:2013:FMd

[Ano13h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 40(4):??, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586691.

Anonymous:2014:BMa

[Ano14a] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 41(1): ??, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586742.

Anonymous:2014:BMb

[Ano14b] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 41(2): ??, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586668.

Anonymous:2014:BMc

[Ano14c] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 41(3):??, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586762.

Anonymous:2014:BMd

[Ano14d] Anonymous. Back matter. Scandinavian Journal of Statistics. Theory and Applications, 41(4):??, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586829.

Anonymous:2014:FMa

[Ano14e] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 41(1): ??, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586719.

Anonymous:2014:FMb

[Ano14f] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 41(2): ??, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586651.

Anonymous:2014:FMc

[Ano14g] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 41(3):??, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586746.

Anonymous:2014:FMd

[Ano14h] Anonymous. Front matter. Scandinavian Journal of Statistics. Theory and Applications, 41(4):??, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586809.

Anonymous:2023:EFC

[Ano23] Anonymous. Errata for “A framework for covariate balance using Bregman distances”. Scandinavian Journal of Statistics. Theory and Applications, 50 (3):1552, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JJCYG21].

Anonymous:2024:CGD

[Ano24] Anonymous. Corrigendum:

Greenland, S. (2023). Divergence vs. decision P - values: A distinction worth making in theory and keeping in practice. Scandinavian Journal of Statistics, 50, 1–35. Scandinavian Journal of Statistics. Theory and Applications, 51(1): 425, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Gre23b].

Antoniadis:1996:SND

[Ant96] Anestis Antoniadis. Smoothing noisy data with tapered coiflets series. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 313–330, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616407.

Andrade:2011:BRM

[AO11] Jose Ailton Alencar Andrade and Anthony O’Hagan. Bayesian robustness modelling of location and scale parameters. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 691–711, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41411145.

Arvastson:2000:ABP

[AOH00] L. Arvastson, H. Olsson, and J. Holst. Asymptotic bias in parameter estimation of AR-processes using recursive least squares with exponential forgetting. Scandinavian Journal of Statistics. Theory and Applications, 27(1):177–192, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616597.

Arjas:2004:CRL

[AP04] Elja Arjas and Jan Parner.

Causal reasoning from longitudinal data*. Scandinavian Journal of Statistics. Theory and Applications, 31(2):171–187, June 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616822. See dis-

cussion [Lau04, Aal04] and replies [Rub04b, Arj04]..

Asci:2007:FCL

[AP07] Claudio Asci and Mauro Piccioni. Functionally compatible local characteristics for the local specification of priors in graphical models. Scandinavian Journal of Statistics. Theory and Applications, 34(4): 829–840, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548583.

Alonso-Pena:2024:NED

[APCG24] María Alonso-Pena, Gerda Claeskens, and Irène Gijbels. Nonparametric estimation of densities on the hypersphere using a parametric guide. Scandinavian Journal of Statistics. Theory and Applications, 51(3): 956–986, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Arfe:2019:RUS

[APM19] Andrea Arfé, Stefano Peluso, and Pietro Muliere. Reinforced urns and the subdistribution beta-Stacy process prior for competing risks analysis. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 706–734, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Akritas:1980:AIC

[AR80] Michael G. Akritas and George G. Roussas. Asymptotic inference in continuous time semi-Markov processes. Scandinavian Journal of Statistics. Theory and Applications, 7(2):73– 79, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615776.

Andhivarothai:1994:PSL

[AR94] N. Andhivarothai and R. V. Ramamoorthi. Partial sufficiency and least favorable experiments. Scandinavian Journal of Statistics. Theory and Applications, 21(4):

493–499, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616333.

Arcones:1998:EEP

[Arc98] Miguel A. Arcones. L_{p}- estimators as estimates of a parameter of location for a sharp–pointed symmetric density. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 693–715, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616533.

Arfi:1998:NPV

[Arf98] Mounir Arfi. Non-parametric variance estimation from ergodic samples. Scandinavian Journal of Statistics. Theory and Applications, 25(1):225–234, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616495.

Arjas:1992:E

[Arj92]  Elja Arjas.  Editorial.

Scandinavian Journal of Statistics. Theory and Applications, 19(1):1, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616221.

Arjas:1994:E

[Arj94]  Elja Arjas.  Editorial.

Scandinavian Journal of Statistics. Theory and Applications, 21(4):i, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616321.

Arjas:2002:D

[Arj02] Elja Arjas. Discussion. Scandinavian Journal of Statistics. Theory and Applications, 29(2):268–269, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [GIA02].

Arjas:2004:RD

[Arj04] Elja Arjas. Reply to discussion. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 198–201, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616823. See [AP04].

Arjas:2011:FDS

[Arj11] Elja Arjas. On future directions in statistical methodologies — some speculations. Scandinavian Journal of Statistics. Theory and Applications, 38(2): 185–194, June 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015488.

Azriel:2023:ODD

[ARP23] David Azriel, Yosef Rinott, and Martin Posch. Optimal designs for the development of personalized treatment rules. Scandinavian Journal of Statistics. Theory and Applications, 50(2):797–824, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Alho:2010:E

[AS10] Juha Alho and Paavo Salminen. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 37(1):1, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000912.

Asmussen:1989:EFG

[Asm89] Søren Asmussen. Exponential families generated by phase-type distributions and other Markov lifetimes. Scandinavian Journal of Statistics. Theory and Applications, 16(4):319–334, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616144.

Asmussen:2000:MAM

[Asm00] Søren Asmussen. Matrixanalytic models and their analysis. Scandinavian Journal of Statistics. Theory and Applications, 27(2): 193–226, June 2000. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616600.

Arnesen:2015:FBB

[AT15] Petter Arnesen and Håkon Tjelmeland. Fully Bayesian binary Markov random field models: Prior specification and posterior simulation. Scandinavian Journal of Statistics. Theory and Applications, 42(4):967–987, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Amiri:2017:EDD

[ATV17] Aboubacar Amiri, Baba Thiam, and Thomas Verdebout. On the estimation of the density of a directional data stream. Scandinavian Journal of Statistics. Theory and Applications, 44 (1):249–267, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Augustin:2004:ECL

[Aug04] Thomas Augustin. An exact corrected log-likelihood function for Cox’s proportional hazards model under measurement error and some extensions. Scandinavian Journal of Statistics. Theory and Applications, 31(1):43–50, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616810.

Akritas:2001:NPE

[AV01] Michael G. Akritas and Ingrid Van Keilegom. Nonparametric estimation of the residual distribution. Scandinavian Journal of Statistics. Theory and Applications, 28(3):549–567, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616678.

Avagyan:2021:SIP

[AV21] Vahe Avagyan and Stijn Vansteelandt. Stable inverse probability weighting estimation for longitudinal studies. Scandinavian Journal of Statistics. Theory and Applications, 48 (3):1046–1067, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Arellano-Valle:2006:UFS

[AVA06] Reinaldo B. Arellano-Valle and Adelchi Azzalini. On the unification of families of skew-normal distributions. Scandinavian Journal of Statistics. Theory and Applications, 33(3):561–574, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616942. See corrigendum [AVA22].

Arellano-Valle:2022:CUF

[AVA22] Reinaldo B. Arellano-Valle and Adelchi Azzalini. Corrigendum to “On the unification of families of skewnormal distributions” published in Scand. J. Stat. vol. 33, pp. 561–574. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1418–1419, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [AVA06].

Arellano-Valle:2013:SEM

[AVCRG13] Reinaldo B. Arellano-Valle, Javier E. Contreras-Reyes, and Marc G. Genton. Shannon entropy and mutual information for multivariate skew-elliptical distributions. Scandinavian Journal of Statistics. Theory and Applications, 40(1):42– 62, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357252.

Aven:1985:TDC

[Ave85] Terje Aven. A theorem for determining the compensator of a counting process. Scandinavian Journal of Statistics. Theory and Applications, 12(1):69– 72, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615974.

Aven:1986:BIP

[Ave86] Terje Aven. Bayesian inference in a parametric counting process model. Scandinavian Journal of Statistics. Theory and Applications, 13(2):87–97, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616015.

Alam:1979:DSO

[AW79] Khursheed Alam and K. T.

Wallenius. Distribution of a sum of order statistics. Scandinavian Journal of Statistics. Theory and Applications, 6(3):123–126, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615746.

Awad:1981:ANP

[Awa81] Adnan M. Awad. Asymptotic normality of the posterior distribution of an average of parameters. Scandinavian Journal of Statistics. Theory and Applications, 8(4):247– 249, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615844.

Azzalini:1985:CDW

[Azz85]  A. Azzalini.  A class

of distributions which includes the normal ones. Scandinavian Journal of Statistics. Theory and Applications, 12(2):171–178, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615982.

Azzalini:2005:DSN

[Azz05a] Adelchi Azzalini. [discussion of “the skew-normal”]: Rejoinder. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 199–200, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616870. See [Azz05b].

Azzalini:2005:SND

[Azz05b] Adelchi Azzalini. The skewnormal distribution and related multivariate families*. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 159–188, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616868. See discussion [Gen05] and reply [Azz05a].

Bachmaier:2011:TEM

[Bac11] Martin Bachmaier. Testing for equivalence of means under heteroskedasticity by an approximate solution of a partial differential equation of infinite order. Scandinavian Journal of Statistics. Theory and Applications, 38(1):147–168, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015388.

Banerjee:2005:LRT

[Ban05] Moulinath Banerjee. Likelihood ratio tests under local and fixed alternatives in monotone function problems. Scandinavian Journal of Statistics. Theory and Applications, 32(4):

507–525, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616895.

Barnard:1976:CII

[Bar76] G. A. Barnard. Conditional inference is not inefficient. Scandinavian Journal of Statistics. Theory and Applications, 3(3):132– 134, September 1976. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615626.

Bergman:1985:RTA

[BAR^{+}85] Bo Bergman, Elja Arjas, Marvin Rausand, Bent Natvig, Kjell A. Doksum, and Tore Schweder. On reliability theory and its applications [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 12(1): 1–41, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615970.

Baringhaus:2003:UEP

[Bar03] Ludwig Baringhaus. On unbiased estimation of positive integral powers of the natural parameter in exponential families. Scandinavian Journal of Statistics. Theory and Applications, 30(3):597–608, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616786.

Bardet:2010:NPE

[BB10] Jean-Marc Bardet and Pierre R. Bertrand. A non-parametric estimator of the spectral density of a continuous-time Gaussian process observed at random times. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 458–476, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000401.

Beutner:2011:EBD

[BB11] Eric Beutner and Laurent Bordes. Estimators based on data-driven generalized weighted Cramér– von Mises distances under censoring — with applications to mixture models. Scandinavian Journal of Statistics. Theory and Applications, 38(1): 108–129, March 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015386.

Butler:2014:MSM

[BB14] Ronald W. Butler and Douglas A. Bronson. Multistate survival models as transient electrical networks. Scandinavian Journal of Statistics. Theory and Applications, 41(1):167–186, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586736.

Belouni:2015:ORD

[BB15] Mohamad Belouni and Karim Benhenni. Optimal and robust designs for estimating the concentration curve and the AUC. Scandinavian Journal of Statistics. Theory and Applications, 42(2):453–470, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bauder:2019:BEE

[BBBS19] David Bauder, Rostyslav Bodnar, Taras Bodnar, and Wolfgang Schmid. Bayesian estimation of the efficient frontier. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 802–830, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Betken:2021:OPL

[BBD^{+}21] Annika Betken, Jannis Buchsteiner, Herold Dehling, Ines Münker, Alexander Schnurr, and Jeannette H. C. Woerner. Ordinal patterns in long-range dependent time series. Scandinavian Journal of Statistics. Theory and Applications, 48(3):969–1000, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Buitendag:2020:CIE

[BBdW20] Sven Buitendag, Jan Beirlant, and Tertius de Wet. Confidence intervals for extreme Pareto-type quantiles. Scandinavian Journal of Statistics. Theory and Applications, 47(1):36– 55, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bentkus:1997:BEB

[BBG97] V. Bentkus, M. Bloznelis, and F. Götze. A Berry– Esséen bound for Mestimators. Scandinavian Journal of Statistics. Theory and Applications, 24(4): 485–502, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616471.

Banerjee:2006:SBR

[BBG06] Moulinath Banerjee, Pinaki Biswas, and Debashis Ghosh. A semiparametric binary regression model involving monotonicity constraints. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 673–697, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616952.

Benth:2007:PPT

[BBK07] Fred Espen Benth, Jūratė Šaltytė Benth, and Steen Koekebakker. Putting a price on temperature*. Scandinavian Journal of Statistics. Theory and Applications, 34(4):746–767, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548579.

Bie:1987:CIC

[BBL87] Ole Bie, Ørnulf Borgan, and Knut Liestøl. Confidence intervals and confidence bands for the cumulative hazard rate function and their small sample properties. Scandinavian Journal of Statistics. Theory and Applications, 14(3): 221–233, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616065.

Bianco:2006:RTS

[BBM06] Ana Bianco, Graciela Boente, and Elena Martínez. Robust tests in semiparametric partly linear models. Scandinavian Journal of Statistics. Theory and Applications, 33(3): 435–450, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616935.

Bissantz:2021:NDC

[BBP21] Kathrin Bissantz, Nicolai Bissantz, and Katharina Proksch. Nonparametric detection of changes over time in image data from fluorescence microscopy of living cells. Scandinavian Journal of Statistics. Theory and Applications, 48 (3):1001–1017, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bahraoui:2018:FGF

[BBQ18] Tarik Bahraoui, Taoufik Bouezmarni, and Jean-François Quessy. A family of goodness-of-fit tests for copulas based on characteristic functions. Scandinavian Journal of Statistics. Theory and Applications, 45(2):301–323, June 2018.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bodnar:2023:SMH

[BBS23] Rostyslav Bodnar, Taras Bodnar, and Wolfgang Schmid. Sequential monitoring of high-dimensional time series. Scandinavian Journal of Statistics. Theory and Applications, 50(3): 962–992, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Brown:1999:BBS

[BC99] Bruce M. Brown and Song Xi Chen. Beta-Bernstein smoothing for regression curves with compact support. Scandinavian Journal of Statistics. Theory and Applications, 26(1):47–59, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616540.

Boruvka:2015:CAM

[BC15] Audrey Boruvka and Richard J. Cook. A Coxaalen model for intervalcensored data. Scandinavian Journal of Statistics. Theory and Applications, 42(2):414–426, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ba:2023:ILH

[BC23] Ismaïla Ba and Jean-François Coeurjolly. Inference for low- and highdimensional inhomogeneous Gibbs point processes. Scandinavian Journal of Statistics. Theory and Applications, 50(3):993–1021, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bertail:2017:EPS

[BCC17] Patrice Bertail, Emilie Chautru, and Stephan Clémençon. Empirical processes in survey sampling with (conditional) Poisson designs. Scandinavian Journal of Statistics. Theory and Applications, 44 (1):97–111, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Baillo:2011:SCF

[BCCA11] Amparo Baíllo, Antonio Cuevas, and Juan Antonio Cuesta-Albertos. Supervised classification for a family of Gaussian functional models. Scandinavian Journal of Statistics. Theory and Applications, 38(3):480–498, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015576.

Bevilacqua:2021:NGG

[BCCAVMO21] Moreno Bevilacqua, Christian Caamaño-Carrillo, Reinaldo B. Arellano-Valle, and Víctor Morales-Oñate. Non-Gaussian geostatistical modeling using (skew) t processes. Scandinavian Journal of Statistics. Theory and Applications, 48(1): 212–245, March 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bai:2019:GQR

[BCCH19] Fangfang Bai, Xuerong Chen, Yan Chen, and Tao Huang. A general quantile residual life model for length-biased rightcensored data. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1191–1205, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Brunel:2008:ESC

[BCG08] Elodie Brunel, Fabienne Comte, and Agathe Guilloux. Estimation strategies for censored lifetimes with a Lexis-diagram type model. Scandinavian Journal of Statistics. Theory and Applications, 35(3): 557–576, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000287.

Boistard:2016:DRI

[BCH16] Helene Boistard, Guillaume Chauvet, and David Haziza. Doubly robust inference for the distribution function in the presence of missing survey data. Scandinavian Journal of Statistics. Theory and Applications, 43(3):683–699, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Besse:2000:AFS

[BCS00] Philippe C. Besse, Hervé Cardot, and David B. Stephenson. Autoregressive forecasting of some functional climatic variations. Scandinavian Journal of Statistics. Theory and Applications, 27(4):673–687, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616634.

Baddeley:2013:LID

[BCS13] Adrian Baddeley, Ya-Mei Chang, and Yong Song. Leverage and influence diagnostics for spatial point processes. Scandinavian Journal of Statistics. Theory and Applications, 40 (1):86–104, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357254.

Birke:2007:ECF

[BD07] Melanie Birke and Holger Dette. Estimating a convex function in nonparametric regression. Scandinavian Journal of Statistics. Theory and Applications, 34(2):384–404, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548558.

Bibbona:2013:EDO

[BD13] Enrico Bibbona and Susanne Ditlevsen. Estimation in discretely observed diffusions killed at a threshold. Scandinavian Journal of Statistics. Theory and Applications, 40(2):274–293, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586713.

Bagchi:2020:SLS

[BD20] Pramita Bagchi and Subhra Sankar Dhar. A study on the least squares estimator of multivariate isotonic regression function. Scandinavian Journal of Statistics. Theory and Applications, 47(4):

1192–1221, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Brown:2003:NGS

[BDH03] Patrick E. Brown, Peter J.

Diggle, and Robin Henderson. A non-Gaussian spatial process model for opacity of flocculated paper. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 355–368, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616768.

Brault:2017:ENB

[BDL^{+}17] Vincent Brault, Maud Delattre, Emilie Lebarbier, Tristan Mary-Huard, and Céline Lévy-Leduc. Estimating the number of block boundaries from diagonal blockwise matrices without penalization. Scandinavian Journal of Statistics. Theory and Applications, 44(2):563–580, June 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bissantz:2012:MCI

[BDP12] Nicolai Bissantz, Holger Dette, and Katharina Proksch. Model checks in inverse regression models with convolution-type operators. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 305–322, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679796.

Beyersmann:2013:WCW

[BDP13] Jan Beyersmann, Susanna Di Termini, and Markus Pauly. Weak convergence of the wild bootstrap for the Aalen–Johansen estimator of the cumulative incidence function of a competing risk. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 387–402, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586677.

Benth:2022:CMS

[BDS22] Fred Espen Benth, Giulia Di Nunno, and Dennis Schroers. Copula measures and Sklar’s theorem in arbitrary dimensions. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1144–1183, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bordes:2006:SET

[BDV06] Laurent Bordes, Céline Delmas, and Pierre Vandekerkhove. Semiparametric estimation of a twocomponent mixture model where one component is known. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 733–752, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616955.

Becheri:2016:AIJ

[BDW16] I. Gaia Becheri, Feike C.

Drost, and Bas J. M. Werker. Asymptotic inference for jump diffusions with state-dependent intensity. Scandinavian Journal of Statistics. Theory and Applications, 43(2): 520–542, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bjerve:1985:UCB

[BDY85] Steinar Bjerve, Kjell A.

Doksum, and Brian S. Yandell. Uniform confidence bounds for regression based on a simple moving average. Scandinavian Journal of Statistics. Theory and Applications, 12(2): 159–169, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615981.

Beirlant:2010:AHI

[BE10] Jan Beirlant and John H. J. Einmahl. Asymptotics for the Hirsch index. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 355–364, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000395.

Bednarski:1993:REC

[Bed93] Tadeusz Bednarski. Robust estimation in Cox’s regression model. Scandinavian Journal of Statistics. Theory and Applications, 20(3): 213–225, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616277.

Basu:1983:MHN

[BEK83] A. P. Basu, Nader Ebrahimi, and Bengt Klefsjö. Multivariate harmonic new better than used in expectation distributions. Scandinavian Journal of Statistics. Theory and Applications, 10(1):19–25, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615896.

Bellio:2003:LMC

[Bel03] Ruggero Bellio. Likelihood methods for controlled calibration. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 339–353, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616767.

Burke:2020:SMR

[BEP20] Kevin Burke, Frank Eriksson, and C. B. Pipper. Semiparametric multiparameter regression survival modeling. Scandinavian Journal of Statistics. Theory and Applications, 47 (2):555–571, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Berg:1974:FSD

[Ber74] Sven Berg. Factorial series distributions, with applications to capture–recapture problems. Scandinavian Journal of Statistics. Theory and Applications, 1 (3):145–152, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615566.

Berg:1975:SPA

[Ber75] Sven Berg. Some properties and applications of a ratio of Stirling numbers of the second kind. Scandinavian Journal of Statistics. Theory and Applications, 2(2):91–94, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615582.

Berg:1976:NUE

[Ber76] Sven Berg. A note on the UMVU estimate in a multiple-recapture census. Scandinavian Journal of Statistics. Theory and Applications, 3(2):86– 88, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615615.

Berg:1977:CPM

[Ber77a] Sven Berg. Certain properties of the multivariate factorial series distributions. Scandinavian Journal of Statistics. Theory and Applications, 4(1):25– 30, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615644.

Bergman:1977:CTT

[Ber77b] Bo Bergman. Crossings in the total time on test plot. Scandinavian Journal of Statistics. Theory and Applications, 4(4):171–177, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615674.

Bergman:1979:ART

[Ber79a]  Bo Bergman.  On age

replacement and the total time on test concept. Scandinavian Journal of Statistics. Theory and Applications, 6(4):161–168, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615755.

Berry:1979:DTA

[Ber79b] Donald A. Berry. Detecting trends in arrangements of ordered objects: a likelihood approach. Scandinavian Journal of Statistics. Theory and Applications, 6 (4):169–174, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615756.

Bergman:1981:TAC

[Ber81] Bo Bergman. Tests against certain ordered multinomial alternatives. Scandinavian Journal of Statistics. Theory and Applications, 8(4):218–226, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615840.

Berk:1982:AOS

[Ber82] Robert H. Berk. On an asymptotically optimal sequential test. Scandinavian Journal of Statistics. Theory and Applications, 9(3): 159–163, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615872.

Berod:1994:CBC

[Bér94] Annick Clerc Bérod. Conditional behavior of confidence intervals. Scandinavian Journal of Statistics. Theory and Applications, 21(2):159–167, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616307.

Berger:2016:ELI

[Ber16] Yves G. Berger. Empirical likelihood inference for the Rao–Hartley–Cochran sampling design. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 721–735, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Berger:2023:UEL

[Ber23]  Yves G. Berger.  Un-

conditional empirical likelihood approach for analytic use of public survey data. Scandinavian Journal of Statistics. Theory and Applications, 50 (1):383–410, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bee:2002:PDN

[BF02] Marco Bee and Bernard Flury. A problem of dimensionality in normal mixture analysis. Scandinavian Journal of Statistics. Theory and Applications, 29(3): 485–500, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616728.

Bondesson:2003:MVV

[BF03] Lennart Bondesson and Jessica Fahlén. Mean and variance of vacancy for hard-core disc processes and applications. Scandinavian Journal of Statistics. Theory and Applications, 30(4):797–816, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616803.

Berger:1980:MSS

[BG80] Roger L. Berger and Shanti S. Gupta. Minimax subset selection rules with applications to unequal variance (unequal sample size) problems. Scandinavian Journal of Statistics. Theory and Applications, 7 (1):21–26, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615766.

Beran:1998:RCE

[BG98] Jan Beran and Sucharita Ghosh. Root-n-consistent estimation in partial linear models with long-memory errors. Scandinavian Journal of Statistics. Theory and Applications, 25(2): 345–357, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616506.

Billiot:2001:EMP

[BG01] Jean-Michel Billiot and Michel Goulard. An estimation method of the pair potential function for Gibbs point processes on spheres. Scandinavian Journal of Statistics. Theory and Applications, 28(1):185–203, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616651.

Bondesson:2011:ESM

[BG11] Lennart Bondesson and Anton Grafström. An extension of Sampford’s method for unequal probability sampling. Scandinavian Journal of Statistics. Theory and Applications, 38(2):377–392, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015500.

Byrne:2013:GMC

[BG13] Simon Byrne and Mark Girolami. Geodesic Monte Carlo on embedded manifolds. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 825–845, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586702.  See

discussion [BG14b, Dry14, Ken14, SLS14, Sim14] and rejoinder [BG14c].

Bortot:2014:LPM

[BG14a] Paola Bortot and Carlo Gaetan. A latent process model for temporal extremes. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 606–621, September 2014.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586749.

Byrne:2014:DAG

[BG14b] Simon Byrne and Mark Girolami. Discussion of the article ‘Geodesic Monte Carlo on Embedded Manifolds’ by Simon Byrne and Mark Girolami. Scandinavian Journal of Statistics. Theory and Applications, 41(1):1–2, March 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

24586720. See [BG13].

Byrne:2014:RGM

[BG14c] Simon Byrne and Mark Girolami. Rejoinder: Geodesic Monte Carlo on embedded manifolds. Scandinavian Journal of Statistics. Theory and Applications, 41(1):19–21, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

24586727. See [BG13].

Bellio:2016:ILI

[BG16] Ruggero Bellio and Annamaria Guolo. Integrated likelihood inference in small sample metaanalysis for continuous outcomes. Scandinavian Journal of Statistics. Theory and Applications, 43(1): 191–201, March 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Borgan:2024:CNL

[BG24a] Ørnulf Borgan and Ingrid K. Glad. A conversation with Nils Lid Hjort. Scandinavian Journal of Statistics. Theory and Applications, 51(3):914–935, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Borgan:2024:SSH

[BG24b] Ørnulf Borgan and Ingrid K. Glad. A special section honoring Nils Lid Hjort. Scandinavian Journal of Statistics. Theory and Applications, 51 (3):913, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Balabdaoui:2019:SEM

[BGH19] Fadoua Balabdaoui, Piet Groeneboom, and Kim Hendrickx. Score estimation in the monotone singleindex model. Scandinavian Journal of Statistics. Theory and Applications, 46(2):517–544, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Branco:2013:OBA

[BGL13] Marcia D’Elia Branco, Marc G. Genton, and Brunero Liseo. Objective Bayesian analysis of skewt distributions. Scandinavian Journal of Statistics. Theory and Applications, 40(1):63–85, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357253.

Beirlant:1984:AOS

[BH84] Jan Beirlant and Lajos Horváth. Approximations of m-overlapping spacings processes. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 225–245, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615961.

Benjamini:1997:MHT

[BH97] Yoav Benjamini and Yosef Hochberg. Multiple hypotheses testing with weights. Scandinavian Journal of Statistics. Theory and Applications, 24(3):407–418, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616464.

Bedrick:1999:PAG

[BH99] Edward J. Bedrick and Joe R. Hill. Properties and applications of the generalized likelihood as a summary function for prediction problems. Scandinavian Journal of Statistics. Theory and Applications, 26(4):593–609, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616581.

Bergman:2014:PPC

[BH14] Jakob Bergman and Björn Holmquist. Poll of polls: a compositional loess model. Scandinavian Journal of Statistics. Theory and Applications, 41(2):301–310, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586654.

Bretagnolle:1988:EOC

[BHC88] Jean Bretagnolle and Catherine Huber-Carol. Effects of omitting covariates in Cox’s model for survival data. Scandinavian Journal of Statistics. Theory and Applications, 15(2): 125–138, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616093.

Bennedsen:2019:LFB

[BHLP19] Mikkel Bennedsen, Ulrich Hounyo, Asger Lunde, and Mikko S. Pakkanen. The local fractional bootstrap. Scandinavian Journal of Statistics. Theory and Applications, 46(1):329–359, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bickel:1976:ALR

[BHR^{+}76] Peter J. Bickel, Sture Holm, Bengt Rosén, Emil Spjøtvoll, Steffen Lauritzen, Søren Johansen, and Ole Barndorff-Nielsen. Another look at robustness: a review of reviews and some new developments [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 3(4):145– 168, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615633.

Bibinger:2011:ECE

[Bib11] Markus Bibinger. Efficient covariance estimation for asynchronous noisy highfrequency data. Scandinavian Journal of Statistics.

Theory and Applications, 38(1):23–45, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015382.

Bickel:2023:SES

[Bic23] David R. Bickel. Statistical evidence and surprise unified under possibility theory. Scandinavian Journal of Statistics. Theory and Applications, 50(3): 923–928, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Biernacki:2007:DML

[Bie07] Christophe Biernacki. Degeneracy in the maximum likelihood estimation of univariate Gaussian mixtures for grouped data and behaviour of the EM algorithm. Scandinavian Journal of Statistics. Theory and Applications, 34(3): 569–586, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548570.

Barczy:2014:ABC

[BIP14] Mátyás Barczy, Márton Ispány, and Gyula Pap. Asymptotic behavior of conditional least squares estimators for unstable integer-valued autoregressive models of order 2. Scandinavian Journal of Statistics. Theory and Applications, 41(4):866–892, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586811.

Baraldo:2013:OPH

[BIPV13] Stefano Baraldo, Francesca Ieva, Anna Maria Paganoni, and Valeria Vitelli. Outcome prediction for heart failure telemonitoring via generalized linear models with functional covariates. Scandinavian Journal of Statistics. Theory and Applications, 40(3):403–416, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586678.

Berk:1978:ROC

[BJ78] Robert H. Berk and Douglas H. Jones. Relatively optimal combinations of test statistics. Scandinavian Journal of Statistics. Theory and Applications, 5(3):158–162, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615706.

Blaesild:1985:SFR

[BJ85] Preben Blæsild and Jens Ledet Jensen. Saddlepoint formulas for reproductive exponential models. Scandinavian Journal of Statistics. Theory and Applications, 12(3):193–202, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615988.

Barbour:1989:LTA

[BJ89] A. D. Barbour and J. L.

Jensen. Local and tail approximations near the Poisson limit. Scandinavian Journal of Statistics. Theory and Applications, 16 (1):75–87, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616123.

Bjork:1993:TFT

[BJ93] Tomas Björk and Björn Johansson. On theorems of De Finetti type for continuous time stochastic processes. Scandinavian Journal of Statistics. Theory and Applications, 20(4): 289–312, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616286.

Branden:2012:NDS

[BJ12] Petter Brändén and Johan Jonasson. Negative dependence in sampling. Scandinavian Journal of Statistics. Theory and Applications, 39(4):830–838, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23357228.

Bolviken:1982:MPR

[BJD82] Erik Bølviken and Kumar Joag-Dev. Monotonicity of the probability of a rectangular region under a multivariate normal distribution. Scandinavian Journal of Statistics. Theory and Applications, 9(3):171–174, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615874.

Bak:2015:CET

[BJFG15] Britta Anker Bak, Jens Ledet Jensen, and Morten Fenger-Grøn. Classification error of the thresholded independence rule. Scandinavian Journal of Statistics. Theory and Applications, 42(1):32–42, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Brendel:2014:WLP

[BJMP14] Michael Brendel, Arnold Janssen, Claus-Dieter Mayer, and Markus Pauly. Weighted logrank permutation tests for randomly right censored life science data. Scandinavian Journal of Statistics. Theory and Applications, 41(3):742–761, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586756.

Bjorkstrom:2010:KST

[Bjö10] Anders Björkström. Krylov sequences as a tool for analysing iterated regression algorithms. Scandinavian Journal of Statistics. Theory and Applications, 37(1):166–175, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000921.

Buhlmann:1995:BBG

[BK95] Peter Bühlmann and Hans R.

Künsch. The blockwise bootstrap for general parameters of a stationary time series. Scandinavian Journal of Statistics. Theory and Applications, 22(1): 35–54, March 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616339.

Bhattacharya:2023:ODB

[BKB23] Shrijita Bhattacharya, Francois Kamper, and Jan Beirlant. Outlier detection based on extreme value theory and applications. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1466–1502, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Boege:2023:GGD

[BKKR23] Tobias Boege, Thomas Kahle, Andreas Kretschmer, and Frank Röttger. The geometry of Gaussian double Markovian distributions. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 665–696, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Baek:2018:OSM

[BKM18] Seungchul Baek, Osamu Komori, and Yanyuan Ma. An optimal semiparametric method for two-group classification. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 806–846, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Borgan:2023:UMI

[BKN23] Ørnulf Borgan, Ruth H.

Keogh, and Aleksander Njøs. Use of multiple imputation in supersampled nested case-control and case-cohort studies. Scandinavian Journal of Statistics. Theory and Applications, 50(1):13–37, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Britton:2011:IET

[BKO11] Tom Britton, Theodore Kypraios, and Philip D. O’Neill. Inference for epidemics with three levels of mixing: Methodology and application to a measles outbreak. Scandinavian Journal of Statistics. Theory and Applications, 38(3): 578–599, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015582.

Beyer:1976:EBM

[BKS76] Jan E. Beyer, Niels Keiding, and W. Simonsen. The exact behaviour of the maximum likelihood estimator in the pure birth process and the pure death process. Scandinavian Journal of Statistics. Theory and Applications, 3(2):61– 72, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615612.

BenAlaya:2020:LAP

[BKT20] Mohamed Ben Alaya, Ahmed Kebaier, and Ngoc Khue Tran. Local asymptotic properties for Cox– Ingersoll–Ross process with discrete observations. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1401–1464, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Buser:2010:BUC

[BKW10] Christoph M. Buser, Hans R. Künsch, and Alain Weber. Biases and uncertainty in climate projections. Scandinavian Journal of Statistics. Theory and Applications, 37(2): 179–199, June 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000874.

Bar-Lev:1983:APM

[BL83] Shaul K. Bar-Lev. On the asymptotic properties of the maximum conditional likelihood estimate for the exponential family of distributions. Scandinavian Journal of Statistics. Theory and Applications, 10 (2):135–140, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615912.

Borgan:1990:NCI

[BL90] Ørnulf Borgan and Knut Liestøl. A note on confidence intervals and bands for the survival function based on transformations. Scandinavian Journal of Statistics. Theory and Applications, 17(1):35–41, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616153.

Baxter:1994:NPC

[BL94] Laurence A. Baxter and Linxiong Li. Non-parametric confidence intervals for the renewal function and the point availability. Scandinavian Journal of Statistics. Theory and Applications, 21(3):277–287, September 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616317.

Borgan:2007:E

[BL07] Ørnulf Borgan and Bo Lindqvist.

Editorial. Scandinavian Journal of Statistics. Theory and Applications, 34(1):

1–2, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548534.

Bartolucci:2008:FIC

[BL08] Francesco Bartolucci and Monia Lupparelli. Focused information criterion for capture–recapture models for closed populations. Scandinavian Journal of Statistics. Theory and Applications, 35(4):629–649, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000293.

Biscio:2017:CEP

[BL17] Christophe Ange Napoléon Biscio and Frédéric Lavancier. Contrast estimation for parametric stationary determinantal point processes. Scandinavian Journal of Statistics. Theory and Applications, 44 (1):204–229, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Blaker:1999:SOD

[Bla99] Helge Blaker. Shrinkage and orthogonal decomposition. Scandinavian Journal of Statistics. Theory and Applications, 26(1):1– 15, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616537.

Blaker:2001:MLS

[Bla01] Helge Blaker. Minimax linear smoothers. Scandinavian Journal of Statistics. Theory and Applications, 28(1):151–160, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616649.

Blaesild:1978:GED

[Blæ78] P. Blæsild. A generalization of the exponential distribution to convex sets in R^{k}. Scandinavian Journal of Statistics. Theory and Applications, 5(4):189– 194, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615714.

Bar-Lev:1992:CPI

[BLBEO92] Shaul K. Bar-Lev, Daoud Bshouty, Peter Enis, and Avital Y. Ohayon. Compositions and products of infinitely divisible variance functions. Scandinavian Journal of Statistics. Theory and Applications, 19 (1):83–89, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616227.

Beran:2020:ASD

[BLG20] Jan Beran, Haiyan Liu, and Sucharita Ghosh. On aggregation of strongly dependent time series. Scandinavian Journal of Statistics. Theory and Applications, 47(3):690–710, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bouveyron:2020:EDS

[BLM20] Charles Bouveyron, Pierre Latouche, and Pierre-Alexandre Mattei. Exact dimensionality selection for Bayesian PCA. Scandinavian Journal of Statistics. Theory and Applications, 47(1):196–211, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Blomqvist:1974:SDG

[Blo74] Nils Blomqvist. A simple derivation of the GI/G/1 waiting time distribution in heavy traffic. Scandinavian Journal of Statistics. Theory and Applications, 1(1):39–40, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615549.

Blom:1975:R

[Blo75] Gunnar Blom. On raoblackwellization. Scandinavian Journal of Statistics. Theory and Applications, 2 (3):145–148, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615594.

Blyth:1993:NCC

[Bly93] Stephen Blyth. A note on correlation curves and Chernoff’s inequality. Scandinavian Journal of Statistics. Theory and Applications, 20(4):375–377, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616292.

Bailey:2001:ESL

[BM01a] R. A. Bailey and H. Monod.

Efficient semi-Latin rectangles: Designs for plant disease experiments. Scandinavian Journal of Statistics. Theory and Applications, 28(2):257–270, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616657.

Brix:2001:STM

[BM01b] Anders Brix and Jesper Møller. Space-time multi type log Gaussian Cox processes with a view to modelling weeds. Scandinavian Journal of Statistics. Theory and Applications, 28(3): 471–488, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616673.

Berthelsen:2003:LNP

[BM03] Kasper K. Berthelsen and Jesper Møller. Likelihood and non-parametric Bayesian MCMC inference for spatial point processes based on perfect simulation and path sampling. Scandinavian Journal of Statistics. Theory and Applications, 30(3):549–564, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616783.

Baker:2015:DES

[BM15] Rose D. Baker and Ian G. McHale. Deterministic evolution of strength in multiple comparisons models: Who is the greatest golfer? Scandinavian Journal of Statistics. Theory and Applications, 42(1):180–196, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bibinger:2016:IMD

[BM16] Markus Bibinger and Per A.

Mykland. Inference for multi-dimensional highfrequency data with an application to conditional independence testing. Scandinavian Journal of Statistics. Theory and Applications, 43(4):1078–1102, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Borgan:1982:IBN

[BMG82] Ørnulf Borgan, Torkil Menné, and Anders Green. Interaction between nickel allergy and hand eczema in the Danish female population. Scandinavian Journal of Statistics. Theory and Applications, 9(3):183–185, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615877.

Bodnar:2019:CLT

[BMP19] Taras Bodnar, Stepan Mazur, and Nestor Parolya. Central limit theorems for functionals of large sample covariance matrix and mean vector in matrixvariate location mixture of normal distributions. Scandinavian Journal of Statistics. Theory and Applications, 46(2):636–660, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bi:2020:IDA

[BMXT20] Nan Bi, Jelena Markovic, Lucy Xia, and Jonathan Taylor. Inferactive data analysis. Scandinavian Journal of Statistics. Theory and Applications, 47 (1):212–249, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Barndorff-Nielsen:1978:HDD

[BN78] O. Barndorff-Nielsen. Hyperbolic distributions and distributions on hyperbolae. Scandinavian Journal of Statistics. Theory and Applications, 5(3):151–157, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615705.

Barndorff-Nielsen:1982:HDS

[BN82] O. Barndorff-Nielsen. The hyperbolic distribution in statistical physics. Scandinavian Journal of Statistics. Theory and Applications, 9(1):43–46, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615853.

Barndorff-Nielsen:1984:CRL

[BN84] O. E. Barndorff-Nielsen.

On conditionality resolution and the likelihood ratio for curved exponential models. Scandinavian Journal of Statistics. Theory and Applications, 11 (3):157–170, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615953. See cor-

rection [BN85b].

Barndorff-Nielsen:1985:CLS

[BN85a] O. E. Barndorff-Nielsen.

Confidence limits from c|ĵ|^{1/2}\overline{L} in the singleparameter case. Scandinavian Journal of Statistics. Theory and Applications, 12(1):83–87, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615976.

Barndorff-Nielsen:1985:CNC

[BN85b] Ole E. Barndorff-Nielsen.

Correction note: On Conditionality Resolution and the Likelihood Ratio for Curved Exponential Models. Scandinavian Journal of Statistics. Theory and Applications, 12(2): 191, September 1985. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615985. See [BN84].

Barndorff-Nielsen:1990:NSS

[BN90] O. E. Barndorff-Nielsen. A note on the standardized signed log likelihood ratio. Scandinavian Journal of Statistics. Theory and Applications, 17(2):157–160, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616163.

Barndorff-Nielsen:1997:NIG

[BN97] Ole E. Barndorff-Nielsen. Normal inverse Gaussian distributions and stochastic volatility modelling. Scandinavian Journal of Statistics. Theory and Applications, 24(1):1–13, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616433.

Birke:2013:TMR

[BN13] Melanie Birke and Natalie Neumeyer. Testing monotonicity of regression functions — an empirical process approach. Scandinavian Journal of Statistics. Theory and Applications, 40(3):438–454, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586680.

Bibbona:2015:HMP

[BN15] Enrico Bibbona and Ilia Negri. Higher moments and prediction-based estimation for the COGA-RCH(1,1) model. Scandinavian Journal of Statistics. Theory and Applications, 42(4):891–910, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Barndorff-Nielsen:1993:LYS

[BNB93] O. E. Barndorff-Nielsen and P. Blæsild. Likelihood yokes and stable combinations of cumulants. Scandinavian Journal of Statistics. Theory and Applications, 20(1):83– 87, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616263.

Barndorff-Nielsen:1991:AIM

[BNC91] O. E. Barndorff-Nielsen and S. R. Chamberlin. An ancillary invariant modification of the signed log likelihood ratio. Scandinavian Journal of Statistics. Theory and Applications, 18 (4):341–352, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616216.

Barndorff-Nielsen:1995:DEB

[BNHH95] Ole E. Barndorff-Nielsen, Inge S. Helland, and Sture Holm. Diversity of evidence and Birnbaum’s Theorem [with comments and reply]. Scandinavian Journal of Statistics. Theory and Applications, 22(4): 513–522, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616378.

Barndorff-Nielsen:1976:MSL

[BNHJP76] O. Barndorff-Nielsen, J. Hoffmann-Jørgensen, and K. Pedersen. On the minimal sufficiency of the likelihood function. Scandinavian Journal of Statistics. Theory and Applications, 3 (1):37–38, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615606.

Barndorff-Nielsen:1974:NEF

[BNK74] O. Barndorff-Nielsen and H. K. Kvist. Note on exponential families and Mancillarity. Scandinavian Journal of Statistics. Theory and Applications, 1(1):

36–38, March 1974. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615548.

Barndorff-Nielsen:1999:TEM

[BNK99] O. E. Barndorff-Nielsen and C. Klüppelberg. Tail exactness of multivariate saddlepoint approximations. Scandinavian Journal of Statistics. Theory and Applications, 26(2):253–264, June 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616554.

Barndorff-Nielsen:2007:LCD

[BNL07] Ole E. Barndorff-Nielsen and Alexander M. Lindner. Lévy copulas: Dynamics and transforms of upsilon type. Scandinavian Journal of Statistics. Theory and Applications, 34(2):298–316, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548553.

Barndorff-Nielsen:2014:IVT

[BNLSV14] Ole E. Barndorff-Nielsen, Asger Lunde, Neil Shephard, and Almut E. D. Veraart. Integer-valued trawl processes: a class of stationary infinitely divisible processes. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 693–724, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586754.

Biedermann:2006:TCC

[BNM^{+}06] Stefanie Biedermann, Eva Nagel, Axel Munk, Hajo Holzmann, and Ansgar Steland. Tests in a case– control design including relatives. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 621–635, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616949.

Barndorff-Nielsen:1979:BHP

[BNP79] Ole Barndorff-Nielsen and Bo V. Pedersen. The bivariate Hermite polynomials up to order six. Scandinavian Journal of Statistics. Theory and Applications, 6(3): 127–128, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615747.

Baksalary:1992:ECU

[BNP92] Jerzy K. Baksalary, Markku Nurhonen, and Simo Puntanen. Effect of correlations and unequal variances in testing for outliers in linear regression. Scandinavian Journal of Statistics. Theory and Applications, 19 (1):91–95, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616228.

Barndorff-Nielsen:2000:EDR

[BNR00] Ole E. Barndorff-Nielsen and Tina Hviid Rydberg. Exact distributional results for random resistance trees. Scandinavian Journal of Statistics. Theory and Applications, 27(1):129–141, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616595.

Barndorff-Nielsen:2003:IOP

[BNS03] Ole E. Barndorff-Nielsen and Neil Shephard. Integrated OU processes and non-Gaussian OU-based stochastic volatility models. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 277–295, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616764.

Barndorff-Nielsen:2005:AMG

[BNS05] Ole Eiler Barndorff-Nielsen and Robert Stelzer. Absolute moments of generalized hyperbolic distributions and approximate scaling of normal inverse Gaussian Lévy processes. Scandinavian Journal of Statistics. Theory and Applications, 32(4):617–637, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616901.

Borgan:1999:ESC

[BO99] Ørnulf Borgan and Espen F. Olsen. The efficiency of simple and counter-matched nested case-control sampling. Scandinavian Journal of Statistics. Theory and Applications, 26(4):493–509, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616575.

Britton:2002:BIS

[BO02] Tom Britton and Philip D. O’Neill. Bayesian inference for stochastic epidemics in populations with random social structure. Scandinavian Journal of Statistics. Theory and Applications, 29(3):375–390, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616722.

Bodnar:2011:PIW

[BO11] Taras Bodnar and Yarema Okhrin. On the product of inverse Wishart and normal distributions with applications to discriminant analysis and portfolio theory. Scandinavian Journal of Statistics. Theory and Applications, 38(2): 311–331, June 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015496.

Boes:2010:CDM

[Boe10] Stefan Boes. Count data models with correlated unobserved heterogeneity. Scandinavian Journal of Statistics. Theory and Applications, 37(3):382–402, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000397.

Bohning:2010:SGC

[Böh10] Dankmar Böhning. Some general comparative points on Chao’s and Zelterman’s estimators of the population size. Scandinavian Journal of Statistics. Theory and Applications, 37(2): 221–236, June 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000876.

Bolviken:1982:PIM

[Bøl82] Erik Bølviken. Probability inequalities for the multivariate normal with nonnegative partial correlations. Scandinavian Journal of Statistics. Theory and Applications, 9(1):49– 58, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615855.

Bolviken:1983:DCR

[Bøl83a] Erik Bølviken. The distribution of certain rational functions of order statistics from exponential distributions. Scandinavian Journal of Statistics. Theory and Applications, 10 (2):117–123, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615910.

Bolviken:1983:NTS

[Bøl83b] Erik Bølviken. New tests of significance in periodogram analysis. Scandinavian Journal of Statistics. Theory and Applications, 10(1): 1–9, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615894.

Bolviken:1988:SPB

[Bøl88] Erik Bølviken. Some probability bounds relevant for Bonferroni significance levels under Gaussian models. Scandinavian Journal of Statistics. Theory and Applications, 15(4): 281–297, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616112.

Bolin:2014:SMF

[Bol14]  David Bolin.  Spatial

Matérn fields driven by non-Gaussian noise. Scandinavian Journal of Statistics. Theory and Applications, 41(3):557–579, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586747.

Bondesson:1975:NDG

[Bon75] Lennart Bondesson. Normal distribution, gamma distribution, and quadratic polynomial statistics. Scandinavian Journal of Statistics. Theory and Applications, 2(3):138–144, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615593.

Bondesson:1976:WSM

[Bon76] Lennart Bondesson. When is the sample mean BLUE? Scandinavian Journal of Statistics. Theory and Applications, 3(3):116–120, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615623.

Bondesson:1979:BEB

[Bon79] Lennart Bondesson. On best equivariant and best unbiased equivariant estimators in the extended Gauss–Markov model. Scandinavian Journal of Statistics. Theory and Applications, 6(1):22–28, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615724.

Bondesson:1982:RCH

[Bon82] Lennart Bondesson. To reduce a composite hypothesis to a simple one by sampling from the structural or a conditional distribution. Scandinavian Journal of Statistics. Theory and Applications, 9(3): 129–138, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615868.

Bondesson:2001:E

[Bon01] Lennart Bondesson. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 28(1): 1, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616642.

Bondesson:2010:CRP

[Bon10] Lennart Bondesson. Conditional and restricted Pareto sampling: Two new methods for unequal probability sampling. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 514–530, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000404.

Bondesson:2012:SPS

[Bon12] Lennart Bondesson. On sampling with prescribed second-order inclusion probabilities. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 813–829, December 2012.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357227.

Borgan:1984:CNM

[Bor84a] Ørnulf Borgan. Correction note: Maximum Likelihood Estimation in Parametric Counting Process Models, with Applications to Censored Failure Time Data. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 275, September 1984. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615966. See [Bor84b].

Borgan:1984:MLE

[Bor84b]  Ørnulf Borgan.  Maxi-

mum likelihood estimation in parametric counting process models, with applications to censored failure time data. Scandinavian Journal of Statistics. Theory and Applications, 11(1): 1–16, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615936. See correction [Bor84a].

Bordes:1999:SAA

[Bor99] Laurent Bordes. Semiparametric additive accelerated life models. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 345–361, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616561.

Bochynek:1989:TCR

[BP89] Juergen Bochynek and Georg Ch. Pflug. Testing coincidences of renewal processes. Scandinavian Journal of Statistics. Theory and Applications, 16(2): 145–152, June 1989. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616130.

Bohning:2005:ANM

[BP05] Dankmar Böhning and Valentin Patilea. Asymptotic normality in mixtures of power series distributions. Scandinavian Journal of Statistics. Theory and Applications, 32 (1):115–131, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616864.

Borgonovo:2022:ICE

[BPR22] Emanuele Borgonovo, Elmar Plischke, and Giovanni Rabitti. Interactions and computer experiments.

Scandinavian Journal of Statistics. Theory and Applications, 49(3):1274–1303, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Beranger:2017:MED

[BPS17] Boris Beranger, Simone A. Padoan, and Scott A. Sisson. Models for extremal dependence derived from skew-symmetric families. Scandinavian Journal of Statistics. Theory and Applications, 44(1):21– 45, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Baxevani:2014:SPA

[BPW14] Anastassia Baxevani, Krzysztof Podgórski, and Jörg Wegener. Sample path asymmetries in non-Gaussian random processes. Scandinavian Journal of Statistics. Theory and Applications, 41(4):1102–1123, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586823.

Berg:2009:LPA

[BQ09] Daniel Berg and Jean-François Quessy. Local power analyses of goodnessof-fit tests for copulas.

Scandinavian Journal of Statistics. Theory and Applications, 36(3):389–412, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000329.

Bahraoui:2022:TMC

[BQ22] Tarik Bahraoui and Jean-François Quessy. Tests of multivariate copula exchangeability based on Lévy measures. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1215–1243, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Best:1981:NMG

[BR81] D. J. Best and J. C. W.

Rayner. A note on Mineo’s grouping method for the chi-square test of goodness of fit. Scandinavian Journal of Statistics. Theory and Applications, 8(3):185–186, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615832. See [Min79] and rejoinder [Min81].

Bergman:1997:BSR

[BR97] Bo Bergman and Mikael Ringi. Bayesian system reliability prediction.

Scandinavian Journal of Statistics. Theory and Applications, 24(1):137–143, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616442.

Banerjee:2003:DGG

[BR03] Moulinath Banerjee and Thomas Richardson. On a dualization of graphical Gaussian models: a correction note. Scandinavian Journal of Statistics. Theory and Applications, 30(4): 817–820, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616804.

Bibinger:2014:SEC

[BR14] Markus Bibinger and Markus Reiß. Spectral estimation of covolatility from noisy observations using local weights. Scandinavian Journal of Statistics. Theory and Applications, 41(1):23–50, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586729.

Bienvenue:2017:LIM

[BR17] Alexis Bienvenüe and Christian Y. Robert. Likelihood inference for multivariate extreme value distributions whose spectral vectors have known conditional distributions. Scandinavian Journal of Statistics. Theory and Applications, 44(1): 130–149, March 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bochkina:2023:BIP

[BR23] Natalia Bochkina and Jenovah Rodrigues. Bayesian inverse problems with heterogeneous variance. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1116–1151, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bohlin:1983:MSA

[BRH83] Torsten Bohlin, Holger Rootzén, and Jan Holst. Mathematical statistics and automatic control [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 10(3):195–221, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615919.

Britton:1997:TDW

[Bri97] Tom Britton. A test to detect within-family infectivity when the whole epidemic process is observed. Scandinavian Journal of Statistics. Theory and Applications, 24(3):315–330, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616458.

Boente:2014:GFT

[BRM14] Graciela Boente, Daniela Rodriguez, and Wenceslao González Manteiga. Goodness-of-fit test for directional data. Scandinavian Journal of Statistics. Theory and Applications, 41(1):259–275, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586741.

Brown:1980:CCC

[Bro80] P. J. Brown. Coherence and complexity in classification problems. Scandinavian Journal of Statistics. Theory and Applications, 7(2):95–98, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615780.

Broberg:1987:NEP

[Bro87a] Per Broberg. A note on the extinction probability of branching populations.

Scandinavian Journal of Statistics. Theory and Applications, 14(2):125–129, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616056.

Brostrom:1987:IMD

[Bro87b] Göran Broström. The influence of mother’s death on infant mortality: a case study in matched data survival analysis. Scandinavian Journal of Statistics. Theory and Applications, 14(2):113–123, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616055.

Bjorkstrom:1999:GVC

[BS99] Anders Björkström and Rolf Sundberg. A generalized view on continuum regression. Scandinavian Journal of Statistics. Theory and Applications, 26(1):17–30, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616538.

Belyaev:2000:ADA

[BS00] Yuri K. Belyaev and Oleg Seleznjev. Approaching in distribution with applications to resampling of stochastic processes. Scandinavian Journal of Statistics. Theory and Applications, 27(2):371–384, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616609.

Bibby:2001:SEF

[BS01] Bo Martin Bibby and Michael Sørensen. Simplified estimating functions for diffusion models with a high-dimensional parameter. Scandinavian Journal of Statistics. Theory and Applications, 28(1):99– 112, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616646.

Bari:2010:BRR

[BS10] Wasimul Bari and Brajendra C. Sutradhar. On bias reduction in robust inference for generalized linear models. Scandinavian Journal of Statistics. Theory and Applications, 37 (1):109–125, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000918.

Bartroff:2016:RPS

[BS16] Jay Bartroff and Jinlin Song. A rejection principle for sequential tests of multiple hypotheses controlling familywise error rates. Scandinavian Journal of Statistics. Theory and Applications, 43(1):3– 19, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bours:2021:LSA

[BS21] Monika Bours and Ansgar Steland. Largesample approximations and change testing for highdimensional covariance matrices of multivariate linear time series and factor models. Scandinavian Journal of Statistics. Theory and Applications, 48 (2):610–654, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Barreto-Souza:2024:NUI

[BSC24] Wagner Barreto-Souza and Ngai Hang Chan. Nearly unstable integer-valued ARCH process and unit root testing. Scandinavian Journal of Statistics. Theory and Applications, 51(1): 402–424, March 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Barreto-Souza:2022:NBP

[BSO22] Wagner Barreto-Souza and Hernando Ombao. The negative binomial process: a tractable model with composite likelihood-based inference. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 568–592, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Belitser:2013:EPC

[BSV13] Eduard Belitser, Paulo Serra, and Harry Van Zanten. Estimating the period of a cyclic nonhomogeneous Poisson process. Scandinavian Journal of Statistics. Theory and Applications, 40(2): 204–218, June 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586709.

Bondesson:2008:LSS

[BT08] Lennart Bondesson and Daniel Thorburn. A list sequential sampling method suitable for real-time sampling. Scandinavian Journal of Statistics. Theory and Applications, 35(3): 466–483, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000282.

Bondesson:2013:SDO

[BT13] Lennart Bondesson and Imbi Traat. On sampling designs with ordered conditional inclusion probabilities. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 724–733, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586696.

Beran:2021:IMU

[BT21] Jan Beran and Klaus Telkmann. On inference for modes under long memory. Scandinavian Journal of Statistics. Theory and Applications, 48(2): 429–455, June 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Bondesson:2006:PSV

[BTL06] Lennart Bondesson, Imbi Traat, and Anders Lundqvist. Pareto sampling versus Sampford and conditional Poisson sampling. Scandinavian Journal of Statistics. Theory and Applications, 33(4):699–720, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616953.

Buhl:1993:EML

[Buh93]  Søren L. Buhl.  On

the existence of maximum likelihood estimators for graphical Gaussian models. Scandinavian Journal of Statistics. Theory and Applications, 20(3): 263–270, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616281.

Butler:1986:EOF

[But86] Ronald W. Butler. Extendibility and the optimality of F, T^{2} and forward variable selection. Scandinavian Journal of Statistics. Theory and Applications, 13(4):257–262, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616034.

Butler:1998:GIG

[But98] Ronald W. Butler. Generalized inverse Gaussian distributions and their Wishart connections. Scandinavian Journal of Statistics. Theory and Applications, 25(1): 69–75, March 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616485.

Bigot:2009:LDD

[BV09] Jérémie Bigot and Sébastien Van Bellegem. Log-density deconvolution by wavelet thresholding. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 749–763, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000350.

Butucea:2014:SMS

[BV14] Cristina Butucea and Pierre Vandekerkhove. Semiparametric mixtures of symmetric distributions. Scandinavian Journal of Statistics. Theory and Applications, 41(1):227–239, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586739.

Birke:2017:SPE

[BVV17] Melanie Birke, Sebastien Van Bellegem, and Ingrid Van Keilegom. Semiparametric estimation in a single-index model with endogenous variables. Scandinavian Journal of Statistics. Theory and Applications, 44(1):168–191, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Butler:2004:DCN

[BW04] Ronald W. Butler and Andrew T. A. Wood. A dimensional CLT for non-central Wilks’ Lambda in multivariate analysis. Scandinavian Journal of Statistics. Theory and Applications, 31(4):585–601, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616852.

Banerjee:2005:CIC

[BW05] Moulinath Banerjee and Jon A. Wellner. Confidence intervals for current status data. Scandinavian Journal of Statistics. Theory and Applications, 32(3): 405–424, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616887.

Breslow:2007:WLS

[BW07] Norman E. Breslow and Jon A. Wellner. Weighted likelihood for semiparametric models and two-phase stratified samples, with application to Cox regression. Scandinavian Journal of Statistics. Theory and Applications, 34(1):86– 102, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548540.

Breslow:2008:TEN

[BW08] Norman E. Breslow and Jon A. Wellner. A Ztheorem with estimated nuisance parameters and correction note for ’Weighted Likelihood for Semiparametric Models and Twophase Stratified Samples, with Application to Cox Regression’. Scandinavian Journal of Statistics. Theory and Applications, 35 (1):186–192, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000392.

Biscio:2019:GCL

[BW19] Christophe Ange Napoléon Biscio and Rasmus Waagepetersen. A general central limit theorem and a subsampling variance estimator for α-mixing point processes. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1168–1190, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Brown:1982:MRS

[BZ82] P. J. Brown and J. V.

Zidek. Multivariate regression shrinkage estimators with unknown covariance matrix. Scandinavian Journal of Statistics. Theory and Applications, 9 (4):209–215, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615884.

Bai:2008:PLM

[BZF08] Yang Bai, Zhongyi Zhu, and Wing K. Fung. Partial linear models for longitudinal data based on quadratic inference functions. Scandinavian Journal of Statistics. Theory and Applications, 35(1): 104–118, March 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000387.

Cacoullos:1977:BEM

[Cac77] T. Cacoullos. Best estimation for multiply truncated power series distributions. Scandinavian Journal of Statistics. Theory and Applications, 4(4):159– 164, September 1977. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615672.

Carroll:1982:TRP

[Car82] R. J. Carroll. Tests for regression parameters in power transformation models. Scandinavian Journal of Statistics. Theory and Applications, 9(4):217– 222, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615885.

Cardot:2007:CFP

[Car07] HervÉ Cardot. Conditional functional principal components analysis. Scandinavian Journal of Statistics. Theory and Applications, 34(2):317–335, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548554.

Capitanio:2003:GMS

[CAS03] A. Capitanio, A. Azzalini, and E. Stanghellini. Graphical models for skewnormal variates. Scandinavian Journal of Statistics. Theory and Applications, 30(1):129–144, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616753.

Cavicchioli:2016:SAM

[Cav16]  Maddalena  Cavicchioli.

Statistical analysis of mixture vector autoregressive models. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1192–1213, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cavicchioli:2023:SAM

[Cav23]  Maddalena  Cavicchioli.

Spectral analysis of Markov switching GARCH models with statistical inference. Scandinavian Journal of Statistics. Theory and Applications, 50(1):102–119, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Conde-Amboage:2021:NLF

[CAVGM21] Mercedes Conde-Amboage, Ingrid Van Keilegom, and Wenceslao González-Manteiga. A new lack-of-fit test for quantile regression with censored data. Scandinavian Journal of Statistics. Theory and Applications, 48(2):655–688, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chow:1984:NMP

[CB84] Winston Kai-Wan Chow and Steinar Bjerve. A new method of power approximation with application to the Wilcoxon statistics. Scandinavian Journal of Statistics. Theory and Applications, 11(1):29– 37, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615938.

Chang:1998:IWE

[CC98] Myron N. Chang and Daehyun Chung. Isotonic window estimators of the baseline hazard function in Cox’s regression model under order restriction. Scandinavian Journal of Statistics. Theory and Applications, 25(1):151–161, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616491.

Clifford:2012:SAP

[CC12] Peter Clifford and Ioana A.

Cosma. A statistical analysis of probabilistic counting algorithms. Scandinavian Journal of Statistics. Theory and Applications, 39(1): 1–14, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411154.

Chang:1998:CEN

[CCH98] I-Shou Chang, Li-Ru Chen, and Chao A. Hsiung. A class of efficient nonparametric K-sample tests for multivariate counting processes with frailties. Scandinavian Journal of Statistics. Theory and Applications, 25(1):127–141, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616489.

Chang:2001:GFT

[CCH01] I-Shou Chang, Yuan-Chuan Chuang, and Chao A. Hsiung. Goodness-of-fit tests for semi-Markov and Markov survival models with one intermediate state. Scandinavian Journal of Statistics. Theory and Applications, 28(3): 505–525, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616675.

Cereda:2023:LTP

[CCV23] Giulia Cereda, Fabio Corradi, and Cecilia Viscardi. Learning the two parameters of the Poisson–Dirichlet distribution with a forensic application. Scandinavian Journal of Statistics. Theory and Applications, 50 (1):120–141, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2019:ANE

[CCWZ19] Xuerong Chen, Yan Chen, Alan T. K. Wan, and Yong Zhou. On the asymptotic non-equivalence of efficient-GMM and MEL estimators in models with missing data. Scandinavian Journal of Statistics. Theory and Applications, 46(2): 361–388, June 2019. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cao:1996:CSM

[CD96] Ricardo Cao and Luc Devroye. The consistency of a smoothed minimum distance estimate. Scandinavian Journal of Statistics. Theory and Applications, 23(4):405–418, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616417.

Cheng:2001:PAH

[CD01] S. Cheng and L. De Haan.

Penultimate approximation for Hill’s estimator. Scandinavian Journal of Statistics. Theory and Applications, 28(3):569–575, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616679.

Causeur:2003:LRM

[CD03] David Causeur and Thierry Dhorne. Linear regression models under conditional independence restrictions. Scandinavian Journal of Statistics. Theory and Applications, 30(3):637–650, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616790.

Comte:2018:SIR

[CD18] F. Comte and C. Duval. Statistical inference for renewal processes. Scandinavian Journal of Statistics. Theory and Applications, 45(1):164–193, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Coeurjolly:2012:TFM

[CDDL12] Jean-Franois Coeurjolly, David Dereudre, Rémy Drouilhet, and Frédéric Lavancier. Takacs–Fiksel method for stationary marked Gibbs point processes. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 416–443, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41679804.

Cho:2016:APE

[CDG16] Yong Bum Cho, Richard A.

Davis, and Souvik Ghosh. Asymptotic properties of the empirical spatial extremogram. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 757–773, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Comte:2015:EJS

[CDGCK15] Fabienne Comte, Celine Duval, Valentine Genon-Catalot, and Johanna Kappus. Estimation of the jump size density in a mixed compound Poisson process. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1023–1044, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Colubi:2006:TCI

[CDMGR06] Ana Colubi, J. Santos Domínguez-Menchero, and Gil González-Rodríguez. Testing constancy for isotonic regressions. Scandinavian Journal of Statistics. Theory and Applications, 33(3):463–475, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616937.

Cappe:2002:CMC

[CDMR02] Olivier Cappé, Randal Douc, Eric Moulines, and Christian Robert. On the convergence of the Monte Carlo maximum likelihood method for latent variable models. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 615–635, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616738.

Cotsakis:2024:PEP

[CDO24] Ryan Cotsakis, Elena Di Bernardino, and Thomas Opitz. On the perimeter estimation of pixelated excursion sets of two-dimensional anisotropic random fields. Scandinavian Journal of Statistics. Theory and Applications, 51(1):268–301, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2020:PLB

[CDQ20] Xuerong Chen, Guoqing Diao, and Jing Qin. Pseudo likelihood-based estimation and testing of missingness mechanism function in nonignorable missing data problems. Scandinavian Journal of Statistics. Theory and Applications, 47(4): 1377–1400, December 2020.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Croux:2011:OME

[CDY11] Christophe Croux, Catherine Dehon, and Abdelilah Yadine. On the optimality of multivariate Sestimators. Scandinavian Journal of Statistics. Theory and Applications, 38(2): 332–341, June 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015497.

Cheng:2011:EEA

[CDZ11] Yebin Cheng, Jan G. De Gooijer, and Dawit Zerom. Efficient estimation of an additive quantile regression model. Scandinavian Journal of Statistics. Theory and Applications, 38(1):46–62, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015383.

Cereda:2017:IMC

[Cer17] Giulia Cereda. Impact of model choice on LR assessment in case of rare haplotype match (frequentist approach). Scandinavian Journal of Statistics. Theory and Applications, 44 (1):230–248, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ceyhan:2010:NTS

[Cey10] Elvan Ceyhan. New tests of spatial segregation based on nearest neighbour contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 37(1):147–165, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000920.

Christensen:2007:TCI

[CFJP07] Ole F. Christensen, Morten Frydenberg, Jens L. Jensen, and Jørgen G. Pedersen. Tests and confidence intervals for an extended variance component using the modified likelihood ratio statistic. Scandinavian Journal of Statistics. Theory and Applications, 34(2):347–364, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548556.

Cook:2024:EML

[CFL24] R. Dennis Cook, Liliana Forzani, and Lan Liu. Envelopes for multivariate linear regression with linearly constrained coefficients. Scandinavian Journal of Statistics. Theory and Applications, 51(2): 429–446, June 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cardot:2003:THF

[CFMS03] Hervé Cardot, Frédéric Ferraty, André Mas, and Pascal Sarda. Testing hypotheses in the functional linear model. Scandinavian Journal of Statistics. Theory and Applications, 30 (1):241–255, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616760.

Cerioli:2019:WAT

[CFR19] Andrea Cerioli, Alessio Farcomeni, and Marco Riani. Wild adaptive trimming for robust estimation and cluster analysis. Scandinavian Journal of Statistics. Theory and Applications, 46 (1):235–256, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cohen:1995:UTT

[CFS95] Arthur Cohen, Mendel Fygenson, and H. B. Sackrowitz. Unbiased testing of treatments against a control. Scandinavian Journal of Statistics. Theory and Applications, 22(1):95– 103, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616343.

Corcuera:1999:GBR

[CG99] José Manuel Corcuera and Federica Giummolè. A generalized Bayes rule for prediction. Scandinavian Journal of Statistics. Theory and Applications, 26(2): 265–279, June 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616555.

Comte:2006:PPE

[CGC06] F. Comte and V. Genon-Catalot. Penalized projection estimator for volatility density. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 875–893, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616963.

Comte:2023:NAE

[CGC23] Fabienne Comte and Valentine Genon-Catalot. Nonparametric adaptive estimation for interacting particle systems. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1716–1755, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Colombi:2019:HMM

[CGGI19] Roberto Colombi, Sabrina Giordano, Anna Gottard, and Maria Iannario. Hierarchical marginal models with latent uncertainty. Scandinavian Journal of Statistics. Theory and Applications, 46(2):595–620, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cox:1981:SAT

[CGL^{+}81] D. R. Cox, Gudmundur Gudmundsson, Georg Lindgren, Lennart Bondesson, Erik Harsaae, Petter Laake, Katarina Juselius, and Steffen L. Lauritzen. Statistical analysis of time series: Some recent developments [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 8(2):93– 115, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615819.

Cardot:2014:VEA

[CGL14] HervÉ Cardot, Camelia Goga, and Pauline Lardin. Variance estimation and asymptotic confidence bands for the mean estimator of sampled functional data with high entropy unequal probability sampling designs. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 516–534, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586666.

Commenges:2007:LGC

[CGP07] Daniel Commenges and Anne Gégout-Petit. Likelihood for generally coarsened observations from multistate or counting process models. Scandinavian Journal of Statistics. Theory and Applications, 34(2):432–450, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548561.

Csorgo:1982:CHP

[CH82] Sándor Csörgő and Lajos Horváth. On cumulative hazard processes under random censorship. Scandinavian Journal of Statistics. Theory and Applications, 9 (1):13–21, June 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615849.

Chang:1996:EEP

[CH96] I-Shou Chang and Chao A.

Hsiung. An efficient estimator for proportional hazards models with frailties. Scandinavian Journal of Statistics. Theory and Applications, 23(1):13– 26, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616382.

Claeskens:2004:GFN

[CH04] Gerda Claeskens and Nils Lid Hjort. Goodness of fit via non-parametric likelihood ratios. Scandinavian Journal of Statistics. Theory and Applications, 31(4): 487–513, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616847.

Cunen:2022:CIA

[CH22] Céline Cunen and Nils Lid Hjort. Combining information across diverse sources: The II-CC-FF paradigm. Scandinavian Journal of Statistics. Theory and Applications, 49 (2):625–656, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2023:GPM

[CH23] Sixia Chen and David Haziza. General purpose multiply robust data integration procedures for handling nonprobability samples. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 697–724, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chandra:1977:MPD

[Cha77] Satish Chandra. On the mixtures of probability distributions. Scandinavian Journal of Statistics. Theory and Applications, 4(3): 105–112, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615660.

Chang:1984:ADM

[Cha84a] Der-Shin Chang. On the asymptotic distribution of the multivariate regression estimator. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 247–255, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615962.

Chao:1984:NEN

[Cha84b] Anne Chao. Nonparametric estimation of the number of classes in a population. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 265–270, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615964.

Chagny:2015:AWK

[Cha15]  Gaëlle Chagny.  Adap-

tive warped kernel estimators. Scandinavian Journal of Statistics. Theory and Applications, 42(2): 336–360, June 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cheng:1991:EUJ

[Che91] K. F. Cheng. M-estimator using jack-knife pseudovalues. Scandinavian Journal of Statistics. Theory and Applications, 18(1):51–61, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616188.

Chen:2009:EIB

[Che09] Hua Yun Chen. Estimation and inference based on Neumann series approximation to locally efficient score in missing data problems. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 713–734, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000348.

Cheng:2013:HMI

[Che13] Guang Cheng. How many iterations are sufficient for efficient semiparametric estimation? Scandinavian Journal of Statistics. Theory and Applications, 40(3): 592–618, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586688.

Chen:2015:STS

[Che15a] Yining Chen. Semiparametric time series models with log-concave innovations: Maximum likelihood estimation and its consistency. Scandinavian Journal of Statistics. Theory and Applications, 42(1):1– 31, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cheng:2015:MCE

[Che15b] Guang Cheng. Moment consistency of the exchangeably weighted bootstrap for semiparametric M-estimation. Scandinavian Journal of Statistics. Theory and Applications, 42(3):665–684, September 2015. CODEN SJSADG.

ISSN 0303-6898 (print), 1467-9469 (electronic).

Chang:2023:SLM

[CHI23] Chih-Hao Chang, Hsin-Cheng Huang, and Ching-Kang Ing. Selection of linear mixed-effects models for clustered data. Scandinavian Journal of Statistics. Theory and Applications, 50(2):875–897, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Christiansen:1974:DFR

[Chr74] H. Dalgas Christiansen.

Distribution-free reliability analysis of two-component systems by means of Rasch’s item analysis model. Scandinavian Journal of Statistics. Theory and Applications, 1(1):33–35, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615547.

Christensen:1989:SPP

[Chr89] E. Susanne Christensen.

Statistical properties of Iprojections within exponential families. Scandinavian Journal of Statistics. Theory and Applications, 16(4):307–318, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616143.

Chang:2007:NPM

[CHW^{+}07] I-Shou Chang, Chao A. Hsiung, Chi-Chung Wen, Yuh-Jenn Wu, and Che-Chi Yang. Non-parametric maximum-likelihood estimation in a semiparametric mixture model for competing-risks data. Scandinavian Journal of Statistics. Theory and Applications, 34(4):870–895, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548585.

Chang:2005:BSA

[CHWY05] I-Shou Chang, Chao A. Hsiung, Yuh-Jenn Wu, and Che-Chi Yang. Bayesian survival analysis using Bernstein polynomials. Scandinavian Journal of Statistics. Theory and Applications, 32(3):447–466, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616889.

Critchley:2008:AGA

[CJ08] Frank Critchley and M. C. Jones. Asymmetry and gradient asymmetry functions: Density-based skewness and kurtosis. Scandinavian Journal of Statistics. Theory and Applications, 35(3): 415–437, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000280.

Commenges:2007:CBS

[CJGPL07] Daniel Commenges, Pierre Joly, Anne Gégout-Petit, and Benoit Liquet. Choice between semi-parametric estimators of Markov and non-Markov multi-state models from coarsened observations. Scandinavian Journal of Statistics. Theory and Applications, 34(1): 33–52, March 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548537.

Christensen:1994:LCS

[CK94] Bent Jesper Christensen and Nicholas M. Kiefer. Local cuts and separate inference. Scandinavian Journal of Statistics. Theory and Applications, 21(4): 389–401, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616325.

Crowder:1997:STM

[CK97] Martin Crowder and Alan Kimber. A score test for the multivariate Burr and other Weibull mixture distributions. Scandinavian Journal of Statistics. Theory and Applications, 24(3): 419–432, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616465.

Cui:2006:ELC

[CK06] Hengjian Cui and Efang Kong. Empirical likelihood confidence region for parameters in semi-linear errors-in-variables models. Scandinavian Journal of Statistics. Theory and Applications, 33(1):153–168, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616914.

Chikamatsu:2023:BLS

[CK23] Kentaro Chikamatsu and Tatsuya Kubokawa. Benchmarked linear shrinkage prediction in the Fay– Herriot small area model. Scandinavian Journal of Statistics. Theory and Applications, 50(2):572–588, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cheng:2001:CEF

[CL01a] R. C. H. Cheng and W. B.

Liu. The consistency of estimators in finite mixture models. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 603–616, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616684.

Chung:2001:OBS

[CL01b] Kam-Hin Chung and Stephen M. S. Lee. Optimal bootstrap sample size in construction of percentile confidence bounds. Scandinavian Journal of Statistics. Theory and Applications, 28(1):225–239, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616653.

Cao:2005:GFT

[CL05] Ricardo Cao and Gábor Lugosi. Goodness-of-fit tests based on the kernel density estimator. Scandinavian Journal of Statistics. Theory and Applications, 32(4): 599–616, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616900.

Consonni:2012:OBF

[CL12] Guido Consonni and Luca La Rocca. Objective Bayes factors for Gaussian directed acyclic graphical models. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 743–756, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357222.

Conn:2019:OPN

[CL19] Daniel Conn and Gang Li. An oracle property of the Nadaraya–Watson kernel estimator for highdimensional nonparametric regression. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 735–764, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cook:2020:FTS

[CL20] Richard J. Cook and Jerald F. Lawless. Failure time studies with intermittent observation and losses to follow-up. Scandinavian Journal of Statistics. Theory and Applications, 47(4): 1035–1063, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chauvet:2021:IUP

[CL21] Guillaume Chauvet and Ronan Le Gleut. Inference under pivotal sampling: Properties, variance estimation, and application to tesselation for spatial sampling. Scandinavian Journal of Statistics. Theory and Applications, 48 (1):108–131, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Clement:1997:EDP

[Cle97] Emmanuelle Clement. Estimation of diffusion processes by simulated moment methods. Scandinavian Journal of Statistics. Theory and Applications, 24(3): 353–369, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616461.

Consonni:2017:OBC

[CLP17] Guido Consonni, Luca La Rocca, and Stefano Peluso. Objective Bayes covariateadjusted sparse graphical model selection. Scandinavian Journal of Statistics. Theory and Applications, 44(3):741–764, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Camerlenghi:2018:BNI

[CLP18] Federico Camerlenghi, Antonio Lijoi, and Igor Prünster. Bayesian nonparametric inference beyond the Gibbs-type framework. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 1062–1091, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cheng:2019:TET

[CLP^{+}19] Guanghui Cheng, Baisen Liu, Liuhua Peng, Baoxue Zhang, and Shurong Zheng. Testing the equality of two high-dimensional spatial sign covariance matrices. Scandinavian Journal of Statistics. Theory and Applications, 46(1): 257–271, March 2019. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Collilieux:2019:FMA

[CLR19] Xavier Collilieux, Emilie Lebarbier, and Stéphane Robin. A factor model approach for the joint segmentation with between-series correlation. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 686–705, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2016:SQR

[CLSZ16] Xuerong Chen, Yeqian Liu, Jianguo Sun, and Yong Zhou. Semiparametric quantile regression analysis of right-censored and length-biased failure time data with partially linear varying effects. Scandinavian Journal of Statistics. Theory and Applications, 43(4):921–938, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chan:1982:TAP

[CM82] Lai K. Chan and Tak K.

Mak. Two adaptive procedures for the estimation of a linear structural relationship. Scandinavian Journal of Statistics. Theory and Applications, 9(4):223– 228, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615886.

Carvalho:1984:ABT

[CM84a] Maria Lucilia Carvalho and Daniel Muller. Asymptotic bilateral tests for branching processes. Scandinavian Journal of Statistics. Theory and Applications, 11(1):39–44, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615939.

Chan:1984:MLE

[CM84b] Lai K. Chan and Tak K.

Mak. Maximum likelihood estimation in multivariate structural relationships. Scandinavian Journal of Statistics. Theory and Applications, 11(1):45– 50, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615940.

Chan:2001:HRM

[CM01] Lai K. Chan and T. K. Mak. Heteroscedastic regression models and applications to off-line quality control. Scandinavian Journal of Statistics. Theory and Applications, 28(3): 445–454, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616671.

Chiou:2004:QLR

[CM04] Jeng-Min Chiou and Hans-Georg Müller. Quasilikelihood regression with multiple indices and smooth link and variance functions. Scandinavian Journal of Statistics. Theory and Applications, 31(3):367–386, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616836.

Conti:2015:IQF

[CM15] Pier Luigi Conti and Daniela Marella. Inference for quantiles of a finite population: Asymptotic versus resampling results. Scandinavian Journal of Statistics. Theory and Applications, 42(2):545–561, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2017:ADS

[CM17a] Kun Chen and Yanyuan Ma. Analysis of double single index models. Scandinavian Journal of Statistics. Theory and Applications, 44(1):1–20, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See correction [CM17b].

Chen:2017:CNA

[CM17b] Kun Chen and Yanyuan Ma. Correction note to ‘Analysis of Double Single Index Models’. Scandinavian Journal of Statistics. Theory and Applications, 44(3):815–816, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [CM17a].

Cai:2020:EME

[CM20a] Juan-Juan Cai and Eni Musta. Estimation of the marginal expected shortfall under asymptotic independence. Scandinavian Journal of Statistics. Theory and Applications, 47(1):56– 83, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cai:2020:VSE

[CM20b] Liqian Cai and Tapabrata Maiti. Variable selection and estimation for highdimensional spatial autoregressive models. Scandinavian Journal of Statistics. Theory and Applications, 47(2):587–607, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ciuperca:2020:CPD

[CM20c] Gabriela Ciuperca and Matúš Maciak. Changepoint detection in a linear model by adaptive fused quantile method. Scandinavian Journal of Statistics. Theory and Applications, 47(2):425–463, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cornuet:2012:AMI

[CMMR12] Jean-Marie Cornuet, Jean-Michel Marin, Antonietta Mira, and Christian P. Robert. Adaptive multiple importance sampling. Scandinavian Journal of Statistics. Theory and Applications, 39(4):798–812, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357226.

Chacon:2008:BBS

[CMN08] José E. Chacón, Jesús Montanero, and Agustín G. Nogales. Bootstrap bandwidth selection using an hdependent pilot bandwidth. Scandinavian Journal of Statistics. Theory and Applications, 35(1):139–157, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000389.

Coeurjolly:2017:PDL

[CMW17] Jean-François Coeurjolly, Jesper Møller, and Rasmus Waagepetersen. Palm distributions for log Gaussian Cox processes. Scandinavian Journal of Statistics. Theory and Applications, 44(1):192–203, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2024:EAS

[CMY24] Haoyu Chen, Tiantian Mao, and Fan Yang. Estimation of the adjusted standard-deviatile for extreme risks. Scandinavian Journal of Statistics.

Theory and Applications, 51(2):643–671, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2016:CSR

[CN16] Fuqi Chen and Sévérien Nkurunziza. A class of Stein-rules in multivariate regression model with structural changes. Scandinavian Journal of Statistics. Theory and Applications, 43(1):83–102, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Clancy:2007:EBI

[CO07] Damian Clancy and Philip D. O’Neill. Exact Bayesian inference and model selection for stochastic models of epidemics among a community of households. Scandinavian Journal of Statistics. Theory and Applications, 34(2):259–274, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548551.

Corander:2003:LGM

[Cor03]  Jukka Corander.  La-

belled graphical models. Scandinavian Journal of Statistics. Theory and Applications, 30(3):493–508, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616779.

Corander:2023:CEA

[Cor23] Jukka Corander. A conversation with Elja Arjas (Helsinki, November 2021 and March 2022). Scandinavian Journal of Statistics. Theory and Applications, 50(1):3–12, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cheng:1998:AEE

[CP98] Shihong Cheng and Jiazhu Pan. Asymptotic expansions of estimators for the tail index with applications. Scandinavian Journal of Statistics. Theory and Applications, 25(4):717–728, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616534.

Charnigo:2007:SMG

[CP07] Richard Charnigo and Ramani S. Pilla. Semiparametric mixtures of generalized exponential families. Scandinavian Journal of Statistics. Theory and Applications, 34(3):535–551, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548568.

Charlier:2020:MOQ

[CPS20] Isabelle Charlier, Davy Paindaveine, and Jérôme Saracco. Multiple-output quantile regression through optimal quantization. Scandinavian Journal of Statistics. Theory and Applications, 47(1):250–278, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cui:2013:CSA

[CPWZ13] Xia Cui, Heng Peng, Songqiao Wen, and Lixing Zhu. Component selection in the additive regression model. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 491–510, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586683.

Chen:2002:CIB

[CQ02] Song Xi Chen and Yong Song Qin. Confidence intervals based on local linear smoother. Scandinavian Journal of Statistics. Theory and Applications, 29(1):89–99, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616701.

Cai:1998:EED

[CR98] Zongwu Cai and George G.

Roussas. Efficient estimation of a distribution function under quadrant dependence. Scandinavian Journal of Statistics. Theory and Applications, 25 (1):211–224, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616494.

Coeurjolly:2013:FCE

[CR13] Jean-François Coeurjolly and Ege Rubak. Fast covariance estimation for innovations computed from a spatial Gibbs point process. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 669–684, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586693.

Cabras:2012:MPS

[CRCV12] Stefano Cabras, Walter Racugno, María Eugenia Castellanos, and Laura Ventura. A matching prior for the shape parameter of the skew-normal distribution. Scandinavian Journal of Statistics. Theory and Applications, 39(2):

236–247, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679792.

Ciuperca:2003:PML

[CRI03] Gabriela Ciuperca, Andrea Ridolfi, and Jérôme Idier. Penalized maximum likelihood estimator for normal mixtures. Scandinavian Journal of Statistics. Theory and Applications, 30(1):45–59, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616748.

Crowder:1991:ICC

[Cro91]  Martin Crowder.  On

the identifiability crisis in competing risks analysis. Scandinavian Journal of Statistics. Theory and Applications, 18(3):223–233, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616205.

Crowder:1998:MMR

[Cro98] Martin Crowder. A multivariate model for repeated failure time measurements. Scandinavian Journal of Statistics. Theory and Applications, 25(1):

53–67, March 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616484.

Crowder:2000:CCR

[Cro00] Martin Crowder. Characterizations of competing risks in terms of independent-risks proxy models. Scandinavian Journal of Statistics. Theory and Applications, 27(1):57– 64, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616590.

Carter:2024:PCG

[CRS24] Jack Storror Carter, David Rossell, and Jim Q. Smith. Partial correlation graphical LASSO. Scandinavian Journal of Statistics. Theory and Applications, 51(1):32–63, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cox:1982:PMG

[CS82] D. R. Cox and E. Spjøtvoll.

On partitioning means into groups. Scandinavian Journal of Statistics. Theory and Applications, 9(3):147– 152, September 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615870.

Currie:1990:TFU

[CS90] Iain D. Currie and Michael A. Stephens. Tests of fit using sample moments. Scandinavian Journal of Statistics. Theory and Applications, 17(2):147–156, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616162.

Cai:2003:LLE

[CS03] Zongwu Cai and Yanqing Sun. Local linear estimation for time-dependent coefficients in Cox’s regression models. Scandinavian Journal of Statistics. Theory and Applications, 30(1):93–111, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616751.

Cox:1977:RST

[CSJ^{+}77] D. R. Cox, Emil Spjøtvoll, Søren Johansen, Willem R. van Zwet, J. F. Bithell, Ole Barndorff-Nielsen, and M. Keuls. The role of significance tests [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 4(2): 49–70, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615652.

Comte:2014:DEO

[CSS14] Fabienne Comte, Adeline Samson, and Julien J. Stirnemann. Deconvolution estimation of onset of pregnancy with replicate observations. Scandinavian Journal of Statistics. Theory and Applications, 41(2):325–345, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586656.

Cassel:1979:PTF

[CSW79] Claes-Magnus Cassel, Carl-Erik Särndal, and Jan H. Wretman. Prediction theory for finite populations when model-based and design-based principles are combined: With an application to a study of choice of transportation mode across the Öresund Straits. Scandinavian Journal of Statistics. Theory and Applications, 6(3):97– 106, December 1979. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615742.

Chen:2014:NRV

[CTGS14] Ziqi Chen, Man-Lai Tang, Wei Gao, and Ning-Zhong Shi. New robust variable selection methods for linear regression models. Scandinavian Journal of Statistics. Theory and Applications, 41(3):725–741, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586755.

Cao:2013:ESH

[CTYF13] Guanqun Cao, David Todem, Lijian Yang, and Jason P. Fine. Evaluating statistical hypotheses using weakly-identifiable estimating functions. Scandinavian Journal of Statistics. Theory and Applications, 40(2):256–273, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586712.

Cucala:2008:IES

[Cuc08] Lionel Cucala. Intensity estimation for spatial point processes observed with noise. Scandinavian Journal of Statistics. Theory and Applications, 35(2): 322–334, June 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548596.

Currie:1980:DSR

[Cur80a] Iain D. Currie. On the distribution of the studentized range in a single normal sample. Scandinavian Journal of Statistics. Theory and Applications, 7(3): 150–154, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615791.

Currie:1980:UTD

[Cur80b] Iain D. Currie. The upper tail of the distribution of W -exponential. Scandinavian Journal of Statistics. Theory and Applications, 7(3):147–149, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615790.

Consonni:2001:CRN

[CV01] Guido Consonni and Piero Veronese. Conditionally reducible natural exponential families and enriched conjugate priors. Scandinavian Journal of Statistics. Theory and Applications, 28(2):377–406, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616665.

Chauveau:2002:ICH

[CV02] Didier Chauveau and Pierre Vandekerkhove. Improving convergence of the Hastings–Metropolis algorithm with an adaptive proposal. Scandinavian Journal of Statistics. Theory and Applications, 29(1):13– 29, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616696.

Chambaz:2014:ITG

[CV14] Antoine Chambaz and Mark J. Van Der Laan. Inference in targeted groupsequential covariate-adjusted randomized clinical trials. Scandinavian Journal of Statistics. Theory and Applications, 41(1):104–140, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586733.

Cronie:2015:FIS

[CV15] O. Cronie and M. N. M. Van Lieshout. A Jfunction for inhomogeneous spatio-temporal point processes. Scandinavian Journal of Statistics. Theory and Applications, 42(2):

562–579, June 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Courgeau:2022:ATI

[CV22] Valentin Courgeau and Almut E. D. Veraart. Asymptotic theory for the inference of the latent trawl model for extreme values. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1448–1495, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cox:1999:LFM

[CW99] D. R. Cox and Nanny Wermuth. Likelihood factorizations for mixed discrete and continuous variables. Scandinavian Journal of Statistics. Theory and Applications, 26(2): 209–220, June 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616551.

Croux:2016:DAT

[CW16] Christophe Croux and Ines Wilms. Discussion of ‘Asymptotic Theory of Outlier Detection Algorithms for Linear Time Series Regression Models’. Scandinavian Journal of Statistics. Theory and Applications, 43(2): 353–356, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JN16a].

Cai:2019:EEP

[CW19] Quan Cai and Suojin Wang. Efficient estimation in partially linear singleindex models for longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 46 (1):116–141, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chiang:2005:KER

[CWH05] Chin-Tsang Chiang, Mei-Cheng Wang, and Chiung-Yu Huang. Kernel estimation of rate function for recurrent event data. Scandinavian Journal of Statistics. Theory and Applications, 32(1):77– 91, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616862.

Cui:2021:ECP

[CWZ21] Yunwei Cui, Rongning Wu, and Qi Zheng. Estimation of change-point for a class of count time series models. Scandinavian Journal of Statistics. Theory and Applications, 48(4):

1277–1313, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Cui:2023:NRK

[CXW23] Wenquan Cui, Jianjun Xu, and Yuehua Wu. A new reproducing kernel-based nonlinear dimension reduction method for survival data. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1365– 1390, September 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chan:2017:HOC

[CY17a] Kin Wai Chan and Chun Yip Yau. High-order corrected estimator of asymptotic variance with optimal bandwidth. Scandinavian Journal of Statistics. Theory and Applications, 44(4): 866–898, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2017:UIS

[CY17b] Yixin Chen and Weixin Yao. Unified inference for sparse and dense longitudinal data in time-varying coefficient models. Scandinavian Journal of Statistics. Theory and Applications, 44(1):268–284, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chen:2011:LPE

[CYL11] Feng Chen, Paul S. F. Yip, and K. F. Lam. On the local polynomial estimators of the counting process intensity function and its derivatives. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 631–649, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41411142.

Clarke:1993:LAT

[CYM93] Brenton R. Clarke, Geoffrey F. Yeo, and Robin K. Milne. Local asymptotic theory for multiple solutions of likelihood equations, with application to a single ion channel model. Scandinavian Journal of Statistics. Theory and Applications, 20(2):133–146, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616269.

Cai:2024:AIA

[CZL24] Bibi Cai, Enwen Zhu, and Shiqing Ling. Asymptotic inference of the ARMA model with time-functional variance noises. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1230–1258, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Chu:2020:SFE

[CZT20] Chenghao Chu, Ying Zhang, and Wanzhu Tu. Stochastic functional estimates in longitudinal models with interval-censored anchoring events. Scandinavian Journal of Statistics. Theory and Applications, 47(3): 638–661, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dabrowska:1987:NPR

[Dab87] Dorota M. Dabrowska.

Non-parametric regression with censored survival time data. Scandinavian Journal of Statistics. Theory and Applications, 14(3): 181–197, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616062.

Dabrowska:1992:VBC

[Dab92] Dorota M. Dabrowska.

Variable bandwidth conditional Kaplan–Meier estimate. Scandinavian Journal of Statistics. Theory and Applications, 19(4): 351–361, June 1992. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616252.

Dabrowska:1996:WCP

[Dab96] Dorota M. Dabrowska. Weak convergence of a product integral dependence measure. Scandinavian Journal of Statistics. Theory and Applications, 23(4):551–580, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616425.

Daley:1977:TBA

[Dal77] D. J. Daley. Tighter bounds for the absolute third moment. Scandinavian Journal of Statistics. Theory and Applications, 4(4):183– 184, September 1977. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615677.

Damsleth:1975:CCG

[Dam75] Eivind Damsleth. Conjugate classes for gamma distributions. Scandinavian Journal of Statistics. Theory and Applications, 2(2):80–84, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615580.

Damsleth:1980:IMV

[Dam80] Eivind Damsleth. Interpolating missing values in a time series. Scandinavian Journal of Statistics. Theory and Applications, 7 (1):33–39, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615768.

Dass:2003:UCF

[DB03] Sarat C. Dass and James O.

Berger. Unified conditional frequentist and Bayesian testing of composite hypotheses. Scandinavian Journal of Statistics. Theory and Applications, 30 (1):193–210, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616757.

Dhar:2018:TIC

[DBD18] Subhra Sankar Dhar, Wicher Bergsma, and Angelos Dassios. Testing independence of covariates and errors in non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 45(3):421–443, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Descloux:2022:RLZ

[DBJ^{+}22] Pascaline Descloux, Claire Boyer, Julie Josse, Aude Sportisse, and Sylvain Sardy. Robust Lasso-Zero for sparse corruption and model selection with missing covariates. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1605–1635, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Durand:2020:PHF

[DBNR20] Guillermo Durand, Gilles Blanchard, Pierre Neuvial, and Etienne Roquain. Post hoc false positive control for structured hypotheses. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1114–1148, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Datta:2010:IPC

[DBS10] Somnath Datta, Dipankar Bandyopadhyay, and Glen A. Satten. Inverse probability of censoring weighted Ustatistics for right-censored data with an application to testing hypotheses. Scandinavian Journal of Statistics. Theory and Applications, 37(4):680–700, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000415.

Durban:2000:APL

[DC00] Maria Durban and Iain D.

Currie. Adjustment of the profile likelihood for a class of normal regression models. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 535–542, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616621.

delCampo:2017:EGF

[dCCU17] Abraham Martín del Campo, Sarah Cepeda, and Caroline Uhler. Exact goodnessof-fit testing for the Ising model. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 285–306, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

DeCastro:2014:BTM

[DCIK14] Mário De Castro, Ming-Hui Chen, Joseph G. Ibrahim, and John P. Klein. Bayesian transformation models for multivariate survival data. Scandinavian Journal of Statistics. Theory and Applications, 41 (1):187–199, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586737.

deCarvalho:1982:TEC

[dCJV82] A. de Carvalho, B. Jørgensen, and M. Væth. Treatment evaluation in chronical diseases: a non-parametric approach using simple scores. Scandinavian Journal of Statistics. Theory and Applications, 9(3):175–177, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615875.

Dabrowska:1988:PLT

[DD88] Dorota M. Dabrowska and Kjell A. Doksum. Partial likelihood in transformation models with censored data. Scandinavian Journal of Statistics. Theory and Applications, 15(1): 1–23, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616080.

Bernardino:2022:SGR

[DD22] Elena Di Bernardino and Céline Duval. Statistics for Gaussian random fields with unknown location and scale using Lipschitz–Killing curvatures. Scandinavian Journal of Statistics. Theory and Applications, 49(1): 143–184, March 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Denis:2024:NPC

[DDBEMT24] Christophe Denis, Charlotte Dion-Blanc, Eddy Ella-Mintsa, and Viet Chi Tran. Nonparametric plugin classifier for multiclass classification of S.D.E. paths. Scandinavian Journal of Statistics. Theory and Applications, 51 (3):1103–1160, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dewan:2004:TDB

[DDK04] Isha Dewan, J. V. Deshpande, and S. B. Kulathinal. On testing dependence between time to failure and cause of failure via conditional probabilities. Scandinavian Journal of Statistics. Theory and Applications, 31(1):79– 91, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616813.

DelCastillo:2014:MDB

[DDL14] Joan Del Castillo, Jalila Daoudi, and Richard Lockhart. Methods to distinguish between polynomial and exponential tails. Scandinavian Journal of Statistics. Theory and Applications, 41(2):382–393, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586659.

Denis:2020:CPM

[DDM20] Christophe Denis, Charlotte Dion, and Miguel Martinez. Consistent procedures for multiclass classification of discrete diffusion paths. Scandinavian Journal of Statistics. Theory and Applications, 47(2):516–554, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

DeOliveira:2006:OPB

[De 06] Victor De Oliveira. On optimal point and block prediction in log-Gaussian random fields. Scandinavian Journal of Statistics. Theory and Applications, 33(3): 523–540, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616940.

Dahmstrom:1982:EFI

[DE82] Karin Dahmström and Gunnar Eklund. Estimation of the frequency of initially hidden distant metastases in patients with nodular malignant melanoma. Scandinavian Journal of Statistics. Theory and Applications, 9(2): 117–120, June 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615863.

Didelez:2004:CGC

[DE04] Vanessa Didelez and David Edwards. Collapsibility of graphical CG-regression models. Scandinavian Journal of Statistics. Theory and Applications, 31(4): 535–551, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616849.

Drton:2006:MLE

[DE06] Mathias Drton and Michael Eichler. Maximum likelihood estimation in Gaussian chain graph models under the alternative Markov property. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 247–257, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616923.

Degenhardt:1996:SLT

[Deg96] H. J. A. Degenhardt. Strong limit theorems for the difference of the perturbed empirical distribution function and the classical empirical distribution function. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 331–351, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616408.

Dorr:2021:TMN

[DEH21] Philip Dörr, Bruno Ebner, and Norbert Henze. Testing multivariate normality by zeros of the harmonic oscillator in characteristic function spaces. Scandinavian Journal of Statistics. Theory and Applications, 48(2):456–501, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Diaconis:1992:FFT

[DEL92] Persi W. Diaconis, Morris L. Eaton, and Steffen L. Lauritzen. Finite De Finetti theorems in linear models and multivariate analysis. Scandinavian Journal of Statistics. Theory and Applications, 19(4):289–315, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616249.

Demidenko:2017:EAS

[Dem17] Eugene Demidenko. Exact and approximate statistical inference for nonlinear regression and the estimating equation approach. Scandinavian Journal of Statistics. Theory and Applications, 44(3):636–665, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Deschatre:2023:AEI

[Des23] Thomas Deschatre. Adaptive estimation of intensity in a doubly stochastic Poisson process. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1756–1794, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dette:2004:RED

[Det04] Holger Dette. On robust and efficient designs for risk estimation in epidemiological studies. Scandinavian Journal of Statistics. Theory and Applications, 31(3): 319–331, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616833.

DeBacker:2020:LCQ

[DEV20a] Mickaël De Backer, Anouar El Ghouch, and Ingrid Van Keilegom. Linear censored quantile regression: A novel minimum-distance approach. Scandinavian Journal of Statistics. Theory and Applications, 47(4): 1275–1306, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dette:2020:MCP

[DEV20b] Holger Dette, Theresa Eckle, and Mathias Vetter. Multiscale change point detection for dependent data. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1243–1274, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dalenius:1974:SPP

[DF74] Tore Dalenius and Ove Frank. Sampling populations with partial variate patterns. Scandinavian Journal of Statistics. Theory and Applications, 1(1):19–27, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615545.

Diaconis:1990:CEF

[DF90] P. Diaconis and D. A.

Freedman. Cauchy’s equation and De Finetti’s theorem. Scandinavian Journal of Statistics. Theory and Applications, 17(3): 235–249, March 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616171.

Ducharme:2003:QMP

[DF03] Gilles R. Ducharme and Benoît Frichot. Quasi most powerful invariant goodness-of-fit tests. Scandinavian Journal of Statistics. Theory and Applications, 30(2):399–414, June 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616771.

Divino:2000:PPI

[DFG00] Fabio Divino, Arnoldo Frigessi, and Peter J. Green. Penalized pseudolikelihood inference in spatial interaction models with covariates. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 445–458, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616616.

Davidov:2014:SIT

[DFI14] Ori Davidov, Konstantinos Fokianos, and George Iliopoulos. Semiparametric inference for the twoway layout under order restrictions. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 622–638, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586750.

Delattre:2013:MLE

[DGCS13] Maud Delattre, Valentine Genon-Catalot, and Adeline Samson. Maximum likelihood estimation for stochastic differential equations with random effects. Scandinavian Journal of Statistics. Theory and Applications, 40(2):322–343, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586715.

Dostal:2016:FCM

[DGGM16] Lucie Dostál, Siegfried Gabler, Matthias Ganninger, and Ralf Münnich. Frame correction modelling with applications to the German register-assisted Census 2011. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 904–920, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dahl:2007:RCM

[DGN07] Fredrik A. Dahl, Jørund Gåsemyr, and Bent Natvig. A robust conflict measure of inconsistencies in Bayesian hierarchical models. Scandinavian Journal of Statistics. Theory and Applications, 34(4):816–828, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548582.

Datta:2002:ATC

[DGSL02] Gauri Sankar Datta, Malay Ghosh, David Daniel Smith, and Parthasarathi Lahiri. On an asymptotic theory of conditional and unconditional coverage probabilities of empirical Bayes confidence intervals. Scandinavian Journal of Statistics. Theory and Applications, 29(1):139–152, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616705.

Dahmstrom:1978:MSP

[DH78] Per Dahmström and Mats Hagnell. Multivariate stratification of primary sampling units in a general purpose survey. Scandinavian Journal of Statistics. Theory and Applications, 5 (1):53–56, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615683.

Duong:2005:CVB

[DH05] Tarn Duong and Martin L. Hazelton. Cross-validation bandwidth matrices for multivariate kernel density estimation. Scandinavian Journal of Statistics. Theory and Applications, 32(3): 485–506, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616891.

DeBlasi:2007:BSA

[DH07] Pierpaolo De Blasi and Nils Lid Hjort. Bayesian survival analysis in proportional hazard models with logistic relative risk. Scandinavian Journal of Statistics. Theory and Applications, 34(1):229–257, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548548.

Dannemann:2008:LRT

[DH08] Jörn Dannemann and Hajo Holzmann. Likelihood ratio testing for hidden Markov models under non-standard conditions. Scandinavian Journal of Statistics. Theory and Applications, 35(2): 309–321, June 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548595.

Dette:2012:SCC

[DH12] Holger Dette and Cedric Heuchenne. Scale checks in censored regression. Scandinavian Journal of Statistics. Theory and Applications, 39(2):323–339, June 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679797.

Doornik:2016:OMS

[DH16] Jurgen A. Doornik and David F. Hendry. Outliers and model selection: Discussion of the paper by Søren Johansen and Bent Nielsen. Scandinavian Journal of Statistics. Theory and Applications, 43(2):360–365, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

DeOliveira:2023:ARP

[DH23] Victor De Oliveira and Zifei Han. Approximate reference priors for Gaussian random fields. Scandinavian Journal of Statistics. Theory and Applications, 50(1):296–326, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dhar:2016:TMI

[Dha16] Subhra Sankar Dhar. Trimmed mean isotonic regression. Scandinavian Journal of Statistics. Theory and Applications, 43(1):202–212, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Daehlen:2024:ABE

[DHH24] Ingrid Dæhlen, Nils Lid Hjort, and Ingrid Hobæk Haff. Accurate bias estimation with applications to focused model selection. Scandinavian Journal of Statistics. Theory and Applications, 51(2): 724–759, June 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Diaconis:2023:AEF

[Dia23]  Persi Diaconis.  Ap-

proximate exchangeability and de Finetti priors in 2022. Scandinavian Journal of Statistics. Theory and Applications, 50(1):38– 53, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Didelez:2007:GMC

[Did07] Vanessa Didelez. Graphical models for composable finite Markov processes. Scandinavian Journal of Statistics. Theory and Applications, 34(1): 169–185, March 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548545.

Dietz:1992:NMR

[Die92] Hans M. Dietz. A non-Markovian relative of the Ornstein–Uhlenbeck process and some of its local statistical properties. Scandinavian Journal of Statistics. Theory and Applications, 19(4):363–379, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616253.

Diggle:1980:EWF

[DK80] Peter J. Diggle and Hans Knutell. Estimating the width of a forest striproad. Scandinavian Journal of Statistics. Theory and Applications, 7(2):85– 89, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615778.

Dachian:2006:HTP

[DK06] Sergueï Dachian and Yury A. Kutoyants. Hypotheses testing: Poisson versus self-exciting. Scandinavian Journal of Statistics. Theory and Applications, 33(2):391–408, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616931.

Davis:1989:LCR

[DL89] T. P. Davis and A. J. Lawrance. The likelihood for competing risk survival analysis. Scandinavian Journal of Statistics. Theory and Applications, 16 (1):23–28, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616119.

Doukhan:2001:FED

[DL01] Paul Doukhan and Sana Louhichi. Functional estimation of a density under a new weak dependence condition. Scandinavian Journal of Statistics. Theory and Applications, 28(2):325–341, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616662.

Diggle:2006:BGD

[DL06] Peter Diggle and Søren Lophaven. Bayesian geostatistical design. Scandinavian Journal of Statistics. Theory and Applications, 33(1):53–64, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616908.

Dereudre:2014:EIP

[DLH14] David Dereudre, Frédéric Lavancier, and Kateřina Staňková Helisová. Estimation of the intensity parameter of the germ-grain Quermassinteraction model when the number of germs is not observed. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 809–829, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586760.

Delecroix:2008:NCR

[DLP08] Michel Delecroix, Olivier Lopez, and Valentin Patilea. Nonlinear censored regression using synthetic data. Scandinavian Journal of Statistics. Theory and Applications, 35(2):248–265, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548591.

Dette:2018:NNI

[DLR18] Holger Dette, Christophe Ley, and Francisco Rubio. Natural (non-)informative priors for skew-symmetric distributions. Scandinavian Journal of Statistics. Theory and Applications, 45(2):405–420, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Damien:1996:IBN

[DLS96] Paul Damien, Purushottam W. Laud, and Adrian F. M. Smith. Implementation of Bayesian non-parametric inference based on beta processes. Scandinavian Journal of Statistics. Theory and Applications, 23(1):27–36, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616383.

Duan:2024:TMR

[DLS^{+}24] Rui Duan, C. Jason Liang, Pamela A. Shaw, Cheng Yong Tang, and Yong Chen. Testing the missing at random assumption in generalized linear models in the presence of instrumental variables. Scandinavian Journal of Statistics. Theory and Applications, 51(1):334–354, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Diggle:1980:SDS

[DM80] Peter J. Diggle and Bertil Matérn. On sampling designs for the study of point-event nearest neighbour distributions in R^{2}. Scandinavian Journal of Statistics. Theory and Applications, 7(2):80–84, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615777.

Diggle:1983:NBQ

[DM83] Peter J. Diggle and Robin K.

Milne. Negative binomial quadrat counts and point processes. Scandinavian Journal of Statistics. Theory and Applications, 10(4): 257–267, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615928.

Ding:2019:MFC

[DM19] Peng Ding and Luke W. Miratrix. Model-free causal inference of binary experimental data. Scandinavian Journal of Statistics. Theory and Applications, 46 (1):200–214, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dasgupta:2024:GOG

[DMK24] Subhadra Dasgupta, Siuli Mukhopadhyay, and Jonathan Keith. G-optimal grid designs for kriging models. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1061–1085, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dawid:2002:PES

[DMPV02] A. P. Dawid, J. Mortera, V. L. Pascali, and D. Van Boxel. Probabilistic expert systems for forensic inference from genetic markers. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 577–595, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616736.

deMare:1988:PFM

[dMR88] Jacques de Maré and Björn Rosander. The production flow in the motor industry: a case study of multistate reliability theory. Scandinavian Journal of Statistics. Theory and Applications, 15(1):51–53, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616084.

Dawid:2016:MSR

[DMV16] A. Philip Dawid, Monica Musio, and Laura Ventura. Minimum scoring rule inference. Scandinavian Journal of Statistics. Theory and Applications, 43(1):123–138, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ding:2015:EMS

[DN15] Ying Ding and Bin Nan.

Estimating mean survival time: When is it possible? Scandinavian Journal of Statistics. Theory and Applications, 42(2): 397–413, June 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dang:2021:NGC

[DNCZ21] Xin Dang, Dao Nguyen, Yixin Chen, and Junying Zhang. A new Gini correlation between quantitative and qualitative variables. Scandinavian Journal of Statistics. Theory and Applications, 48(4):1314–1343, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

DoHa:2010:BRL

[DNL10] Il Do Ha, Maengseok Noh, and Youngjo Lee. Bias reduction of likelihood estimators in semiparametric frailty models. Scandinavian Journal of Statistics. Theory and Applications, 37(2):307–320, June 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000881.

Demiris:2005:BIE

[DO05] Nikolaos Demiris and Philip D. O’Neill. Bayesian inference for epidemics with two levels of mixing. Scandinavian Journal of Statistics. Theory and Applications, 32(2):265–280, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616877.

Doksum:1975:MLA

[Dok75] Kjell A. Doksum. Measures of location and asymmetry. Scandinavian Journal of Statistics. Theory and Applications, 2(1):11– 22, December 1975. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615571.

Doksum:1980:RTM

[Dok80] Kjell Doksum. Rank tests for the matched pair problem with life distributions. Scandinavian Journal of Statistics. Theory and Applications, 7(2):67– 72, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615775. See correction [Dok82].

Doksum:1982:CNR

[Dok82] Kjell Doksum. Correction note: Rank Tests for the Matched Pair Problem with Life Distributions. Scandinavian Journal of Statistics. Theory and Applications, 9(2):121, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615864.  See

[Dok80].

Doornik:2016:EID

[Doo16] Jurgen A. Doornik. An example of instability: Discussion of the paper by Søren Johansen and Bent Nielsen. Scandinavian Journal of Statistics. Theory and Applications, 43 (2):357–359, June 2016.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Doornik:2018:AES

[Doo18] Jurgen A. Doornik. Accelerated estimation of switching algorithms: The cointegrated VAR model and other applications. Scandinavian Journal of Statistics. Theory and Applications, 45(2):283–300, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Damilano:2004:ELC

[DP04] Gabriela Damilano and Pedro Puig. Efficiency of a linear combination of the median and the sample mean: The double truncated normal distribution. Scandinavian Journal of Statistics. Theory and Applications, 31(4):629–637, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616855.

DeCanditiis:2006:SWD

[DP06] Daniela De Canditiis and Marianna Pensky. Simultaneous wavelet deconvolution in periodic setting. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 293–306, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616926.

Daouia:2013:PTE

[DP13] Abdelaati Daouia and Byeong U. Park. On projection-type estimators of multivariate isotonic functions. Scandinavian Journal of Statistics. Theory and Applications, 40(2): 363–386, June 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586717.

Dvorak:2016:PEI

[DP16] Jiří Dvořák and Michaela Prokešová. Parameter estimation for inhomogeneous space–time shot-noise Cox point processes. Scandinavian Journal of Statistics. Theory and Applications, 43(4):939–961, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dai:2018:JMS

[DP18] Hongsheng Dai and Jianxin Pan. Joint modelling of survival and longitudinal data with informative observation times. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 571–589, September 2018.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dette:2009:GFT

[DPFV09] Holger Dette, Juan Carlos Pardo-Fernández, and Ingrid Van Keilegom. Goodnessof-fit tests for multiplicative models with dependent data. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 782–799, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000352.

DiMarzio:2013:NPR

[DPT13] Marco Di Marzio, Agnese Panzera, and Charles C. Taylor. Non-parametric regression for circular responses. Scandinavian Journal of Statistics. Theory and Applications, 40(2): 238–255, June 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586711.

Dette:2006:EIV

[DPV06] Holger Dette, Mark Podolskij, and Mathias Vetter. Estimation of integrated volatility in continuoustime financial models with applications to goodnessof-fit testing. Scandinavian Journal of Statistics.

Theory and Applications, 33(2):259–278, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616924.

Dumitrescu:2021:ILD

[DQR21] Laura Dumitrescu, Wei Qian, and J. N. K. Rao. Inference for longitudinal data from complex sampling surveys: An approach based on quadratic inference functions. Scandinavian Journal of Statistics. Theory and Applications, 48(1):246–274, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Du:2023:TVM

[DQYZ23] Yuqing Du, Lianqiang Qu, Ting Yan, and Yuan Zhang. Time-varying β-model for dynamic directed networks. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1687–1715, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dette:1996:OPD

[DR96] Holger Dette and Ingo Röder. Optimal product designs for multivariate regression with missing terms. Scandinavian Journal of Statistics. Theory and Applications, 23(2): 195–208, June 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616396.

Dunsmore:1997:PSP

[DR97] Ian R. Dunsmore and David J. Robson. Prediction in some Poisson errors in variables models. Scandinavian Journal of Statistics. Theory and Applications, 24(4):543–554, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616475.

Draghici:2000:NAC

[DR00] L. Drăghici and R. V. Ramamoorthi. A note on the absolute continuity and singularity of Polya tree priors and posteriors. Scandinavian Journal of Statistics. Theory and Applications, 27(2):299–303, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616604.

Durot:2010:GFT

[DR10] CÉCile Durot and Laurence Reboul. Goodnessof-fit test for monotone functions. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 422–441, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000399.

Diallo:2018:SAE

[DR18] Mamadou S. Diallo and J. N. K. Rao. Small area estimation of complex parameters under unitlevel models with skewnormal errors. Scandinavian Journal of Statistics. Theory and Applications, 45(4):1092–1116, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Deleamont:2022:RIC

[DR22] Pierre-Yves Deléamont and Elvezio Ronchetti. Robust inference with censored survival data. Scandinavian Journal of Statistics. Theory and Applications, 49(4): 1496–1533, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Drees:1998:SST

[Dre98] Holger Drees. On smooth statistical tail functionals. Scandinavian Journal of Statistics. Theory and Applications, 25(1):187–210, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616493.

Dey:1996:CVP

[DRM96] A. K. Dey, F. H. Ruymgaart, and B. A. Mair. Cross-validation for parameter selection in inverse estimation problems. Scandinavian Journal of Statistics. Theory and Applications, 23(4):609–620, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616428.

DiSerio:2009:SPS

[DRS09] Clelia Di Serio, Yosef Rinott, and Marco Scarsini. Simpson’s Paradox in survival models. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 463–480, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000333.

Sousa:2019:IIL

[dRSHK19] Thiago do Rêgo Sousa, Stephan Haug, and Claudia Klüppelberg. Indirect inference for Lévy-driven continuous-time GARCH models. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 765–801, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sousa:2022:MBE

[dRSS22] Thiago do Rêgo Sousa and Robert Stelzer. Momentbased estimation for the multivariate COGARCH(1,1) process. Scandinavian Journal of Statistics. Theory and Applications, 49 (2):681–717, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dehling:2013:NPC

[DRT13] Herold Dehling, Aeneas Rooch, and Murad S. Taqqu. Non-parametric change-point tests for longrange dependent data. Scandinavian Journal of Statistics. Theory and Applications, 40(1):153–173, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357258.

Dryden:2014:CGM

[Dry14]  Ian L. Dryden.  Com-

ment on Geodesic Monte Carlo on Embedded Manifolds by Byrne and Girolami. Scandinavian Journal of Statistics. Theory and Applications, 41(1):8– 9, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

24586722. See [BG13].

Deshpande:1990:NMA

[DS90] Jayant V. Deshpande and R. P. Suresh. Nonmonotonic ageing. Scandinavian Journal of Statistics. Theory and Applications, 17(3):257–262, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616173.

Diciccio:1994:CAS

[DS94] Thomas J. Diciccio and Steven E. Stern. Constructing approximately standard normal pivots from signed roots of adjusted likelihood ratio statistics. Scandinavian Journal of Statistics. Theory and Applications, 21(4):447–460, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616329.

DeLuna:2003:CMS

[DS03a] Xavier De Luna and Kostas Skouras. Choosing a model selection strategy. Scandinavian Journal of Statistics. Theory and Applications, 30(1):113–128, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616752.

Dette:2003:NST

[DS03b] Holger Dette and Ingrid Spreckelsen. A note on a specification test for time series models based on spectral density estimation. Scandinavian Journal of Statistics. Theory and Applications, 30(3):481–491, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616778.

Ditlevsen:2004:IOI

[DS04] Susanne Ditlevsen and Michael Sørensen. Inference for observations of integrated diffusion processes. Scandinavian Journal of Statistics. Theory and Applications, 31(3):417–429, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616839.

Dauxois:2009:NPT

[DS09] Jean-Yves Dauxois and Sophie Sencey. Nonparametric tests for recurrent events under competing risks. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 649–670, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000345.

Diao:2014:ACI

[DSD^{+}14] Lixia Diao, David D. Smith, Gauri Sankar Datta, Tapabrata Maiti, and Jean D. Opsomer. Accurate confidence interval estimation of small area parameters under the Fay– Herriot model. Scandinavian Journal of Statistics. Theory and Applications, 41(2):497–515, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586665.

Diaconis:2014:CED

[DSH14] Persi Diaconis, Christof Seiler, and Susan Holmes. Connections and extensions: a discussion of the paper by Girolami and Byrne. Scandinavian Journal of Statistics. Theory and Applications, 41(1):3– 7, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586721.

Dean:2014:PEH

[DSJP14] Thomas A. Dean, Sumeetpal S. Singh, Ajay Jasra, and Gareth W. Peters. Parameter estimation for hidden Markov models with intractable likelihoods. Scandinavian Journal of Statistics. Theory and Applications, 41(4):970–987, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586816.

Dette:2013:CBN

[DSS13] Holger Dette, Karl F.

Siburg, and Pavel A. Stoimenov. A copulabased non-parametric measure of regression dependence. Scandinavian Journal of Statistics. Theory and Applications, 40(1):21– 41, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357251.

Ditlevsen:2014:SPD

[DSS14a] Susanne Ditlevsen, Thomas Scheike, and Michael Søensen. Six papers on dynamic statistical models. Scandinavian Journal of Statistics. Theory and Applications, 41(1):22, March 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ditlevsen:2014:ESP

[DSS14b] Susanne Ditlevsen, Thomas Scheike, and Michael Sørensen. Editorial: Six papers on dynamic statistical models. Scandinavian Journal of Statistics. Theory and Applications, 41(1): 22, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586728.

Ding:2009:MRA

[DSWH09] A. Adam Ding, Guangkai Shi, Weijing Wang, and Jin-Jian Hsieh. Marginal regression analysis for semicompeting risks data under dependent censoring. Scandinavian Journal of Statistics. Theory and Applications, 36(3):481–500, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000334.

Dumbgen:2005:BPS

[DT05] Lutz Dümbgen and David E. Tyler. On the breakdown properties of some multivariate M-functionals*. Scandinavian Journal of Statistics. Theory and Applications, 32(2):247–264, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616876.

Dobler:2020:DIN

[DT20] Dennis Dobler and Andrew Titman. Dynamic inference for non-Markov transition probabilities under random right censoring. Scandinavian Journal of Statistics. Theory and Applications, 47(2):572–586, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dette:1995:RMF

[DW95] Holger Dette and Weng Kee Wong. Recurrence moment formulas for D-optimal designs. Scandinavian Journal of Statistics. Theory and Applications, 22(4): 505–512, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616377.

Diebolt:1997:NPT

[DW97] Jean Diebolt and Joseph Ngatchou Wandji. A non-parametric test for generalized firstorder autoregressive models. Scandinavian Journal of Statistics. Theory and Applications, 24(2): 241–259, June 1997. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616451.

Damien:2002:BNP

[DW02] Paul Damien and Stephen Walker. A Bayesian nonparametric comparison of two treatments. Scandinavian Journal of Statistics. Theory and Applications, 29(1):51–56, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616698.

Drton:2016:GIL

[DW16] Mathias Drton and Luca Weihs. Generic identifiability of linear structural equation models by ancestor decomposition. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1035–1045, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Dette:2011:CCQ

[DWV11] Holger Dette, Jens Wagener, and Stanislav Volgushev. Comparing conditional quantile curves. Scandinavian Journal of Statistics. Theory and Applications, 38(1):63–88, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015384.

Duan:2017:EAE

[DY17] Xiaogang Duan and Guosheng Yin. Ensemble approaches to estimating the population mean with missing response. Scandinavian Journal of Statistics. Theory and Applications, 44(4): 899–917, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

DaSilva:2021:CIM

[DZ21] Damião N. Da Silva and Li-Chun Zhang. A calibrated imputation method for secondary data analysis of survey data. Scandinavian Journal of Statistics. Theory and Applications, 48(1):25–41, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ehlers:2008:APC

[EB08] Ricardo S. Ehlers and Stephen P. Brooks. Adaptive proposal construction for reversible jump MCMC. Scandinavian Journal of Statistics. Theory and Applications, 35(4):677–690, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000296.

Emmenegger:2023:PML

[EB23] Corinne Emmenegger and Peter Bühlmann. Plug-in machine learning for partially linear mixed-effects models with repeated measurements. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1553–1567, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Escobar-Bach:2018:LEC

[EBGG18] Mikael Escobar-Bach, Yuri Goegebeur, and Armelle Guillou. Local estimation of the conditional stable tail dependence function. Scandinavian Journal of Statistics. Theory and Applications, 45(3):590–617, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Edwards:1978:CAT

[Edw78] A. W. F. Edwards. Commentary on the arguments of Thomas Bayes. Scandinavian Journal of Statistics. Theory and Applications, 5(2):116–118, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615697.

Edwards:1980:LST

[Edw80] David G. Edwards. Large sample tests for stationarity and reversibility in finite Markov chains. Scandinavian Journal of Statistics. Theory and Applications, 7(4):203–206, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615801.

Efromovich:2005:SBW

[Efr05] Sam Efromovich. A study of blockwise wavelet estimates via lower bounds for a spike function. Scandinavian Journal of Statistics. Theory and Applications, 32(1):133–158, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616865.

Efromovich:2008:OSD

[Efr08] Sam Efromovich. Optimal sequential design in a controlled non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 35(2): 266–285, June 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548592.

Efromovich:2016:ESD

[Efr16] Sam Efromovich. Estimation of the spectral density with assigned risk. Scandinavian Journal of Statistics. Theory and Applications, 43(1):70–82, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Eichelsbacher:2002:MDB

[EG02] Peter Eichelsbacher and Ayalvadi Ganesh. Moderate deviations for Bayes posteriors. Scandinavian Journal of Statistics. Theory and Applications, 29 (1):153–167, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616706.

ElGhouch:2013:MDP

[EGB13] Anouar El Ghouch, Marc G. Genton, and Taoufik Bouezmarni. Measuring the discrepancy of a parametric model via local polynomial smoothing. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 455–470, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586681.

Egeland:1992:MDS

[Ege92] Thore Egeland. Mixing dependence and system reliability. Scandinavian Journal of Statistics. Theory and Applications, 19(2): 173–183, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616236.

ElMethni:2014:NPE

[EGG14] Jonathan El Methni, Laurent Gardes, and Stéphane Girard. Non-parametric estimation of extreme risk measures from conditional heavy-tailed distributions. Scandinavian Journal of Statistics. Theory and Applications, 41(4):988–1012, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586817.

Eerola:2003:JMR

[EGM^{+}03] Mervi Eerola, Dario Gasbarra, P. Helena Mäkelä, Henri Linden, and Andrei Andreev. Joint modelling of recurrent infections and antibody response by Bayesian data augmentation. Scandinavian Journal of Statistics. Theory and Applications, 30(4): 677–698, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616796.

Efstathiou:1998:LAN

[EGPS98] M. Efstathiou, E. Gutiérrez-Peña, and A. F. M. Smith. Laplace approximations for natural exponential families with cuts. Scandinavian Journal of Statistics. Theory and Applications, 25(1):77–92, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616486.

Englund:1988:REL

[EHR88] Jan-Eric Englund, Ulla Holst, and David Ruppert. Recursive M-estimators of location and scale for dependent sequences. Scandinavian Journal of Statistics. Theory and Applications, 15(2):147–159, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616095.

Eie:1983:GBP

[Eie83] Birgit Eie. The generalized Bessel process corresponding to an Ornstein– Uhlenbeck process. Scandinavian Journal of Statistics. Theory and Applications, 10(3):247–250, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615922.

Eberl:2022:EBM

[EK22] Andreas Eberl and Bernhard Klar. Expectilebased measures of skewness. Scandinavian Journal of Statistics. Theory and Applications, 49(1): 373–399, March 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ekstrom:2001:CGM

[Eks01] Magnus Ekström. Consistency of generalized maximum spacing estimates. Scandinavian Journal of Statistics. Theory and Applications, 28(2):343–354, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616663.

Ekstrom:2008:SVE

[Eks08] Magnus Ekström. Subsampling variance estimation for non-stationary spatial lattice data. Scandinavian Journal of Statistics. Theory and Applications, 35(1):38–63, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000384.

Ekstrom:2013:PPT

[Eks13] Magnus Ekström. Powerful parametric tests based on sum-functions of spacings. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 886–898, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586705.

Eggermont:1996:SEB

[EL96] P. P. B. Eggermont and V. N. Lariccia. A simple and effective bandwidth selector for kernel density estimation. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 285–301, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616405.

Esteban:2022:EBP

[ELLV^{+}22] María Dolores Esteban, María José Lombardía, Esther López-Vizcaíno, Domingo Morales, and Agustín Pérez. Empirical best prediction of small area bivariate parameters. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1699–1727, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Erlemann:2022:CMT

[ELY22] Rasmus Erlemann, Richard Lockhart, and Rihan Yao. Cramér–von Mises tests for change points. Scandinavian Journal of Statistics. Theory and Applications, 49(2):802–830, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Egeland:2002:SGG

[EM02] Thore Egeland and Petter F. Mostad. Statistical genetics and genetical statistics: a forensic perspective*. Scandinavian Journal of Statistics. Theory and Applications, 29(2):297–307, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616716.

Eidsvik:2009:ABI

[EMR09] Jo Eidsvik, Sara Martino, and Håvard Rue. Approximate Bayesian inference in spatial generalized linear mixed models. Scandinavian Journal of Statistics. Theory and Applications, 36(1):1–22, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000304.

Eriksson:2015:CSD

[EMS15] Frank Eriksson, Torben Martinussen, and Thomas H. Scheike. Clustered survival data with left-truncation. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1149–1166, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Englund:1980:RTE

[Eng80] Gunnar Englund. Remainder term estimates for the asymptotic normality of order statistics. Scandinavian Journal of Statistics. Theory and Applications, 7(4): 197–202, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615800.

Erto:2015:SEB

[EPM15] Pasquale Erto, Giuliana Pallotta, and Christina M. Mastrangelo. A semiempirical Bayesian chart to monitor Weibull percentiles. Scandinavian Journal of Statistics. Theory and Applications, 42(3): 701–712, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Erhardsson:2008:NPB

[Erh08] Torkel Erhardsson. Nonparametric Bayesian inference for integrals with respect to an unknown finite measure. Scandinavian Journal of Statistics. Theory and Applications, 35(2):369–384, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548599.

Eriksson:1978:AVC

[Eri78] Bo Eriksson. An approximation of the variance of counts for a stationary stochastic point process. Scandinavian Journal of Statistics. Theory and Applications, 5(2):111–115, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615696.

Eriksen:1984:KET

[Eri84] Poul Svante Eriksen. (k, 1) exponential transformation models [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 11(3):129–145, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615951.

Eriksen:1996:TCS

[Eri96] P. Svante Eriksen. Tests in covariance selection models. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 275–284, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616404.

Eriksson:2004:SCA

[Eri04] Kimmo Eriksson. Statistical and combinatorial aspects of comparative genomics*. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 203–216, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616824.

Erlandsen:1981:AHM

[Erl81] Mogens Erlandsen. Affine hypotheses in the mean and in the covariance in multivariate normal families. Scandinavian Journal of Statistics. Theory and Applications, 8(1):10– 16, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615807.

Eubank:1991:BRT

[ES91] R. L. Eubank and P. L.

Speckman. A bias reduction theorem with applications in nonparametric regression. Scandinavian Journal of Statistics. Theory and Applications, 18(3): 211–222, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616204.

Efromovich:2000:AEI

[ES00] Sam Efromovich and Alexander Samarov. Adaptive estimation of the integral of squared regression derivatives. Scandinavian Journal of Statistics. Theory and Applications, 27(2): 335–351, June 2000. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616607.

Esseen:1975:BAT

[Ess75] Carl-Gustav Esseen. Bounds for the absolute third moment. Scandinavian Journal of Statistics. Theory and Applications, 2(3):149– 152, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615595.

Eguchi:2021:STM

[EU21] Shoichi Eguchi and Yuma Uehara. Schwartz-type model selection for ergodic stochastic differential equation models. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 950–968, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Eubank:2000:TNE

[Eub00] R. L. Eubank. Testing for no effect by cosine series methods. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 747–763, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616639.

ElGhouch:2008:NPR

[EV08] Anouar El Ghouch and Ingrid Van Keilegom. Nonparametric regression with dependent censored data. Scandinavian Journal of Statistics. Theory and Applications, 35(2):228–247, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548590.

Evans:2016:GMB

[Eva16] Robin J. Evans. Graphs for margins of Bayesian networks. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 625–648, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Echelard:2015:SEC

[EVP15] Antoine Echelard, Jacques Lévy Véhel, and Anne Philippe. Statistical estimation for a class of self-regulating processes. Scandinavian Journal of Statistics. Theory and Applications, 42 (2):485–503, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Eubank:1994:CRN

[EW94] R. L. Eubank and Suojin Wang. Confidence regions in non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 21(2): 147–158, June 1994. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616306.

Fang:2019:SSI

[Fan19] Yixin Fang. Scalable statistical inference for averaged implicit stochastic gradient descent. Scandinavian Journal of Statistics. Theory and Applications, 46(4):

987–1002, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Farcomeni:2007:SRC

[Far07] Alessio Farcomeni. Some results on the control of the false discovery rate under dependence. Scandinavian Journal of Statistics. Theory and Applications, 34(2):275–297, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548552.

Farcomeni:2009:GAC

[Far09] Alessio Farcomeni. Generalized augmentation to control the false discovery exceedance in multiple testing. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 501–517, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000335.

Farcomeni:2015:GLM

[Far15] Alessio Farcomeni. Generalized linear mixed models based on latent Markov heterogeneity structures. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1127–1135, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fasen:2016:DEH

[Fas16]  Vicky Fasen.  Depen-

dence estimation for highfrequency sampled multivariate CARMA models. Scandinavian Journal of Statistics. Theory and Applications, 43(1):292–320, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fernando:2020:CIV

[FB20] Mahesh N. Fernando and Ronald W. Butler. Confidence intervals for variance component ratios in unbalanced linear mixed models. Scandinavian Journal of Statistics. Theory and Applications, 47(3):817–838, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fernholz:1991:ASC

[Fer91] Luisa Turrin Fernholz. Almost sure convergence of smoothed empirical distribution functions. Scandinavian Journal of Statistics. Theory and Applications, 18(3):255–262, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616208.

Foster:1996:SAP

[FG96] Dean P. Foster and Edward I. George. A simple ancillarity paradox. Scandinavian Journal of Statistics. Theory and Applications, 23(2):233–242, June 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616399.

Finner:2012:FDR

[FGD12] Helmut Finner, Veronika Gontscharuk, and Thorsten Dickhaus. False discovery rate control of stepup-down tests with special emphasis on the asymptotically optimal rejection curve. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 382–397, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679800.

Feron:2020:EVE

[FGH20] Olivier Féron, Pierre Gruet, and Marc Hoffmann. Efficient volatility estimation in a two-factor model. Scandinavian Journal of Statistics. Theory and Applications, 47(3):862–898, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fujimori:2023:SPC

[FGLT23] Kou Fujimori, Yuichi Goto, Yan Liu, and Masanobu Taniguchi. Sparse principal component analysis for high-dimensional stationary time series. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1953–1983, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fan:2023:SCB

[FGY23] Yiwei Fan, Jiaqi Gu, and Guosheng Yin. Sparse concordance-based ordinal classification. Scandinavian Journal of Statistics. Theory and Applications, 50(3): 934–961, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fung:2004:IDM

[FH04] Wing K. Fung and Yue-Qing Hu. Interpreting DNA mixtures with related contributors in subdivided populations. Scandinavian Journal of Statistics. Theory and Applications, 31 (1):115–130, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616815.

Fan:2019:SLA

[FHSZ19] Yanqin Fan, Ming He, Liangjun Su, and Xiao-Hua Zhou. A smoothed Qlearning algorithm for estimating optimal dynamic treatment regimes. Scandinavian Journal of Statistics. Theory and Applications, 46(2):446–469, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fan:1994:RNP

[FHT94] Jianqing Fan, Tien-Chung Hu, and Young K. Truong. Robust non-parametric function estimation. Scandinavian Journal of Statistics. Theory and Applications, 21(4):433–446, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616328.

Facevicova:2016:CTA

[FHTT16] Kamila Fačevicová, Karel Hron, Valentin Todorov, and Matthias Templ. Compositional tables analysis in coordinates. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 962–977, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Facevicova:2018:GAC

[FHTT18] Kamila Fačevicová, Karel Hron, Valentin Todorov, and Matthias Templ. General approach to coordinate representation of compositional tables. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 879–899, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fu:2024:SEE

[FHY24] Xinyu Fu, Mian Huang, and Weixin Yao. Semiparametric efficient estimation in high-dimensional partial linear regression models. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1259–1287, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fan:1998:ALS

[FK98] Jianqing Fan and Eva Kreutzberger. Automatic local smoothing for spectral density estimation. Scandinavian Journal of Statistics. Theory and Applications, 25(2):359–369, June 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616507.

Fushiki:2004:PBB

[FKA04] Tadayoshi Fushiki, Fumiyasu Komaki, and Kazuyuki Aihara. On parametric bootstrapping and Bayesian prediction. Scandinavian Journal of Statistics. Theory and Applications, 31(3):403–416, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616838.

Figueroa-Lopez:2011:CLT

[FL11] José E. Figueroa-López. Central limit theorems for the non-parametric estimation of time-changed Lévy models. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 748–765, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41411148.

Fang:2005:MLE

[FLS05] Hong-Bin Fang, Gang Li, and Jianguo Sun. Maximum likelihood estimation in a semiparametric logistic/proportional-hazards mixture model. Scandinavian Journal of Statistics. Theory and Applications, 32(1):59–75, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616861.

Fraser:1989:MPD

[FM89] D. A. S. Fraser and H. Massam. A mixed primaldual bases algorithm for regression under inequality constraints. application to concave regression. Scandinavian Journal of Statistics. Theory and Applications, 16(1):65–74, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616122.

Franz:1990:AES

[FM90] Jürgen Franz and Ryszard Magiera. Admissible estimators in sequential plans for exponential-type processes. Scandinavian Journal of Statistics. Theory and Applications, 17(4): 275–285, March 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616178.

Fan:2022:CDR

[FM22] Jianing Fan and Hans-Georg Müller. Conditional distribution regression for functional responses. Scandinavian Journal of Statistics. Theory and Applications, 49(2):502–524, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Favre-Martinoz:2016:RIT

[FMHB16] Cyril Favre-Martinoz, David Haziza, and Jean-François Beaumont. Robust inference in two-phase sampling designs with application to unit nonresponse. Scandinavian Journal of Statistics. Theory and Applications, 43(4):1019–1034, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Forman:2011:GFB

[FMS11] Julie L. Forman, Bo Markussen, and Helle Sørensen. Goodnessof-fit based on downsampling with applications to linear drift diffusions. Scandinavian Journal of Statistics. Theory and Applications, 38(2):288–310, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015495.

Feng:2015:RAC

[FMS15] Yanqin Feng, Ling Ma, and Jianguo Sun. Regression analysis of current status data under the additive hazards model with auxiliary covariates. Scandinavian Journal of Statistics.

Theory and Applications, 42(1):118–136, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fougeres:2009:MDE

[FNR09] Anne-Laure Fougères, John P.

Nolan, and Holger Rootzén. Models for dependent extremes using stable mixtures. Scandinavian Journal of Statistics. Theory and Applications, 36(1):42– 59, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000306.

Fokianos:2001:TPR

[Fok01]  Konstantinos  Fokianos.

Truncated Poisson regression for time series of counts. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 645–659, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616687.

Fischer:2014:GMW

[FOS^{+}14] Daniel Fischer, Hannu Oja, Johanna Schleutker, Pranab K. Sen, and Tiina Wahlfors. Generalized Mann–Whitney type tests for microarray experiments. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 672–692, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586753.

Favaro:2011:CRP

[FPW11] Stefano Favaro, Igor Prünster, and Stephen G. Walker. On a class of random probability measures with general predictive structure. Scandinavian Journal of Statistics. Theory and Applications, 38(2): 359–376, June 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015499.

Finner:2000:CVB

[FR00] H. Finner and M. Roters. On the critical value behaviour of multiple decision procedures. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 563–573, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616623.

Fernandez:2021:RKH

[FR21] Tamara Fernández and Nicolás Rivera. A reproducing kernel Hilbert space log-rank test for the twosample problem. Scandinavian Journal of Statistics. Theory and Applications, 48(4):1384–1432, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Frank:1977:EGT

[Fra77a] Ove Frank. Estimation of graph totals. Scandinavian Journal of Statistics. Theory and Applications, 4(2): 81–89, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615654.

Frank:1977:NBS

[Fra77b]  Ove Frank.  A note

on Bernoulli sampling in graphs and Horvitz– Thompson estimation. Scandinavian Journal of Statistics. Theory and Applications, 4(4):178–180, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615675.

Frank:1978:ENC

[Fra78] Ove Frank. Estimation of the number of connected components in a graph by using a sampled subgraph. Scandinavian Journal of Statistics. Theory and Applications, 5(4):177– 188, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615713.

Frees:1989:IOS

[Fre89] Edward W. Frees. Infinite order U-statistics. Scandinavian Journal of Statistics. Theory and Applications, 16(1):29–45, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616120.

Fernandez:1999:LEE

[FRS99] Miguel A. Fernández, Cristina Rueda, and Bonifacio Salvador. The loss of efficiency estimating linear functions under restrictions. Scandinavian Journal of Statistics. Theory and Applications, 26(4):579–592, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616580.

Frydenberg:1990:CGM

[Fry90] Morten Frydenberg. The chain graph Markov property. Scandinavian Journal of Statistics. Theory and Applications, 17(4): 333–353, March 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616181.

Franke:2016:MMD

[FRZ16] Jürgen Franke, Claudia Redenbach, and Na Zhang. On a mixture model for directional data on the sphere. Scandinavian Journal of Statistics. Theory and Applications, 43(1): 139–155, March 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Forman:2008:PDC

[FS08] Julie Lyng Forman and Michael Sørensen. The Pearson diffusions: a class of statistically tractable diffusion processes. Scandinavian Journal of Statistics. Theory and Applications, 35(3):438–465, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000281.

Favetto:2010:PEB

[FS10] Benjamin Favetto and Adeline Samson. Parameter estimation for a bidimensional partially observed Ornstein–Uhlenbeck process with biological application. Scandinavian Journal of Statistics. Theory and Applications, 37(2): 200–220, June 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000875.

Florens:2012:RPL

[FS12] Jean-Pierre Florens and Anna Simoni. Regularized posteriors in linear ill-posed inverse problems. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 214–235, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679791.

Fuentes-Santos:2016:CSB

[FSGMM16] Isabel Fuentes-Santos, Wenceslao González-Manteiga, and Jorge Mateu. Consistent smooth bootstrap kernel intensity estimation for inhomogeneous spatial Poisson point processes. Scandinavian Journal of Statistics. Theory and Applications, 43(2): 416–435, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fremdt:2013:TEC

[FSHK13] Stefan Fremdt, Josef G. Steinebach, Lajos Horváth, and Piotr Kokoszka. Testing the equality of covariance operators in functional samples. Scandinavian Journal of Statistics. Theory and Applications, 40(1):138–152, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357257.

Fabrizi:2016:BCM

[FT16] Enrico Fabrizi and Carlo Trivisano. Bayesian conditional mean estimation in log-normal linear regression models with finite quadratic expected loss. Scandinavian Journal of Statistics. Theory and Applications, 43(4):1064–1077, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ferre:2006:MPF

[FV06] Louis Ferré and Nathalie Villa. Multilayer perceptron with functional inputs: an inverse regression approach. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 807–823, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616959.

Ferraty:2010:VBN

[FVV10] Frédéric Ferraty, Ingrid Van Keilegom, and Philippe Vieu. On the validity of the bootstrap in nonparametric functional regression. Scandinavian Journal of Statistics. Theory and Applications, 37(2): 286–306, June 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000880.

Ferrari:2003:EGS

[FW03] Fiorenzo Ferrari and Abraham Wyner. Estimation of general stationary processes by variable length Markov chains. Scandinavian Journal of Statistics. Theory and Applications, 30(3):459–480, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616777.

Fwu:1985:HP

[Fwu85] Chih-Chy Fwu. Homogeneous parametrization. Scandinavian Journal of Statistics. Theory and Applications, 12(2):179–181, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615983.

Fenstad:1977:TTR

[FWW77] G. U. Fenstad, L. Walløe, and S. Ø. Wille. Three tests for regression compared by stochastic simulation under normal and heavy tailed distribution of errors. Scandinavian Journal of Statistics. Theory and Applications, 4(1):31– 34, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615645.

Wu:2016:SAT

[fWZY16] Ruo fan Wu, Ming Zheng, and Wen Yu. Subgroup analysis with time-to-event data under a logistic-Cox mixture model. Scandinavian Journal of Statistics. Theory and Applications, 43(3):863–878, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Fan:2000:SCB

[FZ00] Jianqing Fan and Wenyang Zhang. Simultaneous confidence bands and hypothesis testing in varyingcoefficient models. Scandinavian Journal of Statistics. Theory and Applications, 27(4):715–731, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616637.

Francq:2006:LRB

[FZ06] Christian Francq and Jean-Michel Zakoïan. Linearrepresentation based estimation of stochastic volatility models. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 785–806, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616958.

Gomes:1986:IMG

[GA86] M. Ivette Gomes and M. Teresa Alpuim. Inference in a multivariate generalized extreme value model: Asymptotic properties of two test statistics. Scandinavian Journal of Statistics. Theory and Applications, 13(4): 291–300, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616038.

Garcia-Angulo:2023:EUM

[GAC23] Andrea C. Garcia-Angulo and Gerda Claeskens. Exact uniformly most powerful postselection confidence distributions. Scandinavian Journal of Statistics. Theory and Applications, 50 (1):358–382, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gadziev:1985:DRW

[Gad85] A. G. Gadẑiev. Delays reducing the waiting time in queueing systems with cyclic services. Scandinavian Journal of Statistics. Theory and Applications, 12(4):301–307, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616002.

Gao:1998:SRS

[Gao98] Jiti Gao. Semiparametric regression smoothing of non-linear time series. Scandinavian Journal of Statistics. Theory and Applications, 25(3):521–539, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616518.

Gardiner:1982:ADM

[Gar82] Joseph C. Gardiner. The asymptotic distribution of mortality rates in competing risks analyses. Scandinavian Journal of Statistics. Theory and Applications, 9(1):31–36, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615851.

Gartner:2003:EGR

[Gär03] Matthias Gärtner. Estimation in a general repair model based on lefttruncated data. Scandinavian Journal of Statistics. Theory and Applications, 30(1):61–73, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616749.

Gaasemyr:2003:AVM

[Gås03]  Jørund Gåsemyr.  On

an adaptive version of the Metropolis–Hastings algorithm with independent proposal distribution. Scandinavian Journal of Statistics. Theory and Applications, 30(1):159–173, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616755.

Gaasemyr:2016:UNL

[Gås16] Jørund Gåsemyr. Uniformity of node level conflict measures in Bayesian hierarchical models based on directed acyclic graphs. Scandinavian Journal of Statistics. Theory and Applications, 43(1):20–34, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gasbarra:2023:DSI

[Gas23] Dario Gasbarra. Discussion on the SJS invited paper by Sander Greenland, Divergence vs. Decision P -values: a Distinction worth making in theory and keeping in Practice. Scandinavian Journal of Statistics. Theory and Applications, 50(3): 929–930, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Gre23b].

Gundersen:2024:TTV

[GBB^{+}24] Kristian Gundersen, Timothée Bacri, Jan Bulla, Sondre Hølleland, Antonello Maruotti, and Bård Støve. Testing for timevarying nonlinear dependence structures: Regimeswitching and local Gaussian correlation. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1012–1060, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ghosh:2005:MDR

[GC05] Anil K. Ghosh and Probal Chaudhuri. On maximum depth and related classifiers. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 327–350, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616881.

Gueuning:2018:HDF

[GC18] Thomas Gueuning and Gerda Claeskens. A highdimensional focused information criterion. Scandinavian Journal of Statistics. Theory and Applications, 45(1):34–61, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Genon-Catalot:1994:EDC

[GCJ94] V. Genon-Catalot and J. Jacod. Estimation of the diffusion coefficient for diffusion processes: Random sampling. Scandinavian Journal of Statistics. Theory and Applications, 21(3): 193–221, September 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616313.

Genon-Catalot:2003:CLE

[GCJL03] Valentine Genon-Catalot, Thierry Jeantheau, and Catherine Laredo. Conditional likelihood estimators for hidden Markov models and stochastic volatility models. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 297–316, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616765.

Genon-Catalot:1987:LTF

[GCL87] V. Genon-Catalot and C. Laredo. Limit theorems for the first hitting times process of a diffusion and statistical applications. Scandinavian Journal of Statistics. Theory and Applications, 14(2): 143–160, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616058.

Genon-Catalot:1992:NPE

[GCLP92] V. Genon-Catalot, C. Laredo, and D. Picard. Nonparametric estimation of the diffusion coefficient by wavelets methods. Scandinavian Journal of Statistics. Theory and Applications, 19(4):317–335, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616250.

Glonek:1988:EML

[GDS88] G. F. V. Glonek, J. N.

Darroch, and T. P. Speed. On the existence of maximum likelihood estimators for hierarchical loglinear models. Scandinavian Journal of Statistics. Theory and Applications, 15(3): 187–193, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616100.

Genton:2005:DSN

[Gen05] Marc G. Genton. Discussion of “The Skewnormal”. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 189–198, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616869. See [Azz05b].

Guilbaud:2001:ENP

[GG01] Olivier Guilbaud and Oliver Guilbaud. Exact nonparametric confidence intervals for quantiles with progressive Type-II censoring. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 699–713, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616690.

Goegebeur:2013:AUE

[GG13] Yuri Goegebeur and Armelle Guillou. Asymptotically unbiased estimation of the coefficient of tail dependence. Scandinavian Journal of Statistics. Theory and Applications, 40(1): 174–189, March 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357259.

Gao:2023:PAR

[GG23] Fengnan Gao and Richard D. Gill. Pitfalls of amateur regression: The Dutch New Herring controversies. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1901–1918, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See commentary [VV24].

Galves:2013:JEI

[GGG13] Antonio Galves, Aurélien Garivier, and Elisabeth Gassiat. Joint estimation of intersecting context tree models. Scandinavian Journal of Statistics. Theory and Applications, 40(2):344–362, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586716.

Gardes:2020:BTM

[GGS20] Laurent Gardes, Stéphane Girard, and Gilles Stupfler. Beyond tail median and conditional tail expectation: Extreme risk estimation using tail L^{p}- optimization. Scandinavian Journal of Statistics. Theory and Applications, 47(3): 922–949, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gross:1987:MPE

[GH87] Shulamith T. Gross and Catherine Huber. Matched pair experiments: Cox and maximum likelihood estimation. Scandinavian Journal of Statistics. Theory and Applications, 14(1): 27–41, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616046.

Gaudard:1989:SAS

[GH89] Marie Gaudard and Donald Hadwin. Sigma-algebras on spaces of probability measures. Scandinavian Journal of Statistics. Theory and Applications, 16(2): 169–175, June 1989. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616132.

Gilula:2000:DAS

[GH00] Zvi Gilula and Shelby J. Haberman. Density approximation by summary statistics: an informationtheoretic approach. Scandinavian Journal of Statistics. Theory and Applications, 27(3):521–534, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616620. See correction [GH08].

Gustafsson:2002:SIW

[GH02] Nils Gustafsson and Gudmund Høst. Statistical issues in weather forecasting*. Scandinavian Journal of Statistics. Theory and Applications, 29(2): 219–239, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616713. See discussion [Høs02] and reply [Gus02].

Gilula:2008:CND

[GH08] Zvi Gilula and Shelby J.

Haberman. Correction note to “Density Approximation by Summary Statistics: an Information-theoretic Approach”. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 762, December 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

41000301. See [GH00].

Grunewald:2012:GSF

[GH12] Maria Grünewald and Ola Hössjer. A general statistical framework for multistage designs. Scandinavian Journal of Statistics. Theory and Applications, 39 (1):131–152, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41411161.

Gandy:2014:MSA

[GH14a] Axel Gandy and Georg Hahn. MMCTest — a safe algorithm for implementing multiple Monte Carlo tests. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1083–1101, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586822.

Gronneberg:2014:CIC

[GH14b] Steffen Grønneberg and Nils Lid Hjort. The copula information criteria. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 436–459, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586662.

Gandy:2016:FMC

[GH16] Axel Gandy and Georg Hahn. A framework for Monte Carlo based multiple testing. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1046–1063, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Groeneboom:2018:CIC

[GH18] Piet Groeneboom and Kim Hendrickx. Confidence intervals for the current status model. Scandinavian Journal of Statistics. Theory and Applications, 45 (1):135–163, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Grun:2021:IGD

[GH21] Bettina Grün and Paul Hofmarcher. Identifying groups of determinants in Bayesian model averaging using Dirichlet process clustering. Scandinavian Journal of Statistics. Theory and Applications, 48 (3):1018–1045, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gaasemyr:2023:PDE

[GH23] Jørund Gåsemyr and Aliaksandr Hubin. Prior distributions expressing ignorance about convex increasing failure rates. Scandinavian Journal of Statistics. Theory and Applications, 50(2):452–476, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gross:1992:RMT

[GHC92] Shulamith T. Gross and Catherine Huber-Carol. Regression models for truncated survival data. Scandinavian Journal of Statistics. Theory and Applications, 19(3):193–213, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616240.

Gandy:2020:IMC

[GHD20] Axel Gandy, Georg Hahn, and Dong Ding. Implementing Monte Carlo tests with p-value buckets. Scandinavian Journal of Statistics. Theory and Applications, 47(3):950–967, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gotze:1995:DRF

[GHH95] F. Götze, L. Heinrich, and C. Hipp. m-dependent random fields with analytic cumulant generating function. Scandinavian Journal of Statistics. Theory and Applications, 22(2): 183–195, June 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616352.

Ghosh:2006:SGC

[Gho06] Debashis Ghosh. Semiparametric global cross-ratio models for bivariate censored data. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 609–619, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616948.

Glad:2003:CDE

[GHU03] Ingrid K. Glad, Nils Lid Hjort, and Nikolai G. Ushakov. Correction of density estimators that are not densities. Scandinavian Journal of Statistics. Theory and Applications, 30(2):415–427, June 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616772.

Gavaris:2002:RD

[GI02] Stratis Gavaris and James N.

Ianelli. Reply to discussion. Scandinavian Journal of Statistics. Theory and Applications, 29(2):

270–271, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [GIA02].

Gavaris:2002:SIF

[GIA02] Stratis Gavaris, James N. Ianelli, and Elja Arjas. Statistical issues in fisheries’ stock assessments*. Scandinavian Journal of Statistics. Theory and Applications, 29(2):245–267, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616714. See [Arj02] and reply [GI02].

Gill:1986:ETI

[Gil86] Richard D. Gill. On estimating transition intensities of a Markov process with aggregate data of a certain type: “occurrences but no exposures”. Scandinavian Journal of Statistics. Theory and Applications, 13(2):113–134, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616017.

Gill:1992:MPL

[Gil92] Richard D. Gill. Marginal partial likelihood. Scandinavian Journal of Statistics. Theory and Applications, 19(2):133–137, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616233.

Gut:1983:LBC

[GJ83] Allan Gut and Svante Janson. The limiting behaviour of certain stopped sums and some applications. Scandinavian Journal of Statistics. Theory and Applications, 10(4): 281–292, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615930.

Girard:2003:EVH

[GJ03] Stéphane Girard and Pierre Jacob. Extreme values and Haar series estimates of point process boundaries. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 369–384, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616769.

Gandy:2005:GFT

[GJ05] Axel Gandy and Uwe Jensen. On goodness-of-fit tests for Aalen’s additive risk model. Scandinavian Journal of Statistics. Theory and Applications, 32(3):

425–445, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616888.

Goddard:2016:RMP

[GJ16] Scott D. Goddard and Valen E. Johnson. Restricted most powerful Bayesian tests for linear models. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1162–1177, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Groeneboom:2008:SRA

[GJW08] Piet Groeneboom, Geurt Jongbloed, and Jon A. Wellner. The support reduction algorithm for computing non-parametric function estimates in mixture models. Scandinavian Journal of Statistics. Theory and Applications, 35(3): 385–399, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000278.

Groeneboom:2012:SPI

[GJW12] Piet Groeneboom, Geurt Jongbloed, and Birgit I. Witte. Smooth plug-in inverse estimators in the current status continuous mark model. Scandinavian Journal of Statistics. Theory and Applications, 39(1):15– 33, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411155.

Gupta:1986:RRM

[GK86] Ramesh C. Gupta and Jerome P. Keating. Relations for reliability measures under length biased sampling. Scandinavian Journal of Statistics. Theory and Applications, 13 (1):49–56, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616010.

Gastwirth:1991:BGD

[GK91] Joseph L. Gastwirth and Abba M. Krieger. On bounding P (X_{2} ≤ X_{1}) from grouped data. Scandinavian Journal of Statistics. Theory and Applications, 18(2):111–117, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616195.

Gasbarra:2000:ACR

[GK00] Dario Gasbarra and S. R.

Karia. Analysis of competing risks by using Bayesian smoothing. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 605–617, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616630.

Gelfand:2003:BSR

[GK03] Alan E. Gelfand and Athanasios Kottas. Bayesian semiparametric regression for median residual life. Scandinavian Journal of Statistics. Theory and Applications, 30(4):651–665, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616794.

Gandy:2013:GCP

[GK13] Axel Gandy and Jan Terje Kvaløy. Guaranteed conditional performance of control charts via bootstrap methods. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 647–668, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586692.

Goh:2021:AMU

[GK21] Gyuhyeong Goh and Jae Kwang Kim. Accounting for model uncertainty in multiple imputation under complex sampling. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 930–949, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gissibl:2021:IER

[GKL21] Nadine Gissibl, Claudia Klüppelberg, and Steffen Lauritzen. Identifiability and estimation of recursive max-linear models. Scandinavian Journal of Statistics. Theory and Applications, 48(1):188–211, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Geng:2002:CNC

[GL02] Zhi Geng and Guangwei Li. Conditions for nonconfounding and collapsibility without knowledge of completely constructed causal diagrams. Scandinavian Journal of Statistics. Theory and Applications, 29(1):169–181, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616707.

Goldstein:2007:CSS

[GL07] Larry Goldstein and Bryan Langholz. Cohort sampling schemes for the Mantel–Haenszel estimator. Scandinavian Journal of Statistics. Theory and Applications, 34(1): 137–154, March 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548543.

Gribkova:2015:NPC

[GL15] Svetlana Gribkova and Olivier Lopez. Nonparametric copula estimation under bivariate censoring. Scandinavian Journal of Statistics. Theory and Applications, 42(4): 925–946, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Glad:1998:PGN

[Gla98] Ingrid K. Glad. Parametrically guided nonparametric regression. Scandinavian Journal of Statistics. Theory and Applications, 25(4):649–668, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616530.

Gloter:2006:PED

[Glo06] Arnaud Gloter. Parameter estimation for a discretely observed integrated diffusion process. Scandinavian Journal of Statistics. Theory and Applications, 33(1):83–104, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616910.

Glombek:2014:SIH

[Glo14] Konstantin Glombek. Statistical inference for highdimensional global minimum variance portfolios. Scandinavian Journal of Statistics. Theory and Applications, 41(4):845–865, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586810.

Guan:2018:SML

[GLQ18] Zhong Guan, Denis H. Y.

Leung, and Jing Qin. Semiparametric maximum likelihood inference for nonignorable nonresponse with callbacks. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 962–984, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Groeneveld:1982:PFL

[GM82] Richard A. Groeneveld and Glen Meeden. Properties of a family of location parameters for skewed distributions. Scandinavian Journal of Statistics. Theory and Applications, 9(4):237– 240, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615889.

Govindarajulu:1983:SRL

[GM83] Z. Govindarajulu and David M. Mason. A strong representation for linear combinations of order statistics with application to fixed-width confidence intervals for location and scale parameters. Scandinavian Journal of Statistics. Theory and Applications, 10(2):97– 115, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615909. See correction [Ano83i].

Gasser:1984:ERF

[GM84] Theo Gasser and Hans-Georg Müller. Estimating regression functions and their derivatives by the kernel method. Scandinavian Journal of Statistics. Theory and Applications, 11 (3):171–185, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615954.

Geyer:1994:SPL

[GM94] Charles J. Geyer and Jesper Møller. Simulation procedures and likelihood inference for spatial point processes. Scandinavian Journal of Statistics. Theory and Applications, 21(4): 359–373, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616323.

Gijbels:1998:LAK

[GM98] I. Gijbels and E. Mammen. Local adaptivity of kernel estimates with plugin local bandwidth selectors. Scandinavian Journal of Statistics. Theory and Applications, 25(3): 503–520, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616517.

Ghosh:2008:EBC

[GM08a] Malay Ghosh and Tapabrata Maiti. Empirical Bayes confidence intervals for means of natural exponential family-quadratic variance function distributions with application to small area estimation. Scandinavian Journal of Statistics. Theory and Applications, 35(3): 484–495, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000283.

Gine:2008:UBE

[GM08b] Evarist Giné and David M. Mason. Uniform in bandwidth estimation of integral functionals of the density function. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 739–761, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000300.

Garcia:2016:OEL

[GM16] Tanya P. Garcia and Yanyuan Ma. Optimal estimator for logistic model with distribution-free random intercept. Scandinavian Journal of Statistics. Theory and Applications, 43(1):156–171, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Grafstrom:2018:SBS

[GM18] Anton Grafström and Alina Matei. Spatially balanced sampling of continuous populations. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 792–805, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gardes:2023:NAC

[GM23] Laurent Gardes and Samuel Maistre. Nonparametric asymptotic confidence intervals for extreme quantiles. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 825–841, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gottard:2011:QSG

[GMA11] Anna Gottard, Giovanni Maria Marchetti, and Alan Agresti. Quasi-symmetric graphical log-linear models. Scandinavian Journal of Statistics. Theory and Applications, 38(3):447–465, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015574.

Garthwaite:2024:LCC

[GME24] Paul H. Garthwaite, Maha W.

Moustafa, and Fadlalla G. Elfadaly. Locally correct confidence intervals for a binomial proportion: A new criteria for an interval estimator. Scandinavian Journal of Statistics. Theory and Applications, 51 (1):220–244, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Giron:2006:OTP

[GMMT06] F. Javier Girón, M. Lina Martínez, Elías Moreno, and Francisco Torres. Objective testing procedures in linear models: Calibration of the p-values. Scandinavian Journal of Statistics. Theory and Applications, 33(4):765–784, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616957.

Gonzalez-Manteiga:2011:RCN

[GMPFV11] Wenceslao González-Manteiga, Juan Carlos Pardo-Fernández, and Ingrid Van Keilegom. ROC curves in non-parametric locationscale regression models. Scandinavian Journal of Statistics. Theory and Applications, 38(1):169–184, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015389.

George:1993:CLD

[GMS93] E. I. George, U. E. Makov, and A. F. M. Smith. Conjugate likelihood distributions. Scandinavian Journal of Statistics. Theory and Applications, 20(2): 147–156, March 1993. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616270.

Gugushvili:2020:DDD

[GMvdM20] Shota Gugushvili, Ester Mariucci, and Frank van der Meulen. Decompounding discrete distributions: a nonparametric Bayesian approach. Scandinavian Journal of Statistics. Theory and Applications, 47(2):464–492, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gaasemyr:1995:SAR

[GN95] Jørund Gåsemyr and Bent Natvig. Some aspects of reliability analysis in shock models. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 385–393, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616367.

Gaasemyr:1998:PDP

[GN98] Jørund Gåsemyr and Bent Natvig. The posterior distribution of the parameters of component lifetimes based on autopsy data in a shock model. Scandinavian Journal of Statistics. Theory and Applications, 25(2):271–292, June 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616501.

Gaasemyr:2009:ECM

[GN09] Jørund Gåsemyr and Bent Natvig. Extensions of a conflict measure of inconsistencies in Bayesian hierarchical models. Scandinavian Journal of Statistics. Theory and Applications, 36(4):822–838, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000354.

Guillen:2007:IEN

[GNPM07] Montserrat Guillen, Jens P. Nielsen, and Ana M. Perez-Marin. Improving the efficiency of the Nelson– Aalen estimator: the naive local constant estimator. Scandinavian Journal of Statistics. Theory and Applications, 34(2):419–431, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548560.

Gottschau:1994:MCM

[Got94] Adam Gottschau. Markov chain models for multivariate binary panel data. Scandinavian Journal of Statistics. Theory and Applications, 21(1):57–71, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616298.

Gijbels:2015:ECW

[GOV15] Irène Gijbels, Marek Omelka, and Noël Veraverbeke. Estimation of a copula when a covariate affects only marginal distributions. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1109–1126, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Guegan:1989:CNN

[GP89a] D. Guegan and D. T.

Pham. Correction note: A Note on the Estimation of the Parameters of the Diagonal Bilinear Model by the Method of Least Squares. Scandinavian Journal of Statistics. Theory and Applications, 16(4): 363, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616147. See [GP89b].

Guegan:1989:NEP

[GP89b] D. Guegan and D. T.

Pham. A note on the estimation of the parameters of the diagonal bilinear model by the method of least squares. Scandinavian Journal of Statistics. Theory and Applications, 16(2):129–136, June 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616128. See cor-

rection [GP89a].

Garcia-Portugues:2021:GFT

[GPÁLÁPGM21] Eduardo García-Portugués, Javier Álvarez-Liébana, Gonzalo Álvarez-Pérez, and Wenceslao González-Manteiga. A goodness-of-fit test for the functional linear model with functional response. Scandinavian Journal of Statistics. Theory and Applications, 48(2):502–528, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gutierrez-Pena:2004:CPR

[GPM04] Eduardo Gutiérrez-Peña and Pietro Muliere. Conjugate priors represent strong pre-experimental assumptions. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 235–246, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616826.

Gather:1996:UMT

[GPP96] Ursula Gather, Jörg Pawlitschko, and Iris Pigeot. Unbiasedness of multiple tests. Scandinavian Journal of Statistics. Theory and Applications, 23(1): 117–127, March 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616390.

Ghosh:2023:RSI

[GPST23] Abhik Ghosh, Erica Ponzi, Torkjel Sandanger, and Magne Thoresen. Robust sure independence screening for nonpolynomial dimensional generalized linear models. Scandinavian Journal of Statistics. Theory and Applications, 50 (3):1232–1262, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Garcia-Portugues:2016:TPM

[GPVCGM16] Eduardo García-Portugués, Ingrid Van Keilegom, Rosa M. Crujeiras, and Wenceslao González-Manteiga. Testing parametric models in linear-directional regression. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1178–1191, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Genest:2006:GFP

[GQR06] Christian Genest, Jean-François Quessy, and Bruno Rémillard. Goodness-of-fit procedures for copula models based on the probability integral transformation. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 337–366, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616928.

Green:2001:MHD

[GR01] Peter J. Green and Sylvia Richardson. Modelling heterogeneity with and without the Dirichlet process. Scandinavian Journal of Statistics. Theory and Applications, 28(2): 355–375, June 2001. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616664.

Gayraud:2005:RCB

[GR05] Ghislaine Gayraud and Judith Rousseau. Rates of convergence for a Bayesian level set estimation. Scandinavian Journal of Statistics. Theory and Applications, 32(4):639–660, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616902.

Germain:2010:WCR

[GR10] Jean-François Germain and François Roueff. Weak convergence of the regularization path in penalized Mestimation. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 477–495, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000402.

Gregoire:1993:LSC

[Gré93] Gérard Grégoire. Least squares cross-validation for counting process intensities. Scandinavian Journal of Statistics. Theory and Applications, 20(4): 343–360, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616289.

Grenander:2011:RFD

[Gre11] Ulf Grenander. Reflections on future directions. Scandinavian Journal of Statistics. Theory and Applications, 38(2):195–196, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015489.

Greenland:2023:CSP

[Gre23a] Sander Greenland. Connecting simple and precise P -values to complex and ambiguous realities (includes rejoinder to comments on “Divergence vs. decision P -values”). Scandinavian Journal of Statistics. Theory and Applications, 50(3):899–914, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Vos23, Gre23b].

Greenland:2023:DVD

[Gre23b] Sander Greenland. Divergence versus decision P - values: a distinction worth making in theory and keeping in practice: Or, how divergence P -values measure evidence even when decision P -values do not. Scandinavian Journal of Statistics. Theory and Applications, 50(1):54–88, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Lav23, Gre23a, Gas23] and corrigendum [Ano24].

Grigelionis:1980:MAS

[Gri80] B. Grigelionis. A martingale approach to the statistical problems of point processes. Scandinavian Journal of Statistics. Theory and Applications, 7(4): 190–196, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615799.

Grillenzoni:2009:KLI

[Gri09] Carlo Grillenzoni. Kernel likelihood inference for time series. Scandinavian Journal of Statistics. Theory and Applications, 36 (1):127–140, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000311.

Gross:1996:WEL

[Gro96] Shulamith T. Gross. Weighted estimation in linear regression for truncated survival data. Scandinavian Journal of Statistics. Theory and Applications, 23(2): 179–193, June 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616395.

Gronnesby:1997:TCE

[Grø97] Jon Ketil Grønnesby. Testing covariate effects in Aalen’s linear hazard model. Scandinavian Journal of Statistics. Theory and Applications, 24(1):125–135, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616441.

Groeneboom:2012:LRT

[Gro12] Piet Groeneboom. Likelihood ratio type two-sample tests for current status data. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 645–662, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357216.

Groeneboom:2021:GFC

[Gro21] Piet Groeneboom. Grenander functionals and Cauchy’s formula. Scandinavian Journal of Statistics. Theory and Applications, 48 (1):275–294, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gluck:2022:SLI

[GRS22] Jochen Glück, Stefan Roth, and Evgeny Spodarev. A solution to a linear integral equation with an application to statistics of infinitely divisible moving averages. Scandinavian Journal of Statistics. Theory and Applications, 49 (3):1244–1273, September 2022. CODEN SJSADG.

ISSN 0303-6898 (print), 1467-9469 (electronic).

Gustavsson:1976:PRM

[GS76] Jan Gustavsson and Åke Svensson. A Poisson regression model applied to classes of road accidents with small frequencies. Scandinavian Journal of Statistics. Theory and Applications, 3(2):49– 60, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615611.

Gawronski:1980:DEM

[GS80] Wolfgang Gawronski and Ulrich Stadtmüller. On density estimation by means of Poisson’s distribution. Scandinavian Journal of Statistics. Theory and Applications, 7(2):90–94, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615779.

Glidden:1999:SLE

[GS99] D. V. Glidden and S. G.

Self. Semiparametric likelihood estimation in the Clayton–Oakes failure time model. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 363–372, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616562.

Gut:2002:TSC

[GS02] Allan Gut and Josef Steinebach. Truncated sequential change-point detection based on renewal counting processes. Scandinavian Journal of Statistics. Theory and Applications, 29(4):693–719, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616742.

Grafstrom:2014:HSR

[GS14] Anton Grafström and Lina Schelin. How to select representative samples. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 277–290, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586652.

Gaasemyr:2019:APP

[GS19a] Jørund Gåsemyr and Ida Scheel. Alternatives to post-processing posterior predictive p values. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1252–1273, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gjessing:2019:E

[GS19b] Håkon K. Gjessing and Hans J. Skaug. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 46(1):1, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Goulard:1996:PEM

[GSG96] Michel Goulard, Aila Särkkä, and Pavel Grabarnik. Parameter estimation for marked Gibbs point processes through the maximum pseudo-likelihood method. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 365–379, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616410.

Ghosh:2006:EHB

[GSK06] Malay Ghosh, Karabi Sinha, and Dalho Kim. Empirical and hierarchical Bayesian estimation in finite population sampling under structural measurement error models. Scandinavian Journal of Statistics. Theory and Applications, 33(3):591–608, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616944.

Girard:2022:NEC

[GSUC22] Stéphane Girard, Gilles Stupfler, and Antoine Usseglio-Carleve. Nonparametric extreme conditional expectile estimation. Scandinavian Journal of Statistics. Theory and Applications, 49(1):78–115, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Gannoun:2005:NPQ

[GSYB05] Ali Gannoun, Jérôme Saracco, Ao Yuan, and George E. Bonney. Nonparametric quantile regression with censored data. Scandinavian Journal of Statistics. Theory and Applications, 32(4):527–550, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616896.

Greenshtein:1998:OSS

[GT98] Eitan Greenshtein and Erik Torgersen. Ordering of sequentially sampled exponential experiments. Scandinavian Journal of Statistics. Theory and Applications, 25(2):

325–329, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616504.

Guan:2007:CLC

[Gua07] Yongtao Guan. A composite likelihood crossvalidation approach in selecting bandwidth for the estimation of the pair correlation function. Scandinavian Journal of Statistics. Theory and Applications, 34(2):336–346, June 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548555.

Guilbaud:1977:EPE

[Gui77] Olivier Guilbaud. Estimating the probability eigenvector and related characteristics of an ergodic transition matrix. Scandinavian Journal of Statistics. Theory and Applications, 4 (3):97–104, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615659.

Guilbaud:1979:IEM

[Gui79] Olivier Guilbaud. Interval estimation of the median of a general distribution. Scandinavian Journal of Statistics. Theory and Applications, 6(1):29– 36, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615725.

Guilbaud:1980:ABE

[Gui80] Olivier Guilbaud. Asymptotic behavior of the empirical distribution function at a random point and some applications. Scandinavian Journal of Statistics. Theory and Applications, 7(4): 181–189, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615798.

Guilbaud:1982:FNI

[Gui82] Olivier Guilbaud. Functions of non-IID random vectors expressed as functions of IID random vectors. Scandinavian Journal of Statistics. Theory and Applications, 9(4):229– 233, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615887.

Guilbaud:1986:SIK

[Gui86] Olivier Guilbaud. Stochastic inequalities for Kolmogorov and similar statistics, with confidence region applications. Scandinavian Journal of Statistics. Theory and Applications, 13(4): 301–305, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616039.

Guilbaud:2004:ENP

[Gui04]  Olivier Guilbaud.  Ex-

act non-parametric confidence, prediction and tolerance intervals with progressive Type-II censoring. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 265–281, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616828.

Guolo:2011:MEC

[Guo11] Annamaria Guolo. Measurement error correction by exploiting gene — environment independence in family-based case — control studies. Scandinavian Journal of Statistics. Theory and Applications, 38(2):237–251, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015492.

Gupta:1976:SCD

[Gup76] Ramesh C. Gupta. Some characterizations of distributions by properties of their forward and backward recurrence times in a renewal process. Scandinavian Journal of Statistics. Theory and Applications, 3 (4):215–216, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615637.

Gustafsson:2002:RD

[Gus02] Nils Gustafsson. Reply to discussion. Scandinavian Journal of Statistics. Theory and Applications, 29(2):242–243, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Høs02].

Gill:1993:NSP

[GV93] Richard D. Gill and Aad W. Van Der Vaart. Non- and semi-parametric maximum likelihood estimators and the von Mises method: II. Scandinavian Journal of Statistics. Theory and Applications, 20(4):271–288, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616285. See also Part 1 [GWP89].

Gao:2024:CEL

[GW24] Junzhuo Gao and Lei Wang. Communicationefficient low-dimensional parameter estimation and inference for high-dimensional -quantile regression. Scandinavian Journal of Statistics. Theory and Applications, 51(1):302–333, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Groendyke:2011:BIC

[GWH11] Chris Groendyke, David Welch, and David R. Hunter. Bayesian inference for contact networks given epidemic data. Scandinavian Journal of Statistics. Theory and Applications, 38(3):600–616, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015583.

Gill:1989:NSP

[GWP89] Richard D. Gill, Jon A.

Wellner, and Jens Præstgaard. Non- and semi-parametric maximum likelihood estimators and the von Mises method (part 1) [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 16(2):97– 128, June 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616127. See also Part 2 [GV93].

Geng:2000:MGM

[GWT00] Zhi Geng, Kang Wan, and Feng Tao. Mixed graphical models with missing data and the partial imputation EM algorithm. Scandinavian Journal of Statistics. Theory and Applications, 27(3):433–444, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616615.

Gao:2023:FMA

[GZZM23] Ziwen Gao, Jiahui Zou, Xinyu Zhang, and Yanyuan Ma. Frequentist model averaging for envelope models. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1325– 1364, September 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Heikkinen:1998:NPB

[HA98] Juha Heikkinen and Elja Arjas. Non-parametric Bayesian estimation of a spatial Poisson intensity. Scandinavian Journal of Statistics. Theory and Applications, 25(3):435–450, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616514.

Haag:2008:NPR

[Haa08] Berthold R. Haag. Nonparametric regression tests using dimension reduction techniques. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 719–738, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000299.

Haggstrom:2007:PSO

[Häg07] Olle Häggström. Problem solving is often a matter of cooking up an appropriate Markov chain*. Scandinavian Journal of Statistics. Theory and Applications, 34(4):768–780, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548580.

Hald:2001:HCG

[Hal01]  Anders Hald.  On the

history of the correction for grouping, 1873–1922. Scandinavian Journal of Statistics. Theory and Applications, 28(3):417–428, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616669.

Hamdy:1988:RAT

[Ham88] H. I. Hamdy. Remarks on the asymptotic theory of triple stage estimation of the normal mean. Scandinavian Journal of Statistics. Theory and Applications, 15(4):303–310, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616114.

Han:2016:CIP

[Han16] Peisong Han. Combining inverse probability weighting and multiple imputation to improve robustness of estimation. Scandinavian Journal of Statistics. Theory and Applications, 43 (1):246–260, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hartvig:2002:SGM

[Har02] Niels Væver Hartvig. A stochastic geometry model for functional magnetic resonance images. Scandinavian Journal of Statistics. Theory and Applications, 29(3):333–353, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616720.

Hazelton:1996:BSL

[Haz96] Martin Hazelton. Bandwidth selection for local density estimators. Scandinavian Journal of Statistics. Theory and Applications, 23(2):221–232, June 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616398.

Holzmann:2006:ISE

[HB06] Hajo Holzmann and Leif Boysen. Integrated square error asymptotics for supersmooth deconvolution. Scandinavian Journal of Statistics. Theory and Applications, 33(4):849–860, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616961.

Halle:2020:CEF

[HBD^{+}20] Kari Krizak Halle, Øyvind Bakke, Srdjan Djurovic, Anja Bye, Einar Ryeng, Ulrik Wisløff, Ole A. Andreassen, and Mette Langaas. Computationally efficient familywise error rate control in genome-wide association studies using score tests for generalized linear models. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1090–1113, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Harkanen:2017:NPB

[HBH17] Tommi Härkänen, Anna But, and Jari Haukka. Nonparametric Bayesian intensity model: Exploring timeto-event data on two time scales. Scandinavian Journal of Statistics. Theory and Applications, 44(3): 798–814, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hung:2010:OCM

[HC10] Hung Hung and Chin-Tsang Chiang. Optimal composite markers for timedependent receiver operating characteristic curves with censored survival data. Scandinavian Journal of Statistics. Theory and Applications, 37(4):664–679, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000414.

Huang:2017:EIP

[HC17] Ming-Yueh Huang and Chin-Tsang Chiang. Estimation and inference procedures for semiparametric distribution models with varying linear-index. Scandinavian Journal of Statistics. Theory and Applications, 44(2):396–424, June 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hu:2015:RAL

[HCS15] Na Hu, Xuerong Chen, and Jianguo Sun. Regression analysis of lengthbiased and right-censored failure time data with missing covariates. Scandinavian Journal of Statistics. Theory and Applications, 42(2):438–452, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hansen:2016:E

[HD16] Niels Richard Hansen and Peter Dalgaard. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 43(1):1– 2, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hazelton:2022:PCT

[HD22] Martin L. Hazelton and Tilman M. Davies. Pointwise comparison of two multivariate density functions. Scandinavian Journal of Statistics. Theory and Applications, 49(4): 1791–1810, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Heikkila:2019:NGO

[Hei19]  Matias Heikkilä.  Non-

parametric geometric outlier detection. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 1300–1314, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Helland:1982:CLT

[Hel82] Inge S. Helland. Central limit theorems for martingales with discrete or continuous time. Scandinavian Journal of Statistics. Theory and Applications, 9(2): 79–94, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615860.

Helland:1990:PLS

[Hel90]  Inge S. Helland.  Par-

tial least squares regression and statistical models. Scandinavian Journal of Statistics. Theory and Applications, 17(2):97–114, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616159.

Helland:1998:PAA

[Hel98] Inge S. Helland. A population approach to analysis of variance models. Scandinavian Journal of Statistics. Theory and Applications, 25(1):3–15, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616480.

Helland:2000:MRP

[Hel00] Inge S. Helland. Model reduction for prediction in regression models. Scandinavian Journal of Statistics. Theory and Applications, 27(1):1–20, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616586.

Hellton:2023:PAR

[Hel23] Kristoffer H. Hellton. Penalized angular regression for personalized predictions. Scandinavian Journal of Statistics. Theory and Applications, 50(1): 184–212, March 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Henze:1986:PRS

[Hen86] Norbert Henze. A probabilistic representation of the ’skew-normal’ distribution. Scandinavian Journal of Statistics. Theory and Applications, 13(4): 271–275, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616036.

He:2016:PHR

[HESZ16] Peng He, Frank Eriksson, Thomas H. Scheike, and Mei-Jie Zhang. A proportional hazards regression model for the subdistribution with covariatesadjusted censoring weight for competing risks data. Scandinavian Journal of Statistics. Theory and Applications, 43(1):103–122, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Heuch:2005:DFM

[Heu05]  Ivar Heuch.  Discus-

sion of “Functional Modelling”. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 241, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616874. See [Mül05b].

Hayati:2024:KME

[HFP24] Saeed Hayati, Kenji Fukumizu, and Afshin Parvardeh. Kernel mean embedding of probability measures and its applications to functional data analysis. Scandinavian Journal of Statistics. Theory and Applications, 51(2): 447–484, June 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

He:2023:VSH

[HFS23] Yifan He, Yang Feng, and Xinyuan Song. Variable selection for high-dimensional generalized linear model with block-missing data. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1279–1297, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hardle:1985:RKE

[HG85] Wolfgang Härdle and Theo Gasser. On robust kernel estimation of derivatives of regression functions. Scandinavian Journal of Statistics. Theory and Applications, 12(3): 233–240, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615992.

Hansen:1996:KMT

[HGB96] Martin B. Hansen, Richard D. Gill, and Adrian Baddeley. Kaplan–Meier type estimators for linear contact distributions. Scandinavian Journal of Statistics. Theory and Applications, 23(2):129–155, June 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616393.

Hjorth:1982:MSF

[HH82] Urban Hjorth and U. Hjort. Model selection and forward validation. Scandinavian Journal of Statistics. Theory and Applications, 9 (2):95–105, June 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615861.

Hwang:2016:EAP

[HH16] Wen-Han Hwang and Richard Huggins. Estimating abundance from presence– absence maps via a paired negative-binomial model. Scandinavian Journal of Statistics. Theory and Applications, 43(2):573–586, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Han:2023:RIH

[HHHL23] Dongxiao Han, Miao Han, Jian Huang, and Yuanyuan Lin. Robust inference for high-dimensional single index models. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1590–1615, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hopfner:2002:NPE

[HHL02] R. Höpfner, M. Hoffmann, and E. Löcherbach. Nonparametric estimation of the death rate in branching diffusions. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 665–692, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616741.

Hudecova:2017:TSC

[HHM17] Šárka Hudecová, Marie Hušková, and Simos G. Meintanis. Tests for structural changes in time series of counts. Scandinavian Journal of Statistics. Theory and Applications, 44(4): 843–865, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Harkanen:2003:NPF

[HHVA03] Tommi Härkänen, Hannu Hausen, Jorma I. Virtanen, and Elja Arjas. A nonparametric frailty model for temporally clustered multivariate failure times. Scandinavian Journal of Statistics. Theory and Applications, 30(3):523–533, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616781.

Hinkley:1979:NWM

[Hin79] David V. Hinkley. A note on the weighted means problem. Scandinavian Journal of Statistics. Theory and Applications, 6(1): 37–40, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615726.

Herberts:2004:ODC

[HJ04] Tina Herberts and Uwe Jensen. Optimal detection of a change point in a Poisson process for different observation schemes. Scandinavian Journal of Statistics. Theory and Applications, 31(3):347–366, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616835.

Hahn:2016:HSO

[HJ16] Ute Hahn and Eva B. Vedel Jensen. Hidden secondorder stationary spatial point processes. Scandinavian Journal of Statistics. Theory and Applications, 43(2):455–475, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Henze:2021:TGH

[HJG21] Norbert Henze and María Dolores Jiménez-Gamero. A test for Gaussianity in Hilbert spaces via the empirical characteristic functional. Scandinavian Journal of Statistics. Theory and Applications, 48(2): 406–428, June 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hobert:2018:CAM

[HJKQ18] James P. Hobert, Yeun Ji Jung, Kshitij Khare, and Qian Qin. Convergence analysis of MCMC algorithms for Bayesian multivariate linear regression with non-Gaussian errors. Scandinavian Journal of Statistics. Theory and Applications, 45(3):513–533, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hjort:1986:BEA

[Hjo86a] Nils Lid Hjort. Bayes estimators and asymptotic efficiency in parametric counting process models. Scandinavian Journal of Statistics. Theory and Applications, 13(2):63–85, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616014. See correction [Hjo86b].

Hjort:1986:CNB

[Hjo86b] Nils Lid Hjort. Correction note: Bayes Estimators and Asymptotic Efficiency in Parametric Counting Process Models. Scandinavian Journal of Statistics. Theory and Applications, 13(4): 313, September 1986. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616041. See [Hjo86a].

Hjort:1988:LSM

[Hjo88] Nils Lid Hjort. On largesample multiple comparison methods. Scandinavian Journal of Statistics. Theory and Applications, 15(4): 259–271, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616110.

Hatefi:2015:MMA

[HJO15] Armin Hatefi, Mohammad Jafari Jozani, and Omer Ozturk. Mixture model analysis of partially rank-ordered set samples: Age groups of fish from length-frequency data. Scandinavian Journal of Statistics. Theory and Applications, 42(3):848–871, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hobert:2006:UMC

[HJR06] James P. Hobert, Galin L.

Jones, and Christian P. Robert. Using a Markov chain to construct a tractable approximation of an intractable probability distribution. Scandinavian Journal of Statistics. Theory and Applications, 33(1):37– 51, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616907.

Helmers:1990:BEB

[HJS90] R. Helmers, P. Janssen, and R. Serfling. Berry– Esséen and bootstrap results for generalized Lstatistics. Scandinavian Journal of Statistics. Theory and Applications, 17 (1):65–77, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616155.

Hopfner:1997:MDE

[HK97] R. Höpfner and Yu. Kutoyants. On minimum distance estimation in recurrent Markov step processes. I. Scandinavian Journal of Statistics. Theory and Applications, 24(1):61–79, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616437.

Hardle:1999:TRM

[HK99] Wolfgang Härdle and Alois Kneip. Testing a regression model when we have smooth alternatives in mind. Scandinavian Journal of Statistics. Theory and Applications, 26(2): 221–238, June 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616552.

Hormann:2015:NEH

[HK15] Siegfried Hörmann and Łukasz Kidziński. A note on estimation in Hilbertian linear models. Scandinavian Journal of Statistics. Theory and Applications, 42(1):43–62, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hillier:2024:EEM

[HK24] Grant Hillier and Raymond M. Kan. On the expectations of equivariant matrix-valued functions of Wishart and inverse Wishart matrices. Scandinavian Journal of Statistics. Theory and Applications, 51(2):697–723, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hobolth:2002:RBE

[HKD02] Asger Hobolth, John T.

Kent, and Ian L. Dryden. On the relation between edge and vertex modelling in shape analysis. Scandinavian Journal of Statistics. Theory and Applications, 29(3):355–374, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616721.

Hansen:2011:IBE

[HKJ11] Linda V. Hansen, Markus Kiderlen, and Eva B. Vedel Jensen. Image-based empirical importance sampling: an efficient way of estimating intensities. Scandinavian Journal of Statistics. Theory and Applications, 38(3):393–408, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015571.

Hoem:1976:STD

[HKK^{+}76] Jan M. Hoem, Niels Keiding, Hannu Kulokari, Bent Natvig, Ole Barndorff-Nielsen, and Jørgen Hilden. The statistical theory of demographic rates: a review of current developments [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 3(4):169– 185, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615634.

Hlubinka:2022:MQB

[HKŠ22] Daniel Hlubinka, Lukáš Kotík, and Miroslav Šiman. Multivariate quantiles with both overall and directional probability interpretation. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1586–1604, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hopfner:1999:SMB

[HL99] R. Höpfner and E. Löcherbach. Statistical models for birth and death on a flow: Local absolute continuity and likelihood ratio processes. Scandinavian Journal of Statistics. Theory and Applications, 26(1):107–128, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616544.

Huebner:2000:AAS

[HL00] M. Huebner and S. Lototsky. Asymptotic analysis of the sieve estimator for a class of parabolic SPDEs. Scandinavian Journal of Statistics. Theory and Applications, 27(2): 353–370, June 2000. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616608.

Hu:2002:CRC

[HL02] Chengcheng Hu and D. Y.

Lin. Cox regression with covariate measurement error. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 637–655, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616739.

Ho:2008:IBC

[HL08] Yvonne H. S. Ho and Stephen M. S. Lee. Iterated bootstrap- t confidence intervals for density functions. Scandinavian Journal of Statistics. Theory and Applications, 35(2):295–308, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548594.

Huang:2023:DIT

[HLP23] Bingyao Huang, Yanyan Liu, and Liuhua Peng. Distributed inference for twosample U-statistics in massive data analysis. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1090–1115, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Haeusler:1999:WAC

[HM99] Erich Haeusler and David M. Mason. Weighted approximations to continuous time martingales with applications. Scandinavian Journal of Statistics. Theory and Applications, 26(2): 281–295, June 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616556.

Hirotsu:2002:CAM

[HM02] Chihiro Hirotsu and Kohei Marumo. Changepoint analysis as a method for isotonic inference. Scandinavian Journal of Statistics.

Theory and Applications, 29(1):125–138, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616704.

Herbei:2009:HSI

[HM09] Radu Herbei and Ian W.

Mckeague. Hybrid samplers for ill-posed inverse problems. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 839–853, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000355.

Hothorn:2018:MLT

[HMB18] Torsten Hothorn, Lisa Möst, and Peter Bühlmann. Most likely transformations. Scandinavian Journal of Statistics. Theory and Applications, 45(1): 110–134, March 2018. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Holzmann:2006:IFM

[HMG06] Hajo Holzmann, Axel Munk, and Tilmann Gneiting. Identifiability of finite mixtures of elliptical distributions. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 753–763, December 2006.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616956.

Hirukawa:2022:UCR

[HMP22] Masayuki Hirukawa, Irina Murtazashvili, and Artem Prokhorov. Uniform convergence rates for nonparametric estimators smoothed by the beta kernel. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1353–1382, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Horvath:2021:DEL

[HMR21] Lajos Horváth, Curtis Miller, and Gregory Rice. Detecting early or late changes in linear models with heteroscedastic errors. Scandinavian Journal of Statistics. Theory and Applications, 48(2): 577–609, June 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Haggstrom:1999:ESM

[HN99] Olle Häggström and Karin Nelander. On exact simulation of Markov random fields using coupling from the past. Scandinavian Journal of Statistics. Theory and Applications, 26(3):

395–411, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616564.

Huskova:2019:STN

[HNNS19] Marie Hušková, Natalie Neumeyer, Tobias Niebuhr, and Leonie Selk. Specification testing in nonparametric AR-ARCH models. Scandinavian Journal of Statistics. Theory and Applications, 46(1):26– 58, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hoang:2022:NEF

[HNRT22] Van Ha Hoang, Thanh Mai Pham Ngoc, Vincent Rivoirard, and Viet Chi Tran. Nonparametric estimation of the fragmentation kernel based on a partial differential equation stationary distribution approximation. Scandinavian Journal of Statistics. Theory and Applications, 49(1): 4–43, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Henderson:1993:ILH

[HO93] Robin Henderson and Paul Oman. Influence in linear hazard models. Scandinavian Journal of Statistics. Theory and Applications, 20(3):195–212, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616276.

Hoem:1976:OMM

[Hoe76] Jan M. Hoem. On the optimality of modified minimum chi-square analytic graduation. Scandinavian Journal of Statistics. Theory and Applications, 3(2): 89–92, March 1976. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615616.

Hoem:1978:CND

[Hoe78] Jan M. Hoem. Correction note: Design-Based and Model-Based Inference in Survey Sampling [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 5(4):214, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615719.  See

[STH^{+}78].

Hjort:1994:TSS

[HOF^{+}94] Nils Lid Hjort, Henning Omre, Marianne Frisén, Fred Godtliebsen, Jon Helgeland, Jesper Møller, Eva B. Vedel Jensen, Mats Rudemo, and Henrik Stryhn. Topics in spatial statistics [with discussion, comments and rejoinder]. Scandinavian Journal of Statistics. Theory and Applications, 21(4): 289–357, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616322.

Hoglund:1978:SFP

[Hög78] Thomas Höglund. Sampling from a finite population. a remainder term estimate. Scandinavian Journal of Statistics. Theory and Applications, 5(1):69– 71, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615687.

Hogfeldt:1979:LTV

[Hög79] Peter Högfeldt. On low F - test values in linear models. Scandinavian Journal of Statistics. Theory and Applications, 6(4):175– 178, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615757.

Hojsgaard:2004:SIC

[Høj04] Søren Højsgaard. Statistical inference in context specific interaction models for contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 31(1):143–158, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616817.

Hokstad:1975:SVT

[Hok75] Per Hokstad. A supplementary variable technique applied to the M—G—1 queue. Scandinavian Journal of Statistics. Theory and Applications, 2(2):95– 98, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615583.

Hokstad:1976:UWH

[Hok76] Per Hokstad. The use of Wiener–Hopf decomposition in the study of waiting time and busy period distribution for the G—G—1 queue. Scandinavian Journal of Statistics. Theory and Applications, 3(2):79– 85, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615614.

Holm:1975:AMC

[Hol75a] Sture Holm. Asymptotic minimax character of SPR tests in one-parameter exponential classes. Scandinavian Journal of Statistics. Theory and Applications, 2(2):49–60, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615577.

Holm:1975:SIS

[Hol75b]  Sture Holm.  Sequen-

tial inversion sum tests. Scandinavian Journal of Statistics. Theory and Applications, 2(1):1–10, December 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615570.

Holm:1979:SSR

[Hol79] Sture Holm. A simple sequentially rejective multiple test procedure. Scandinavian Journal of Statistics. Theory and Applications, 6(2):65–70, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615733.

Holst:1980:MOC

[Hol80a]  Lars Holst.  On ma-

trix occupancy, committee, and capture–recapture problems. Scandinavian Journal of Statistics. Theory and Applications, 7(3): 139–146, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615789.

Holst:1980:CRS

[Hol80b] Ulla Holst. Convergence of a recursive stochastic algorithm with m-dependent observations. Scandinavian Journal of Statistics. Theory and Applications, 7(4): 207–215, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615802. See cor-

rection [Hol81b].

Holst:1981:SAR

[Hol81a] Lars Holst. Some asymptotic results for incomplete multinomial or Poisson samples. Scandinavian Journal of Statistics. Theory and Applications, 8(4):243–246, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615843.

Holst:1981:CNC

[Hol81b]  Ulla Holst.  Correction

note: Convergence of a Recursive Stochastic Algorithm with m-Dependent Observations. Scandinavian Journal of Statistics. Theory and Applications, 8(2): 124, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615822. See [Hol80b].

Holm:1993:ABC

[Hol93] Sture Holm. Abstract bootstrap confidence intervals in linear models. Scandinavian Journal of Statistics. Theory and Applications, 20(2):157–170, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616271.

Hopfner:1987:RST

[Höp87] Reinhard Höpfner. A remark on stopping times associated with continuoustime ergodic Markov chains and an application to statistical inference in birthand-death processes. Scandinavian Journal of Statistics. Theory and Applications, 14(3):211–219, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616064.

Hopfner:1990:NRB

[Höp90] Reinhard Höpfner. Null recurrent birth-and-death processes, limits of certain martingales, and local asymptotic mixed normality. Scandinavian Journal of Statistics. Theory and Applications, 17(3): 201–215, March 1990. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616169.

Hopfner:1999:SMD

[Höp99] Reinhard Höpfner. On statistical models for ddimensional stable processes, and some generalizations. Scandinavian Journal of Statistics. Theory and Applications, 26(4): 611–620, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616582.

Horvath:1985:SNL

[Hor85] Lajos Horváth. A strong non-linear renewal theorem with applications to sequential analysis. Scandinavian Journal of Statistics. Theory and Applications, 12(4):271–280, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615999.

Host:2002:D

[Høs02] Gudmund Høst. Discussion. Scandinavian Journal of Statistics. Theory and Applications, 29(2): 239–241, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [GH02].

Huser:2021:MID

[HOT21] Raphaël Huser, Thomas Opitz, and Emeric Thibaud. Max-infinitely divisible models and inference for spatial extremes. Scandinavian Journal of Statistics. Theory and Applications, 48 (1):321–348, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hougaard:1986:CST

[Hou86] Philip Hougaard. Covariance stabilizing transformations in non-linear regression. Scandinavian Journal of Statistics. Theory and Applications, 13(3): 207–210, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616026.

Hougaard:1987:MMS

[Hou87] Philip Hougaard. Modelling multivariate survival.

Scandinavian Journal of Statistics. Theory and Applications, 14(4):291–304, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616072.

Hougaard:2012:DCD

[Hou12] Philip Hougaard. Discussion of ‘The Current Duration Approach to Estimating Time to Pregnancy’. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 207–209, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

41679789. See [KHSS12a].

Huang:2000:QLE

[HP00] Jian Huang and Yudi Pawitan. Quasi-likelihood estimation of non-invertible moving average processes. Scandinavian Journal of Statistics. Theory and Applications, 27(4):689–702, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616635.

Huang:2009:ART

[HP09] Yijian Huang and Limin Peng. Accelerated recurrence time models. Scandinavian Journal of Statistics. Theory and Applications, 36(4):636–648, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000344.

Hashorva:2021:MER

[HPR21] Enkelejd Hashorva, Simone A. Padoan, and Stefano Rizzelli. Multivariate extremes over a random number of observations. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 845–880, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hu:2024:MME

[HPS24] Shuang Hu, Zuoxiang Peng, and Johan Segers. Modeling multivariate extreme value distributions via Markov trees. Scandinavian Journal of Statistics. Theory and Applications, 51(2):760–800, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hall:2008:TEC

[HQR08] Peter Hall, Peihua Qiu, and Christian Rau. Tracking edges, corners and vertices in an image. Scandinavian Journal of Statistics. Theory and Applications, 35(1):

1–17, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000382.

Helvik:1987:MCE

[HS87] Bjarne E. Helvik and Anders Rygh Swensen. Modelling of clustering effects in point processes. an application to failures in SPCsystems. Scandinavian Journal of Statistics. Theory and Applications, 14(1): 57–66, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616048.

Horvath:1995:TCM

[HS95] Lajos Horváth and Monika Serbinowska. Testing for changes in multinomial observations: The lindisfarne scribes problem. Scandinavian Journal of Statistics. Theory and Applications, 22(3):371–384, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616366.

Huang:1998:RCE

[HS98] Jianhua Z. Huang and Charles J. Stone. The L_{2} rate of convergence for event history regression with time-dependent covariates. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 603–620, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616527.

Huang:2004:FCR

[HS04] Jianhua Z. Huang and Haipeng Shen. Functional coefficient regression models for non-linear time series: a polynomial spline approach. Scandinavian Journal of Statistics. Theory and Applications, 31(4): 515–534, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616848.

Ho:2006:TSS

[HS06] Hwai-Chung Ho and Grace S.

Shieh. Two-stage Ustatistics for hypothesis testing. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 861–873, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616962.

Hall:2010:REZ

[HS10] Daniel B. Hall and Jing Shen. Robust estimation for zero-inflated Poisson regression. Scandinavian Journal of Statistics. Theory and Applications, 37(2):237–252, June 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000877.

Hering:2012:EAC

[HS12] Christian Hering and Ulrich Stadtmüller. Estimating Archimedean copulas in high dimensions. Scandinavian Journal of Statistics. Theory and Applications, 39(3):461–479, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679806.

Horrmann:2017:LDA

[HS17] Julia Hörrmann and Anne Marie Svane. Local digital algorithms applied to Boolean models. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 369–395, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Holme:1974:LRC

[HST74] Ingar Holme, Emil Spjøtvoll, and Jens O. Tveter. Likelihood ratio confidence bands for a normal distribution function. Scandinavian Journal of Statistics. Theory and Applications, 1(2): 87–92, March 1974. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615557.

Helland:2012:NOP

[HST12] Inge S. Helland, Solve Sæbø, and Håkon Tjelmeland. Near optimal prediction from relevant components. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 695–713, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357219.

Hu:2003:RPE

[HSW03] X. Joan Hu, Jianguo Sun, and Lee-Jen Wei. Regression parameter estimation from panel counts. Scandinavian Journal of Statistics. Theory and Applications, 30(1):25–43, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616747.

Hammer:2008:CVM

[HT08] Hugo Hammer and Håkon Tjelmeland. Control variates for the Metropolis– Hastings algorithm. Scandinavian Journal of Statistics. Theory and Applications, 35(3):400–414, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000279.

Hazelton:2010:SDD

[HT10] Martin L. Hazelton and Berwin A. Turlach. Semiparametric density deconvolution. Scandinavian Journal of Statistics. Theory and Applications, 37 (1):91–108, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000917.

Hellton:2014:IME

[HT14] Kristoffer Herland Hellton and Magne Thoresen. The impact of measurement error on principal component analysis. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1051–1063, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586820.

Hellton:2017:WWP

[HT17] Kristoffer H. Hellton and Magne Thoresen. When and why are principal component scores a good tool for visualizing highdimensional data? Scandinavian Journal of Statistics. Theory and Applications, 44(3):581–597, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hsu:2015:ASS

[HTK15] Wei-Wen Hsu, David Todem, and Kyungmann Kim. Adjusted supremum scoretype statistics for evaluating non-standard hypotheses. Scandinavian Journal of Statistics. Theory and Applications, 42(3): 746–759, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Huang:2013:FCL

[Hua13] Yijian Huang. Fast censored linear regression. Scandinavian Journal of Statistics. Theory and Applications, 40(4):789–806, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586700.

Huang:2024:DOA

[Hua24]  Hengzhen Huang.  De-

sign for order-of-addition experiments with two-level components. Scandinavian Journal of Statistics. Theory and Applications, 51 (1):142–170, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Huckemann:2011:IPM

[Huc11] Stephan Huckemann. Inference on 3D Procrustes means: Tree bole growth, rank deficient diffusion tensors and perturbation models. Scandinavian Journal of Statistics. Theory and Applications, 38(3): 424–446, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23015573.

Huzurbazar:1999:FMG

[Huz99] Aparna V. Huzurbazar. Flowgraph models for generalized phase type distributions having nonexponential waiting times. Scandinavian Journal of Statistics. Theory and Applications, 26(1):145–157, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616546.

Hjorth:2005:SDD

[HV05] Urban Hjorth and Anna Vadeby. Subsample distribution distance and MCMC convergence. Scandinavian Journal of Statistics. Theory and Applications, 32(2):313–326, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616880.

Hamid:2006:REG

[HV06] Jemila Seid Hamid and Dietrich Von Rosen. Residuals in the extended growth curve model. Scandinavian Journal of Statistics. Theory and Applications, 33 (1):121–138, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616912.

Hjort:2008:MPQ

[HV08] Nils Lid Hjort and Cristiano Varin. ML, PL, QL in Markov chain models. Scandinavian Journal of Statistics. Theory and Applications, 35(1):64– 82, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000385.

Heinen:2022:KSR

[HV22] Andréas Heinen and Alfonso Valdesogo. The Kendall and Spearman rank correlations of the bivariate skew normal distribution. Scandinavian Journal of Statistics. Theory and Applications, 49(4):

1669–1698, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Harkanen:2000:CPT

[HVA00] Tommi Härkänen, Jorma I. Virtanen, and Elja Arjas. Caries on permanent teeth: a non-parametric Bayesian analysis. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 577–588, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616628.

Hao:2014:LIA

[HVV14] Chengcheng Hao, Dietrich Von Rosen, and Tatjana Von Rosen. Local influence analysis in AB–BA crossover designs. Scandinavian Journal of Statistics. Theory and Applications, 41(4):1153–1166, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586826.

Huang:1995:EMD

[HW95] Jian Huang and Jon A. Wellner. Estimation of a monotone density or monotone hazard under random censoring. Scandinavian Journal of Statistics. Theory and Applications, 22(1):

3–33, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616338.

Helmers:1998:WBF

[HW98] Roelof Helmers and Marten Wegkamp. Wild bootstrapping in finite populations with auxiliary information. Scandinavian Journal of Statistics. Theory and Applications, 25(2): 383–399, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616509.

Huang:2017:MLE

[HW17] Xianzheng Huang and Md Shamim Sarker Warasi. Maximum likelihood estimators in regression models for error-prone group testing data. Scandinavian Journal of Statistics. Theory and Applications, 44(4): 918–931, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Heine:2020:PPF

[HWC20] Kari Heine, Nick Whiteley, and A. Taylan Cemgil. Parallelizing particle filters with butterfly interactions. Scandinavian Journal of Statistics. Theory and Applications, 47(2):

361–396, June 2020. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Huang:2015:ECS

[HWH15] Yih-Huei Huang, Chi-Chung Wen, and Yu-Hua Hsu. The extensively corrected score for measurement error models. Scandinavian Journal of Statistics. Theory and Applications, 42(4):911–924, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Hu:2023:RDD

[HYLT23] Zongliang Hu, Yiping Yang, Gaorong Li, and Tiejun Tong. Regularized distribution: definition, properties, and applications. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1884–1900, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Huang:2018:SIA

[HYWC18] Mian Huang, Weixin Yao, Shaoli Wang, and Yixin Chen. Statistical inference and applications of mixture of varying coefficient models. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 618–643, September 2018.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Huang:2022:EIL

[HYZ22] Qian Huang, Jinhong You, and Liwen Zhang. Efficient inference of longitudinal/functional data models with time-varying additive structure. Scandinavian Journal of Statistics. Theory and Applications, 49(2):744–771, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Huang:2007:EEM

[HZZ07] Jianhua Z. Huang, Liangyue Zhang, and Lan Zhou. Efficient estimation in marginal partially linear models for longitudinal/ clustered data using splines. Scandinavian Journal of Statistics. Theory and Applications, 34(3):451–477, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548564.

Irincheeva:2012:GLL

[ICG12] Irina Irincheeva, Eva Cantoni, and Marc G. Genton. Generalized linear latent variable models with flexible distribution of latent variables. Scandinavian Journal of Statistics. Theory and Applications, 39(4):

663–680, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357217.

Iftimi:2019:SOA

[ICM19] A. Iftimi, O. Cronie, and F. Montes. Second-order analysis of marked inhomogeneous spatiotemporal point processes: Applications to earthquake data. Scandinavian Journal of Statistics. Theory and Applications, 46(3):661–685, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Inglot:1994:PAP

[IKL94] Tadeusz Inglot, Wilbert C. M. Kallenberg, and Teresa Ledwina. Power approximations to and power comparison of smooth goodness-of-fit tests. Scandinavian Journal of Statistics. Theory and Applications, 21(2):131–145, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616305.

Imori:2015:MSC

[Imo15]  Shinpei Imori.  Model

selection criterion based on the multivariate quasilikelihood for generalized estimating equations. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1214–1224, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Irle:1990:AOG

[Irl90] A. Irle. Asymptotic optimality of general sequential probability ratio tests. Scandinavian Journal of Statistics. Theory and Applications, 17(4):321–332, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616180.

Ibragimov:1999:AKM

[IS99] R. Ibragimov and Sh.

Sharakhmetov. Analogues of Khintchine, Marcinkiewicz– Zygmund and Rosenthal inequalities for symmetric statistics. Scandinavian Journal of Statistics. Theory and Applications, 26(4): 621–633, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616583.

Istas:1996:ESF

[Ist96] Jacques Istas. Estimating the singularity function of a Gaussian process with applications. Scandinavian Journal of Statistics. Theory and Applications, 23(4): 581–595, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616426.

Imori:2014:SFC

[IYW14] Shinpei Imori, Hirokazu Yanagihara, and Hirofumi Wakaki. Simple formula for calculating bias-corrected AIC in generalized linear models. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 535–555, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586667.

Izem:2005:DFM

[Ize05] Rima Izem. Discussion of “Functional Modelling”. Scandinavian Journal of Statistics. Theory and Applications, 32(2):241–242, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616874. See [Mül05b].

Jacobsen:1989:EUM

[Jac89] Martin Jacobsen. Existence and unicity of MLEs in discrete exponential family distributions. Scandinavian Journal of Statistics.

Theory and Applications, 16(4):335–349, March 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616145.

Jacod:2000:NPK

[Jac00] Jean Jacod. Non-parametric kernel estimation of the coefficient of a diffusion. Scandinavian Journal of Statistics. Theory and Applications, 27(1):83– 96, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616592.

Jacobsen:2001:DOD

[Jac01]  Martin Jacobsen.  Dis-

cretely observed diffusions: Classes of estimating functions and small Δoptimality. Scandinavian Journal of Statistics. Theory and Applications, 28 (1):123–149, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616648.

Jagers:1977:DLF

[Jag77]  Peter Jagers.  A de-

cent likelihood function is certainly minimal sufficient. Scandinavian Journal of Statistics. Theory and Applications, 4(1):35– 36, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615646.

Joreskog:1981:ACS

[JAL^{+}81] Karl G. Jöreskog, Erling B. Andersen, Petter Laake, D. R. Cox, and Tore Schweder. Analysis of covariance structures [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 8(2):65– 92, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615818.

Janssen:1991:OST

[Jan91] Arnold Janssen. Optimal ksample tests for randomly censored data. Scandinavian Journal of Statistics. Theory and Applications, 18(2):135–152, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616197.

Jansen:2002:RD

[Jan02] Eilev S. Jansen. Reply to discussion. Scandinavian Journal of Statistics. Theory and Applications, 29(2):216–217, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JJ02].

Jha:2020:ORF

[JB20] Jayant Jha and Atanu Biswas. Orientation relationship in finite dimensional space. Scandinavian Journal of Statistics. Theory and Applications, 47 (3):1011–1034, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jensen:1978:CPI

[Jen78] Eva Bjørn Jensen. Conditional plausibility inference in contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 5(3):129–140, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615703.

Jensen:1979:ADB

[Jen79] Eva Bjørn Jensen. The asymptotic difference between the likelihood and plausibility functions in exponential families. Scandinavian Journal of Statistics. Theory and Applications, 6(1):41–44, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615727.

Jensen:1981:PIT

[Jen81a] D. R. Jensen. Power of invariant tests for linear hypotheses under spherical symmetry. Scandinavian Journal of Statistics. Theory and Applications, 8(3): 169–174, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615829.

Jensen:1981:HD

[Jen81b] Jens Ledet Jensen. On the hyperboloid distribution. Scandinavian Journal of Statistics. Theory and Applications, 8(4):193– 206, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615838.

Jensen:1986:IMG

[Jen86] J. L. Jensen. Inference for the mean of a gamma distribution with unknown shape parameter. Scandinavian Journal of Statistics. Theory and Applications, 13(2): 135–151, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616018.

Jensen:1987:DMB

[Jen87a] E. B. Jensen. Designand model-based stereological analysis of arbitrarily shaped particles. Scandinavian Journal of Statistics. Theory and Applications, 14(3):161–180, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616061.

Jensen:1987:AEN

[Jen87b] J. L. Jensen. On asymptotic expansions in non-ergodic models. Scandinavian Journal of Statistics. Theory and Applications, 14(4): 305–318, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616073.

Jensen:1989:AES

[Jen89] J. L. Jensen. Asymptotic expansions for strongly mixing Harris recurrent Markov chains. Scandinavian Journal of Statistics. Theory and Applications, 16(1):47–63, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616121.

Jensen:1993:ANE

[Jen93a] J. L. Jensen. Asymptotic normality of estimates in spatial point processes. Scandinavian Journal of Statistics. Theory and Applications, 20(2):97– 109, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616267.

Jensen:1993:HSS

[Jen93b] J. L. Jensen. A historical sketch and some new results on the improved log likelihood ratio statistic. Scandinavian Journal of Statistics. Theory and Applications, 20(1):1– 15, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616256.

Jensen:1997:SDC

[Jen97] Jens Ledet Jensen. A simple derivation of r^{∗} for curved exponential families. Scandinavian Journal of Statistics. Theory and Applications, 24(1):33– 46, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616435.

Jiang:2005:PSC

[JFKC05] Hongyu Jiang, Jason P.

Fine, Michael R. Kosorok, and Rick Chappell. Pseudo self-consistent estimation of a copula model with informative censoring. Scandinavian Journal of Statistics. Theory and Applications, 32(1):1–20, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616858.

Jiao:2023:BPD

[JFO23] Shuhao Jiao, Ron D.

Frostig, and Hernando Ombao. Break point detection for functional covariance. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 477–512, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jensen:1979:RMT

[JGØ79] E. B. Jensen, H. J. G. Gundersen, and R. Østerby. Reconstruction of membrane thickness distribution from orthogonal intercept lengths in a plane section. Scandinavian Journal of Statistics. Theory and Applications, 6(4):182– 183, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615760.

Jalilian:2013:DVS

[JGW13] Abdollah Jalilian, Yongtao Guan, and Rasmus Waagepetersen. Decomposition of variance for spatial Cox processes. Scandinavian Journal of Statistics. Theory and Applications, 40(1):119–137, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357256.

Johansson:2005:SNM

[JH05] Jan-Olof Johansson and Ola Hössjer. A shot-noise model for paper fibres with non-uniform random orientations. Scandinavian Journal of Statistics. Theory and Applications, 32(3): 351–363, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616884.

Jiang:2017:SPT

[JH17] Fei Jiang and Sebastien Haneuse. A semi-parametric transformation frailty model for semi-competing risks survival data. Scandinavian Journal of Statistics. Theory and Applications, 44 (1):112–129, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jansen:2002:SIM

[JJ02] Eilev S. Jansen and Søren Johansen. Statistical issues in macroeconomic modelling*. Scandinavian Journal of Statistics. Theory and Applications, 29(2): 193–213, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616712. See [Joh02] and reply [Jan02].

Josey:2021:FCB

[JJCYG21] Kevin P. Josey, Elizabeth Juarez-Colunga, Fan Yang, and Debashis Ghosh. A framework for covariate balance using Bregman distances. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 790–816, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See errata [Ano23].

Jones:1992:CKD

[JK92] M. C. Jones and R. F. Kappenman. On a class of kernel density estimate bandwidth selectors. Scandinavian Journal of Statistics. Theory and Applications, 19(4):337–349, June 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616251.

Jorgensen:2004:POB

[JK04] Bent Jørgensen and Sven Jesper Knudsen. Parameter orthogonality and bias adjustment for estimating functions. Scandinavian Journal of Statistics. Theory and Applications, 31 (1):93–114, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616814.

Jasra:2019:BIS

[JKM19] Ajay Jasra, Kengo Kamatani, and Hiroki Masuda. Bayesian inference for stable Lévy-driven stochastic differential equations with high-frequency data. Scandinavian Journal of Statistics. Theory and Applications, 46(2): 545–574, June 2019. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jing:2012:PHM

[JKN12] Junmei Jing, Inge Koch, and Kanta Naito. Polynomial histograms for multivariate density and mode estimation. Scandinavian Journal of Statistics. Theory and Applications, 39(1): 75–96, March 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411158.

Jing:2002:PSA

[JKR02] Bing-Yi Jing, John E. Kolassa, and John Robinson. Partial saddlepoint approximations for transformed means. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 721–731, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616743.

James:2006:CDF

[JLP06] Lancelot F. James, Antonio Lijoi, and Igor Prünster. Conjugacy as a distinctive feature of the Dirichlet process. Scandinavian Journal of Statistics. Theory and Applications, 33 (1):105–120, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616911.

James:2009:PAN

[JLP09] Lancelot F. James, Antonio Lijoi, and Igor Prünster. Posterior analysis for normalized random measures with independent increments. Scandinavian Journal of Statistics. Theory and Applications, 36(1):76– 97, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000308.

Jain:2019:ITC

[JLRT19] Shantanu Jain, Michael Levine, Predrag Radivojac, and Michael W. Trosset. Identifiability of twocomponent skew normal mixtures with one known component. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 955–986, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jin:2006:RRA

[JLY06] Z. Jin, D. Y. Lin, and Z. Ying. Rank regression analysis of multivariate failure time data based on marginal linear models. Scandinavian Journal of Statistics. Theory and Applications, 33(1):1– 23, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616905.

Jupp:1983:NME

[JM83] P. E. Jupp and K. V. Mardia. A note on the maximum-entropy principle. Scandinavian Journal of Statistics. Theory and Applications, 10(1):45– 47, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615900.

Janssen:1993:RTG

[JM93] Arnold Janssen and Hartmut Milbrodt. Rényi type goodness of fit tests with adjusted principal direction of alternatives. Scandinavian Journal of Statistics. Theory and Applications, 20(3):177–194, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616275.

Janssen:2001:CSS

[JM01] Arnold Janssen and Claus-Dieter Mayer. Conditional studentized survival tests for randomly censored models. Scandinavian Journal of Statistics. Theory and Applications, 28(2):283–293, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616659.

Jacobsen:2016:NLS

[JM16] Martin Jacobsen and Torben Martinussen. A note on the large sample properties of estimators based on generalized linear models for correlated pseudoobservations. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 845–862, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jorgensen:1994:ABV

[JMT94] Bent Jørgensen, José Raúl Martínez, and Min Tsao. Asymptotic behaviour of the variance function. Scandinavian Journal of Statistics. Theory and Applications, 21(3):223–243, September 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616314.

Johansen:2016:ATO

[JN16a] Søren Johansen and Bent Nielsen. Asymptotic theory of outlier detection algorithms for linear time series regression models. Scandinavian Journal of Statistics. Theory and Applications, 43(2):321–348, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See discussion [ACR16, CW16, CW16, Oja16, Ron16, Zwa16] and rejoinder [JN16b].

Johansen:2016:RAT

[JN16b] Søren Johansen and Bent Nielsen. Rejoinder: Asymptotic Theory of Outlier Detection Algorithms for Linear Time Series Regression Models. Scandinavian Journal of Statistics. Theory and Applications, 43(2):374–381, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JN16a].

Jiongo:2019:BMS

[JN19] Valery D. Jiongo and Pierre Nguimkeu. Bootstrapping mean-squared errors of robust small-area estimators: Application to the methodof-payments surveys data. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1274–1299, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Johansen:1983:STR

[JNS^{+}83] Søren Johansen, Ragnar Norberg, Ib Skovgaard, Emil Spjötvoll, Rolf Sundberg, Poul Thyregod, and Aage Vølund. Some topics in regression [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 10(3):161–194, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615918.

Johansen:1977:SRA

[Joh77]  Søren Johansen.  Some

remarks on M-ancillarity. Scandinavian Journal of Statistics. Theory and Applications, 4(4):181–182, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615676.

Johansen:1978:PLE

[Joh78] Søren Johansen. The product limit estimator as maximum likelihood estimator. Scandinavian Journal of Statistics. Theory and Applications, 5(4):195– 199, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615715.

Johansen:1982:AIR

[Joh82] Søren Johansen. Asymptotic inference in random coefficient regression models. Scandinavian Journal of Statistics. Theory and Applications, 9(4):201– 207, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615883.

Johansson:1990:UPP

[Joh90] Björn Johansson. Unbiased prediction in the Poisson and Yule processes. Scandinavian Journal of Statistics. Theory and Applications, 17(2):135–145, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616161.

Johansen:1997:LAM

[Joh97] Søren Johansen. Likelihood analysis of the i (2) model. Scandinavian Journal of Statistics. Theory and Applications, 24(4): 433–462, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616469.

Johansen:2002:D

[Joh02] Søren Johansen. Discussion. Scandinavian Journal of Statistics. Theory and Applications, 29(2): 213–216, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JJ02].

Johnson:2008:PBF

[Joh08] Valen E. Johnson. Properties of Bayes factors based on test statistics. Scandinavian Journal of Statistics. Theory and Applications, 35(2):354–368, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548598.

Johnson:2017:NTS

[Joh17] Brent A. Johnson. Nonparametric two-sample tests of the marginal mark distribution with censored marks. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 545–562, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jones:1978:LPE

[Jon78] Douglas H. Jones. Limiting Pitman efficient combinations of test statistiscs. Scandinavian Journal of Statistics. Theory and Applications, 5(4):209– 212, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615718.

Jones:1991:RTS

[Jon91] Michael P. Jones. Robust tests for survival data involving a single continuous covariate. Scandinavian Journal of Statistics. Theory and Applications, 18 (4):323–332, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616214.

Jones:2001:UTS

[Jon01a] Michael P. Jones. Unmasking the trend sought by Jonckheere-type tests for right-censored data. Scandinavian Journal of Statistics. Theory and Applications, 28(3):527–535, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616676.

Jongbloed:2001:SML

[Jon01b] Geurt Jongbloed. Sieved maximum likelihood estimation in Wicksell’s problem and related deconvolution problems. Scandinavian Journal of Statistics. Theory and Applications, 28(1):161–183, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616650.

Jorgensen:1986:SPE

[Jør86]  Bent Jørgensen.  Some

properties of exponential dispersion models. Scandinavian Journal of Statistics. Theory and Applications, 13(3):187–197, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616024.

Jorgensen:1992:FRL

[Jør92] Bent Jørgensen. Finding rank leverage subsets in regression. Scandinavian Journal of Statistics. Theory and Applications, 19(2): 139–156, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616234.

Jeong:2006:LSA

[JP06] S.-O. Jeong and B. U. Park. Large sample approximation of the distribution for convex-hull estimators of boundaries. Scandinavian Journal of Statistics. Theory and Applications, 33 (1):139–151, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616913.

Jiang:2015:LRT

[JQ15] Tiefeng Jiang and Yongcheng Qi. Likelihood ratio tests for high-dimensional normal distributions. Scandinavian Journal of Statistics. Theory and Applications, 42(4):988–1009, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jonrup:1976:RAS

[JR76] Hans Jönrup and Bo Rennermalm. Regression analysis in samples from finite populations. Scandinavian Journal of Statistics. Theory and Applications, 3 (1):33–36, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615605.

Jarner:2007:CHT

[JR07] Søren F. Jarner and Gareth O.

Roberts. Convergence of heavy-tailed Monte Carlo Markov chain algorithms. Scandinavian Journal of Statistics. Theory and Applications, 34(4):781–815, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548581.

Jankowski:2012:CRM

[JS12] Hanna Jankowski and Larissa Stanberry. Confidence regions for means of random sets using oriented distance functions. Scandinavian Journal of Statistics. Theory and Applications, 39(2):340–357, June 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679798.

Jasra:2011:ILD

[JSDT11] Ajay Jasra, David A. Stephens, Arnaud Doucet, and Theodoros Tsagaris. Inference for Lévy-driven stochastic volatility models via adaptive sequential Monte Carlo. Scandinavian Journal of Statistics. Theory and Applications, 38(1): 1–22, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23015381.

Jennen-Steinmetz:1986:APR

[JSG86] C. Jennen-Steinmetz and Th. Gasser. The asymptotic power of runs tests. Scandinavian Journal of Statistics. Theory and Applications, 13(4):263–269, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616035.

Jorgensen:1991:MIG

[JSW91] B. Jørgensen, V. Seshadri, and G. A. Whitmore. On the mixture of the inverse Gaussian distribution with its complementary reciprocal. Scandinavian Journal of Statistics. Theory and Applications, 18(1):77–89, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616190.

Jensen:2007:STM

[JT07] Eva B. Vedel Jensen and Thordis L. Thorarinsdottir. A spatio-temporal model for functional magnetic resonance imaging data — with a view to resting state networks. Scandinavian Journal of Statistics. Theory and Applications, 34(3): 587–614, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548571.

Jammalamadaka:2021:ATS

[JTT21] Sreenivasa Rao Jammalamadaka, Emanuele Taufer, and György H. Terdik. Asymptotic theory for statistics based on cumulant vectors with applications. Scandinavian Journal of Statistics. Theory and Applications, 48(2): 708–728, June 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jung:2008:LIG

[Jun08] Kang-Mo Jung. Local influence in generalized estimating equations. Scandinavian Journal of Statistics. Theory and Applications, 35(2):286–294, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548593.

Jun:2011:NSC

[Jun11]  Mikyoung Jun.  Non-

stationary cross-covariance models for multivariate processes on a globe. Scandinavian Journal of Statistics. Theory and Applications, 38(4):726–747, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41411147.

Jongbloed:2006:PES

[JV06] Geurt Jongbloed and Frank H. Van Der Meulen. Parametric estimation for subordinators and induced OU processes. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 825–847, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616960.

Jeon:2024:DER

[JV24] Jeong Min Jeon and Ingrid Van Keilegom. Density estimation and regression analysis on hyperspheres in the presence of measurement error. Scandinavian Journal of Statistics. Theory and Applications, 51(2):513–556, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jalilian:2011:RAI

[JVA11] Abdollah H. Jalilian and Mohammad Q. Vahidi-Asl. Residual analysis for inhomogeneous Neyman–Scott processes. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 617–630, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41411141.

Jongbloed:2022:BNE

[JvdMP22] Geurt Jongbloed, Frank H.

van der Meulen, and Lixue Pang. Bayesian nonparametric estimation in the current status continuous mark model. Scandinavian Journal of Statistics. Theory and Applications, 49 (3):1329–1352, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jin:2010:MVP

[JW10] Lei Jin and Suojin Wang.

A model validation procedure when covariate data are missing at random.

Scandinavian Journal of Statistics. Theory and Applications, 37(3):403–421, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000398.

Janic-Wroblewska:2000:DDR

[JWL00] Alicja Janic-Wróblewska and Teresa Ledwina. Data driven rank test for twosample problem. Scandinavian Journal of Statistics. Theory and Applications, 27(2):281–297, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616603.

Jaaskinen:2014:SMC

[JXCK14] Väinö Jääskinen, Jie Xiong, Jukka Corander, and Timo Koski. Sparse Markov chains for sequence data. Scandinavian Journal of Statistics. Theory and Applications, 41(3):639–655, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586751.

Kulathinal:2006:BIC

[KA06] Sangita Kulathinal and Elja Arjas. Bayesian inference from case–cohort data with multiple endpoints. Scandinavian Journal of Statistics. Theory and Applications, 33(1):25– 36, March 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616906.

Kabe:1978:BVM

[Kab78] D. G. Kabe. On binary variables multiple regression. Scandinavian Journal of Statistics. Theory and Applications, 5(1):72– 74, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615688.

Karvanen:2015:SDC

[Kar15] Juha Karvanen. Study design in causal models. Scandinavian Journal of Statistics. Theory and Applications, 42(2):361–377, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Karttunen:2020:AMB

[Kar20] Henri Karttunen. An autoregressive model based on the generalized hyperbolic distribution. Scandinavian Journal of Statistics. Theory and Applications, 47(3): 787–816, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kauermann:1996:DGG

[Kau96] Göran Kauermann. On a dualization of graphical Gaussian models. Scandinavian Journal of Statistics. Theory and Applications, 23(1):105–116, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616389.

Kim:2004:VES

[KB04] Hyon-Jung Kim and Dennis D. Boos. Variance estimation in spatial regression using a non-parametric semivariogram based on residuals. Scandinavian Journal of Statistics. Theory and Applications, 31(3): 387–401, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616837.

Kuroki:2011:SAP

[KC11] Manabu Kuroki and Zhihong Cai. Statistical analysis of ‘probabilities of causation’ using co-variate information. Scandinavian Journal of Statistics. Theory and Applications, 38(3): 564–577, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015581.

Komanska:1984:UEV

[KD84] Henryka Komanska and S. W. Dharmadhikari. Unbiased estimation of the variance of an estimator of lanke. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 271–274, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615965.

Kent:2014:CDP

[Ken14] John T. Kent. Contribution to the discussion of the paper Geodesic Monte Carlo on Embedded Manifolds. Scandinavian Journal of Statistics. Theory and Applications, 41(1):10– 11, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

24586723. See [BG13].

Kessler:1997:EED

[Kes97] Mathieu Kessler. Estimation of an ergodic diffusion from discrete observations. Scandinavian Journal of Statistics. Theory and Applications, 24(2): 211–229, June 1997. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616449.

Kessler:2000:SEE

[Kes00]  Mathieu Kessler.  Sim-

ple and explicit estimating functions for a discretely observed diffusion process. Scandinavian Journal of Statistics. Theory and Applications, 27(1): 65–82, March 2000. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616591.

Kneib:2007:MMA

[KF07] Thomas Kneib and Ludwig Fahrmeir. A mixed model approach for geoadditive hazard regression. Scandinavian Journal of Statistics. Theory and Applications, 34(1):207–228, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548547.

Krupskii:2018:LFC

[KG18] Pavel Krupskii and Marc G. Genton. Linear factor copula models and their properties. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 861–878, December 2018.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Knorr-Held:1999:CPP

[KH99] Leonhard Knorr-Held. Conditional prior proposals in dynamic models. Scandinavian Journal of Statistics. Theory and Applications, 26(1):129–144, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616545.

Kuroki:2016:EAC

[KH16] Manabu Kuroki and Takahiro Hayashi. On the estimation accuracy of causal effects using supplementary variables. Scandinavian Journal of Statistics. Theory and Applications, 43(2):505–519, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Krebs:2022:FCL

[KH22] Johannes Krebs and Christian Hirsch. Functional central limit theorems for persistent Betti numbers on cylindrical networks. Scandinavian Journal of Statistics. Theory and Applications, 49(1):427–454, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Klein:2022:MCT

[KHBK22] Nadja Klein, Torsten Hothorn, Luisa Barbanti, and Thomas Kneib. Multivariate conditional transformation models. Scandinavian Journal of Statistics. Theory and Applications, 49(1): 116–142, March 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ko:2019:FIC

[KHH19] Vinnie Ko, Nils Lid Hjort, and Ingrid Hobæk Haff. Focused information criteria for copulas. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 1117–1140, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kong:1998:ESC

[KHL98] Fan Hui Kong, Weihua Huang, and Xiaoming Li. Estimating survival curves under proportional hazards model with covariate measurement errors. Scandinavian Journal of Statistics. Theory and Applications, 25(4):573–587, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616525.

Knorr-Held:2002:BUM

[KHR02] Leonhard Knorr-Held and Håvard Rue. On block updating in Markov random field models for disease mapping. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 597–614, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616737.

Kim:2019:SSD

[KHSJ19] Byungwon Kim, Stephan Huckemann, Jörn Schulz, and Sungkyu Jung. Smallsphere distributions for directional data with application to medical imaging. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1047–1071, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Keiding:2012:CDA

[KHSS12a] Niels Keiding, Oluf K. Højbjerg Hansen, Ditte Nørbo Sørensen, and Rémy Slama. The current duration approach to estimating time to pregnancy. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 185–204, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

41679787. See discussion [Aal12, Hou12] and reply [KHSS12b].

Keiding:2012:RCD

[KHSS12b] Niels Keiding, Oluf K. Højbjerg Hansen, Ditte Nørbo Sørensen, and Rémy Slama. Reply to the comments by Drs Aalen and Hougaard on ‘The Current Duration Approach to Estimating Time to Pregnancy’ by Niels Keiding et al. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 210–213, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

41679790. See [KHSS12a].

Kubokawa:2014:MUB

[KHT14] Tatsuya Kubokawa, Mana Hasukawa, and Kunihiko Takahashi. On measuring uncertainty of benchmarked predictors with application to disease risk estimate. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 394–413, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586660.

Kim:1997:EST

[Kim97] Gie-Whan Kim. Exact slopes of test statistics for the multivariate exponential family. Scandinavian Journal of Statistics. Theory and Applications, 24(3): 387–406, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616463.

Kim:2003:PCM

[Kim03] Yongdai Kim. On the posterior consistency of mixtures of Dirichlet process priors with censored data. Scandinavian Journal of Statistics. Theory and Applications, 30(3):535–547, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616782.

Kim:2016:NCS

[KJH16] Yongdai Kim, Jong-June Jeon, and Sangmi Han. A necessary condition for the strong oracle property. Scandinavian Journal of Statistics. Theory and Applications, 43(2): 610–624, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kuchler:2000:DES

[KK00] Uwe Küchler and Yury A.

Kutoyants. Delay estimation for some stationary diffusion-type processes.

Scandinavian Journal of Statistics. Theory and Applications, 27(3):405–414, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616613.

Kim:2006:NCD

[KK06] Sung-Ho Kim and Seong-Ho Kim. A note on collapsibility in DAG models of contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 33(3):575–590, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616943.

Kottas:2009:BSM

[KK09] Athanasios Kottas and Milovan KrnjajiĆ. Bayesian semiparametric modelling in quantile regression. Scandinavian Journal of Statistics. Theory and Applications, 36(2):297–319, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000322.

Kabaila:2019:LSC

[KK19] Paul Kabaila and Rupert E. H. Kuveke. The large sample coverage probability of confidence intervals in general regression models after a preliminary hypothesis test. Scandinavian Journal of Statistics. Theory and Applications, 46(2):432–445, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kono:2023:CBI

[KK23] Haruki Kono and Tatsuya Kubokawa. Consistent Bayesian information criterion based on a mixture prior for possibly highdimensional multivariate linear regression models. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1022–1047, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kim:2017:BAP

[KKC17] Gwangsu Kim, Yongdai Kim, and Taeryon Choi. Bayesian analysis of the proportional hazards model with time-varying coefficients. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 524–544, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kohler:2018:UQC

[KKMP18] Michael Kohler, Adam Krzyżak, Shashidhar Mallapur, and Roland Platz. Uncertainty quantification in case of imperfect models: a non-Bayesian approach. Scandinavian Journal of Statistics. Theory and Applications, 45(3):729–752, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Koo:1999:LDE

[KKP99] Ja-Yong Koo, Charles Kooperberg, and Jinho Park. Logspline density estimation under censoring and truncation. Scandinavian Journal of Statistics. Theory and Applications, 26(1):87–105, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616543.

Kluppelberg:2008:SPM

[KKP08] Claudia Klüppelberg, Gabriel Kuhn, and Liang Peng. Semi-parametric models for the multivariate tail dependence function — the asymptotically dependent case. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 701–718, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000298.

Kutzker:2024:FST

[KKW24] Tim Kutzker, Nadja Klein, and Dominik Wied. Flexible specification testing in quantile regression models. Scandinavian Journal of Statistics. Theory and Applications, 51(1):355–383, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Keiding:1978:MML

[KL78] Niels Keiding and Steffen L.

Lauritzen. Marginal maximum likelihood estimates and estimation of the offspring mean in a branching process. Scandinavian Journal of Statistics. Theory and Applications, 5(2): 106–110, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615695.

Kuchler:1989:EFE

[KL89] Uwe Küchler and Steffen L. Lauritzen. Exponential families, extreme point models and minimal space– time invariant functions for stochastic processes with stationary and independent increments. Scandinavian Journal of Statistics. Theory and Applications, 16(3): 237–261, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616137.

Kang:2014:PCT

[KL14] Jiwon Kang and Sangyeol Lee. Parameter change test for Poisson autoregressive models. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1136–1152, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586825.

Kim:2022:MCL

[KL22] Moosup Kim and Sangyeol Lee. Maximum composite likelihood estimation for spatial extremes models of Brown–Resnick type with application to precipitation data. Scandinavian Journal of Statistics. Theory and Applications, 49 (3):1023–1059, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Klefsjo:1981:HSU

[Kle81] Bengt Klefsjö. HNBUE survival under some shock models. Scandinavian Journal of Statistics. Theory and Applications, 8(1):39– 47, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615811.

Klefsjo:1982:APT

[Kle82] Bengt Klefsjö. On aging properties and total time on test transforms. Scandinavian Journal of Statistics. Theory and Applications, 9 (1):37–41, June 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615852.

Klefsjo:1983:TEA

[Kle83] Bengt Klefsjö. Testing exponentiality against HN-BUE. Scandinavian Journal of Statistics. Theory and Applications, 10(2):65– 75, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615906.

Klein:1991:SSM

[Kle91] John P. Klein. Small sample moments of some estimators of the variance of the Kaplan–Meier and Nelson–Aalen estimators. Scandinavian Journal of Statistics. Theory and Applications, 18(4):333–340, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616215.

Klemela:1999:AMR

[Kle99] Jussi Klemelä. Asymptotic minimax risk for the white noise model on the sphere. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 465–473, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616568.

Kleppe:2016:ASS

[Kle16] Tore Selland Kleppe. Adaptive step size selection for Hessian-based manifold Langevin samplers. Scandinavian Journal of Statistics. Theory and Applications, 43(3):788–805, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kleppe:2024:LDG

[Kle24] Tore Selland Kleppe. Logdensity gradient covariance and automatic metric tensors for Riemann manifold Monte Carlo methods. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1206–1229, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kobilinsky:1991:EDG

[KM91] A. Kobilinsky and H. Monod.

Experimental design generated by group morphisms: an introduction. Scandinavian Journal of Statistics. Theory and Applications, 18(2):119–134, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616196.

Kluppelberg:1994:SLT

[KM94] Claudia Klüppelberg and Thomas Mikosch. Some limit theory for the selfnormalised periodogram of stable processes. Scandinavian Journal of Statistics. Theory and Applications, 21(4):485–491, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616332.

Kobilinsky:1995:JRF

[KM95a] A. Kobilinsky and H. Monod.

Juxtaposition of regular factorial designs and the complex linear model. Scandinavian Journal of Statistics. Theory and Applications, 22(2):223–254, June 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616355.

Kundu:1995:CME

[KM95b] Debasis Kundu and Amit Mitra. Consistent method of estimating superimposed exponential signals. Scandinavian Journal of Statistics. Theory and Applications, 22(1):73– 82, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616341.

Klaassen:2000:CES

[KM00] Chris A. J. Klaassen and Robert M. Mnatsakanov. Consistent estimation of the structural distribution function. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 733–746, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616638.

Kundu:2021:BVS

[KMG21] Debamita Kundu, Riten Mitra, and Jeremy T. Gaskins. Bayesian variable selection for multioutcome models through shared shrinkage. Scandinavian Journal of Statistics. Theory and Applications, 48(1): 295–320, March 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kiwitt:2012:ECE

[KN12] Sebastian Kiwitt and Natalie Neumeyer. Estimating the conditional error distribution in non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 259–281, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679794.

Knight:1998:DMA

[Kni98] Keith Knight. A “delta method” approach to Bahadur– Kiefer theorems. Scandinavian Journal of Statistics. Theory and Applications, 25(3):555–568, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616520.

Kauermann:2003:LLE

[KO03] Göran Kauermann and J. D. Opsomer. Local likelihood estimation in generalized additive models. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 317–337, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616766.

Koike:2014:ECC

[Koi14] Yuta Koike. An estimator for the cumulative covolatility of asynchronously observed semimartingales with jumps. Scandinavian Journal of Statistics. Theory and Applications, 41(2):460–481, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586663.

Kolehmainen:1981:BME

[Kol81] Osmo Kolehmainen. Bayesian models in estimating the total of a finite population: Towards a general theory. Scandinavian Journal of Statistics. Theory and Applications, 8(1):27– 32, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615809.

Kolassa:1997:IPE

[Kol97] John E. Kolassa. Infinite parameter estimates in logistic regression, with application to approximate conditional inference. Scandinavian Journal of Statistics. Theory and Applications, 24(4):523–530, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616473.

Koo:1999:LDB

[Koo99] Ja-Yong Koo. Logspline deconvolution in Besov space. Scandinavian Journal of Statistics. Theory and Applications, 26(1):73– 86, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616542.

Kopczewska:2023:SBM

[Kop23] Katarzyna Kopczewska.

Spatial bootstrapped microeconometrics: Forecasting for out-of-sample geolocations in big data. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1391–1419, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Korhonen:1982:PSM

[Kor82] Markku P. Korhonen. On the performance of some multiple comparison procedures with unequal variances. Scandinavian Journal of Statistics. Theory and Applications, 9(4):241– 247, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615890.

Korsholm:2000:SNV

[Kor00] Lars Korsholm. The semiparametric normal variancemean mixture model. Scandinavian Journal of Statistics. Theory and Applications, 27(2):227–261, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616601.

Koster:1999:VMI

[Kos99] Jan T. A. Koster. On the validity of the Markov interpretation of path diagrams of Gaussian structural equations systems with correlated errors. Scandinavian Journal of Statistics. Theory and Applications, 26(3):413–431, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616565.

Kousgaard:1976:MPC

[Kou76] Nils Kousgaard. Models for paired comparisons with ties. Scandinavian Journal of Statistics. Theory and Applications, 3(1):1– 14, September 1976. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615602.

Kousgaard:1979:CAA

[Kou79] Nils Kousgaard. A conditional approach to the analysis of data from paired comparison experiments incorporating within-pair order effects. Scandinavian Journal of Statistics. Theory and Applications, 6 (4):154–160, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615754.

Kousgaard:1984:ASF

[Kou84] Nils Kousgaard. Analysis of a sound field experiment by a model for paired comparisons with explanatory variables. Scandinavian Journal of Statistics. Theory and Applications, 11(1):51–57, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615941.

Koutrouvelis:1985:DFP

[Kou85] Ioannis A. Koutrouvelis.

Distribution-free procedures for location and symmetry inference problems based on the empirical characteristic function. Scandinavian Journal of Statistics. Theory and Applications, 12(4): 257–269, June 1985. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615998.

Kumar:1977:STG

[KP77a] A. Kumar and P. K. Pathak. Sufficiency and tests of goodness of fit. Scandinavian Journal of Statistics. Theory and Applications, 4(1):39–43, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615648.

Kumar:1977:TAB

[KP77b] A. Kumar and P. K. Pathak. Two applications of Basu’s Lemma. Scandinavian Journal of Statistics. Theory and Applications, 4(1):37–38, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615647.

Kessler:2002:CAR

[KP02] Mathieu Kessler and Silvestre Paredes. Computational aspects related to martingale estimating functions for a discretely observed diffusion. Scandinavian Journal of Statistics. Theory and Applications, 29(3):425–440, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616725.

Kluppelberg:2021:ECC

[KP21] Claudia Klüppelberg and Viet Son Pham. Estimation of causal continuoustime autoregressive moving average random fields. Scandinavian Journal of Statistics. Theory and Applications, 48(1):132–163, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kulathinal:2022:E

[KPS22] Sangita Kulathinal, Jaakko Peltonen, and Mikko J. Sillanpää. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 49(1):1–3, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kulathinal:2023:PEA

[KPS23] Sangita Kulathinal, Jaakko Peltonen, and Mikko J. Sillanpää. Professor Elja Arjas: a prominent figure in establishing statistics in Finland. Scandinavian Journal of Statistics. Theory and Applications, 50(1): 1–2, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kulathinal:2024:E

[KPS24] Sangita Kulathinal, Jaakko Peltonen, and Mikko J. Sillanpää. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 51(2):427–428, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kessler:2001:ALB

[KR01] Mathieu Kessler and Anders Rahbek. Asymptotic likelihood based inference for co-integrated homogeneous Gaussian diffusions. Scandinavian Journal of Statistics. Theory and Applications, 28(3):455–470, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616672.

Klimova:2015:ISC

[KR15a] Anna Klimova and Tamás Rudas. Iterative scaling in curved exponential families. Scandinavian Journal of Statistics. Theory and Applications, 42(3):832–847, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kuhlmann:2015:EFL

[KR15b] Malte Kuhlmann and Claudia Redenbach. Estimation of fibre length distributions from fibre endpoints. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1010–1022, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kuljus:2015:GMS

[KR15c] Kristi Kuljus and Bo Ranneby. Generalized maximum spacing estimation for multivariate observations. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1092–1108, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kuljus:2020:ANG

[KR20] Kristi Kuljus and Bo Ranneby. Asymptotic normality of generalized maximum spacing estimators for multivariate observations. Scandinavian Journal of Statistics. Theory and Applications, 47(3):968–989, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kreiner:1987:AMC

[Kre87] Svend Kreiner. Analysis of multidimensional contingency tables by exact conditional tests: Techniques and strategies. Scandinavian Journal of Statistics. Theory and Applications, 14(2):97–112, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616054.

Kristiansen:1995:CCC

[Kri95] Gundorph K. Kristiansen. A counterexample to a conjecture concerning infinitely divisible distributions. Scandinavian Journal of Statistics. Theory and Applications, 22(1): 139–141, March 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616347.

Kollo:2007:ADL

[KRV07] Tõnu Kollo, Anu Roos, and Dietrich Von Rosen. Approximation of the distribution of the location parameter in the growth curve model. Scandinavian Journal of Statistics. Theory and Applications, 34(3): 499–510, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548566.

Kowalski:1988:LPN

[KS88] Aleksander Kowalski and Dominik Szynal. On linear predictors for nonstationary ARMA models. Scandinavian Journal of Statistics. Theory and Applications, 15(2):111–116, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616091.

Kuchler:1994:EFS

[KS94] Uwe Küchler and Michael Sørensen. Exponential families of stochastic processes with time-continuous likelihood functions. Scandinavian Journal of Statistics. Theory and Applications, 21(4):421–431, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616327.

Kotz:1999:BDE

[KS99] Samuel Kotz and Nozer D.

Singpurwalla. On a bivariate distribution with exponential marginals. Scandinavian Journal of Statistics. Theory and Applications, 26(3):451–464, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616567.

Kvam:2001:NPE

[KS01] Paul H. Kvam and Harshinder Singh. On nonparametric estimation of the survival function with competing risks. Scandinavian Journal of Statistics. Theory and Applications, 28(4):715–724, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616691.

Kleppe:2008:BFN

[KS08] Tore Selland Kleppe and Hans J. Skaug. Building and fitting non-Gaussian latent variable models via the moment-generating function. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 664–676, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000295.

Kirch:2022:SCP

[KS22] Claudia Kirch and Christina Stoehr. Sequential change point tests based on Ustatistics. Scandinavian Journal of Statistics. Theory and Applications, 49 (3):1184–1214, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Koul:1987:ESP

[KSM87] Hira L. Koul, Gerald L.

Sievers, and Joseph McKean. An estimator of the scale parameter for the rank analysis of linear models under general score functions. Scandinavian Journal of Statistics. Theory and Applications, 14(2): 131–141, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616057.

Koplowitz:1995:CCE

[KSN95] Jack Koplowitz, Jeffrey E.

Steif, and Olle Nerman. On Cover’s consistent estimator. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 395–397, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616368.

Kauermann:2013:FCD

[KSR13] Göran Kauermann, Christian Schellhase, and David Ruppert. Flexible copula density estimation with penalized hierarchical Bsplines. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 685–705, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586694.

Klebanov:2021:OPE

[KSSR21] Ilja Klebanov, Alexander Sikorski, Christof Schütte, and Susanna Röblitz. Objective priors in the empirical Bayes framework. Scandinavian Journal of Statistics. Theory and Applications, 48(4):1212–1233, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kooperberg:1995:RCH

[KST95] Charles Kooperberg, Charles J. Stone, and Young K. Truong. The L_{2} rate of convergence for hazard regression. Scandinavian Journal of Statistics. Theory and Applications, 22(2): 143–157, June 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616350.

Knautz:1995:SBB

[KT95] Henning Knautz and Götz Trenkler. Some bounds for bias and variance of S^{2} under dependence. Scandinavian Journal of Statistics. Theory and Applications, 22(1):121–128, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616345.

Karavias:2019:GFP

[KT19] Yiannis Karavias and Elias Tzavalis. Generalized fixed-T panel unit root tests. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1227–1251, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kuhnt:2004:OIP

[Kuh04] Sonja Kuhnt. Outlier identification procedures for contingency tables using maximum likelihood and L_{1} estimates. Scandinavian Journal of Statistics. Theory and Applications, 31(3): 431–442, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616840.

Kunsch:1983:AML

[Kün83] H. Künsch. Approximations to the maximum likelihood equations for some Gaussian random fields. Scandinavian Journal of Statistics. Theory and Applications, 10(3):239–246, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615921.

Kuroki:2016:IDI

[Kur16] Manabu Kuroki. The identification of direct and indirect effects in studies with an unmeasured intermediate variable. Scandinavian Journal of Statistics. Theory and Applications, 43 (1):228–245, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kurisu:2018:PVT

[Kur18] Daisuke Kurisu. Power variations and testing for co-jumps: The small noise approach. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 482–512, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kutoyants:2019:CLE

[Kut19] Yury A. Kutoyants. On cusp location estimation for perturbed dynamical systems. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 1206–1226, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kollo:1998:UAA

[KV98] Tõnu Kollo and Dietrich Von Rosen. A unified approach to the approximation of multivariate densities. Scandinavian Journal of Statistics. Theory and Applications, 25(1):93– 109, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616487.

Kalogridis:2023:ROE

[KV23] Ioannis Kalogridis and Stefan Van Aelst. Robust optimal estimation of location from discretely sampled functional data. Scandinavian Journal of Statistics. Theory and Applications, 50(2):411–451, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kabaila:2016:MAC

[KWA16] Paul Kabaila, A. H. Welsh, and Waruni Abeysekera. Model-averaged confidence intervals. Scandinavian Journal of Statistics. Theory and Applications, 43(1): 35–48, March 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kim:2020:SPS

[KXZA20] Soyoung Kim, Yayun Xu, Mei-Jie Zhang, and Kwang-Woo Ahn. Stratified proportional subdistribution hazards model with covariate-adjusted censoring weight for case-cohort studies. Scandinavian Journal of Statistics. Theory and Applications, 47(4): 1222–1242, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kojadinovic:2012:NPT

[KY12] Ivan Kojadinovic and Jun Yan. A non-parametric test of exchangeability for extreme-value and left-tail decreasing bivariate copulas. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 480–496, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41679807.

Kong:2021:HDR

[KYZC21] Shengchun Kong, Zhuqing Yu, Xianyang Zhang, and Guang Cheng. Highdimensional robust inference for Cox regression models using desparsified lasso. Scandinavian Journal of Statistics. Theory and Applications, 48(3):1068– 1095, September 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Kratschmer:2017:SIE

[KZ17] Volker Krätschmer and Henryk Zähle. Statistical inference for expectilebased risk measures. Scandinavian Journal of Statistics. Theory and Applications, 44(2):425–454, June 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

LaRocca:2008:BNP

[La 08] Luca La Rocca. Bayesian non-parametric estimation of smooth hazard rates for seismic hazard assessment. Scandinavian Journal of Statistics. Theory and Applications, 35(3):524–539, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000285.

Lunardon:2016:SOA

[LA16] Nicola Lunardon and Gianfranco Adimari. Secondorder accurate confidence regions based on members of the generalized power divergence family. Scandinavian Journal of Statistics. Theory and Applications, 43(1):213–227, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Laake:1975:OAO

[Laa75] Petter Laake. On the optimal allocation of observations in experiments with mixtures. Scandinavian Journal of Statistics.

Theory and Applications, 2 (3):153–157, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615596.

Laake:1978:ESE

[Laa78] Petter Laake. An evaluation of synthetic estimates of employment. Scandinavian Journal of Statistics. Theory and Applications, 5 (1):57–60, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615684.

Laake:1988:EAD

[Laa88] Petter Laake. Exact and asymptotic distributions of the maximum likelihood estimate in a simple factor analysis model. Scandinavian Journal of Statistics. Theory and Applications, 15(3):195–201, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616101.

Lauritzen:1989:MGA

[LAE^{+}89] Steffen L. Lauritzen, Anders Holst Andersen, David Edwards, Karl G. Jöreskog, and Søren Johansen. Mixed graphical association models [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 16(4): 273–306, March 1989. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616142.

Laippala:1979:EBA

[Lai79] Pekka Laippala. The empirical Bayes approach with floating optimal sample size in binomial experimentation. Scandinavian Journal of Statistics. Theory and Applications, 6(3):113–118, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615744. See correction [Lai80].

Laippala:1980:CNE

[Lai80] Pekka Laippala. Correction note: The Empirical Bayes Approach with Floating Optimal Sample Size in Binomial Experimentation. Scandinavian Journal of Statistics. Theory and Applications, 7(2): 105, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615783. See [Lai79].

Laippala:1983:SRC

[Lai83] Pekka Laippala. Simulation results concerning the empirical Bayes two-action rules with floating optimal sample size. Scandinavian Journal of Statistics. Theory and Applications, 10(2):87–96, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615908.

Liu:2012:SCB

[LAKZ12] Wei Liu, Pascal Ah-Kine, and Sanyu Zhou. Simultaneous confidence bands for linear regression with covariates constrained in intervals. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 543–553, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41679811.

Lanke:1974:IEM

[Lan74a] Jan Lanke. Interval estimation of a median. Scandinavian Journal of Statistics. Theory and Applications, 1(1):28–32, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615546.

Lanke:1974:IHR

[Lan74b] Jan Lanke. On an inequality of Hoeffding and Rosén. Scandinavian Journal of Statistics. Theory and Applications, 1(2):84– 86, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615556.

Langholz:2007:UCI

[Lan07] Bryan Langholz. Use of cohort information in the design and analysis of casecontrol studies. Scandinavian Journal of Statistics. Theory and Applications, 34(1):120–136, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548542.

Lang:2013:MLE

[Lan13] Joseph B. Lang. Maximum likelihood estimation for multinomial-Poisson models: a generalization of Birch’s numerical invariance results. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 530–548, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586685.

Lando:2023:TOS

[LAO23] Tommaso Lando, Idir Arab, and Paulo Eduardo Oliveira. Transform orders and stochastic monotonicity of statistical functionals. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1183–1200, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lauritzen:1974:SPE

[Lau74] Steffen L. Lauritzen. Sufficiency, prediction and extreme models. Scandinavian Journal of Statistics. Theory and Applications, 1 (3):128–134, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615564.

Lauritzen:1975:GEM

[Lau75] Steffen L. Lauritzen. General exponential models for discrete observations. Scandinavian Journal of Statistics. Theory and Applications, 2(1):23–33, December 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615572.

Lauritzen:1976:NCS

[Lau76] Steffen L. Lauritzen. Note on a case-study in Box– Jenkins seasonal forecasting of time series. Scandinavian Journal of Statistics. Theory and Applications, 3(2):93–94, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615617.

Lauritzen:1998:E

[Lau98] Steffen L. Lauritzen. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 25(1): 1, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616479.

Lauritzen:2004:DC

[Lau04] Steffen L. Lauritzen. Discussion on causality. Scandinavian Journal of Statistics. Theory and Applications, 31(2):189–192, June 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616823. See [AP04].

Lavine:2023:CDV

[Lav23]  Michael Lavine.  Com-

ments on Divergence vs. decision P -values: a distinction worth making in theory and keeping in practice — or, how divergence P -values measure evidence even when decision P - values do not by Greenland in Scandinavian Journal of Statistics, 2023. Scandinavian Journal of Statistics. Theory and Applications, 50(3):915–919, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Gre23a].

Lawrance:1982:IDG

[Law82] A. J. Lawrance. The innovation distribution of a gamma distributed autoregressive process. Scandinavian Journal of Statistics. Theory and Applications, 9 (4):234–236, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615888.

Ljung:1980:AVA

[LB80] Greta M. Ljung and George E. P. Box. Analysis of variance with autocorrelated observations. Scandinavian Journal of Statistics. Theory and Applications, 7(4): 172–180, September 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615797.

Li:1988:MJS

[LB88] Tze Fen Li and Dinesh S. Bhoj. A modified James– Stein estimator with application to multiple regression analysis. Scandinavian Journal of Statistics.

Theory and Applications, 15(1):33–37, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616082.

Linder:1994:BLF

[LB94] Ernst Linder and Gutti Jogesh Babu. Bootstrapping the linear functional relationship with known error variance ratio. Scandinavian Journal of Statistics. Theory and Applications, 21(1):21–39, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616296.

Lindkvist:1998:CNP

[LB98] Håkan Lindkvist and Yuri Belyaev. A class of nonparametric tests in the competing risks model for comparing two samples. Scandinavian Journal of Statistics. Theory and Applications, 25(1):143–150, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616490.

Lauritzen:1984:EPM

[LBND^{+}84] Steffen L. Lauritzen, Ole E.

Barndorff-Nielsen, A. P. Dawid, Persi Diaconis, and Søren Johansen. Extreme point models in statistics [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 11(2):65– 91, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615945.

Lindley:1978:BAD

[LBNE^{+}78] D. V. Lindley, O. Barndorff-Nielsen, Gustav Elfving, Erik Harsaae, Daniel Thorburn, Anders Hald, and Emil Spjötvoll. The Bayesian approach [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 5(1):1–26, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615681.

Lee:2000:BAG

[LC00a] J. C. Lee and C. H. Chang. Bayesian analysis of a growth curve model with a general autoregressive covariance structure. Scandinavian Journal of Statistics. Theory and Applications, 27(4):703–713, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616636.

Lingjaerde:2000:SSP

[LC00b] Ole C. Lingjærde and Nils Christophersen. Shrinkage structure of partial least squares. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 459–473, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616617.

Liquet:2011:CEB

[LC11a] Benoit Liquet and Daniel Commenges. Choice of estimators based on different observations: Modified AIC and LCV criteria. Scandinavian Journal of Statistics. Theory and Applications, 38(2): 268–287, June 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015494.

Liu:2011:GAM

[LC11b] Hai Liu and Kung-Sik Chan. Generalized additive models for zeroinflated data with partial constraints. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 650–665, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41411143.

Liu:2022:IET

[LC22] Xialu Liu and Elynn Y. Chen. Identification and estimation of threshold matrix-variate factor models. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1383– 1417, September 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Leonardi:2024:SRP

[LCF24] Florencia Leonardi, Rodrigo Carvalho, and Iara Frondana. Structure recovery for partially observed discrete Markov random fields on graphs under not necessarily positive distributions. Scandinavian Journal of Statistics. Theory and Applications, 51(1):64–88, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lin:2009:ATS

[LCZ09] Lu Lin, Xia Cui, and Lixing Zhu. An adaptive twostage estimation method for additive models. Scandinavian Journal of Statistics. Theory and Applications, 36(2):248–269, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000320.

Li:2014:VCT

[LCZ14] Zaixing Li, Fei Chen, and Lixing Zhu. Variance components testing in ANOVA-type mixed models. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 482–496, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586664.

Liu:2022:LSD

[LCZW22] Jialuo Liu, Tingjin Chu, Jun Zhu, and Haonan Wang. Large spatial data modeling and analysis: a Krylov subspace approach. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1115–1143, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lukas:2012:PRG

[LDA12] Mark A. Lukas, Frank R.

De Hoog, and Robert S. Anderssen. Performance of robust GCV and modified GCV for spline smoothing. Scandinavian Journal of Statistics. Theory and Applications, 39(1):97– 115, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411159.

Lundberg:1980:IES

[LdM80] Erik Lundberg and Jacques de Maré. Interval estimates in the spectroscopy calibration problem. Scandinavian Journal of Statistics. Theory and Applications, 7 (1):40–42, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615769.

Lindsten:2015:UEP

[LDM15] Fredrik Lindsten, Randal Douc, and Eric Moulines. Uniform ergodicity of the particle Gibbs sampler. Scandinavian Journal of Statistics. Theory and Applications, 42(3):775–797, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liang:2015:CLT

[LdUád15] Han-Ying Liang, Jacobo de Uña-álvarez, and María del Carmen Iglesias-Pérez. A central limit theorem in non-parametric regression with truncated, censored and dependent data. Scandinavian Journal of Statistics. Theory and Applications, 42(1):256–269, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Laud:2006:CAR

[LDW06] Purushottam W. Laud, Paul Damien, and Stephen G. Walker. Computations via auxiliary random functions for survival models. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 219–226, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616920.

Li:2016:GMM

[LDY16] Hui Li, Xiaogang Duan, and Guosheng Yin. Generalized method of moments for additive hazards model with clustered dental survival data. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1124–1139, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lee:1997:ENC

[Lee97] Chung-Bow Lee. Estimating the number of change points in exponential families distributions. Scandinavian Journal of Statistics. Theory and Applications, 24(2):201–210, June 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616448.

Lindskou:2020:ODC

[LET20] Mads Lindskou, Poul Svante Eriksen, and Torben Tvedebrink. Outlier detection in contingency tables using decomposable graphical models. Scandinavian Journal of Statistics. Theory and Applications, 47(2):347–360, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lindqvist:2022:CMC

[LET22] Bo H. Lindqvist, Rasmus Erlemann, and Gunnar Taraldsen. Conditional Monte Carlo revisited. Scandinavian Journal of Statistics. Theory and Applications, 49(3): 943–968, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lin:2016:SEE

[LFL16] Huazhen Lin, Zhe Fei, and Yi Li. A semiparametrically efficient estimator of the time-varying effects for survival data with time-dependent treatment. Scandinavian Journal of Statistics. Theory and Applications, 43(3):649–663, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2009:SIA

[LG09] Yehua Li and Marc G. Genton. Single-index additive vector autoregressive time series models. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 369–388, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000328.

Liu:2013:CCG

[LG13] Binghui Liu and Jianhua Guo. Collapsibility of conditional graphical models. Scandinavian Journal of Statistics. Theory and Applications, 40(2): 191–203, June 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586708.

Li:2019:WSB

[LGL19] Shuang Li, James Gleaton, and James Lynch. What is the shape of a bundle? an analysis of Rosen’s fibrous composites experiments using the chain-of-bundles model. Scandinavian Journal of Statistics. Theory and Applications, 46(1):59–86, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2011:ELI

[LGP11] Zhouping Li, Yun Gong, and Liang Peng. Empirical likelihood intervals for conditional value-at-risk in heteroscedastic regression models. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 781–787, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41411150.

Loland:2013:SCI

[LHHF13] Anders Løland, Ragnar Bang Huseby, Nils Lid Hjort, and Arnoldo Frigessi. Statistical corrections of invalid correlation matrices. Scandinavian Journal of Statistics. Theory and Applications, 40(4):807–824, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586701.

Li:2016:CVI

[LHML16] Jialiang Li, Zhipeng Huang, Shuangge Ma, and Mei-Ling Ting Lee. Collective versus individual effects in survival analysis of multiple failures. Scandinavian Journal of Statistics. Theory and Applications, 43(2):543–557, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lee:2003:CTP

[LHNN03] Sangyeol Lee, Jeongcheol Ha, Okyoung Na, and Seongryong Na. The Cusum test for parameter change in time series models. Scandinavian Journal of Statistics. Theory and Applications, 30(4):781–796, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616802.

Liu:2016:SCT

[LHW^{+}16] Wei Liu, Yang Han, Fang Wan, Frank Bretz, and Anthony J. Hayter. Simultaneous confidence tubes in multivariate linear regression. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 879–885, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2018:CST

[LHWS18] Shuwei Li, Tao Hu, Peijie Wang, and Jianguo Sun. A class of semiparametric transformation models for doubly censored failure time data. Scandinavian Journal of Statistics. Theory and Applications, 45(3):

682–698, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2001:QLE

[Li01] Bing Li. On quasi likelihood equations with nonparametric weights. Scandinavian Journal of Statistics. Theory and Applications, 28(4):577–602, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616683.

Lindvall:1976:MBP

[Lin76] Torgny Lindvall. On the maximum of a branching process. Scandinavian Journal of Statistics. Theory and Applications, 3 (4):209–214, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615636.

Lindqvist:1977:HFD

[Lin77] Bo Lindqvist. How fast does a Markov chain forget the initial state? a decision theoretical approach. Scandinavian Journal of Statistics. Theory and Applications, 4(4):145–152, September 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615670.

Lindgren:1978:MRM

[Lin78a] Georg Lindgren. Markov regime models for mixed distributions and switching regressions. Scandinavian Journal of Statistics. Theory and Applications, 5 (2):81–91, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615692.

Lindqvist:1978:NBT

[Lin78b] Bo Lindqvist. A note on Bernoulli trials with dependence. Scandinavian Journal of Statistics. Theory and Applications, 5(4):205– 208, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615717.

Lindqvist:1978:LII

[Lin78c] Bo Lindqvist. On the loss of information incurred by lumping states of a Markov chain. Scandinavian Journal of Statistics. Theory and Applications, 5(2):92– 98, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615693.

Lindqvist:1988:ETM

[Lin88] Bo Henry Lindqvist. Estimation and testing in Markov models for repairable systems with positively dependent components. Scandinavian Journal of Statistics. Theory and Applications, 15(4): 243–257, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616109.

Lindqvist:1994:MPU

[Lin94] Bo Lindqvist. The minimal path upper bound for the moments of a reliability function. Scandinavian Journal of Statistics. Theory and Applications, 21(1):83–90, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616300.

Lin:2000:NKA

[Lin00] Yan-Xia Lin. A new kind of asymptotic quasi-score estimating function. Scandinavian Journal of Statistics. Theory and Applications, 27(1):97–109, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616593.

Lin:2014:GSS

[Lin14] Pei-Sheng Lin. Generalized scan statistics for disease surveillance. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 791–808, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586759.

Lv:2018:EHD

[LJZ^{+}18] Shaogao Lv, Jiakun Jiang, Fanyin Zhou, Jian Huang, and Huazhen Lin. Estimating high-dimensional additive Cox model with time-dependent covariate processes. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 900–922, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Leucht:2015:MST

[LKN15] Anne Leucht, Jens-Peter Kreiss, and Michael H. Neumann. A model specification test for GARCH(1,1) processes. Scandinavian Journal of Statistics. Theory and Applications, 42(4): 1167–1193, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2023:GIC

[LKT^{+}23] Fan Li, Jessica Kasza, Elizabeth L. Turner, Paul J. Rathouz, Andrew B. Forbes, and John S. Preisser. Generalizing the information content for stepped wedge designs: a marginal modeling approach. Scandinavian Journal of Statistics. Theory and Applications, 50 (3):1048–1067, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lehmann:1990:PVU

[LL90] E. L. Lehmann and Wei-Yin Loh. Pointwise versus uniform robustness of some large-sample tests and confidence intervals. Scandinavian Journal of Statistics. Theory and Applications, 17(3):177–187, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616167.

Li:1996:CST

[LL96] Bing Li and Bruce G. Lindsay. Chi-square tests for generalized estimating equations with possibly misspecified weights. Scandinavian Journal of Statistics. Theory and Applications, 23(4):489–509, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616422.

Ludena:1999:WES

[LL99] Carenne Ludeña and Marc Lavielle. The Whittle estimator for strongly dependent stationary Gaussian fields. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 433–450, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616566.

Li:2006:RBD

[LL06] Qian H. Li and Stephen W.

Lagakos. On the relationship between directional and omnibus statistical tests. Scandinavian Journal of Statistics. Theory and Applications, 33(2):239–246, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616922.

Lee:2009:NMQ

[LL09] Taewook Lee and Sangyeol Lee. Normal mixture quasimaximum likelihood estimator for GARCH models. Scandinavian Journal of Statistics. Theory and Applications, 36(1): 157–170, March 2009. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000313.

Lee:2012:IBC

[LL12] Sangyeol Lee and Taewook Lee. Inference for Box–Cox transformed threshold GARCH models with nuisance parameters. Scandinavian Journal of Statistics. Theory and Applications, 39(3):568–589, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679813.

Lageraas:2020:HAS

[LL20] Andreas Lagerås and Mathias Lindholm. How to ask sensitive multiple-choice questions. Scandinavian Journal of Statistics. Theory and Applications, 47 (2):397–424, June 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lin:2021:GSR

[LLCW21] Lu Lin, Lili Liu, Xia Cui, and Kangning Wang. A generalized semiparametric regression and its efficient estimation. Scandinavian Journal of Statistics. Theory and Applications, 48(1): 1–24, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liu:2020:HTQ

[LLLP20] Guanfu Liu, Pengfei Li, Yukun Liu, and Xiaolong Pu. Hypothesis testing for quantitative trait locus effects in both location and scale in genetic backcross studies. Scandinavian Journal of Statistics. Theory and Applications, 47(4): 1064–1089, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lloyd:1988:BNT

[Llo88] Christopher J. Lloyd. Bivariate normal transformation models. Scandinavian Journal of Statistics. Theory and Applications, 15(3): 177–185, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616099.

Liu:2022:EEE

[LLS^{+}22] Lan Liu, Wei Li, Zhihua Su, Dennis Cook, Luca Vizioli, and Essa Yacoub. Efficient estimation via envelope chain in magnetic resonance imaging-based studies. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 481–501, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liu:2019:CRC

[LLXH19] Li Liu, Yanyan Liu, Yi Xiong, and X. Joan Hu. Cox regression of clustered event times with covariates missing not at random. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1315–1346, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2017:SRM

[LLY17] Rui Li, Chenlei Leng, and Jinhong You. A semiparametric regression model for longitudinal data with nonstationary errors. Scandinavian Journal of Statistics. Theory and Applications, 44(4):932–950, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:2018:JCL

[LLY18] Weiming Li, Zeng Li, and Jianfeng Yao. Joint central limit theorem for eigenvalue statistics from several dependent large dimensional sample covariance matrices with application. Scandinavian Journal of Statistics. Theory and Applications, 45(3):699–728, September 2018. CODEN SJSADG.

ISSN 0303-6898 (print), 1467-9469 (electronic).

Liu:2022:SEM

[LLYC22] Guannan Liu, Wei Long, Bingduo Yang, and Zongwu Cai. Semiparametric estimation and model selection for conditional mixture copula models. Scandinavian Journal of Statistics. Theory and Applications, 49 (1):287–330, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Letac:2004:AIM

[LM04] Gérard Letac and Hélène Massam. All invariant moments of the Wishart distribution. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 295–318, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616830.

Lavancier:2016:MAL

[LM16] Frédéric Lavancier and Jesper Møller. Modelling aggregation on the large scale and regularity on the small scale in spatial point pattern datasets. Scandinavian Journal of Statistics. Theory and Applications, 43(2):587–609, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lopuhaa:2018:SIE

[LM18] Hendrik P. Lopuhaä and Eni Musta. Smoothed isotonic estimators of a monotone baseline hazard in the Cox model. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 753–791, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lingohr:2023:CTT

[LM23] Daniel Lingohr and Gernot Müller. Continuoustime threshold autoregressions with jumps: Properties, estimation, and application to electricity markets. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 638–664, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lejay:2024:MLE

[LM24] Antoine Lejay and Sara Mazzonetto. Maximum likelihood estimator for skew Brownian motion: The convergence rate. Scandinavian Journal of Statistics. Theory and Applications, 51(2):612–642, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lupparelli:2009:PFB

[LMB09] Monia Lupparelli, Giovanni M. Marchetti, and Wicher P. Bergsma. Parameterizations and fitting of Bi-directed graph models to categorical data. Scandinavian Journal of Statistics. Theory and Applications, 36(3):559–576, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000338.

Liu:2014:IAC

[LMH14] Jin Liu, Shuangge Ma, and Jian Huang. Integrative analysis of cancer diagnosis studies with composite penalization. Scandinavian Journal of Statistics. Theory and Applications, 41(1):87–103, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586732.

Lund:2022:SME

[LMH22] Adam Lund, Søren Wengel Mogensen, and Niels Richard Hansen. Soft maximin estimation for heterogeneous data. Scandinavian Journal of Statistics. Theory and Applications, 49(4): 1761–1790, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lejay:2014:BMS

[LMT14] Antoine Lejay, Ernesto Mordecki, and Soledad Torres. Is a Brownian motion skew? Scandinavian Journal of Statistics. Theory and Applications, 41(2):346–364, June 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586657.

Liski:1995:PTS

[LN95] Erkki P. Liski and Tapio Nummi. Prediction of tree stems to improve efficiency in automatized harvesting of forests. Scandinavian Journal of Statistics. Theory and Applications, 22(2): 255–269, June 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616356.

Lee:2013:QRE

[LN13a] Sangyeol Lee and Jungsik Noh. Quantile regression estimator for GARCH models. Scandinavian Journal of Statistics. Theory and Applications, 40(1):2– 20, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357250.

Lopuhaa:2013:SCN

[LN13b] Hendrik P. Lopuhaä and Gabriela F. Nane. Shape constrained non-parametric estimators of the baseline distribution in Cox proportional hazards model. Scandinavian Journal of Statistics. Theory and Applications, 40(3):619–646, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586689.

Lo:1981:BNS

[Lo81] Albert Y. Lo. Bayesian nonparametric statistical inference for shock models and wear processes. Scandinavian Journal of Statistics. Theory and Applications, 8(4):237–242, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615842.

Li:2016:FCM

[LO16] Zhiguo Li and Kouros Owzar. Fitting Cox models with doubly censored data using spline-based sieve marginal likelihood. Scandinavian Journal of Statistics. Theory and Applications, 43(2):476–486, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lok:2007:SNM

[Lok07] Judith J. Lok. Structural nested models and standard software: a mathematical foundation through partial likelihood. Scandinavian Journal of Statistics. Theory and Applications, 34 (1):186–206, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548546.

Longford:2012:CNR

[Lon12] Nicholas T. Longford. Comparing normal random samples, with uncertainty about the priors and utilities. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 729–742, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357221.

Louani:1998:LDL

[Lou98] Djamal Louani. Large deviations limit theorems for the kernel density estimator. Scandinavian Journal of Statistics. Theory and Applications, 25(1): 243–253, March 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616497.

Lee:2001:CST

[LP01] Sangyeol Lee and Siyun Park. The Cusum of squares test for scale changes in infinite order moving average processes. Scandinavian Journal of Statistics. Theory and Applications, 28(4):625–644, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616686.

Lejay:2020:MLD

[LP20] Antoine Lejay and Paolo Pigato. Maximum likelihood drift estimation for a threshold diffusion. Scandinavian Journal of Statistics. Theory and Applications, 47(3):609–637, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ljungdahl:2022:MPE

[LP22] Mathias Mørck Ljungdahl and Mark Podolskij. Multidimensional parameter estimation of heavy-tailed moving averages. Scandinavian Journal of Statistics. Theory and Applications, 49(2):593–624, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lin:2015:NBD

[LPB15] Lizhen Lin, Walter W. Piegorsch, and Rabi Bhattacharya. Nonparametric benchmark dose estimation with continuous doseresponse data. Scandinavian Journal of Statistics. Theory and Applications, 42(3):713–731, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lacayo:1982:OSD

[LPPS82] Herbert Lacayo, Carlos A. B. Pereira, Frank Proschan, and Carl Erik Särndal. Optimal sample depends on optimality criterion. Scandinavian Journal of Statistics. Theory and Applications, 9(1):47– 48, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615854.

Lu:2022:AML

[LPPW22] Kevin W. Lu, Phillip J. Paine, Simon P. Preston, and Andrew T. A. Wood. Approximate maximum likelihood estimation for one-dimensional diffusions observed on a fine grid. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1085– 1114, September 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lijoi:2023:FCH

[LPR23] Antonio Lijoi, Igor Prünster, and Giovanni Rebaudo. Flexible clustering via hidden hierarchical Dirichlet priors. Scandinavian Journal of Statistics. Theory and Applications, 50 (1):213–234, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liu:2003:THG

[LPS03] Xin Liu, Cristian Pasarica, and Yongzhao Shao. Testing homogeneity in gamma mixture models. Scandinavian Journal of Statistics. Theory and Applications, 30(1):227–239, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616759.

Lavancier:2021:AEF

[LPW21] Frédéric Lavancier, Arnaud Poinas, and Rasmus Waagepetersen. Adaptive estimating function inference for nonstationary determinantal point processes. Scandinavian Journal of Statistics. Theory and Applications, 48(1):87– 107, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liu:2024:NCM

[LPYZ24] Yiming Liu, Guangming Pan, Guangren Yang, and Wang Zhou. Nonparametric conditional mean testing via an extreme-type statistic in high dimension. Scandinavian Journal of Statistics. Theory and Applications, 51(2):801–831, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Luo:2017:AOS

[LQ17] Ruiyan Luo and Xin Qi. Asymptotic optimality of sparse linear discriminant analysis with arbitrary number of classes. Scandinavian Journal of Statistics. Theory and Applications, 44(3):598–616, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liang:2009:ELB

[LQZR09] Hua Liang, Yongsong Qin, Xinyu Zhang, and David Ruppert. Empirical likelihood-based inferences for generalized partially linear models. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 433–443, September 2009.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000331.

Landers:1976:NC

[LR76] D. Landers and L. Rogge.

A note on completeness. Scandinavian Journal of Statistics. Theory and Applications, 3(3):139, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615628.

Li:2006:ISD

[LR06] Yi Li and Louise Ryan.

Inference on survival data with covariate measurement error — an imputationbased approach. Scandinavian Journal of Statistics. Theory and Applications, 33(2):169–190, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616917.

Lindgren:2008:SOR

[LR08] Finn Lindgren and Håvard Rue. On the second-order random walk model for irregular locations. Scandinavian Journal of Statistics. Theory and Applications, 35(4):691–700, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000297.

Lie:2023:DIM

[LRSS23] Han Cheng Lie, Daniel Rudolf, Björn Sprungk, and T. J. Sullivan. Dimensionindependent Markov chain Monte Carlo on the sphere. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1818–1858, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lindgren:1987:EVT

[LRT^{+}87] Georg Lindgren, Holger Rootzén, Dag Tjøstheim, Richard A. Davis, L. Nilsson, S. Uvell, Anders Milhøj, and Jacques de Maré. Extreme values: Theory and technical applications [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 14(4):241–279, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616070.

Lu:2023:DAI

[LRT23] Xiaoyu Lu, Tom Rainforth, and Yee Whye Teh. Daisee: Adaptive importance sampling by balancing exploration and exploitation.

Scandinavian Journal of Statistics. Theory and Applications, 50(3):1298–1324, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lin:1996:MCT

[LS96] D. Y. Lin and C. F. Spiekerman. Model checking techniques for parametric regression with censored data. Scandinavian Journal of Statistics. Theory and Applications, 23(2): 157–177, June 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616394.

Lawrance:1998:CCM

[LS98] A. J. Lawrance and N. M.

Spencer. Curved chaotic map time series models and their stochastic reversals. Scandinavian Journal of Statistics. Theory and Applications, 25(2): 371–382, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616508.

Lu:2015:EEP

[LS15] Xuewen Lu and Peter X.-K.

Song. Efficient estimation of the partly linear additive hazards model with current status data. Scandinavian Journal of Statistics. Theory and Applications, 42 (1):306–328, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lonnstedt:2023:RUV

[LS23] Ingrid M. Lönnstedt and Terence P. Speed. Remove unwanted variation retrieves unknown experimental designs. Scandinavian Journal of Statistics. Theory and Applications, 50(1):89–101, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Luukkonen:1988:TLU

[LST88] Ritva Luukkonen, Pentti Saikkonen, and Timo Teräsvirta. Testing linearity in univariate time series models. Scandinavian Journal of Statistics. Theory and Applications, 15(3):161–175, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616098.

Leskinen:1977:EDP

[LT77] Esko Leskinen and Timo Teräsvirta. On the estimation of dynamic price elasticities with Box–Jenkins methods: a case study. Scandinavian Journal of Statistics. Theory and Applications, 4(1):1–10, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615641.

Li:2008:NPE

[LT08] Gang Li and Chi-Hong Tseng. Non-parametric estimation of a survival function with two-stage design studies. Scandinavian Journal of Statistics. Theory and Applications, 35(2):193–211, June 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548588.

Loe:2021:EUB

[LT21] Margrethe Kvale Loe and Håkon Tjelmeland. Ensemble updating of binary state vectors by maximizing the expected number of unchanged components. Scandinavian Journal of Statistics. Theory and Applications, 48(4):1148–1185, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lu:2021:PDT

[Lu21]  Yang Lu.  The pre-

dictive distributions of thinning-based count processes. Scandinavian Journal of Statistics. Theory and Applications, 48(1):42– 67, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ludena:2004:MCE

[Lud04] Carenne Ludeña. Minimum contrast estimation for fractional diffusions. Scandinavian Journal of Statistics. Theory and Applications, 31(4):613–628, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616854.

Lue:2015:IRM

[Lue15] Heng-Hui Lue. An inverseregression method of dependent variable transformation for dimension reduction with non-linear confounding. Scandinavian Journal of Statistics. Theory and Applications, 42(3): 760–774, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lund:2000:SBP

[Lun00] Jens Lund. Sampling bias in population studies — how to use the Lexis diagram. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 589–604, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616629.

Luschgy:1994:ABN

[Lus94] Harald Luschgy. Asymptotic behavior of Neyman– Pearson tests for autoregressive processes. Scandinavian Journal of Statistics. Theory and Applications, 21(4):461–473, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616330.

Li:2002:LRC

[LV02] Gang Li and Ingrid Van Keilegom. Likelihood ratio confidence bands in non– parametric regression with censored data. Scandinavian Journal of Statistics. Theory and Applications, 29(3):547–562, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616731.

Li:2013:EBM

[LV13] Feng Li and Mattias Villani. Efficient Bayesian multivariate surface regression. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 706–723, December 2013.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586695.

Lopez:2009:ELN

[LVV09] Elisa M. Molanes Lopez, Ingrid Van Keilegom, and Noël Veraverbeke. Empirical likelihood for nonsmooth criterion functions. Scandinavian Journal of Statistics. Theory and Applications, 36(3):413–432, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000330.

Ledwina:2012:TST

[LW12] Teresa Ledwina and Grzegorz Wyłupek. Two-sample test against one-sided alternatives. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 358–381, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679799.

Liang:2023:RMA

[LW23] Zhongqi Liang and Qihua Wang. A robust model averaging approach for partially linear models with responses missing at random. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1933–1952, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Li:1997:EAS

[LWY97] Linxiong Li, Terry Watkins, and Qiqing Yu. An EM algorithm for smoothing the self-consistent estimator of survival functions with interval-censored data. Scandinavian Journal of Statistics. Theory and Applications, 24(4):531–542, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616474.

Li:2016:NEF

[LXZ16] Zhouping Li, Jinfeng Xu, and Wang Zhou. On nonsmooth estimating functions via jackknife empirical likelihood. Scandinavian Journal of Statistics. Theory and Applications, 43(1):49– 69, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Luna:2003:BKP

[LY03] Sara Sjöstedt-De Luna and Alastair Young. The bootstrap and kriging prediction intervals. Scandinavian Journal of Statistics. Theory and Applications, 30 (1):175–192, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616756.

Lau:2008:EBR

[LY08] Eric H. Y. Lau and Paul S. F. Yip. Estimating the basic reproductive number in the general epidemic model with an unknown initial number of susceptible individuals. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 650–663, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000294.

Lynch:1988:MGC

[Lyn88] James Lynch. Mixtures, generalized convexity and balayages. Scandinavian Journal of Statistics. Theory and Applications, 15(3): 203–210, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616102.

Liu:2022:ETC

[LYW22] Xin Liu, Rong-Xian Yue, and Weng Kee Wong. Equivalence theorems for c and D_{A}-optimality for linear mixed effects models with applications to multitreatment group assignments in health care. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1842–1859, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Leung:2015:ESS

[LYZ15] Denis Heng Yan Leung, Ken Yamada, and Biao Zhang. Enriching surveys with supplementary data and its application to studying wage regression. Scandinavian Journal of Statistics. Theory and Applications, 42(1):155–179, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Loh:1997:EMD

[LZ97] Wei-Liem Loh and Cun-Hui Zhang. Estimating mixing densities in exponential family models for discrete variables. Scandinavian Journal of Statistics. Theory and Applications, 24(1):15–32, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616434.

Liese:1999:NEP

[LZ99] F. Liese and K. Ziegler.

A note on empirical process methods in the theory of Poisson point processes. Scandinavian Journal of Statistics. Theory and Applications, 26(4): 533–537, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616577.

Liang:2008:SIR

[LZ08] Hua Liang and Yong Zhou. Semiparametric inference for ROC curves with censoring. Scandinavian Journal of Statistics. Theory and Applications, 35(2): 212–227, June 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548589.

Li:2010:VCS

[LZ10] Zaixing Li and Lixing Zhu. On variance components in semiparametric mixed models for longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 442–457, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000400.

Lee:2023:RQR

[LZC23] Danhyang Lee, Li-Chun Zhang, and Sixia Chen. Robust quasi-randomization-based estimation with ensemble learning for missing data. Scandinavian Journal of Statistics. Theory and Applications, 50 (3):1263–1278, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Liu:2024:NPH

[LZL^{+}24] Jinyuan Liu, Xinlian Zhang, Tuo Lin, Ruohui Chen, Yuan Zhong, Tian Chen, Tsungchin Wu, Chenyu Liu, Anna Huang, Tanya T. Nguyen, Ellen E. Lee, Dilip V. Jeste, and Xin M. Tu. A new paradigm for high-dimensional data: Distance-based semiparametric feature aggregation framework via betweensubject attributes. Scandinavian Journal of Statistics. Theory and Applications, 51(2):672–696, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Lin:2014:SRA

[LZZ14] Huazhen Lin, Ling Zhou, and Xiaohua Zhou. Semiparametric regression analysis of longitudinal skewed data. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1031–1050, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586819.

Mabon:2017:ADN

[Mab17] Gwennaëlle Mabon. Adaptive deconvolution on the non-negative real line. Scandinavian Journal of Statistics. Theory and Applications, 44(3):707–740, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mack:1982:GFP

[Mac82] Y. P. Mack. On a goodnessof-fit problem of some nonparametric density estimates. Scandinavian Journal of Statistics. Theory and Applications, 9(3):179– 182, September 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615876.

MacEachern:1993:CSC

[Mac93] Steven N. MacEachern. A characterization of some conjugate prior distributions for exponential families. Scandinavian Journal of Statistics. Theory and Applications, 20(1):77– 82, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616262.

Madsen:1976:SAM

[Mad76] Mette Madsen. Statistical analysis of multiple contingency tables. two examples. Scandinavian Journal of Statistics. Theory and Applications, 3(3):97–106, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615621.

Makowski:2005:DT

[Mak05] Armand M. Makowski.

Discussion of “Teletraffic”. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 217–218, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616872. See [Nor05].

Mammen:1992:HOA

[Mam92] E. Mammen. Higher order accuracy of the bootstrap method for smooth functionals. Scandinavian Journal of Statistics. Theory and Applications, 19(3): 255–269, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616243.

Mancini:2009:NPT

[Man09]  Cecilia Mancini.  Non-

parametric threshold estimation for models with stochastic diffusion coefficient and jumps. Scandinavian Journal of Statistics. Theory and Applications, 36(2):270–296, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000321.

Marohn:1998:AET

[Mar98] Frank Marohn. An adaptive efficient test for Gumbel domain of attraction. Scandinavian Journal of Statistics. Theory and Applications, 25(2):311–324, June 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616503.

Martinussen:1999:CRI

[Mar99] Torben Martinussen. Cox regression with incomplete covariate measurements using the EM-algorithm. Scandinavian Journal of Statistics. Theory and Applications, 26(4):479–491, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616574.

Martino:2011:ABI

[MAR11] Sara Martino, Rupali Akerkar, and Håvard Rue. Approximate Bayesian inference for survival models. Scandinavian Journal of Statistics. Theory and Applications, 38(3):514–528, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015578.

Mathiasen:1979:PF

[Mat79] P. E. Mathiasen. Prediction functions. Scandinavian Journal of Statistics. Theory and Applications, 6(1):1–21, June 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615723.

Moyeed:1991:SAM

[MB91] R. A. Moyeed and A. J.

Baddeley. Stochastic approximation of the MLE for a spatial point pattern. Scandinavian Journal of Statistics. Theory and Applications, 18(1):39–50, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616187.

Munch:2023:TES

[MBMG23] Anders Munch, Marie Skov Breum, Torben Martinussen, and Thomas A. Gerds. Targeted estimation of state occupation probabilities for the non-Markov illness-death model. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1532–1551, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

McSwiggan:2017:KDE

[MBN17] Greg McSwiggan, Adrian Baddeley, and Gopalan Nair. Kernel density estimation on a linear network. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 324–345, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Moreno:2003:BIU

[MBR03] Elias Moreno, Francesco Bertolino, and Walter Racugno. Bayesian inference under partial prior information. Scandinavian Journal of Statistics. Theory and Applications, 30(3): 565–580, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616784.

Morgan:1997:CCE

[MC97] Peter B. Morgan and Noel Cressie. A comparison of the cost-efficiencies of the sequential, groupsequential, and variablesample-size-sequential probability ratio tests. Scandinavian Journal of Statistics. Theory and Applications, 24(2):181–200, June 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616447.

Mukerjee:2003:ELP

[MC03] Rahul Mukerjee and Zehua Chen. On expected lengths of predictive intervals. Scandinavian Journal of Statistics. Theory and Applications, 30(4): 757–766, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616800.

McGregor:1988:ANM

[McG88] J. R. McGregor. Asymptotic normality of the munch numbers. Scandinavian Journal of Statistics. Theory and Applications, 15(4):299–301, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616113.

McKenzie:1987:IDG

[McK87] Ed McKenzie. Innovation distributions for gamma and negative binomial autoregressions. Scandinavian Journal of Statistics. Theory and Applications, 14(1): 79–85, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616050.

Moller:2010:SST

[MDA10] Jesper Møller and Carlos Díaz-Avalos. Structured spatio-temporal shot-noise Cox point process models, with a view to modelling forest fires. Scandinavian Journal of Statistics. Theory and Applications, 37(1): 2–25, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000913.

Mhalla:2019:RTM

[MdCCD19] L. Mhalla, M. de Carvalho, and V. Chavez-Demoulin. Regression-type models for extremal dependence. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 1141–1167, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Meinshausen:2006:FDC

[Mei06] Nicolai Meinshausen. False discovery control for multiple tests of association under general dependence. Scandinavian Journal of Statistics. Theory and Applications, 33(2):227–237, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616921.

Meitz:2024:SIG

[Mei24]  Mika Meitz.  Statisti-

cal inference for generative adversarial networks and other minimax problems. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1323–1356, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mejza:1985:THM

[Mej85] Stanisław Mejza. On testing hypotheses in a mixed linear model for incomplete block designs. Scandinavian Journal of Statistics. Theory and Applications, 12(3): 241–247, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615993.

Messer:2022:BCP

[Mes22] Michael Messer. Bivariate change point detection: Joint detection of changes in expectation and variance. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 886–916, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Masry:1997:LPE

[MF97] Elias Masry and Jianqing Fan. Local polynomial estimation of regression functions for mixing processes. Scandinavian Journal of Statistics. Theory and Applications, 24(2): 165–179, June 1997. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616446.

Mangin:1995:LAP

[MG95] B. Mangin and B. Goffinet. Local asymptotic power of maximum likelihood ratio tests when the information matrix is singular, with an application in genetics. Scandinavian Journal of Statistics. Theory and Applications, 22(4):465–475, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616374.

Murdoch:1998:ESC

[MG98] D. J. Murdoch and P. J.

Green. Exact sampling from a continuous state space. Scandinavian Journal of Statistics. Theory and Applications, 25(3): 483–502, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616516.

Ma:2004:FCS

[MG04] Yanyuan Ma and Marc G.

Genton. Flexible class of skew-symmetric distributions. Scandinavian Journal of Statistics. Theory and Applications, 31(3): 459–468, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616842.

Menezes:2008:KVE

[MGSFB08] Raquel Menezes, Pilar Garcia-Soidán, and Manuel Febrero-Bande. A kernel variogram estimator for clustered data. Scandinavian Journal of Statistics. Theory and Applications, 35(1):18–37, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000383.

Miao:1997:EML

[MH97] Weiwen Miao and Marjorie Hahn. Existence of maximum likelihood estimates for multi-dimensional exponential families. Scandinavian Journal of Statistics. Theory and Applications, 24(3):371–386, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616462.

Meintanis:2010:GFT

[MH10a] Simos G. Meintanis and Zdeněk Hlávka. Goodnessof-fit tests for bivariate and multivariate skew-normal distributions. Scandinavian Journal of Statistics. Theory and Applications, 37(4): 701–714, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000416.

Moller:2010:LIU

[MH10b] Jesper Møller and Kateřina Helisová. Likelihood inference for unions of interacting discs. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 365–381, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000396.

Michel:2009:PEP

[Mic09] RenÉ Michel. Parametric estimation procedures in multivariate generalized Pareto models. Scandinavian Journal of Statistics. Theory and Applications, 36(1):60–75, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000307.

Milhoj:1985:MSA

[Mil85] Anders Milhøj. The moment structure of ARCH processes. Scandinavian Journal of Statistics. Theory and Applications, 12(4): 281–292, June 1985. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616000.

Mineo:1979:NGM

[Min79] A. Mineo. A new grouping method for the right evaluation of the chi-square test of goodness-of-fit. Scandinavian Journal of Statistics. Theory and Applications, 6 (4):145–153, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615753.  See

note [BR81] and rejoinder [Min81].

Mineo:1981:RBR

[Min81] A. Mineo. Rejoinder to Best and Rayner’s: A Note on Mineo’s Grouping Method for the Chi-Square Test of Goodness of Fit. Scandinavian Journal of Statistics. Theory and Applications, 8(3):187–188, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4615833. See [BR81].

Miura:1978:TMI

[Miu78] Carole K. Miura. Tests for the mean of the inverse Gaussian distribution. Scandinavian Journal of Statistics. Theory and Applications, 5(4):200– 204, June 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615716.

Miura:1981:SEA

[Miu81]  Ryozo Miura.  Spacing

estimation of the asymptotic variance of trimmed rank estimators of location. Scandinavian Journal of Statistics. Theory and Applications, 8(1):48– 54, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615812.

Martin-Lof:1974:NRU

[ML74] Per Martin-Löf. The notion of redundancy and its use as a quantitative measure of the discrepancy between a statistical hypothesis and a set of observational data [with discussion]. Scandinavian Journal of Statistics. Theory and Applications, 1(1):3– 18, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615544.

Martin-Lof:1975:RSP

[ML75] Per Martin-Löf. Reply to Sverdrup’s polemical article Tests without Power. Scandinavian Journal of Statistics. Theory and Applications, 2(3):161–165, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615598. See [Sve75].

Martin-Lof:1986:EEF

[ML86] Anders Martin-Löf. Entropy estimates for the first passage time of a random walk to a time dependent barrier. Scandinavian Journal of Statistics. Theory and Applications, 13(3): 221–229, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616028.

Marzec:1993:GFI

[MM93] Leszek Marzec and Paweł Marzec. Goodness of fit inference based on stratification in Cox’s regression model. Scandinavian Journal of Statistics. Theory and Applications, 20(3): 227–238, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616278.

Masarotto:2024:CBS

[MM24] Valentina Masarotto and Guido Masarotto. Covariancebased soft clustering of functional data based on the Wasserstein–Procrustes metric. Scandinavian Journal of Statistics. Theory and Applications, 51(2): 485–512, June 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Martin:2023:MGA

[MMO23] James S. Martin, David J. Murrell, and Sofia C. Olhede. Multivariate geometric anisotropic Cox processes. Scandinavian Journal of Statistics. Theory and Applications, 50 (3):1420–1465, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mansourvar:2016:AMR

[MMS16] Zahra Mansourvar, Torben Martinussen, and Thomas H. Scheike. An additive–multiplicative restricted mean residual life model. Scandinavian Journal of Statistics. Theory and Applications, 43(2): 487–504, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mohr:2021:NVC

[MN21] Maria Mohr and Natalie Neumeyer. Nonparametric volatility change detection. Scandinavian Journal of Statistics. Theory and Applications, 48(2): 529–548, June 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Munk:2007:NPA

[MNS07] Axel Munk, Natalie Neumeyer, and Achim Scholz. Nonparametric analysis of covariance — the case of inhomogeneous and heteroscedastic noise. Scandinavian Journal of Statistics. Theory and Applications, 34(3):511–534, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548567.

Moller:1976:RWP

[Møl76] Svend Kreiner Møller. The Rasch–Weibull process. Scandinavian Journal of Statistics. Theory and Applications, 3(3):107–115, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615622.

Moller:1986:BAS

[Møl86] J. Møller. Bartlett adjustments for structured covariances. Scandinavian Journal of Statistics. Theory and Applications, 13 (1):1–15, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616006.

Molchanov:1994:SAB

[Mol94] Ilya S. Molchanov. On statistical analysis of Boolean models with non-random grains. Scandinavian Journal of Statistics. Theory and Applications, 21(1):73– 82, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616299.

Molchanov:1998:LTS

[Mol98] Ilya S. Molchanov. A limit theorem for solutions of inequalities. Scandinavian Journal of Statistics. Theory and Applications, 25 (1):235–242, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616496.

Maczynski:1980:IS

[MP80] M. J. Ma̧czyński and P. K.

Pāthak. Integration of surveys. Scandinavian Journal of Statistics. Theory and Applications, 7(3):130–138, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615788.

Mitra:1984:AOI

[MP84] S. K. Mitra and P. K.

Pathak. Algorithms for optimal integration of two or three surveys. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 257–263, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615963.

Martinussen:2014:ECO

[MP14] Torben Martinussen and Christian Bressen Pipper. Estimation of causal odds of concordance using the Aalen additive model. Scandinavian Journal of Statistics. Theory and Applications, 41(1):141–151, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586734.

Mazo:2021:PVN

[MP21] Gildas Mazo and François Portier. Parametric versus nonparametric: The fitness coefficient. Scandinavian Journal of Statistics. Theory and Applications, 48(4): 1344–1383, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ma:2022:MAM

[MP22] Chuoxin Ma and Jianxin Pan. Multistate analysis of multitype recurrent event and failure time data with event feedbacks in biomarkers. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 864–885, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Montoril:2019:WBE

[MPV19] Michel H. Montoril, Aluísio Pinheiro, and Brani Vidakovic. Wavelet-based estimators for mixture regression. Scandinavian Journal of Statistics. Theory and Applications, 46 (1):215–234, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Musta:2024:TSE

[MPV24] Eni Musta, Valentin Patilea, and Ingrid Van Keilegom. A two-step estimation procedure for semiparametric mixture cure models. Scandinavian Journal of Statistics. Theory and Applications, 51(3):987–1011, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Moodie:2010:EOD

[MR10] Erica E. M. Moodie and Thomas S. Richardson. Estimating optimal dynamic regimes: Correcting bias under the null. Scandinavian Journal of Statistics. Theory and Applications, 37(1):126–146, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000919.

Moller:2012:SPP

[MR12] Jesper Møller and Jakob Gulddahl Rasmussen. A sequential point process model and Bayesian inference for spatial point patterns with linear structures. Scandinavian Journal of Statistics. Theory and Applications, 39(4):618–634, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357214.

Martins:2014:EIN

[MR14] Thiago G. Martins and Håvard Rue. Extending integrated nested Laplace approximation to a class of near-Gaussian latent models. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 893–912, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586812.

Marie:2023:NWE

[MR23] Nicolas Marie and Amélie Rosier. Nadaraya–Watson estimator for I.I.D. paths of diffusion processes. Scandinavian Journal of Statistics. Theory and Applications, 50(2):589–637, June 2023. CODEN SJSADG.

ISSN 0303-6898 (print), 1467-9469 (electronic).

Moller:2024:CPD

[MR24] Jesper Møller and Jakob G.

Rasmussen. Cox processes driven by transformed Gaussian processes on linear networks — a review and new contributions. Scandinavian Journal of Statistics. Theory and Applications, 51 (3):1288–1322, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mcvinish:2009:BGF

[MRM09] Ross Mcvinish, Judith Rousseau, and Kerrie Mengersen. Bayesian goodness of fit testing with mixtures of triangular distributions. Scandinavian Journal of Statistics. Theory and Applications, 36(2): 337–354, June 2009. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000324.

Maiti:2014:PES

[MRS14] Tapabrata Maiti, Hao Ren, and Samiran Sinha. Prediction error of small area predictors shrinking both means and variances. Scandinavian Journal of Statistics. Theory and Applications, 41(3):775–790, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586758.

Moberg:1978:MLE

[MS78] Lars Moberg and Rolf Sundberg. Maximum likelihood estimation of a linear functional relationship when one of the departure variances is known. Scandinavian Journal of Statistics. Theory and Applications, 5(1):61– 64, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615685.

Marshall:1986:MNB

[MS86] Albert W. Marshall and Moshe Shaked. Multivariate new better than used distributions: a survey. Scandinavian Journal of Statistics. Theory and Applications, 13(4):277–290, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616037.

Miamee:1991:EHP

[MS91] A. G. Miamee and H. Salehi. An example of a harmonizable process whose spectral domain is not complete. Scandinavian Journal of Statistics. Theory and Applications, 18(3): 249–254, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616207.

Moller:1994:SAS

[MS94] Jesper Møller and Michael Sørensen. Statistical analysis of a spatial birth-and-death process model with a view to modelling linear dune fields. Scandinavian Journal of Statistics. Theory and Applications, 21(1): 1–19, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616295.

Martinussen:1998:EPS

[MS98a] Torben Martinussen and Thomas H. Scheike. Estimating the population survival function using additional information recorded over time: a filter based approach. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 621–635, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616528.

McKeague:1998:PLE

[MS98b] Ian W. McKeague and Sundarraman Subramanian. Product-limit estimators and Cox regression with missing censoring information. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 589–601, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616526.

Martinussen:2001:SAA

[MS01] Torben Martinussen and Thomas H. Scheike. Sampling adjusted analysis of dynamic additive regression models for longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 28(2): 303–323, June 2001. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616661.

Martinussen:2009:CSS

[MS09] Torben Martinussen and Thomas H. Scheike. Covariate selection for the semiparametric additive risk model. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 602–619, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000342.

Morgenthaler:2012:AVS

[MS12] Stephan Morgenthaler and Robert G. Staudte. Advantages of variance stabilization. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 714–728, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357220.

Martinussen:2023:EYR

[MS23] Torben Martinussen and Thomas Harder Scheike. Efficient -year risk regression using the logistic model. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1919–1932, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

McGee:2024:MAM

[MS24] Glen McGee and Alex Stringer. Marginal additive models for populationaveraged inference in longitudinal and cluster-correlated data. Scandinavian Journal of Statistics. Theory and Applications, 51(1): 384–401, March 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mselmi:2022:GLM

[Mse22]  Farouk Mselmi.  Gen-

eralized linear model for subordinated Lévy processes. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 772–801, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Marriott:2001:BAS

[MSP01] J. M. Marriott, N. M. Spencer, and A. N. Pettitt. A Bayesian approach to selecting covariates for prediction. Scandinavian Journal of Statistics. Theory and Applications, 28(1):87– 97, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616645.

Maatta:2016:SSL

[MSR16] Jussi Määttä, Daniel F. Schmidt, and Teemu Roos. Subset selection in linear regression using sequentially normalized least squares: Asymptotic theory. Scandinavian Journal of Statistics. Theory and Applications, 43(2):382–395, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Martinussen:2002:EEF

[MSSM02] Torben Martinussen, Thomas H.

Scheike, Ib M. Skovgaard, and T. Matinerssen. Efficient estimation of fixed and time-varying covariate effects in multiplicative intensity models. Scandinavian Journal of Statistics. Theory and Applications, 29(1):57–74, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616699.

Moller:1998:LGC

[MSW98] Jesper Møller, Anne Randi Syversveen, and Rasmus Plenge Waagepetersen. Log Gaussian Cox processes. Scandinavian Journal of Statistics. Theory and Applications, 25(3):451–482, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616515.

Maiti:2016:FME

[MSZ16] Tapabrata Maiti, Samiran Sinha, and Ping-Shou Zhong. Functional mixed effects model for small area estimation. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 886–903, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mira:2002:ECP

[MT02] Antonietta Mira and Luke Tierney. Efficiency and convergence properties of slice samplers. Scandinavian Journal of Statistics. Theory and Applications, 29(1):1–12, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616695.

Madsen:2003:BTA

[MT03] Jesper Madsen and René Tabanera. Bartlett’s test applied to variance component models. Scandinavian Journal of Statistics. Theory and Applications, 30(3): 609–616, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616787.

Moller:2014:GAS

[MT14] Jesper Møller and Håkon Toftaker. Geometric anisotropic spatial point pattern analysis and Cox processes. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 414–435, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586661.

McConville:2019:ASP

[MT19] Kelly S. McConville and Daniell Toth. Automated selection of post-strata using a model-assisted regression tree estimator. Scandinavian Journal of Statistics. Theory and Applications, 46(2):389–413, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mammen:1999:SSS

[MTA99] E. Mammen and C. Thomas-Agnan. Smoothing splines and shape restrictions. Scandinavian Journal of Statistics. Theory and Applications, 26(2):239–252, June 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616553.

McKeague:1991:GFT

[MU91] Ian W. McKeague and Klaus J. Utikal. Goodnessof-fit tests for additive hazards and proportional hazards models. Scandinavian Journal of Statistics. Theory and Applications, 18(3): 177–195, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616202.

Muller:1985:KEZ

[Mül85] Hans-Georg Müller. Kernel estimators of zeros and of location and size of extrema of regression functions. Scandinavian Journal of Statistics. Theory and Applications, 12(3): 221–232, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615991.

Muller:1992:GFD

[Mül92] Hans-Georg Müller. Goodnessof-fit diagnostics for regression models. Scandinavian Journal of Statistics. Theory and Applications, 19(2): 157–172, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616235.

Muller:1993:BKM

[Mül93] Hans-Georg Müller. On the boundary kernel method for non-parametric curve estimation near endpoints. Scandinavian Journal of Statistics. Theory and Applications, 20(4):313–328, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616287.

Muller:2005:DFM

[Mül05a] Hans-Georg Müller. [discussion of “functional modelling”]: Rejoinder. Scandinavian Journal of Statistics. Theory and Applications, 32(2):245–246, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616875. See [Mül05b].

Muller:2005:FMC

[Mül05b] Hans-Georg Müller. Functional modelling and classification of longitudinal data*. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 223–240, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616873. See discussion [Heu05, Ize05, Ram05] and rejoinder [Mül05a].

Munk:2002:TGF

[Mun02] Axel Munk. Testing the goodness of fit of parametric regression models with random Toeplitz forms. Scandinavian Journal of Statistics. Theory and Applications, 29(3):501–533, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616729.

Murphy:1993:TTD

[Mur93] S. A. Murphy. Testing for a time dependent coefficient in Cox’s regression model. Scandinavian Journal of Statistics. Theory and Applications, 20(1): 35–50, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616258.

Murphy:1995:CLT

[Mur95] S. A. Murphy. A central limit theorem for local martingales with applications to the analysis of longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 279–294, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616360.

Mustonen:1981:ISD

[Mus81] Seppo Mustonen. On interactive statistical data processing. Scandinavian Journal of Statistics. Theory and Applications, 8(3): 129–136, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615826.

Moolgavkar:1987:CRP

[MV87] Suresh H. Moolgavkar and David J. Venzon. Confidence regions for parameters of the proportional hazards model: a simulation study. Scandinavian Journal of Statistics. Theory and Applications, 14(1): 43–56, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616047.

Martin:2020:NHC

[MV20] Ole Martin and Mathias Vetter. The null hypothesis of (common) jumps in case of irregular and asynchronous observations. Scandinavian Journal of Statistics. Theory and Applications, 47(3):711–756, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Muller:2015:PLM

[MvdG15] Patric Müller and Sara van de Geer. The partial linear model in high dimensions. Scandinavian Journal of Statistics. Theory and Applications, 42(2): 580–608, June 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mukerjee:1993:NML

[MW93] Hari Mukerjee and Jane-Ling Wang. Nonparametric maximum likelihood estimation of an increasing hazard rate for uncertain cause-of-death data. Scandinavian Journal of Statistics. Theory and Applications, 20(1):17– 33, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616257.

Muliere:1997:BNP

[MW97] Pietro Muliere and Stephen Walker. A Bayesian nonparametric approach to survival analysis using Polya trees. Scandinavian Journal of Statistics. Theory and Applications, 24(3): 331–340, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616459.

Moller:2007:MSS

[MW07] Jesper Møller and Rasmus P. Waagepetersen. Modern statistics for spatial point processes*. Scandinavian Journal of Statistics. Theory and Applications, 34(4):643–684, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

41548576. See discussion [Ano07f].

Maathuis:2008:IMJ

[MW08] Marloes H. Maathuis and Jon A. Wellner. Inconsistency of the MLE for the joint distribution of interval-censored survival times and continuous marks. Scandinavian Journal of Statistics. Theory and Applications, 35(1):83– 103, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000386.

Ma:2010:TSS

[MW10] Yanyuan Ma and Alan H.

Welsh. Transformation and smoothing in sample survey data. Scandinavian Journal of Statistics. Theory and Applications, 37(3): 496–513, September 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000403.

Marra:2012:CPC

[MW12] Giampiero Marra and Simon N. Wood. Coverage properties of confidence intervals for generalized additive model components. Scandinavian Journal of Statistics. Theory and Applications, 39(1):53– 74, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41411157.

Muller:2024:EWC

[MWW24] Samuel Muller, Suojin Wang, and A. H. Welsh. The effect of the working correlation on fitting models to longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 51(2):891–912, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Miao:2015:MDP

[MWY15] Yu Miao, Yanling Wang, and Guangyu Yang. Moderate deviation principles for empirical covariance in the neighbourhood of the unit root. Scandinavian Journal of Statistics. Theory and Applications, 42 (1):234–255, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Mykland:2011:DGA

[MZ11] Per A. Mykland and Lan Zhang. The double Gaussian approximation for high frequency data. Scandinavian Journal of Statistics. Theory and Applications, 38(2):215–236, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015491.

Naes:1982:ADE

[Næs82] Tormod Næs. The asymptotic distribution of the estimator for the regression parameter in Cox’s regression model. Scandinavian Journal of Statistics. Theory and Applications, 9(2):107–115, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615862.

Nair:1982:PCB

[Nai82] Vijayan N. Nair. Q–Q plots with confidence bands for comparing several populations. Scandinavian Journal of Statistics. Theory and Applications, 9(4):193– 200, March 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615882.

Natvig:1975:QMW

[Nat75] Bent Natvig. On a queuing model where potential customers are discouraged by queue length. Scandinavian Journal of Statistics. Theory and Applications, 2 (1):34–42, December 1975.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615573.

Natvig:1985:NLM

[Nat85] Bent Natvig. New light on measures of importance of system components. Scandinavian Journal of Statistics. Theory and Applications, 12(1):43– 54, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615971.

Natvig:1993:SEB

[Nat93] Bent Natvig. Strict and exact bounds for the availabilities in a fixed time interval for multistate monotone systems. Scandinavian Journal of Statistics. Theory and Applications, 20 (2):171–175, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616272.

Nieto-Barajas:2012:RPT

[NBM12] Luis E. Nieto-Barajas and Peter Müller. Rubbery Polya tree. Scandinavian Journal of Statistics. Theory and Applications, 39 (1):166–184, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41411163.

Nieto-Barajas:2002:MBG

[NBW02] Luis E. Nieto-Barajas and Stephen G. Walker. Markov beta and gamma processes for modelling hazard rates. Scandinavian Journal of Statistics. Theory and Applications, 29(3):413–424, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616724.

Nieto-Barajas:2008:BSC

[NBY08] Luis E. Nieto-Barajas and Guosheng Yin. Bayesian semiparametric cure rate model with an unknown threshold. Scandinavian Journal of Statistics. Theory and Applications, 35(3): 540–556, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000286.

Nayak:1992:EUC

[NC92] Tapan K. Nayak and Mary C. Christman. Effect of unequal catchability on estimates of the number of classes in a population. Scandinavian Journal of Statistics. Theory and Applications, 19(3):

281–287, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616245.

Ning:2015:CPR

[NC15] Yang Ning and Yong Chen. A class of pseudolikelihood ratio tests for homogeneity in exponential tilt mixture models. Scandinavian Journal of Statistics. Theory and Applications, 42(2):504–517, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ni:2018:TPS

[NC18] Ai Ni and Jianwen Cai. Tuning parameter selection in Cox proportional hazards model with a diverging number of parameters. Scandinavian Journal of Statistics. Theory and Applications, 45(3):557–570, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Neumeyer:2021:WGF

[NDH^{+}21] Natalie Neumeyer, Miguel A. Delgado, Lajos Horváth, Simos Meintanis, Emanuele Taufer, and Lixing Zhu. 4th workshop on goodnessof-fit, change-point, and related problems, Trento, 2019. Scandinavian Journal of Statistics. Theory and Applications, 48(2): 371–374, June 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Natvig:1987:BES

[NE87] Bent Natvig and Henriette Eide. Bayesian estimation of system reliability. Scandinavian Journal of Statistics. Theory and Applications, 14(4): 319–327, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616074.

Nerman:1977:MGS

[Ner77]  Olle Nerman.  On the

maximal generation size of a non-critical Galton– Watson process. Scandinavian Journal of Statistics. Theory and Applications, 4(3):131–135, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615664.

Nerman:1998:SMC

[Ner98] Olle Nerman. Stochastic monotonicity and conditioning in the limit. Scandinavian Journal of Statistics. Theory and Applications, 25(3):569–572, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616521.

Neumann:1997:OCP

[Neu97] Michael H. Neumann. Optimal change-point estimation in inverse problems. Scandinavian Journal of Statistics. Theory and Applications, 24(4):503–521, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616472.

Neumeyer:2009:SRB

[Neu09] Natalie Neumeyer. Smooth residual bootstrap for empirical processes of nonparametric regression residuals. Scandinavian Journal of Statistics. Theory and Applications, 36(2): 204–228, June 2009. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000318.

Noh:2013:QFM

[NEV13] Hohsuk Noh, Anouar El Ghouch, and Ingrid Van Keilegom. Quality of fit measures in the framework of quantile regression. Scandinavian Journal of Statistics. Theory and Applications, 40(1):

105–118, March 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357255.

Nuno:2022:ENC

[NG22] Michelle M. Nuño and Daniel L. Gillen. On estimation in the nested case-control design under nonproportional hazards. Scandinavian Journal of Statistics. Theory and Applications, 49(1):265–286, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nguyen:2023:FSI

[NG23] Hien Duy Nguyen and Mayetri Gupta. Finite sample inference for empirical Bayesian methods. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1616–1640, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nielsen:1992:CPA

[NGAS92] Gert G. Nielsen, Richard D.

Gill, Per Kragh Andersen, and Thorkild I. A. Sørensen. A counting process approach to maximum likelihood estimation in frailty models. Scandinavian Journal of Statistics. Theory and Applications, 19(1):25–43, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616223.

Nelder:1994:RRR

[NGMS94] J. A. Nelder, E. I. George, U. E. Makov, and A. F. M. Smith. Reader reaction: a re-analysis of the pumpfailure data [with response]. Scandinavian Journal of Statistics. Theory and Applications, 21(2):187–191, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616310.

Niu:2018:ENP

[NGZ18] Cuizhen Niu, Xu Guo, and Lixing Zhu. Enhancements of non-parametric generalized likelihood ratio test: Bias correction and dimension reduction. Scandinavian Journal of Statistics. Theory and Applications, 45(2):217–254, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Naes:1993:RCR

[NH93] Tormod Næs and Inge S. Helland. Relevant components in regression. Scandinavian Journal of Statistics. Theory and Applications, 20(3):239–250, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616279.

Neal:2015:FSM

[NH15] Peter Neal and Chien Lin Terry Huang. Forward simulation Markov chain Monte Carlo with applications to stochastic epidemic models. Scandinavian Journal of Statistics. Theory and Applications, 42(2):378–396, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nghiem:2022:EGM

[NHMW22] Linh H. Nghiem, Francis K. C. Hui, Samuel Müller, and Alan H. Welsh. Estimation of graphical models for skew continuous data. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1811–1841, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Neumann:2019:EGT

[NHS^{+}19] Matthias Neumann, Christian Hirsch, Jakub Staněk, Viktor Beneš, and Volker Schmidt. Estimation of geodesic tortuosity and constrictivity in stationary random closed sets. Scandinavian Journal of Statistics. Theory and Applications, 46(3):848–884, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nicolau:2014:NMM

[Nic14] João Nicolau. A new model for multivariate Markov chains. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1124–1135, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586824.

Niemi:1983:ENN

[Nie83] Hannu Niemi. On the effects of a nonstationary noise on ARMA models. Scandinavian Journal of Statistics. Theory and Applications, 10(1):11– 17, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615895.

Niemi:1984:TPL

[Nie84] Hannu Niemi. Theoretical properties of linear predictors for ARIMA models. Scandinavian Journal of Statistics. Theory and Applications, 11(2):

113–122, March 1984. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615947.

Nielsen:1997:ESC

[Nie97a] Bent Nielsen. Expected survival in the Cox model. Scandinavian Journal of Statistics. Theory and Applications, 24(2):275–287, June 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616453.

Nielsen:1997:IMD

[Nie97b] Søren Feodor Nielsen. Inference and missing data: Asymptotic results. Scandinavian Journal of Statistics. Theory and Applications, 24(2):261–274, June 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616452.

Nielsen:1998:MBC

[Nie98] Jens Perch Nielsen. Multiplicative bias correction in kernel hazard estimation. Scandinavian Journal of Statistics. Theory and Applications, 25(3): 541–553, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616519.

Nielsen:1999:ESC

[Nie99] Bent Nielsen. Expected survival in the Cox model: an addendum. Scandinavian Journal of Statistics. Theory and Applications, 26(1): 159, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616547.

Nipoti:2018:BSP

[NJG18] Bernardo Nipoti, Alejandro Jara, and Michele Guindani. A Bayesian semiparametric partially PH model for clustered time-to-event data. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 1016–1035, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Noh:2016:QRL

[NL16] Jungsik Noh and Sangyeol Lee. Quantile regression for location-scale time series models with conditional heteroscedasticity. Scandinavian Journal of Statistics. Theory and Applications, 43(3):700–720, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nurminen:1987:EBA

[NM87] Markku Nurminen and Pertti Mutanen. Exact Bayesian analysis of two proportions. Scandinavian Journal of Statistics. Theory and Applications, 14(1): 67–77, June 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616049. See corrigenda [Ano96e].

Nguyen:2014:EEP

[NM14] Van Hanh Nguyen and Catherine Matias. On efficient estimators of the proportion of true null hypotheses in a multiple testing setup. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1167–1194, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586827.

Nordland:2024:ETE

[NM24] Andreas Nordland and Torben Martinussen. Estimation of treatment effect among treatment responders with a time-to-event endpoint. Scandinavian Journal of Statistics. Theory and Applications, 51 (3):1161–1180, September 2024. CODEN SJSADG.

ISSN 0303-6898 (print), 1467-9469 (electronic).

Norlen:1975:CMW

[Nor75] Urban Norlén. The covariance matrices for which least squares is best linear unbiased. Scandinavian Journal of Statistics. Theory and Applications, 2(2):85–90, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615581.

Norberg:1977:IRC

[Nor77] Ragnar Norberg. Inference in random coefficient regression models with oneway and nested classifications. Scandinavian Journal of Statistics. Theory and Applications, 4(2):71– 80, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615653.

Nordberg:1980:ANM

[Nor80] Lennart Nordberg. Asymptotic normality of maximum likelihood estimators based on independent, unequally distributed observations in exponential family models. Scandinavian Journal of Statistics. Theory and Applications, 7 (1):27–32, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615767.

Nordberg:1981:SSE

[Nor81] Lennart Nordberg. Stepwise selection of explanatory variables in the binary logit model. Scandinavian Journal of Statistics. Theory and Applications, 8 (1):17–26, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615808.

Norros:1986:CRM

[Nor86] Ilkka Norros. A compensator representation of multivariate life length distributions, with applications. Scandinavian Journal of Statistics. Theory and Applications, 13(2): 99–112, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616016.

Norell:1990:PSN

[Nor90] Lennart Norell. Partitioning a set of normal populations by their locations with respect to a number of controls. Scandinavian Journal of Statistics. Theory and Applications, 17 (3):217–233, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616170.

Norros:2005:TSP

[Nor05]  Ilkka Norros.  Teletraf-

fic as a stochastic playground*. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 201–215, June 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

4616871. See discussion [Mak05, Sze05].

Naieni:2023:RDR

[NR23] Amir-Abbas Mofidian Naieni and Reyhaneh Rikhtehgaran. Regularization in dynamic random-intercepts models for analysis of longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 50(2):513–549, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Neumeyer:2006:CSC

[NS06] Natalie Neumeyer and Stefan Sperlich. Comparison of separable components in different samples. Scandinavian Journal of Statistics. Theory and Applications, 33(3):477–501, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616938.

Nielsen:2001:BBC

[NT01] Jens Perch Nielsen and Carsten Tanggaard. Boundary and bias correction in kernel hazard estimation. Scandinavian Journal of Statistics. Theory and Applications, 28(4):675–698, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616689.

Nakakita:2019:IED

[NU19] Shogo H. Nakakita and Masayuki Uchida. Inference for ergodic diffusions plus noise. Scandinavian Journal of Statistics. Theory and Applications, 46(2): 470–516, June 2019. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nielsen:2004:SIT

[NV04] Linda Stougaard Nielsen and Eva B. Vedel Jensen. Statistical inference for transformation inhomogeneous point processes. Scandinavian Journal of Statistics. Theory and Applications, 31(1):131–142, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616816.

Neumeyer:2009:CPT

[NV09] Natalie Neumeyer and Ingrid Van Keilegom. Changepoint tests for the error distribution in non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 518–541, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000336.

Nguyen:2017:STO

[NV17] Michele Nguyen and Almut E. D. Veraart. Spatiotemporal Ornstein–Uhlenbeck processes: Theory, simulation and statistical inference. Scandinavian Journal of Statistics. Theory and Applications, 44(1):46– 80, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Nicoleris:2006:PPD

[NW06] Theodoros Nicoleris and Stephen G. Walker. On a posterior predictive density sample size criterion. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 209–218, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616919.

Neal:2017:CNC

[NX17] Peter Neal and Fei Xiang.

Collapsing of non-centred parameterized MCMC algorithms with applications to epidemic models. Scandinavian Journal of Statistics. Theory and Applications, 44(1):81–96, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ning:2018:CWE

[NYR18] Yang Ning, Grace Yi, and Nancy Reid. A class of weighted estimating equations for semiparametric transformation models with missing covariates. Scandinavian Journal of Statistics. Theory and Applications, 45(1):87–109, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

OKeeffe:2016:AEL

[OB16] Aidan G. O’Keeffe and Gianluca Baio. Approaches to the estimation of the local average treatment effect in a regression discontinuity design. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 978–995, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Borgan:2007:DAR

[øBFHB07] ørnulf Borgan, Rosemeire L. Fiaccone, Robin Henderson, and Mauricio L. Barreto. Dynamic analysis of recurrent event data with missing observations, with application to infant diarrhoea in Brazil. Scandinavian Journal of Statistics. Theory and Applications, 34(1):53–69, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548538.

Oesting:2018:SMS

[OBL18] Marco Oesting, Liliane Bel, and Christian Lantuéjoul. Sampling from a maxstable process conditional on a homogeneous functional with an application for downscaling climate data. Scandinavian Journal of Statistics. Theory and Applications, 45 (2):382–404, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Opsomer:2012:MBN

[OFFL12] Jean D. Opsomer, Mario Francisco-Fernández, and Xiaoxi Li. Model-based non-parametric variance estimation for systematic sampling. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 528–542, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41679810.

Ogundimu:2016:SSM

[OH16] Emmanuel O. Ogundimu and Jane L. Hutton. A sample selection model with skew-normal distribution. Scandinavian Journal of Statistics. Theory and Applications, 43(1):172–190, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Overgaard:2021:AIR

[OH21] Morten Overgaard and Stefan Nygaard Hansen. On the assumption of independent right censoring. Scandinavian Journal of Statistics. Theory and Applications, 48(4):1234–1255, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Okano:2024:LAB

[OHIS24] Ryo Okano, Yasuyuki Hamura, Kaoru Irie, and Shonosuke Sugasawa. Locally adaptive Bayesian isotonic regression using half shrinkage priors. Scandinavian Journal of Statistics. Theory and Applications, 51(1):109–141, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ohlsson:1986:ANR

[Ohl86] Esbjörn Ohlsson. Asymptotic: Normality of the Rao, Hartley, Cochran estimator: an application of the martingale CLT. Scandinavian Journal of Statistics. Theory and Applications, 13(1):17–28, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616007.

Omelka:2021:MPL

[OHN21] Marek Omelka, Šárka Hudecová, and Natalie Neumeyer. Maximum pseudo-likelihood estimation based on estimated residuals in copula semiparametric models. Scandinavian Journal of Statistics. Theory and Applications, 48(4):1433–1473, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Oja:1981:LSS

[Oja81]  Hannu Oja.  On loca-

tion, scale, skewness and kurtosis of univariate distributions. Scandinavian Journal of Statistics. Theory and Applications, 8(3): 154–168, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615828.

Oja:1999:AIM

[Oja99] Hannu Oja. Affine invariant multivariate sign and rank tests and corresponding estimates: a review. Scandinavian Journal of Statistics. Theory and Applications, 26(3):319–343, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616560.

Oja:2016:DAT

[Oja16] Hannu Oja. Discussion on “Asymptotic Theory of Outlier Detection Algorithms for Linear Time Series Regression Problems” by Søren Johansen and Bent Nielsen. Scandinavian Journal of Statistics. Theory and Applications, 43(2): 366–367, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JN16a].

Ogata:2000:EBA

[OKK^{+}00] Yosihiko Ogata, Koichi Katsura, Niels Keiding, Claus Holst, and Anders Green. Empirical Bayes age–period–cohort analysis of retrospective incidence data. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 415–432, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616614.

Oktaba:1988:ILS

[OKW88] Wiktor Oktaba, Andrzej Kornacki, and Jacek Wawrzosek. Invariant linearly sufficient transformations of the general Gauss– Markoff model. estimation and testing. Scandinavian Journal of Statistics. Theory and Applications, 15(2): 117–124, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616092.

Olsson:1996:EPT

[Ols96] Marita Olsson. Estimation of phase-type distributions from censored data. Scandinavian Journal of Statistics. Theory and Applications, 23(4):443–460, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616419.

Overgaard:2018:EVP

[OPP18] Morten Overgaard, Erik Thorlund Parner, and Jan Pedersen. Estimating the variance in a pseudoobservation scheme with competing risks. Scandinavian Journal of Statistics. Theory and Applications, 45(4):923–940, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Olsson:1994:IMQ

[OR94] Jonny Olsson and Holger Rootzén. An image model for quantal response analysis in perimetry. Scandinavian Journal of Statistics. Theory and Applications, 21(4):375–387, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616324.

Ouadah:2020:DBG

[ORL20] Sarah Ouadah, Stéphane Robin, and Pierre Latouche. Degree-based goodness-of-fit tests for heterogeneous random graph models: Independent and exchangeable cases. Scandinavian Journal of Statistics. Theory and Applications, 47(1):156–181, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Oman:1996:EMS

[OS96] Samuel D. Oman and M. S. Srivastava. Exact mean squared error comparisons of the inverse and classical estimators in multiunivariate linear calibration. Scandinavian Journal of Statistics. Theory and Applications, 23(4): 473–488, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616421.

Ould-Said:1997:NEP

[OS97] Elias Ould-Saïd. A note on ergodic processes prediction via estimation of the conditional mode function. Scandinavian Journal of Statistics. Theory and Applications, 24(2): 231–239, June 1997. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616450.

Ogihara:2024:EDP

[OS24] Teppei Ogihara and Mitja Stadje. Efficient drift parameter estimation for ergodic solutions of backward SDEs. Scandinavian Journal of Statistics. Theory and Applications, 51 (3):1181–1205, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Olsen:2008:SSA

[OSG08] Lena R. Olsen, Sigrunn H.

Sørbye, and Fred Godtliebsen. A scale-space approach for detecting nonstationarities in time series. Scandinavian Journal of Statistics. Theory and Applications, 35(1): 119–138, March 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000388.

Ogata:2009:CRP

[OT09] Hiroaki Ogata and Masanobu Taniguchi. Cressie–Read power-divergence statistics for non-Gaussian vector stationary processes. Scandinavian Journal of Statistics. Theory and Applications, 36(1):141–156, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000312.

Overbeck:1998:ECB

[Ove98] Ludger Overbeck. Estimation for continuous branching processes. Scandinavian Journal of Statistics. Theory and Applications, 25 (1):111–126, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616488.

Palmgren:2004:ICM

[Pal04] Juni Palmgren. Introduction to causal modelling and inference. Scandinavian Journal of Statistics. Theory and Applications, 31(2):159–160, June 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616820.

Pal:2009:EPE

[Pal09] Jayanta Kumar Pal. Endpoint estimation for decreasing densities: Asymptotic behaviour of the penalized likelihood ratio. Scandinavian Journal of Statistics. Theory and Applications, 36(4):764–781, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000351.

Pan:2002:GFT

[Pan02] Wei Pan. Goodness-of-fit tests for GEE with correlated binary data. Scandinavian Journal of Statistics. Theory and Applications, 29(1):101–110, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616702.

Paparoditis:2000:SDB

[Pap00] Efstathios Paparoditis. Spectral density based goodnessof-fit tests for time series models. Scandinavian Journal of Statistics. Theory and Applications, 27 (1):143–176, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616596.

Papathomas:2008:CBV

[Pap08] Michail Papathomas. Correlated binary variables and multi-level probability assessments. Scandinavian Journal of Statistics. Theory and Applications, 35 (1):169–185, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000391.

Parner:2001:CLA

[Par01] E. T. Parner. A composite likelihood approach to multivariate survival data. Scandinavian Journal of Statistics. Theory and Applications, 28(2): 295–302, June 2001. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616660.

Parkinson:2020:CMT

[Par20] Judith H. Parkinson. Combined multiple testing of multivariate survival times by censored empirical likelihood. Scandinavian Journal of Statistics. Theory and Applications, 47(3): 757–786, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Persson:2006:EPA

[PBB06] Anna H. Persson, Lennart Bondesson, and Niclas Börlin. Estimation of polygons and areas. Scandinavian Journal of Statistics. Theory and Applications, 33(3):541–559, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616941.

Pratelli:2012:SAH

[PBBM12] Luca Pratelli, Alberto Baccini, Lucio Barabesi, and Marzia Marcheselli. Statistical analysis of the Hirsch index. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 681–694, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357218.

Picard:2009:MSM

[PBHMC09] Nicolas Picard, Avner Bar-Hen, Frédéric Mortier, and Joël Chadœuf. The multi-scale marked areainteraction point process: a model for the spatial pattern of trees. Scandinavian Journal of Statistics. Theory and Applications, 36(1):23–41, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000305.

Putt:1999:GSC

[PC99] Mary Putt and Vernon M.

Chinchilli. Generalized Lstatistics: Correction to the form of the asymptotic variance presented by Helmers et al. (1990). Scandinavian Journal of Statistics. Theory and Applications, 26(3): 475–476, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616569.

Pitt:2002:CFO

[PCW02] Michael K. Pitt, Chris Chatfield, and Stephen G. Walker. Constructing first order stationary autoregressive models via latent processes. Scandinavian Journal of Statistics. Theory and Applications, 29(4): 657–663, December 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616740.

Pokharel:2022:EBI

[PD22] Gyanendra Pokharel and Rob Deardon. Emulationbased inference for spatial infectious disease transmission models incorporating event time uncertainty. Scandinavian Journal of Statistics. Theory and Applications, 49(1):455–479, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Picchini:2010:SDM

[PDD10] Umberto Picchini, Andrea De Gaetano, and Susanne Ditlevsen. Stochastic differential mixed-effects models. Scandinavian Journal of Statistics. Theory and Applications, 37(1):67–90 176, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/

41000916. See corrections [Ano10f].

Pons:1987:ECP

[PdT87] O. Pons and E. de Turckheim. Estimation in Cox’s periodic model with a histogram-type estimator for the underlying intensity. Scandinavian Journal of Statistics. Theory and Applications, 14(4): 329–345, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616075.

Pons:1991:TIB

[PdT91] Odile Pons and Elisabeth de Turckheim. Tests of independence for bivariate censored data based on the empirical joint hazard function. Scandinavian Journal of Statistics. Theory and Applications, 18(1):21–37, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616186.

Pedersen:1975:SUA

[Ped75a] J. Granfeldt Pedersen. On strong unimodality and Mancillarity with applications to contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 2(3):127– 137, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615592.

Pedersen:1975:SUT

[Ped75b] J. Granfeldt Pedersen. On strong unimodality of twodimensional discrete distributions with applications to M-ancillarity. Scandinavian Journal of Statistics. Theory and Applications, 2(2):99–102, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615584.

Pedersen:1995:NAM

[Ped95] Asger Roer Pedersen. A new approach to maximum likelihood estimation for stochastic differential equations based on discrete observations. Scandinavian Journal of Statistics. Theory and Applications, 22(1): 55–71, March 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616340.

Pedersen:2000:ENO

[Ped00] Asger Roer Pedersen. Estimating the nitrous oxide emission rate from the soil surface by means of a diffusion model. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 385–403, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616612.

Paglia:2022:ESD

[PEK22] Jacopo Paglia, Jo Eidsvik, and Juha Karvanen. Efficient spatial designs using Hausdorff distances and Bayesian optimization. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1060–1084, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Penskaya:1995:MSC

[Pen95] M. Y. Penskaya. Mean square consistent estimation of a ratio. Scandinavian Journal of Statistics. Theory and Applications, 22(1):129–137, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616346.

Persson:1979:NWO

[Per79] Tore Persson. A new way to obtain Watson’s U^{2}. Scandinavian Journal of Statistics. Theory and Applications, 6(3):119–122, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615745.

Pereyra:2014:DPB

[Per14] Marcelo Pereyra. Discussion on the paper by Byrne and Girolami. Scandinavian Journal of Statistics. Theory and Applications, 41(1):12–13, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586724.  See

[BG13].

Petrone:1999:RBP

[Pet99]  Sonia Petrone.  Ran-

dom Bernstein polynomials. Scandinavian Journal of Statistics. Theory and Applications, 26(3): 373–393, September 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616563.

Paige:2008:RIC

[PF08] Robert L. Paige and P. Harshini Fernando. Robust inference in conditionally linear nonlinear regression models. Scandinavian Journal of Statistics. Theory and Applications, 35 (1):158–168, March 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000390.

Pfanzagl:1993:SOU

[Pfa93] J. Pfanzagl. Sequences of optimal unbiased estimators need not be asymptotically optimal. Scandinavian Journal of Statistics. Theory and Applications, 20(1): 73–76, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616261.

Pardo-Fernandez:2015:NPA

[PFJGE15] Juan Carlos Pardo-Fernández, María Dolores Jiménez-Gamero, and Anouar El Ghouch. A non-parametric ANOVA-type test for regression curves based on characteristic functions. Scandinavian Journal of Statistics. Theory and Applications, 42(1):197–213, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pardo-Fernandez:2006:CRC

[PFV06] Juan Carlos Pardo-Fernández and Ingrid Van Keilegom. Comparison of regression curves with censored responses. Scandinavian Journal of Statistics. Theory and Applications, 33(3): 409–434, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616934.

Potgieter:2013:CFB

[PG13] Cornelis J. Potgieter and Marc G. Genton. Characteristic function-based semiparametric inference for skew-symmetric models. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 471–490, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586682.

Piccioni:2000:ISN

[Pic00] Mauro Piccioni. Independence structure of natural conjugate densities to exponential families and the Gibbs’ Sampler. Scandinavian Journal of Statistics. Theory and Applications, 27(1):111–127, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616594.

Puig:2018:NPE

[PK18] Pedro Puig and Célestin C. Kokonendji. Non-parametric estimation of the number of zeros in truncated count distributions. Scandinavian Journal of Statistics. Theory and Applications, 45(2):347–365, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pal:2017:TCM

[PKH17] Subahdip Pal, Kshitij Khare, and James P. Hobert. Trace class Markov chains for Bayesian inference with generalized double Pareto shrinkage priors. Scandinavian Journal of Statistics. Theory and Applications, 44(2): 307–323, June 2017. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Park:1997:STT

[PKR^{+}97] B. U. Park, W. C. Kim, D. Ruppert, M. C. Jones, D. F. Signorini, and R. Kohn. Simple transformation techniques for improved non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 24(2): 145–163, June 1997. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616445.

Poinas:2023:AAL

[PL23] Arnaud Poinas and Frédéric Lavancier. Asymptotic approximation of the likelihood of stationary determinantal point processes. Scandinavian Journal of Statistics. Theory and Applications, 50(2):842–874, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Page:2017:EPP

[PLHS17] Garritt L. Page, Yajun Liu, Zhuoqiong He, and Donchu Sun. Estimation and prediction in the presence of spatial confounding for spatial linear models. Scandinavian Journal of Statistics. Theory and Applications, 44(3): 780–797, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Park:2006:SEE

[PLKP06] Byeong U. Park, Young Kyung Lee, Tae Yoon Kim, and Cheolyong Park. A simple estimator of error correlation in non-parametric regression models. Scandinavian Journal of Statistics. Theory and Applications, 33(3):451–462, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616936.

Pawitan:2023:ECO

[PLL23] Yudi Pawitan, Hangbin Lee, and Youngjo Lee. Epistemic confidence in the observed confidence interval. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1859–1883, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pipper:2003:LBE

[PM03] Christian Bressen Pipper and Torben Martinussen. A likelihood based estimating equation for the Clayton–Oakes model with marginal proportional hazards. Scandinavian Journal of Statistics. Theory and Applications, 30(3): 509–521, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616780.

Pensar:2017:SLC

[PNC17] Johan Pensar, Henrik Nyman, and Jukka Corander. Structure learning of contextual Markov networks using marginal pseudolikelihood. Scandinavian Journal of Statistics. Theory and Applications, 44 (2):455–479, June 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pollard:1995:URL

[Pol95]  David Pollard.  Uni-

form ratio limit theorems for empirical processes. Scandinavian Journal of Statistics. Theory and Applications, 22(3):271–278, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616359.

Pons:1986:TIB

[Pon86] Odile Pons. A test of independence between two censored survival times. Scandinavian Journal of Statistics. Theory and Applications, 13(3):173–185, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616023.

Portier:2016:EPV

[Por16] François Portier. An empirical process view of inverse regression. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 827–844, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Paparoditis:2016:LPP

[PP16] Efstathios Paparoditis and Philip Preuß. On local power properties of frequency domain-based tests for stationarity. Scandinavian Journal of Statistics. Theory and Applications, 43(3):664–682, September 2016. CODEN SJSADG.

ISSN 0303-6898 (print), 1467-9469 (electronic).

Pumi:2021:DMD

[PPS21] Guilherme Pumi, Taiane Schaedler Prass, and Rafael Rigão Souza. A dynamic model for doublebounded time series with chaotic-driven conditional averages. Scandinavian Journal of Statistics. Theory and Applications, 48(1): 68–86, March 2021. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pipper:2007:CGD

[PR07] Christian B. Pipper and Christian Ritz. Checking the grouped data version of the Cox model for interval-grouped survival data. Scandinavian Journal of Statistics. Theory and Applications, 34(2): 405–418, June 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548559.

Praestgaard:1995:PBK

[Præ95] Jens Thomas Præstgaard.

Permutation and bootstrap Kolmogorov–Smirnov tests for the equality of two distributions. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 305–322, September 1995.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616362.

Prewitt:2003:EBS

[Pre03] Kathryn A. Prewitt. Efficient bandwidth selection in non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 30(1):75– 92, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616750.

Prendergast:2005:IFS

[Pre05] L. A. Prendergast. Influence functions for sliced inverse regression. Scandinavian Journal of Statistics. Theory and Applications, 32(3):385–404, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616886.

Pokarowski:2022:ILM

[PRS^{+}22] Piotr Pokarowski, Wojciech Rejchel, Agnieszka Sołtys, Michał Frej, and Jan Mielniczuk. Improving lasso for model selection and prediction. Scandinavian Journal of Statistics. Theory and Applications, 49(2):

831–863, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Paindaveine:2021:PTE

[PRV21] Davy Paindaveine, Joséa Rasoafaraniaina, and Thomas Verdebout. Preliminary test estimation in uniformly locally asymptotically normal models. Scandinavian Journal of Statistics. Theory and Applications, 48(2):689–707, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pukelsheim:1983:CMP

[PS83] Friedrich Pukelsheim and George P. H. Styan. Convexity and monotonicity properties of dispersion matrices of estimators in linear models. Scandinavian Journal of Statistics. Theory and Applications, 10 (2):145–149, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615914.

Penttinen:1989:SAC

[PS89] A. Penttinen and D. Stoyan.

Statistical analysis for a class of line segment processes. Scandinavian Journal of Statistics. Theory and Applications, 16(2): 153–168, June 1989. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616131.

Pace:1992:NCC

[PS92] Luigi Pace and Alessandra Salvan. A note on conditional cumulants in canonical exponential families. Scandinavian Journal of Statistics. Theory and Applications, 19(2):185–191, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616237.

Pawlak:1999:KDE

[PS99] M. Pawlak and U. Stadtmüller. Kernel density estimation with generalized binning. Scandinavian Journal of Statistics. Theory and Applications, 26(4):539–561, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616578.

Prendergast:2010:IFD

[PS10] Luke A. Prendergast and Jodie A. Smith. Influence functions for dimension reduction methods: an example influence study of principal Hessian direction analysis. Scandinavian Journal of Statistics. Theory and Applications, 37(4): 588–611, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000410.

Prendergast:2013:SIR

[PS13] Luke A. Prendergast and Simon J. Sheather. On sensitivity of inverse response plot estimation and the benefits of a robust estimation approach. Scandinavian Journal of Statistics. Theory and Applications, 40(2):219–237, June 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586710.

Park:2020:DRC

[PS20] Jin-Hong Park and S. Yaser Samadi. Dimension reduction for the conditional mean and variance functions in time series. Scandinavian Journal of Statistics. Theory and Applications, 47(1):134–155, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pantle:2010:EIC

[PSS10] Ursa Pantle, Volker Schmidt, and Evgeny Spodarev. On the estimation of integrated covariance functions of stationary random fields. Scandinavian Journal of Statistics. Theory and Applications, 37(1):47– 66, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000915.

Paige:2009:VRS

[PSW09] Robert L. Paige, Shan Sun, and Keyi Wang. Variance reduction in smoothing splines. Scandinavian Journal of Statistics. Theory and Applications, 36 (1):112–126, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000310.

Payne:1992:GBC

[PT92] R. W. Payne and R. D. Tobias. General balance, combination of information and the analysis of covariance. Scandinavian Journal of Statistics. Theory and Applications, 19 (1):3–23, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616222.

Paige:2009:SBB

[PTF09] Robert L. Paige, A. Alexandre Trindade, and P. Harshini Fernando. Saddlepointbased bootstrap inference for quadratic estimating equations. Scandinavian Journal of Statistics. Theory and Applications, 36 (1):98–111, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000309.

Pukkila:1982:ITF

[Puk82] Tarmo Pukkila. On the identification of transfer function noise models with several correlated inputs. Scandinavian Journal of Statistics. Theory and Applications, 9(3):139–146, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615869.

Pons:2000:NSC

[PV00] Odile Pons and Michael Visser. A non-stationary Cox model. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 619–639, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616631.

Preuss:2013:TSH

[PVD13] Philip Preuss, Mathias Vetter, and Holger Dette. Testing semiparametric hypotheses in locally stationary processes. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 417–437, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586679.

Puig:2024:SML

[PVFF24] Pedro Puig, Jordi Valero, and Amanda Fernández-Fontelo. Some mechanisms leading to underdispersion: Old and new proposals. Scandinavian Journal of Statistics. Theory and Applications, 51 (1):245–267, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Pal:2006:DBC

[PW06] Jayanta Kumar Pal and Michael Woodroofe. On the distance between cumulative sum diagram and its greatest convex minorant for unequally spaced design points. Scandinavian Journal of Statistics. Theory and Applications, 33(2):279–291, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616925.

Preston:2010:TSB

[PW10] Simon P. Preston and Andrew T. A. Wood.

Two-sample bootstrap hypothesis tests for threedimensional labelled landmark data. Scandinavian Journal of Statistics. Theory and Applications, 37(4): 568–587, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000409.

Pagani:2022:DRD

[PWN22] Filippo Pagani, Martin Wiegand, and Saralees Nadarajah. An n-dimensional Rosenbrock distribution for Markov chain Monte Carlo testing. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 657–680, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Parzen:1997:SCI

[PWY97] M. I. Parzen, L. J. Wei, and Z. Ying. Simultaneous confidence intervals for the difference of two survival functions. Scandinavian Journal of Statistics. Theory and Applications, 24(3): 309–314, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616457.

Qian:2023:NBS

[QB23] Jing Qian and Rebecca A.

Betensky. Nonparametric bounds for the survivor function under general dependent truncation. Scandinavian Journal of Statistics. Theory and Applications, 50(1):327–357, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Qin:1998:SLB

[Qin98] Jing Qin. Semiparametric likelihood based method for goodness of fit tests and estimation in upgraded mixture models. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 681–691, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616532.

Qin:2001:ELC

[QJ01] Gengsheng Qin and Bing-Yi Jing. Empirical likelihood for censored linear regression. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 661–673, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616688.

Qi:2015:SPC

[QL15] Xin Qi and Ruiyan Luo. Sparse principal component analysis in Hilbert space. Scandinavian Journal of Statistics. Theory and Applications, 42(1):270–289, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Quintana:2015:CSV

[QMP15] Fernando A. Quintana, Peter Müller, and Ana Luisa Papoila. Cluster-specific variable selection for product partition models. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1065–1077, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Qiu:2016:CEE

[QQZ16] Zhiping Qiu, Jing Qin, and Yong Zhou. Composite estimating equation method for the accelerated failure time model with lengthbiased sampling data. Scandinavian Journal of Statistics. Theory and Applications, 43(2):396–415, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Qiu:2008:SME

[QST08] Zhenguo Qiu, Peter X.- K. Song, and Ming Tan. Simplex mixed-effects models for longitudinal proportional data. Scandinavian Journal of Statistics. Theory and Applications, 35(4): 577–596, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000290.

Quessy:2012:TBE

[Que12] Jean-François Quessy. Testing for bivariate extreme dependence using Kendall’s process. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 497–514, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41679808.

Qin:1996:ELS

[QW96] Jing Qin and Augustine Wong. Empirical likelihood in a semi-parametric model. Scandinavian Journal of Statistics. Theory and Applications, 23(2): 209–219, June 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616397.

Qaqish:2012:ORE

[QZP12] Bahjat F. Qaqish, Richard C. Zink, and John S. Preisser. Orthogonalized residuals for estimation of marginally specified association parameters in multivariate binary data. Scandinavian Journal of Statistics. Theory and Applications, 39(3):515–527, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679809.

Rahiala:1986:IPE

[Rah86] Markku Rahiala. Identification and preliminary estimation in linear transfer function models. Scandinavian Journal of Statistics. Theory and Applications, 13(4):239–255, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616033.

Ramsay:2005:DFM

[Ram05] James O. Ramsay. Discussion of “Functional Modelling”. Scandinavian Journal of Statistics. Theory and Applications, 32(2): 243, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616874. See [Mül05b].

Ranneby:1975:APE

[Ran75] Bo Ranneby. Asymptotic properties of the empirical transition frequencies for a stationary chain with discrete time. Scandinavian Journal of Statistics. Theory and Applications, 2 (2):103–105, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615585.

Ranneby:1978:NSC

[Ran78] Bo Ranneby. On necessary and sufficient conditions for consistency of MLE’s in Markov chain models. Scandinavian Journal of Statistics. Theory and Applications, 5(2):99– 105, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615694.

Ranneby:1984:MSM

[Ran84] Bo Ranneby. The maximum spacing method. an estimation method related to the maximum likelihood method. Scandinavian Journal of Statistics. Theory and Applications, 11(2):93–112, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615946.

Rapallo:2003:AMB

[Rap03] Fabio Rapallo. Algebraic Markov bases and MCMC for two-way contingency tables. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 385–397, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616770.

Rapallo:2012:OPO

[Rap12] Fabio Rapallo. Outliers and patterns of outliers in contingency tables with algebraic statistics. Scandinavian Journal of Statistics. Theory and Applications, 39(4):784–797, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23357225.

Randall:2021:SDR

[RAQ21] Hayley Randall, Andreas Artemiou, and Xingye Qiao. Sufficient dimension reduction based on distance-weighted discrimination. Scandinavian Journal of Statistics. Theory and Applications, 48(4): 1186–1211, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rodrigues:2010:CCB

[RD10] Alexandre Rodrigues and Peter J. Diggle. A class of convolution-based models for spatio-temporal processes with non-separable covariance structure. Scandinavian Journal of Statistics. Theory and Applications, 37(4):553–567, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000408.

Romanescu:2017:FIN

[RD17a] Razvan G. Romanescu and Rob Deardon. Fast inference for network models of infectious disease spread. Scandinavian Journal of Statistics. Theory and Applications, 44(3):666–683, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rosen:2017:OR

[RD17b] Sophia Rosen and Ori Davidov. Ordered regressions. Scandinavian Journal of Statistics. Theory and Applications, 44(4):817–842, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Reiss:1981:NES

[Rei81] R.-D. Reiss. Nonparametric estimation of smooth distribution functions. Scandinavian Journal of Statistics.

Theory and Applications, 8(2):116–119, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615820.

Ren:2003:GFT

[Ren03] Jian-Jian Ren. Goodness of fit tests with interval censored data. Scandinavian Journal of Statistics. Theory and Applications, 30 (1):211–226, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616758.

Roy:2022:ENE

[RFK22] Pourab Roy, Jason P. Fine, and Michael R. Kosorok. Efficiency of naive estimators for accelerated failure time models under lengthbiased sampling. Scandinavian Journal of Statistics. Theory and Applications, 49(2):525–541, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rousson:2003:CII

[RGS03] Valentin Rousson, Theo Gasser, and Burkhardt Seifert. Confidence intervals for intraclass correlation in inter-rater reliability. Scandinavian Journal of Statistics. Theory and Applications, 30(3): 617–624, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616788.

Richardson:2003:MPA

[Ric03] Thomas Richardson. Markov properties for acyclic directed mixed graphs. Scandinavian Journal of Statistics. Theory and Applications, 30(1):145–157, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616754.

Rice:2023:DDV

[Ric23] Kenneth Rice. Discussion of “Divergence vs. Decision P -values: a Distinction Worth Making in Theory and Keeping in Practice — or, How Divergence P -values Measure Evidence Even When Decision P - values Do Not” by Sander Greenland. Scandinavian Journal of Statistics. Theory and Applications, 50(3): 931–933, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Risager:1980:SCA

[Ris80] Folmer Risager. Simple correlated autoregressive processes. Scandinavian Journal of Statistics. Theory and Applications, 7(2):49– 60, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615773.

Risager:1981:MCS

[Ris81] Folmer Risager. Model checking of simple correlated autoregressive processes. Scandinavian Journal of Statistics. Theory and Applications, 8(3):137– 153, September 1981. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615827.

Ritz:2004:GFT

[Rit04] Christian Ritz. Goodnessof-fit tests for mixed models. Scandinavian Journal of Statistics. Theory and Applications, 31(3): 443–458, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616841.

Roverato:2006:BOV

[RL06] Alberto Roverato and Luca La Rocca. On block ordering of variables in graphical modelling. Scandinavian Journal of Statistics. Theory and Applications, 33(1):65–81, March 2006.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616909.

Roininen:2018:SAF

[RLOS18] Lassi Roininen, Sari Lasanen, Mikko Orispää, and Simo Särkkä. Sparse approximations of fractional Matérn fields. Scandinavian Journal of Statistics. Theory and Applications, 45 (1):194–216, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rabier:2019:AHD

[RMG19] Charles-Elie Rabier, Brigitte Mangin, and Simona Grusea. On the accuracy in highdimensional linear models and its application to genomic selection. Scandinavian Journal of Statistics. Theory and Applications, 46(1):289–313, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Robinson:1978:SIR

[Rob78] P. M. Robinson. Statistical inference for a random coefficient autoregressive model. Scandinavian Journal of Statistics. Theory and Applications, 5(3): 163–168, December 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615707.

Rojo:1998:EQF

[Roj98]  Javier Rojo.  Estima-

tion of the quantile function of an IFRA distribution. Scandinavian Journal of Statistics. Theory and Applications, 25(2): 293–310, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616502.

Romano:2004:NPT

[Rom04] Joseph P. Romano. On non-parametric testing, the uniform behaviour of the t-test, and related problems. Scandinavian Journal of Statistics. Theory and Applications, 31(4): 567–584, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616851.

Ronchetti:2016:DPA

[Ron16] Elvezio Ronchetti. Discussion of the paper “Asymptotic Theory of Outlier Detection Algorithms for Linear Time Series Regression Models” by Johansen & Nielsen. Scandinavian Journal of Statistics. Theory and Applications, 43(2):

368–370, June 2016. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JN16a].

Rosen:1974:ATRa

[Ros74a] Bengt Rosén. Asymptotic theory for Des Raj’s estimator. I. Scandinavian Journal of Statistics. Theory and Applications, 1(2):71– 83, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615555.

Rosen:1974:ATRb

[Ros74b] Bengt Rosén. Asymptotic theory for Des Raj’s estimator. II. Scandinavian Journal of Statistics. Theory and Applications, 1 (3):135–144, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615565.

Rosenlund:1977:UPT

[Ros77] Stig I. Rosenlund. Upwards passage times in the nonnegative birth-death process. Scandinavian Journal of Statistics. Theory and Applications, 4(2):90– 92, June 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615655.

Rosenlund:1978:TPT

[Ros78] Stig I. Rosenlund. Transition probabilities for a truncated birth–death process. Scandinavian Journal of Statistics. Theory and Applications, 5(2):119–122, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615698.

Rosenbaum:1989:SAM

[Ros89] P. R. Rosenbaum. Sensitivity analysis for matched observational studies with many ordered treatments. Scandinavian Journal of Statistics. Theory and Applications, 16(3):227–236, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616136.

Roverato:2002:HIW

[Rov02] Alberto Roverato. Hyper inverse Wishart distribution for non-decomposable graphs and its application to Bayesian inference for Gaussian graphical models. Scandinavian Journal of Statistics. Theory and Applications, 29(3): 391–411, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616723.

Roverato:2005:UAC

[Rov05] Alberto Roverato. A unified approach to the characterization of equivalence classes of DAGs, chain graphs with no flags and chain graphs. Scandinavian Journal of Statistics. Theory and Applications, 32(2):295–312, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616879.

Roverato:2015:LML

[Rov15] Alberto Roverato. Logmean linear parameterization for discrete graphical models of marginal independence and the analysis of dichotomizations. Scandinavian Journal of Statistics. Theory and Applications, 42(2):627–648, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rotnitzky:1995:SPE

[RR95] Andrea Rotnitzky and James M. Robins. Semiparametric estimation of models for means and covariances in the presence of missing data. Scandinavian Journal of Statistics. Theory and Applications, 22(3):

323–333, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616363.

Roberts:2001:MCI

[RR01] Gareth O. Roberts and Jeffrey S. Rosenthal. Markov chains and De-initializing processes. Scandinavian Journal of Statistics. Theory and Applications, 28(3): 489–504, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616674.

Rootzen:2013:E

[RR13] Holger Rootzén and Mats Rudemo. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 40(1):1, March 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357249.

Rasch:1983:DEE

[RS83] D. Rasch and E. Schimke.

Distribution of estimators in exponential regression: a simulation study. Scandinavian Journal of Statistics. Theory and Applications, 10(4):293–300, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615931.

Rudemo:1994:ADM

[RS94] Mats Rudemo and Henrik Stryhn. Approximating the distribution of maximum likelihood contour estimators in two-region images. Scandinavian Journal of Statistics. Theory and Applications, 21(1):41– 55, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616297.

Robeva:2021:MLE

[RSTU21] Elina Robeva, Bernd Sturmfels, Ngoc Tran, and Caroline Uhler. Maximum likelihood estimation for totally positive log-concave densities. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 817–844, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rue:2002:FGM

[RT02] Håvard Rue and Håkon Tjelmeland. Fitting Gaussian Markov random fields to Gaussian fields. Scandinavian Journal of Statistics. Theory and Applications, 29(1):31–49, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616697.

Rubin:2004:DIC

[Rub04a] Donald B. Rubin. Direct and indirect causal effects via potential outcomes*. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 161–170, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616821.

Rubin:2004:RD

[Rub04b] Donald B. Rubin. Reply to discussion. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 196–198, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616823. See [AP04].

Rudemo:1982:ECH

[Rud82] Mats Rudemo. Empirical choice of histograms and kernel density estimators. Scandinavian Journal of Statistics. Theory and Applications, 9(2):65– 78, June 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615859.

Rue:1997:LFM

[Rue97] Håvard Rue. A loss function model for the restoration of grey level images. Scandinavian Journal of Statistics. Theory and Applications, 24(1):103–114, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616439.

Rivero:2004:MBI

[RV04] Carlos Rivero and Teófilo Valdés. Mean-based iterative procedures in linear models with general errors and grouped data. Scandinavian Journal of Statistics. Theory and Applications, 31(3):469–486, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616843.

Robins:2015:PBI

[RVG15] James M. Robins, Tyler J. VanderWeele, and Richard D. Gill. A proof of Bell’s inequality in quantum mechanics using causal interactions. Scandinavian Journal of Statistics. Theory and Applications, 42(2): 329–335, June 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Rivera:2013:ODJ

[RW13] Camilo Rivera and Guenther Walther. Optimal detection of a jump in the intensity of a Poisson process or in a density with likelihood ratio statistics. Scandinavian Journal of Statistics. Theory and Applications, 40(4):752–769, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586698.

Rychlik:1996:FSL

[Ryc96] Igor Rychlik. Fatigue and stochastic loads. Scandinavian Journal of Statistics. Theory and Applications, 23(4):387–404, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616416.

Ryden:1995:CAN

[Ryd95] Tobias Rydén. Consistent and asymptotically normal parameter estimates for Markov modulated Poisson processes. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 295–303, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616361.

Rougier:2016:VLS

[RZM16] Jonathan Rougier and Andrew Zammit-Mangion. Visualization for large-scale Gaussian updates. Scandinavian Journal of Statistics. Theory and Applications, 43(4):1153–1161, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Spjotvoll:1980:CRE

[SA80] Emil Spjøtvoll and Are Halvor Aastveit. Comparison of robust estimators on data from field experiments. Scandinavian Journal of Statistics. Theory and Applications, 7(1):1–13, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615764.

Saarela:2011:MBM

[SA11] Olli Saarela and Elja Arjas. A method for Bayesian monotonic multiple regression. Scandinavian Journal of Statistics. Theory and Applications, 38(3): 499–513, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015577.

Saarela:2015:NPB

[SA15] Olli Saarela and Elja Arjas. Non-parametric Bayesian hazard regression for chronic disease risk assessment. Scandinavian Journal of Statistics. Theory and Applications, 42 (2):609–626, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Saegusa:2015:VEU

[Sae15] Takumi Saegusa. Variance estimation under two-phase sampling. Scandinavian Journal of Statistics. Theory and Applications, 42(4): 1078–1091, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Saikkonen:1983:APS

[Sai83] Pentti Saikkonen. Asymptotic properties of some tests for cross correlation. Scandinavian Journal of Statistics. Theory and Applications, 10(4):269–280, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615929.

Sakamoto:2019:BRM

[Sak19] Wataru Sakamoto. Biasreduced marginal Akaike information criteria based on a Monte Carlo method for linear mixed-effects models. Scandinavian Journal of Statistics. Theory and Applications, 46(1):87– 115, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Samuelsen:1989:ATN

[Sam89] Sven Ove Samuelsen. Asymptotic theory for nonparametric estimators from doubly censored data. Scandinavian Journal of Statistics. Theory and Applications, 16(1):1–21, September 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616118.

Sansonnet:2014:WTE

[San14] Laure Sansonnet. Wavelet thresholding estimation in a Poissonian interactions model with application to genomic data. Scandinavian Journal of Statistics. Theory and Applications, 41(1):200–226, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586738.

Sardy:2009:APM

[Sar09] Sylvain Sardy. Adaptive posterior mode estimation of a sparse sequence for model selection. Scandinavian Journal of Statistics. Theory and Applications, 36(4):577–601, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000341.

Sasieni:1992:NOP

[Sas92]  Peter Sasieni.  Non-

orthogonal projections and their application to calculating the information in a partly linear Cox model. Scandinavian Journal of Statistics. Theory and Applications, 19(3):215–233, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616241.

Samuelsen:2007:SCC

[SÅS07] Sven Ove Samuelsen, Hallvard Ånestad, and Anders Skrondal. Stratified case-cohort analysis of general cohort sampling designs. Scandinavian Journal of Statistics. Theory and Applications, 34(1): 103–119, March 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548541.

Song:2024:EEN

[SAS24] Weixing Song, Kanwal Ayub, and Jianhong Shi. Extrapolation estimation for nonparametric regression with measurement error. Scandinavian Journal of Statistics. Theory and Applications, 51(1):1– 31, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sato:1996:MSS

[Sat96] Michikazu Sato. A minimal sufficient statistic and representations of the densities. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 381–384, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616411.

Shenton:1985:DAR

[SB85] L. R. Shenton and K. O. Bowman. Distributions arising from reinforcementdepletion urn models. Scandinavian Journal of Statistics. Theory and Applications, 12(4):293–299, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616001.

Singh:1990:AOB

[SB90] Kesar Singh and G. J.

Babu. On asymptotic optimality of the bootstrap. Scandinavian Journal of Statistics. Theory and Applications, 17(1):1–9, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616150.

Schladitz:2000:TOP

[SB00] K. Schladitz and A. J.

Baddeley. A third order point process characteristic. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 657–671, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616633.

Sutradhar:2005:MRB

[SBB05] Brajendra C. Sutradhar, Atanu Biswas, and Wasimul Bari. Marginal regression for binary longitudinal data in adaptive clinical trials. Scandinavian Journal of Statistics. Theory and Applications, 32 (1):93–113, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616863.

Saltyte-Benth:2005:LDA

[ŠBD05] Jūratė Šaltytė-Benth and Kȩstutis Dučinskas. Linear discriminant analysis of multivariate spatial– temporal regressions. Scandinavian Journal of Statistics. Theory and Applications, 32(2):281–294, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616878.

Skare:2003:ISI

[SBH03] Øivind Skare, Erik Bølviken, and Lars Holden. Improved sampling-importance resampling and reduced bias importance sampling. Scandinavian Journal of Statistics. Theory and Applications, 30(4):719–737, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616798.

Sundberg:1999:MCD

[SBM^{+}99] Rolf Sundberg, Philip J. Brown, Harald Martens, Tormod Næs, Samuel D. Oman, and Svante Wold. Multivariate calibration: Direct and indirect regression methodology [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 26(2):161–207, June 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616550.

Sengupta:1998:TPH

[SBR98] D. Sengupta, A. Bhattacharjee, and B. Rajeev. Testing for the proportionality of hazards in two samples against the increasing cumulative hazard ratio alternative. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 637–647, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616529.

Schelldorfer:2011:EHD

[SBV11] Jürg Schelldorfer, Peter Bühlmann, and Sara Van De Geer. Estimation for high-dimensional linear mixed-effects models using ℓ_{1}-penalization. Scandinavian Journal of Statistics. Theory and Applications, 38(2):197–214, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015490.

Schaubel:2006:RMR

[SC06] Douglas Schaubel and Jianwen Cai. Rate/mean regression for multiplesequence recurrent event data with missing event category. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 191–207, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616918.

Schweder:1975:WEA

[Sch75] Tore Schweder. Window estimation of the asymptotic variance of rank estimators of location. Scandinavian Journal of Statistics. Theory and Applications, 2 (3):113–126, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615591. See cor-

rection [Sch81].

Schweder:1979:SAM

[Sch79] Tore Schweder. A statistical analysis of the mating behaviour of Euchaeta norvegica (Copepoda: Calanoida). Scandinavian Journal of Statistics. Theory and Applications, 6(2):71–76, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615734.

Schweder:1980:SSS

[Sch80] Tore Schweder. Scandinavian statistics, some early lines of development. Scandinavian Journal of Statistics. Theory and Applications, 7(3):113–129, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615787.

Schweder:1981:CNW

[Sch81] Tore Schweder. Correction note: Window Estimation of the Asymptotic Variance of Rank Estimators of Location. Scandinavian Journal of Statistics. Theory and Applications, 8(1):55, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4615813. See [Sch75].

Schweder:1982:DM

[Sch82] Tore Schweder. On the dispersion of mixtures. Scandinavian Journal of Statistics. Theory and Applications, 9(3):165–169, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615873.

Scheike:1994:PRL

[Sch94] Thomas H. Scheike. Parametric regression for longitudinal data with counting process measurement times. Scandinavian Journal of Statistics. Theory and Applications, 21(3): 245–263, September 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616315.

Schweder:2002:LSM

[Sch02]  Tore Schweder.  Large

structured models in applied sciences; challenges for statistics. Scandinavian Journal of Statistics. Theory and Applications, 29(2):189–191, June 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616711.

Scheike:2004:E

[Sch04] Thomas Scheike. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 31(1): 1, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616807.

Scricciolo:2007:RCB

[Scr07]  Catia Scricciolo.  On

rates of convergence for Bayesian density estimation. Scandinavian Journal of Statistics. Theory and Applications, 34(3): 626–642, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548573.

Singh:1985:SNA

[SD85] Harshinder Singh and Jayant V. Deshpande. On some new ageing properties. Scandinavian Journal of Statistics. Theory and Applications, 12(3): 213–220, September 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615990.

Seeger:1993:ASM

[See93] Paul Seeger. Application of statistical methods for design and analysis of individual bridge competitions with a small number of players. Scandinavian Journal of Statistics. Theory and Applications, 20(1): 89–95, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

4616264. See corrigenda [Ano96e].

Seeger:1996:NRC

[See96]  Paul Seeger.  Nested

row–column and generalized Howell designs used in bridge competitions. Scandinavian Journal of Statistics. Theory and Applications, 23(2):243–255, June 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616400.

Segers:2002:ATT

[Seg02] Johan Segers. Abelian and Tauberian theorems on the bias of the Hill estimator. Scandinavian Journal of Statistics. Theory and Applications, 29(3):461–483, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616727.

Seifert:2024:CVA

[Sei24] Miriam Isabel Seifert. Characterization of valid auxiliary functions for representations of extreme value distributions and their max-domains of attraction. Scandinavian Journal of Statistics. Theory and Applications, 51 (2):832–860, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sengupta:1988:LPU

[Sen88] Ashis Sengupta. On loss of power under additional information: an example. Scandinavian Journal of Statistics. Theory and Applications, 15(1):25– 31, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616081.

Seaman:2016:LIN

[SFW16] Shaun R. Seaman, Daniel Farewell, and Ian R. White. Linear increments with non-monotone missing data and measurement error. Scandinavian Journal of Statistics. Theory and Applications, 43(4):996–1018, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Singh:1978:PCM

[SG78] Jagbir Singh and R. S.

Gupta. A paired comparison model allowing for ties. Scandinavian Journal of Statistics. Theory and Applications, 5(1):65– 68, September 1978. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615686.

Sjogren:2004:CWS

[SG04] Niclas Sjögren and Olivier Guilbaud. Comparison of within-subject covariance matrices in 2 × 2 crossover trials with multivariate response. Scandinavian Journal of Statistics. Theory and Applications, 31(4): 603–612, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616853.

Sun:2012:ESM

[SG12] Yanqing Sun and Peter B. Gilbert. Estimation of stratified markspecific proportional hazards models with missing marks. Scandinavian Journal of Statistics. Theory and Applications, 39(1):34– 52, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411156.

Shen:2015:ABP

[SG15] Weining Shen and Subhashis Ghosal. Adaptive Bayesian procedures using random series priors. Scandinavian Journal of Statistics. Theory and Applications, 42(4):1194–1213, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Scheel:2011:GDI

[SGR11] Ida Scheel, Peter J. Green, and Jonathan C. Rougier. A graphical diagnostic for identifying influential model choices in Bayesian hierarchical models. Scandinavian Journal of Statistics. Theory and Applications, 38(3):529–550, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015579.

Shao:1996:DFT

[SH96] Yongzhao Shao and Marjorie G. Hahn. On a distribution-free test of fit for continuous distribution functions. Scandinavian Journal of Statistics. Theory and Applications, 23(1): 63–73, March 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616386.

Schweder:2002:CL

[SH02] Tore Schweder and Nils Lid Hjort. Confidence and likelihood*. Scandinavian Journal of Statistics. Theory and Applications, 29(2): 309–332, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616717.

Stoltenberg:2021:MID

[SH21] Emil Aas Stoltenberg and Nils Lid Hjort. Models and inference for on–off data via clipped Ornstein– Uhlenbeck processes. Scandinavian Journal of Statistics. Theory and Applications, 48(3):908–929, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Shao:2012:PIS

[Sha12] Xiaofeng Shao. Parametric inference in stationary time series models with dependent errors. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 772–783, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/23357224.

Sheather:1994:DBB

[SHD94] Simon J. Sheather, Thomas P. Hettmansperger, and Margaret R. Donald. Databased bandwidth selection for kernel estimators of the integral of f^{2}(x). Scandinavian Journal of Statistics. Theory and Applications, 21(3):265–275, September 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616316.

Shimizu:2017:TES

[Shi17] Yasutaka Shimizu. Threshold estimation for stochastic processes with small noise. Scandinavian Journal of Statistics. Theory and Applications, 44(4): 951–988, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sibson:1980:DTA

[Sib80] Robin Sibson. The Dirichlet tessellation as an aid in data analysis. Scandinavian Journal of Statistics. Theory and Applications, 7 (1):14–20, December 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615765.

Simpson:2014:CDP

[Sim14] Daniel Simpson. Contribution to the discussion of the paper P“Geodesic Monte Carlo on Embedded Manifolds”. Scandinavian Journal of Statistics. Theory and Applications, 41(1):16– 18, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

24586726. See [BG13].

Sengupta:1993:IDL

[SJ93] Debasis Sengupta and S. Rao Jammalamadaka. Inference from discrete life history data: a counting process approach. Scandinavian Journal of Statistics. Theory and Applications, 20(1):51–61, June 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616259.

Stockmarr:1994:GDA

[SJ94] Anders Stockmarr and Martin Jacobsen. Gaussian diffusions and autoregressive processes: Weak convergence and statistical inference. Scandinavian Journal of Statistics. Theory and Applications, 21(4): 403–419, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616326.

Sant:2022:CLR

[SJKS22] Jaromir Sant, Paul A. Jenkins, Jere Koskela, and Dario Spanò. Convergence of likelihood ratios and estimators for selection in nonneutral Wright–Fisher diffusions. Scandinavian Journal of Statistics. Theory and Applications, 49(4): 1728–1760, December 2022.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sjostedt:2000:RDR

[Sjö00] Sara Sjöstedt. Resampling m-dependent random variables with applications to forecasting. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 543–561, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616622.

Sutradhar:2008:UGQ

[SJS08] Brajendra C. Sutradhar, Vandna Jowaheer, and Gary Sneddon. On a unified generalized quasilikelihood approach for familial-longitudinal nonstationary count data. Scandinavian Journal of Statistics. Theory and Applications, 35(4):597–612, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000291.

Sun:2019:HTR

[SJZI19] Qiang Sun, Bai Jiang, Hongtu Zhu, and Joseph G. Ibrahim. Hard thresholding regression. Scandinavian Journal of Statistics. Theory and Applications, 46(1):314–328, March 2019.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sugasawa:2019:ATM

[SK19] Shonosuke Sugasawa and Tatsuya Kubokawa. Adaptively transformed mixedmodel prediction of general finite-population parameters. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 1025–1046, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Safken:2020:CCP

[SK20] Benjamin Säfken and Thomas Kneib. Conditional covariance penalties for mixed models. Scandinavian Journal of Statistics. Theory and Applications, 47(3): 990–1010, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Steutel:1979:IDT

[SKBBN79] F. W. Steutel, John T. Kent, Lennart Bondesson, and Ole Barndorff-Nielsen. Infinite divisibility in theory and practice [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 6(2):57–64, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615732.

Skovgaard:1981:EED

[Sko81a] Ib M. Skovgaard. Edgeworth expansions of the distributions of maximum likelihood estimators in the general (non I.I.D.) case. Scandinavian Journal of Statistics. Theory and Applications, 8(4):227– 236, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615841.

Skovgaard:1981:TEE

[Sko81b] Ib M. Skovgaard. Transformation of an Edgeworth expansion by a sequence of smooth functions. Scandinavian Journal of Statistics. Theory and Applications, 8(4):207– 217, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615839.

Skovgaard:1984:RGM

[Sko84] Lene Theil Skovgaard. A Riemannian geometry of the multivariate normal model. Scandinavian Journal of Statistics. Theory and Applications, 11(4): 211–223, September 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615960.

Skovgaard:1986:SMC

[Sko86] Ib Skovgaard. A statistical model for competition experiments. Scandinavian Journal of Statistics. Theory and Applications, 13 (1):29–38, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616008.

Skovgaard:2001:LA

[Sko01] Ib M. Skovgaard. Likelihood asymptotics. Scandinavian Journal of Statistics. Theory and Applications, 28(1):3–32, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616643.

Sugasawa:2017:EUB

[SKO17] Shonosuke Sugasawa, Tatsuya Kubokawa, and Kota Ogasawara. Empirical uncertain Bayes methods in area-level models. Scandinavian Journal of Statistics. Theory and Applications, 44(3):684–706, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sugasawa:2019:HBS

[SKR19] Shonosuke Sugasawa, Tatsuya Kubokawa, and J. N. K. Rao. Hierarchical Bayes small-area estimation with an unknown link function. Scandinavian Journal of Statistics. Theory and Applications, 46(3):885–897, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Saikkonen:1988:LMT

[SL88] Pentti Saikkonen and Ritva Luukkonen. Lagrange multiplier tests for testing nonlinearities in time series models. Scandinavian Journal of Statistics. Theory and Applications, 15(1):55– 68, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616085.

Singh:1990:VJK

[SL90] Kesar Singh and Regina Y.

Liu. On the validity of the jack-knife procedure. Scandinavian Journal of Statistics. Theory and Applications, 17(1):11–21, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616151.

Svensson:2006:EWF

[SLB06] Ingrid Svensson, Sara Sjöstedt-De Luna, and Lennart Bondesson. Estimation of wood fibre length distributions from censored data through an EM algorithm. Scandinavian Journal of Statistics. Theory and Applications, 33(3): 503–522, September 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616939.

Shen:2019:RAL

[SLCN19] Weining Shen, Suyu Liu, Yong Chen, and Jing Ning. Regression analysis of longitudinal data with outcome-dependent sampling and informative censoring. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 831–847, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Staicu:2014:LRT

[SLCR14] Ana-Maria Staicu, Yingxing Li, Ciprian M. Crainiceanu, and David Ruppert. Likelihood ratio tests for dependent data with applications to longitudinal and functional data analysis. Scandinavian Journal of Statistics. Theory and Applications, 41(4):932–949, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586814.

Shahbaba:2014:CDP

[SLS14] Babak Shahbaba, Shiwei Lan, and Jeffrey Streets. Contribution to the discussion of the paper ‘Geodesic Monte Carlo on Embedded Manifolds’. Scandinavian Journal of Statistics. Theory and Applications, 41(1): 14–15, March 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/

24586725. See [BG13].

Slud:1992:PLC

[Slu92]  Eric V. Slud.  Partial

likelihood for continuoustime stochastic processes. Scandinavian Journal of Statistics. Theory and Applications, 19(2):97–109, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616231.

Slud:1997:TID

[Slu97] Eric Slud. Testing for imperfect debugging in software reliability. Scandinavian Journal of Statistics. Theory and Applications, 24(4):555–572, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616476.

Scheike:2004:MLE

[SM04a] Thomas H. Scheike and Torben Martinussen. Maximum likelihood estimation for Cox’s regression model under case–cohort sampling. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 283–293, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616829.

Scheike:2004:ETT

[SM04b] Thomas H. Scheike and Torben Martinussen. On estimation and tests of time-varying effects in the proportional hazards model. Scandinavian Journal of Statistics. Theory and Applications, 31(1):51– 62, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616811.

Senturk:2005:CAC

[ŞM05] Damla Şentürk and Hans-Georg Müller. Covariate adjusted correlation analysis via varying coefficient models. Scandinavian Journal of Statistics. Theory and Applications, 32(3): 365–383, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616885.

Saraiva:2012:CGE

[SM12] Erlandson F. Saraiva and Luis A. Milan. Clustering gene expression data using a posterior split–merge– birth procedure. Scandinavian Journal of Statistics. Theory and Applications, 39(3):399–415, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41679803.

Sardy:2024:SAM

[SM24] Sylvain Sardy and Xiaoyu Ma. Sparse additive models in high dimensions with wavelets. Scandinavian Journal of Statistics. Theory and Applications, 51 (1):89–108, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sharma:2014:CMB

[SMB14] Gaurav Sharma, Thomas Mathew, and Ionut Bebu. Combining multivariate bioassays: Accurate inference using small sample asymptotics. Scandinavian Journal of Statistics. Theory and Applications, 41 (1):152–166, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586735.

Song:2012:RAL

[SMS12] Xinyuan Song, Xiaoyun Mu, and Liuquan Sun. Regression analysis of longitudinal data with timedependent covariates and informative observation times. Scandinavian Journal of Statistics. Theory and Applications, 39(2): 248–258, June 2012. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41679793.

Seillier-Moiseiwitsch:1992:PTM

[SMSD92] F. Seillier-Moiseiwitsch, T. J. Sweeting, and A. Philip Dawid. Prequential tests of model fit. Scandinavian Journal of Statistics. Theory and Applications, 19 (1):45–60, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616224.

Sellero:2005:URP

[SMV05] César Sánchez Sellero, Wenceslao González Manteiga, and Ingrid Van Keilegom. Uniform representation of product–limit integrals with applications. Scandinavian Journal of Statistics. Theory and Applications, 32(4):563–581, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616898.

Scheike:2011:ARM

[SMZ11] Thomas H. Scheike, Torben Martinussen, and Mei-Jie Zhang. The additive risk model for estimation of effect of haplotype match in BMT studies. Scandinavian Journal of Statistics. Theory and Applications, 38(3): 409–423, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015572.

Schweder:1988:SVB

[SN88] Tore Schweder and Ragnar Norberg. A significance version of the basic Neyman–Pearson theory for scientific hypothesis testing [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 15(4): 225–242, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616108.

Selk:2013:TCI

[SN13] Leonie Selk and Natalie Neumeyer. Testing for a change of the innovation distribution in nonparametric autoregression: The sequential empirical process approach. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 770–788, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586699.

Saegusa:2024:CBS

[SN24] Takumi Saegusa and Peter Nandori. Confidence bands for survival curves from outcome-dependent stratified samples. Scandinavian Journal of Statistics. Theory and Applications, 51 (3):1086–1102, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Syversveen:1997:CMP

[SO97] Anne Randi Syversveen and Henning Omre. Conditioning of marked point processes within a Bayesian framework. Scandinavian Journal of Statistics. Theory and Applications, 24(3): 341–352, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616460.

Saetrom:2013:UQE

[SO13] Jon Sætrom and Henning Omre. Uncertainty quantification in the ensemble Kalman filter. Scandinavian Journal of Statistics. Theory and Applications, 40(4):868–885, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586704.

Song:2000:MDM

[Son00] Peter Xue-Kun Song. Multivariate dispersion models generated from Gaussian copula. Scandinavian Journal of Statistics. Theory and Applications, 27(2):305–320, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616605.

Sorensen:1998:CST

[Sør98]  Michael Sørensen.  On

comparison of stopping times in sequential procedures for exponential families of stochastic processes. Scandinavian Journal of Statistics. Theory and Applications, 25(2):

331–343, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616505.

Sorensen:2001:DOD

[Sør01]  Helle Sørensen.  Dis-

cretely observed diffusions: Approximation of the continuous-time score function. Scandinavian Journal of Statistics. Theory and Applications, 28 (1):113–121, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616647.

Sorensen:2003:SLA

[Sør03]  Helle Sørensen.  Simu-

lated likelihood approximations for stochastic volatility models. Scandinavian Journal of Statistics. Theory and Applications, 30(2): 257–276, June 2003. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616763.

Sergides:2009:FDT

[SP09] Marios Sergides and Efstathios Paparoditis. Frequency domain tests of semiparametric hypotheses for locally stationary processes. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 800–821, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000353.

Senoussi:2022:NST

[SP22] Rachid Senoussi and Emilio Porcu. Nonstationary space–time covariance functions induced by dynamical systems. Scandinavian Journal of Statistics. Theory and Applications, 49 (1):211–235, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Spjotvoll:1974:MTA

[Spj74]  Emil Spjøtvoll.  Multi-

ple testing in the analysis of variance. Scandinavian Journal of Statistics. Theory and Applications, 1(3):97–114, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615562.

Schwartz:2023:LNM

[SPK23] William K. Schwartz, Sonja Petrović, and Hemanshu Kaul. Longitudinal network models and permutationuniform Markov chains. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1201–1231, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sermaidis:2013:MCM

[SPR^{+}13] Giorgos Sermaidis, Omiros Papaspiliopoulos, Gareth O. Roberts, Alexandros Beskos, and Paul Fearnhead. Markov chain Monte Carlo for exact inference for diffusions. Scandinavian Journal of Statistics. Theory and Applications, 40(2): 294–321, June 2013. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586714.

Sun:2001:NPT

[SR01] Jianguo Sun and Shesh N. Rai. Non-parametric tests for the comparison of point processes based on incomplete data. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 725–732, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616692.

Skold:2003:DEM

[SR03] M. Sköld and G. O. Roberts. Density estimation for the Metropolis– Hastings algorithm. Scandinavian Journal of Statistics. Theory and Applications, 30(4):699–718, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616797.

Sorbye:2011:SCB

[SR11] Sigrunn H. Sørbye and Håvard Rue. Simultaneous credible bands for latent Gaussian models. Scandinavian Journal of Statistics. Theory and Applications, 38(4):712–725, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411146.

Skrondal:2007:LVM

[SRH07] Anders Skrondal and Sophia Rabe-Hesketh. Latent variable modelling: a survey*. Scandinavian Journal of Statistics. Theory and Applications, 34(4):712–745, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548578.

Srivastava:1997:RRD

[Sri97]  M. S. Srivastava.  Re-

duced rank discrimination. Scandinavian Journal of Statistics. Theory and Applications, 24(1):115–124, March 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616440.

Stene:1979:DSF

[SS79] Jon Stene and Eeva Stene. Down’s Syndrome and father’s age. Scandinavian Journal of Statistics. Theory and Applications, 6 (4):180–181, March 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615759.

Stute:1980:GGC

[SS80] W. Stute and G. Schumann. A general Glivenko– Cantelli theorem for stationary sequences of random observations. Scandinavian Journal of Statistics. Theory and Applications, 7(2):102–104, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615782.

Stoica:1998:PLS

[SS98] Petre Stoica and Torsten Söderström. Partial least squares: a first-order analysis. Scandinavian Journal of Statistics. Theory and Applications, 25(1):17– 24, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616481.

Stoyan:2000:IRE

[SS00] Dietrich Stoyan and Helga Stoyan. Improving ratio estimators of second order point process characteristics. Scandinavian Journal of Statistics. Theory and Applications, 27(4): 641–656, December 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616632.

Sain:2002:ZBL

[SS02] Stephan R. Sain and David W. Scott. Zero-bias locally adaptive density estimators. Scandinavian Journal of Statistics. Theory and Applications, 29(3): 441–460, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616726.

Schmidt:2006:NPE

[SS06] Rafael Schmidt and Ulrich Stadtmüller. Nonparametric estimation of tail dependence. Scandinavian Journal of Statistics. Theory and Applications, 33(2):307–335, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616927.

Steinbakk:2009:PPV

[SS09] Gunnhildur Högnadóttir Steinbakk and Geir Olve Storvik. Posterior predictive p-values in Bayesian hierarchical models. Scandinavian Journal of Statistics. Theory and Applications, 36(2):320–336, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000323.

Savitsky:2018:SBU

[SS18] Terrance D. Savitsky and Sanvesh Srivastava. Scalable Bayes under informative sampling. Scandinavian Journal of Statistics. Theory and Applications, 45(3):534–556, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Salehabadi:2015:PEM

[SSD15] Sedigheh Mirzaei Salehabadi, Debasis Sengupta, and Rituparna Das. Parametric estimation of Menarcheal age distribution based on recall data. Scandinavian Journal of Statistics. Theory and Applications, 42(1):290–305, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Sun:2009:ELI

[SSZ09] Yanqing Sun, Rajeshwari Sundaram, and Yichuan Zhao. Empirical likelihood inference for the Cox model with time-dependent coefficients via local partial likelihood. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 444–462, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000332.

Srivastava:1976:EEP

[ST76] V. K. Srivastava and Ramji Tiwari. Evaluation of expectations of products of stochastic matrices. Scandinavian Journal of Statistics. Theory and Applications, 3(3):135–138, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615627.

Stepniak:1981:CLM

[ST81] Czesław Stȩpniak and Erik Torgersen. Comparison of linear models with partially known covariances with respect to unbiased estimation. Scandinavian Journal of Statistics. Theory and Applications, 8(3):183–184, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615831.

Sardy:2010:DET

[ST10] Sylvain Sardy and Paul Tseng. Density estimation by total variation penalized likelihood driven by the sparsity ℓ_{1} information criterion. Scandinavian Journal of Statistics. Theory and Applications, 37(2):321–337, June 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000882.

Stove:2012:CED

[ST12] Bård Støve and Dag Tjøstheim. A convolution estimator for the density of nonlinear regression observations. Scandinavian Journal of Statistics. Theory and Applications, 39(2):282–304, June 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41679795.

Stefanov:1988:SST

[Ste88] Valeri T. Stefanov. On some stopping times for dependent bemoulli trials. Scandinavian Journal of Statistics. Theory and Applications, 15(1):39– 50, March 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616083.

Stefanov:1991:NEF

[Ste91] Valeri T. Stefanov. Noncurved exponential families associated with observations over finite state Markov chains. Scandinavian Journal of Statistics. Theory and Applications, 18(4):353–356, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616217.

Sarndal:1978:DBM

[STH^{+}78] Carl-Erik Särndal, Ib Thomsen, Jan M. Hoem, D. V. Lindley, O. Barndorff-Nielsen, and Tore Dalenius. Design-based and modelbased inference in survey sampling [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 5 (1):27–52, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615682. See cor-

rection, discussion, and reply [Hoe78].

Stijnen:1982:MEB

[Sti82] Theo Stijnen. A monotone empirical Bayes estimator and test for the one-parameter continuous exponential family based on spacings. Scandinavian Journal of Statistics. Theory and Applications, 9(3): 153–158, September 1982. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615871.

Sugasawa:2017:BES

[STK17] Shonosuke Sugasawa, Hiromasa Tamae, and Tatsuya Kubokawa. Bayesian estimators for small area models shrinking both means and variances. Scandinavian Journal of Statistics. Theory and Applications, 44(1):150–167, March 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Selk:2022:MBR

[STM22] Leonie Selk, Charles Tillier, and Orlando Marigliano. Multivariate boundary regression models. Scandinavian Journal of Statistics. Theory and Applications, 49(1):400–426, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Schmid:2016:ORS

[STMC16] Timo Schmid, Nikos Tzavidis, Ralf Münnich, and Ray Chambers. Outlier robust small-area estimation under spatial correlation. Scandinavian Journal of Statistics. Theory and Applications, 43(3):806–826, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Storvik:2011:FMH

[Sto11]  Geir Storvik.  On the

flexibility of Metropolis– Hastings acceptance probabilities in auxiliary variable proposal generation. Scandinavian Journal of Statistics. Theory and Applications, 38(2):342–358, June 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015498.

Stute:1983:LPT

[Stu83] Winfried Stute. Last passage times of M-estimators. Scandinavian Journal of Statistics. Theory and Applications, 10(4):301–305, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615932.

Stute:1994:BKM

[Stu94] Winfried Stute. The bias of Kaplan–Meier integrals. Scandinavian Journal of Statistics. Theory and Applications, 21(4):475–484, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616331.

Stute:1996:DCU

[Stu96] Winfried Stute. Distributional convergence under random censorship when covariables are present. Scandinavian Journal of Statistics. Theory and Applications, 23(4):461–471, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616420.

Sun:2001:GFT

[STZ01] Yanqing Sun, Ram C. Tiwari, and Jyoti N. Zalkikar. Goodness of fit tests for multivariate counting process models with applications. Scandinavian Journal of Statistics. Theory and Applications, 28 (1):241–256, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616654.

Seppanen:1992:CBL

[SU92] Eila Seppänen and Esa Uusipaikka. Confidence bands for linear regression over restricted regions. Scandinavian Journal of Statistics. Theory and Applications, 19(1):73–81, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616226.

Sundberg:1974:MLT

[Sun74] Rolf Sundberg. Maximum likelihood theory for incomplete data from an exponential family. Scandinavian Journal of Statistics. Theory and Applications, 1(2):49–58, March 1974. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615553.

Sundberg:1975:SRA

[Sun75] Rolf Sundberg. Some results about decomposable (or Markov-type) models for multidimensional contingency tables: Distribution of marginals and partitioning of tests. Scandinavian Journal of Statistics. Theory and Applications, 2(2):71–79, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615579.

Sundberg:1983:PAR

[Sun83] Rolf Sundberg. The predictive approach and randomized population type models for finite population inference from two-stage samples. Scandinavian Journal of Statistics. Theory and Applications, 10(3): 223–238, September 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615920.

Sun:1995:RRM

[Sun95] Li Sun. Risk ratio and minimaxity in estimating the multivariate normal mean with unknown variance. Scandinavian Journal of Statistics. Theory and Applications, 22(1): 105–120, March 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616344.

Sundberg:1996:PEG

[Sun96] Rolf Sundberg. The precision of the estimated generalized least squares estimator in multivariate calibration. Scandinavian Journal of Statistics. Theory and Applications, 23(3): 257–274, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616403.

Sundberg:2010:FML

[Sun10] Rolf Sundberg. Flat and multimodal likelihoods and model lack of fit in curved exponential families. Scandinavian Journal of Statistics. Theory and Applications, 37(4):632–643, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000412.

Spjotvoll:1976:CSL

[SV76] Emil Spjøtvoll and Rolf Volden. On the choice of significance level in linear regression. Scandinavian Journal of Statistics. Theory and Applications, 3 (1):28–32, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615604.

Slud:2004:CNE

[SV04] E. V. Slud and F. Vonta.

Consistency of the NPML estimator in the rightcensored transformation model. Scandinavian Journal of Statistics. Theory and Applications, 31(1):21– 41, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616809.

Swanepoel:2005:NKD

[SV05] Jan W. H. Swanepoel and François C. Van Graan. A new kernel distribution function estimator based on a non-parametric transformation of the data. Scandinavian Journal of Statistics. Theory and Applications, 32(4):551–562, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616897.

Schreiber:2010:DLP

[SV10] Tomasz Schreiber and Marie-Colette Van Lieshout. Disagreement loop and path creation/annihilation algorithms for binary planar Markov fields with applications to image segmentation. Scandinavian Journal of Statistics. Theory and Applications, 37(2): 264–285, June 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000879.

Sverdrup:1975:TP

[Sve75] Erling Sverdrup. Tests without power. Scandinavian Journal of Statistics.

Theory and Applications, 2 (3):158–160, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615597. See reply

[ML75].

Sverdrup:1976:STM

[Sve76]  Erling Sverdrup.  Sig-

nificance testing in multiple statistical inference. Scandinavian Journal of Statistics. Theory and Applications, 3(2):73–78, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615613.

Sverdrup:1977:TPR

[Sve77] Erling Sverdrup. Tests without power: a rejoinder. Scandinavian Journal of Statistics. Theory and Applications, 4(3):136–138, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615665.

Svensson:1986:EEC

[Sve86]  Åke Svensson.  Esti-

mating effects in certain before-and-after studies. Scandinavian Journal of Statistics. Theory and Applications, 13(1):39–47, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616009.

Svensson:1990:AEC

[Sve90a] Åke Svensson. Asymptotic estimation in counting processes with parametric intensities based on one realization. Scandinavian Journal of Statistics. Theory and Applications, 17 (1):23–33, December 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616152.

Sverdrup:1990:DMC

[Sve90b]  Erling Sverdrup.  The

delta multiple comparison method. performance and usefulness. Scandinavian Journal of Statistics. Theory and Applications, 17(2): 115–134, September 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616160.

Stenlund:1975:MCS

[SW75] Hans Stenlund and Anders Westlund. A Monte Carlo study of simple random sampling from a finite population. Scandinavian Journal of Statistics. Theory and Applications, 2 (2):106–108, March 1975.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615586.

Stenlund:1976:ANM

[SW76] Hans Stenlund and Anders Westlund. On the asymptotic normality of the mean estimator: a Monte Carlo study of three sampling designs. Scandinavian Journal of Statistics. Theory and Applications, 3(3): 127–131, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615625.

Sarndal:1984:CFE

[SW84] Carl-Erik Särndal and Roger L. Wright. Cosmetic form of estimators in survey sampling. Scandinavian Journal of Statistics. Theory and Applications, 11 (3):146–156, March 1984. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615952.

Skovlund:1987:SSS

[SW87] Eva Skovlund and Lars Walløe. A simulation study of a sequential t-test developed by Armitage. Scandinavian Journal of Statistics. Theory and Applications, 14(4):347–352, September 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616076.

Sitter:1993:ODB

[SW93a] R. R. Sitter and C. F. J. Wu. Optimal designs for binary response experiments: Fieller, d, and a criteria. Scandinavian Journal of Statistics. Theory and Applications, 20(4):329–341, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616288.

Stute:1993:MSS

[SW93b] Winfried Stute and Jane-Ling Wang. Multi-sample U-statistics for censored data. Scandinavian Journal of Statistics. Theory and Applications, 20(4): 369–374, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616291.

Schick:2004:RCO

[SW04] Anton Schick and Wolfgang Wefelmeyer. Root n consistent and optimal density estimators for moving average processes. Scandinavian Journal of Statistics. Theory and Applications, 31(1):63–78, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616812.

Sun:2005:STV

[SW05] Yanqing Sun and Hulin Wu. Semiparametric timevarying coefficients regression model for longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 32(1):21– 47, March 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616859.

Song:2018:CLT

[SW18] Yuping Song and Hanchao Wang. Central limit theorems of local polynomial threshold estimator for diffusion processes with jumps. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 644–681, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Song:2019:GNC

[SW19] Xiao Song and Ching-Yun Wang. GMM nonparametric correction methods for logistic regression with error-contaminated covariates and partially observed instrumental variables. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 898–919, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Swensen:1983:NCD

[Swe83] Anders Rygh Swensen. A note on convergence of distributions of conditional moments. Scandinavian Journal of Statistics. Theory and Applications, 10(1): 41–44, March 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615899.

Swensen:1988:ECP

[Swe88] Anders Rygh Swensen. Estimating change in a proportion by combining measurements from a true and a fallible classifier. Scandinavian Journal of Statistics. Theory and Applications, 15(2):139–145, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616094.

Sun:2006:EAB

[SWS06] Liuquan Sun, Lianming Wang, and Jianguo Sun. Estimation of the association for bivariate intervalcensored failure time data.

Scandinavian Journal of Statistics. Theory and Applications, 33(4):637–649, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616950.

Singpurwalla:1993:MDI

[SY93] Nozer D. Singpurwalla and Mark A. Youngren. Multivariate distributions induced by dynamic environments. Scandinavian Journal of Statistics. Theory and Applications, 20(3): 251–261, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616280.

Schick:2000:CGM

[SY00] Anton Schick and Qiqing Yu. Consistency of the GMLE with mixed case interval-censored data. Scandinavian Journal of Statistics. Theory and Applications, 27(1):45–55, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616589.

Srinivasan:1995:CLT

[SZ95] C. Srinivasan and Mai Zhou. A central limit theorem for weighted and integrated martingales. Scandinavian Journal of Statistics. Theory and Applications, 22(4):493–504, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616376.

Stute:2002:MCG

[SZ02] Winfried Stute and Li-Xing Zhu. Model checks for generalized linear models. Scandinavian Journal of Statistics. Theory and Applications, 29(3):535–545, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616730.

Scheike:2007:DMR

[SZ07] Thomas H. Scheike and Mei-Jie Zhang. Direct modelling of regression effects for transition probabilities in multistate models. Scandinavian Journal of Statistics. Theory and Applications, 34(1):17– 32, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548536.

Smaga:2020:LHT

[SZ20] Łukasz Smaga and Jin-Ting Zhang. Linear hypothesis testing for weighted functional data with applications. Scandinavian Journal of Statistics. Theory and Applications, 47(2): 493–515, June 2020. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Szekli:2005:DT

[Sze05] Ryszard Szekli. Discussion of “Teletraffic”. Scandinavian Journal of Statistics. Theory and Applications, 32(2):219–221, June 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Nor05].

Scheike:2002:AMC

[SZS02] Thomas H. Scheike, Mei-Jie Zhang, and Mei-Jie Shang. An additive– multiplicative Cox–Aalen regression model. Scandinavian Journal of Statistics. Theory and Applications, 29(1):75–88, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616700.

Sun:2005:GLR

[SZZ05] Jianguo Sun, Qiang Zhao, and Xingqiu Zhao. Generalized log-rank tests for interval-censored failure time data. Scandinavian Journal of Statistics. Theory and Applications, 32(1): 49–57, March 2005. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616860.

Takeishi:2023:APM

[Tak23] Shota Takeishi. Asymptotic properties of the maximum smoothed partial likelihood estimator in the changeplane Cox model. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1503–1531, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tang:1994:UEF

[Tan94] Boxin Tang. Unbiased estimation for finite population parameters. Scandinavian Journal of Statistics. Theory and Applications, 21(1):91–95, March 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616301.

Tanaka:2009:SCM

[Tan09] Kentaro Tanaka. Strong consistency of the maximum likelihood estimator for finite mixtures of location–scale distributions when penalty is imposed on the ratios of the scale parameters. Scandinavian Journal of Statistics. Theory and Applications, 36 (1):171–184, March 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000314.

Taqqu:2002:MED

[Taq02] Murad S. Taqqu. The modelling of Ethernet data and of signals that are heavytailed with infinite variance*. Scandinavian Journal of Statistics. Theory and Applications, 29(2): 273–295, June 2002. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616715.

Tjelmeland:1998:MRF

[TB98] Håkon Tjelmeland and Julian Besag. Markov random fields with higher-order interactions. Scandinavian Journal of Statistics. Theory and Applications, 25(3): 415–433, September 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616513.

Tardivel:2022:SRL

[TB22] Patrick J. C. Tardivel and Małgorzata Bogdan. On the sign recovery by least absolute shrinkage and selection operator, thresholded least absolute shrinkage and selection operator, and thresholded basis pursuit denoising. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1636–1668, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tsai:2005:TAS

[TC05] Henghsiu Tsai and K. S. Chan. Temporal aggregation of stationary and nonstationary continuous-time processes. Scandinavian Journal of Statistics. Theory and Applications, 32(4): 583–597, December 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616899.

Tong:1995:PON

[TCC^{+}95] Howell Tong, K. S. Chan, D. R. Cox, Colleen D. Cutler, D. Guégan, Jens Ledet Jensen, Søren Johansen, A. J. Lawrance, Blake Lebaron, T. Ozaki, Douglas W. Nychka, Stephen Ellner, Barbara A. Bailey, A. R. Gallant, L. R. Smith, R. L. Smith, and Rodney C. L. Wolff. A personal overview of nonlinear time series analysis from a chaos perspective [with discussion and rejoinder]. Scandinavian Journal of Statistics. Theory and Applications, 22(4): 399–445, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616372.

Tian:2023:LSS

[TCK^{+}23] Jinjin Tian, Xu Chen, Eugene Katsevich, Jelle Goeman, and Aaditya Ramdas. Large-scale simultaneous inference under dependence. Scandinavian Journal of Statistics. Theory and Applications, 50(2): 750–796, June 2023. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Torabi:2009:EBE

[TDR09] Mahmoud Torabi, Gauri S.

Datta, and J. N. K. Rao. Empirical Bayes estimation of small area means under a nested error linear regression model with measurement errors in the covariates. Scandinavian Journal of Statistics. Theory and Applications, 36(2): 355–369, June 2009. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000325.

Terasvirta:1977:EFD

[Ter77a] Timo Teräsvirta. Effect of feedback on the distribution of the portmanteau test statistic. Scandinavian Journal of Statistics. Theory and Applications, 4(1):20–24, March 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615643.

Terasvirta:1977:ISD

[Ter77b] Timo Teräsvirta. The invertibility of sums of discrete MA and ARMA processes. Scandinavian Journal of Statistics. Theory and Applications, 4(4):165– 170, September 1977. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615673.

Terasvirta:1981:SRI

[Ter81] Timo Teräsvirta. Some results on improving the least squares estimation of linear models by mixed estimation. Scandinavian Journal of Statistics. Theory and Applications, 8 (1):33–38, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615810.

Terasvirta:1983:RSL

[Ter83]  Timo Teräsvirta.  Re-

stricted superiority of linear homogeneous estimators over ordinary least squares. Scandinavian Journal of Statistics. Theory and Applications, 10(1): 27–33, March 1983. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615897.

Terada:2014:SCR

[Ter14] Yoshikazu Terada. Strong consistency of reduced Kmeans clustering. Scandinavian Journal of Statistics. Theory and Applications, 41(4):913–931, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586813.

Talamakrouni:2015:GCR

[TEV15] Majda Talamakrouni, Anouar El Ghouch, and Ingrid Van Keilegom. Guided censored regression. Scandinavian Journal of Statistics. Theory and Applications, 42 (1):214–233, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Trippa:2012:CNR

[TF12] Lorenzo Trippa and Stefano Favaro. A class of normalized random measures with an exact predictive sampling scheme. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 444–460, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41679805.

Tecuapetla-Gomez:2017:AER

[TGM17] Inder Tecuapetla-Gómez and Axel Munk. Autocovariance estimation in regression with a discontinuous signal and m-dependent errors: a difference-based approach. Scandinavian Journal of Statistics. Theory and Applications, 44 (2):346–368, June 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tjelmeland:2001:MJP

[TH01] Håkon Tjelmeland and Bjørn Kåre Hegstad. Mode jumping proposals in MCMC. Scandinavian Journal of Statistics. Theory and Applications, 28(1):205–223, March 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616652.

Todem:2018:QSS

[THF18] David Todem, Wei-Wen Hsu, and Jason P. Fine. A quasi-score statistic for homogeneity testing against covariate-varying heterogeneity. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 465–481, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Thorburn:1977:ANJ

[Tho77] Daniel Thorburn. On the asymptotic normality of the jackknife. Scandinavian Journal of Statistics. Theory and Applications, 4(3): 113–118, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615661.

Thomsen:1981:UMC

[Tho81] Ib Thomsen. The use of Markov chain models in sampling from finite populations. Scandinavian Journal of Statistics. Theory and Applications, 8(1):1–9, September 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615806.

Thorburn:1983:FAT

[Tho83] Daniel Thorburn. Forecasting aggregate time series using empirical transition matrices. Scandinavian Journal of Statistics. Theory and Applications, 10(1):35–39, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615898.

Thorisson:1995:CMP

[Tho95] Hermann Thorisson. Coupling methods in probability theory. Scandinavian Journal of Statistics. Theory and Applications, 22(2):159–182, June 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616351.

Tong:2009:VSP

[THSS09] Xingwei Tong, Xin He, Liuquan Sun, and Jianguo Sun. Variable selection for panel count data via nonconcave penalized estimating function. Scandinavian Journal of Statistics. Theory and Applications, 36(4): 620–635, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000343.

Thulin:2014:CAC

[Thu14] Måns Thulin. Coverageadjusted confidence intervals for a binomial proportion. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 291–300, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586653.

Thyregod:1975:BSS

[Thy75] P. Thyregod. Bayesian single sampling plans for life-testing with truncation of the number of failures. Scandinavian Journal of Statistics. Theory and Applications, 2(2):61– 70, March 1975. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615578.

Torgersen:1976:CSE

[TJL^{+}76] Erik N. Torgersen, Søren Johansen, Bo Lindqvist, J. Hilden, and O. Barndorff-Nielsen. Comparison of statistical experiments [with discussion and reply]. Scandinavian Journal of Statistics. Theory and Applications, 3(4):186– 208, March 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615635.

Tjostheim:1994:NLT

[Tjø94] Dag Tjøstheim. Non-linear time series: a selective review. Scandinavian Journal of Statistics. Theory and Applications, 21(2):97– 130, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616304.

Tjostheim:1995:E

[Tjø95] Dag Tjøstheim. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 22(1): 1, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616337.

Tjostheim:1996:E

[Tjø96] Dag Tjøstheim. Editorial. Scandinavian Journal of Statistics. Theory and Applications, 23(4):385, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616415.

Tjur:1982:CBR

[Tju82] Tue Tjur. A connection between Rasch’s item analysis model and a multiplicative Poisson model. Scandinavian Journal of Statistics. Theory and Applications, 9 (1):23–30, June 1982. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615850.

Tang:2023:PCS

[TKLM23] Yifu Tang, Claudia Kirch, Jeong Eun Lee, and Renate Meyer. Posterior consistency for the spectral density of non-Gaussian stationary time series. Scandinavian Journal of Statistics. Theory and Applications, 50(3):1152–1182, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tonaki:2023:ECP

[TKU23a] Yozo Tonaki, Yusuke Kaino, and Masayuki Uchida. Estimation for change point of discretely observed ergodic diffusion processes. Scandinavian Journal of Statistics. Theory and Applications, 50(1):142–183, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tonaki:2023:PEL

[TKU23b] Yozo Tonaki, Yusuke Kaino, and Masayuki Uchida. Parameter estimation for linear parabolic SPDEs in two space dimensions based on high frequency data. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1568–1589, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Thoresen:2003:URL

[TL03] Magne Thoresen and Petter Laake. The use of replicates in logistic measurement error modelling. Scandinavian Journal of Statistics. Theory and Applications, 30(3):625–636, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616789.

Terasvirta:1986:MSC

[TM86] Timo Teräsvirta and Ilkka Mellin. Model selection criteria and model selection tests in regression models. Scandinavian Journal of Statistics. Theory and Applications, 13(3):159–171, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616022.

Tocquet:2001:LBI

[Toc01] A. S. Tocquet. Likelihood based inference in non-linear regression models using the p * and r * approach. Scandinavian Journal of Statistics. Theory and Applications, 28(3): 429–443, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616670.

Torgersen:1988:BCR

[Tor88] Erik Torgersen. On Bahadur’s converse of the Rao–Blackwell Theorem. Extension to majorized experiments. Scandinavian Journal of Statistics. Theory and Applications, 15(4): 273–280, December 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616111.

Telschow:2021:FIR

[TPH21] Fabian J. E. Telschow, Michael R. Pierrynowski, and Stephan F. Huckemann. Functional inference on rotational curves under sample-specific group actions and identification of human gait. Scandinavian Journal of Statistics. Theory and Applications, 48(4): 1256–1276, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tran:2011:LRC

[Tra11] Minh-Ngoc Tran. The loss rank criterion for variable selection in linear regression analysis. Scandinavian Journal of Statistics. Theory and Applications, 38(3):

466–479, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015575.

Trenkler:1983:BLE

[Tre83]  Götz Trenkler.  Biased

linear estimators and estimable functions. Scandinavian Journal of Statistics. Theory and Applications, 10(1):53–55, March 1983. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615902.

Tribouley:2003:WMC

[Tri03]  Karine Tribouley.  A

wavelet method for confidence intervals in the density model. Scandinavian Journal of Statistics. Theory and Applications, 30(4): 739–756, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616799.

Tsai:2015:ISL

[TRL15] Henghsiu Tsai, Heiko Rachinger, and Edward M. H. Lin. Inference of seasonal long-memory time series with measurement error. Scandinavian Journal of Statistics. Theory and Applications, 42 (1):137–154, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Thyregod:1991:HNI

[TS91] Poul Thyregod and Henrik Spliid. A hypothesis of no interaction in factorial experiments with a binary response. Scandinavian Journal of Statistics. Theory and Applications, 18(3): 197–209, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616203.

Tjur:1991:AVD

[TSH91] Tue Tjur, Rolf Sundberg, and Inge Helland. Analysis of variance and design of experiments [with discussion and rejoinder]. Scandinavian Journal of Statistics. Theory and Applications, 18(4):273–322, March 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616213.

Tian:2017:ASI

[TT17] Xiaoying Tian and Jonathan Taylor. Asymptotics of selective inference. Scandinavian Journal of Statistics. Theory and Applications, 44(2):480–499, June 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Taraldsen:2022:IPI

[TTL22] Gunnar Taraldsen, Jarle Tufto, and Bo H. Lindqvist. Improper priors and improper posteriors. Scandinavian Journal of Statistics. Theory and Applications, 49(3):969–991, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tang:2018:ERE

[TTZZ18] Man-Lai Tang, Nian-Sheng Tang, Pu-Ying Zhao, and Hongtu Zhu. Efficient robust estimation for linear models with missing response at random. Scandinavian Journal of Statistics. Theory and Applications, 45(2):366–381, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tsybakov:1996:RCE

[TvdM96] A. B. Tsybakov and E. C. van der Meulen. Rootn consistent estimators of entropy for densities with unbounded support. Scandinavian Journal of Statistics. Theory and Applications, 23(1):75– 83, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616387.

Togersen:2004:SMD

[TW04] Frede Aakmann Tøgersen and Rasmus Waagepetersen. Statistical modelling and deconvolution of yield meter data. Scandinavian Journal of Statistics. Theory and Applications, 31(2): 247–264, June 2004. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616827.

Tang:2018:CES

[TWL18] Yanlin Tang, Huixia Judy Wang, and Hua Liang. Composite estimation for single-index models with responses subject to detection limits. Scandinavian Journal of Statistics. Theory and Applications, 45(3): 444–464, September 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Tsai:1995:APN

[TZ95] Wei-Yann Tsai and Cun-Hui Zhang. Asymptotic properties of nonparametric maximum likelihood estimator for intervaltruncated data. Scandinavian Journal of Statistics. Theory and Applications, 22(3):361–370, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616365.

Uchida:2004:EDO

[Uch04] Masayuki Uchida. Estimation for discretely observed small diffusions based on approximate martingale estimating functions. Scandinavian Journal of Statistics. Theory and Applications, 31(4):553–566, December 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616850.

Uehara:2023:SIS

[ULK23] Masatoshi Uehara, Danhyang Lee, and Jae-Kwang Kim. Statistical inference with semiparametric nonignorable nonresponse models. Scandinavian Journal of Statistics. Theory and Applications, 50(4): 1795–1817, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Utazi:2017:BSC

[Uta17] Chigozie E. Utazi. Bayesian single changepoint estimation in a parameter-driven model. Scandinavian Journal of Statistics. Theory and Applications, 44(3): 765–779, September 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Vaeth:1979:NBO

[Væt79] Michael Væth. A note on the behaviour of occurrence/exposure rates when the survival distribution is not exponential. Scandinavian Journal of Statistics. Theory and Applications, 6 (2):77–80, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615735.

Vaillant:1991:NBD

[Vai91]  Jean Vaillant.  Nega-

tive binomial distributions of individuals and spatiotemporal Cox processes. Scandinavian Journal of Statistics. Theory and Applications, 18(3):235–248, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616206.

VanDerLaan:1998:TIL

[Van98] Mark J. Van Der Laan. The two-interval line-segment problem. Scandinavian Journal of Statistics. Theory and Applications, 25 (1):163–186, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616492.

VanZanten:2001:NCE

[Van01] J. H. Van Zanten. A note on consistent estimation of multivariate parameters in ergodic diffusion models. Scandinavian Journal of Statistics. Theory and Applications, 28(4): 617–623, December 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616685.

VanHouwelingen:2007:DPL

[Van07a] Hans C. Van Houwelingen. Dynamic prediction by landmarking in event history analysis. Scandinavian Journal of Statistics. Theory and Applications, 34(1):70–85, March 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548539.

Vansteelandt:2007:CPE

[Van07b] Stijn Vansteelandt. On confounding, prediction and efficiency in the analysis of longitudinal and crosssectional clustered data. Scandinavian Journal of Statistics. Theory and Applications, 34(3):478–498, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41548565.

Vanderweele:2011:CDM

[Van11] Tyler J. Vanderweele. Controlled direct and mediated effects: Definition, identification and bounds. Scandinavian Journal of Statistics. Theory and Applications, 38(3):551–563, September 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015580.

VanLieshout:2013:DMR

[Van13] M. N. M. Van Lieshout.

Discrete multicolour random mosaics with an application to network extraction. Scandinavian Journal of Statistics. Theory and Applications, 40(4): 734–751, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586697.

vandeGeer:2014:WDR

[Van14] Sara Van De Geer. Weakly decomposable regularization penalties and structured sparsity. Scandinavian Journal of Statistics. Theory and Applications, 41(1):72–86, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586731.

Vartia:1976:ILC

[Var76]  Yrjö O. Vartia.  Ideal

log-change index numbers. Scandinavian Journal of Statistics. Theory and Applications, 3(3):121–126, September 1976. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615624.

Vardeman:1979:NAS

[Var79] Stephen B. Vardeman. A note on the applicability of sequence compound decision schemes. Scandinavian Journal of Statistics. Theory and Applications, 6 (2):86–88, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615737.

VanLieshout:1999:IDB

[VB99] M. N. M. Van Lieshout and A. J. Baddeley. Indices of dependence between types in multivariate point patterns. Scandinavian Journal of Statistics. Theory and Applications, 26(4): 511–532, December 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616576.

VanDerLaan:1997:SDC

[VBJ97] Mark J. Van Der Laan, Peter J. Bickel, and Nicholas P. Jewell. Singly and doubly censored current status data: Estimation, asymptotics and regression. Scandinavian Journal of Statistics. Theory and Applications, 24(3):289–307, September 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616456.

Vansteelandt:2018:IRE

[VD18] Stijn Vansteelandt and Vanessa Didelez. Improving the robustness and efficiency of covariate-adjusted linear instrumental variable estimators. Scandinavian Journal of Statistics. Theory and Applications, 45(4): 941–961, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

vanderLaan:1996:ENL

[vdL96] Mark J. van der Laan. Efficiency of the NPMLE in the line-segment problem. Scandinavian Journal of Statistics. Theory and Applications, 23(4):527–550, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616424.

vanderVaart:1991:EHD

[vdV91] Aad van der Vaart. Efficiency and Hadamard differentiability. Scandinavian Journal of Statistics. Theory and Applications, 18 (1):63–75, September 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616189.

vanderVaart:1994:WCS

[vdV94] Aad van der Vaart. Weak convergence of smoothed empirical processes. Scandinavian Journal of Statistics. Theory and Applications, 21(4):501–504, December 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616334.

vandeWiel:2019:LLE

[vdWBM19] Mark A. van de Wiel, Dennis E. Te Beest, and Magnus M. Münch. Learning from a lot: Empirical Bayes for highdimensional model-based prediction. Scandinavian Journal of Statistics. Theory and Applications, 46(1): 2–25, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

vanEs:1992:EFR

[vE92] Bert van Es. Estimating functionals related to a density by a class of statistics based on spacings. Scandinavian Journal of Statistics. Theory and Applications, 19(1):61–72, December 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616225.

Vellaisamy:2012:ACD

[Vel12] P. Vellaisamy. Average collapsibility of distribution dependence and quantile regression coefficients. Scandinavian Journal of Statistics. Theory and Applications, 39(1):153–165, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41411162.

Vetter:2012:ECC

[Vet12] Mathias Vetter. Estimation of correlation for continuous semimartingales. Scandinavian Journal of Statistics. Theory and Applications, 39(4):757–771, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        23357223.

vanEs:1996:CME

[vEvZ96] A. J. van Es and M. C. A. van Zuijlen. Convex minorant estimators of distributions in non-parametric deconvolution problems. Scandinavian Journal of Statistics. Theory and Applications, 23(1):85–104, March 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616388.

vanHouwelingen:1980:RCR

[vH80] Hans van Houwelingen. Risk convergence rates in empirical Bayes classification. Scandinavian Journal of Statistics. Theory and Applications, 7(2):99– 101, March 1980. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4615781.

Vihola:2020:IST

[VHF20] Matti Vihola, Jouni Helske, and Jordan Franks. Importance sampling type estimators based on approximate marginal Markov chain Monte Carlo. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1339–1376, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

VanWesel:2011:CPC

[VHK11] Floryt Van Wesel, Herbert Hoijtink, and Irene Klugkist. Choosing priors for constrained analysis of variance: Methods based on training data. Scandinavian Journal of Statistics. Theory and Applications, 38(4): 666–690, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41411144.

vanHouwelingen:1985:CMS

[vHV85] Hans C. van Houwelingen and Albert Verbeek. On the construction of monotone symmetric decision rules for distributions with monotone likelihood ratio. Scandinavian Journal of Statistics. Theory and Applications, 12(1):73– 81, June 1985. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615975.

Vidoni:2001:PDS

[Vid01] Paolo Vidoni. Proper dispersion state space models for stochastic volatility. Scandinavian Journal of Statistics. Theory and Applications, 28(2):

271–281, June 2001. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616658.

Vidoni:2009:IPI

[Vid09] Paolo Vidoni. Improved prediction intervals and distribution functions. Scandinavian Journal of Statistics. Theory and Applications, 36(4):735–748, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000349.

Vidoni:2021:BMM

[Vid21] Paolo Vidoni. Boosting multiplicative model combination. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 761–789, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Vieu:1999:MKP

[Vie99]  Philippe Vieu.  Multi-

ple kernel procedure: an asymptotic support. Scandinavian Journal of Statistics. Theory and Applications, 26(1):61–72, March 1999. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616541.

VanDerLaan:2001:NDC

[VJ01] Mark J. Van Der Laan and Nicholas P. Jewell. The NPMLE for doubly censored current status data. Scandinavian Journal of Statistics. Theory and Applications, 28(3):537–547, September 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616677.

Vannman:1995:SCI

[VK95] Kerstin Vännman and Samuel Kotz. A superstructure of capability indices: Distributional properties and implications. Scandinavian Journal of Statistics. Theory and Applications, 22(4):477–491, December 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616375.

Vexler:2014:CCV

[VKY^{+}14] Albert Vexler, Young Min Kim, Jihnhee Yu, Nicole A. Lazar, and Alan D. Hutson. Computing critical values of exact tests by incorporating Monte Carlo simulations combined with statistical tables. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1013–1030, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586818.

vanLieshout:2018:NID

[vL18] M. N. M. van Lieshout. Nonparametric indices of dependence between components for inhomogeneous multivariate random measures and marked sets. Scandinavian Journal of Statistics. Theory and Applications, 45(4):985–1015, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Virta:2021:FTJ

[VLIN21] Joni Virta, Niko Lietzén, Pauliina Ilmonen, and Klaus Nordhausen. Fast tensorial JADE. Scandinavian Journal of Statistics. Theory and Applications, 48(1):164–187, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

vanLieshout:2023:SEA

[vLM23] Maria N. M. van Lieshout and Robin L. Markwitz. State estimation for aoristic models. Scandinavian Journal of Statistics. Theory and Applications, 50 (3):1068–1089, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

VonSachs:2000:NPC

[VM00] Rainer Von Sachs and Brenda Macgibbon. Nonparametric curve estimation by wavelet thresholding with locally stationary errors. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 475–499, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616618.

Veronese:2015:FCD

[VM15] Piero Veronese and Eugenio Melilli. Fiducial and confidence distributions for real exponential families. Scandinavian Journal of Statistics. Theory and Applications, 42(2):471–484, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Vihrs:2022:ABI

[VMG22] Ninna Vihrs, Jesper Møller, and Alan E. Gelfand. Approximate Bayesian inference for a spatial point process model exhibiting regularity and random aggregation. Scandinavian Journal of Statistics. Theory and Applications, 49 (1):185–210, March 2022.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Veraverbeke:2011:ECC

[VOG11] Noël Veraverbeke, Marek Omelka, and Irène Gijbels. Estimation of a conditional copula and association measures. Scandinavian Journal of Statistics. Theory and Applications, 38(4):766–780, December 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41411149.

Vonta:1996:EEN

[Von96] Filia Vonta. Efficient estimation in a non-proportional hazards model in survival analysis. Scandinavian Journal of Statistics. Theory and Applications, 23(1): 49–61, March 1996. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616385.

Vos:2023:CDV

[Vos23] Paul W. Vos. Comments on Divergence vs. Decision P -values. Scandinavian Journal of Statistics. Theory and Applications, 50(3): 920–922, September 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [Gre23a].

vanPul:1992:APC

[vP92] Mark C. van Pul. Asymptotic properties of a class of statistical models in software reliability. Scandinavian Journal of Statistics. Theory and Applications, 19(3):235–253, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616242.

vonRosen:1988:MIW

[vR88] Dietrich von Rosen. Moments for the inverted Wishart distribution. Scandinavian Journal of Statistics. Theory and Applications, 15(2):97–109, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616090.

vonRosen:1994:PLM

[vR94] Dietrich von Rosen. PLS, linear models and invariant spaces. Scandinavian Journal of Statistics. Theory and Applications, 21(2): 179–186, June 1994. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616309.

vonRosen:1995:IOM

[vR95] Dietrich von Rosen. Influential observations in multivariate linear models. Scandinavian Journal of Statistics. Theory and Applications, 22(2): 207–222, June 1995. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616354.

Varadhan:2008:SGC

[VR08] Ravi Varadhan and Christophe Roland. Simple and globally convergent methods for accelerating the convergence of any EM algorithm. Scandinavian Journal of Statistics. Theory and Applications, 35(2): 335–353, June 2008. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548597.

VanLieshout:2007:PSL

[VS07] M. N. M. Van Lieshout and T. Schreiber. Perfect simulation for length-interacting polygonal Markov fields in the plane. Scandinavian Journal of Statistics. Theory and Applications, 34(3): 615–625, September 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41548572.

Vogt:2021:CSE

[VS21] Michael Vogt and Matthias Schmid. Clustering with statistical error control. Scandinavian Journal of Statistics. Theory and Applications, 48(3):729–760, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

VanEs:2005:ANK

[VU05] Bert Van Es and Hae-Won Uh. Asymptotic normality of kernel-type deconvolution estimators. Scandinavian Journal of Statistics. Theory and Applications, 32(3):467–483, September 2005. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616890.

VanOurs:2024:CPA

[VV24] Jan C. Van Ours and Ben Vollaard. Commentary on “Pitfalls of amateur regression: The Dutch New Herring controversies”. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1388–1389, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [GG23].

Voutilainen:2022:VVG

[VVI^{+}22] Marko Voutilainen, Lauri Viitasaari, Pauliina Ilmonen, Soledad Torres, and Ciprian Tudor. Vectorvalued generalized Ornstein– Uhlenbeck processes: Properties and parameter estimation. Scandinavian Journal of Statistics. Theory and Applications, 49(3): 992–1022, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Vexler:2009:ORC

[VW09] Albert Vexler and Chengqing Wu. An optimal retrospective change point detection policy. Scandinavian Journal of Statistics. Theory and Applications, 36(3): 542–558, September 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000337.

Villa:2015:OBC

[VW15] Cristiano Villa and Stephen Walker. An objective Bayesian criterion to determine model prior probabilities. Scandinavian Journal of Statistics. Theory and Applications, 42(4): 947–966, December 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Vogt:2019:ENA

[VW19] Michael Vogt and Christopher Walsh. Estimating nonlinear additive models with nonstationarities and correlated errors. Scandinavian Journal of Statistics. Theory and Applications, 46(1):160–199, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

vanZanten:2003:EPE

[vZ03] Harry van Zanten. On empirical processes for ergodic diffusions and rates of convergence of M-estimators. Scandinavian Journal of Statistics. Theory and Applications, 30(3):443–458, September 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616776.

Waagepetersen:2006:SBG

[Waa06] Rasmus Waagepetersen. A simulation-based goodnessof-fit test for random effects in generalized linear mixed models. Scandinavian Journal of Statistics. Theory and Applications, 33(4): 721–731, December 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616954.

Walley:1997:BDM

[Wal97] Peter Walley. A bounded derivative model for prior ignorance about a realvalued parameter. Scandinavian Journal of Statistics. Theory and Applications, 24(4):463–483, December 1997. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616470.

Walker:2000:NID

[Wal00] Stephen G. Walker. A note on the innovation distribution of a gamma distributed autoregressive process. Scandinavian Journal of Statistics. Theory and Applications, 27(3): 575–576, September 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616624.

Walker:2024:MPD

[Wal24] Stephen G. Walker. Martingale posterior distributions for cumulative hazard functions. Scandinavian Journal of Statistics. Theory and Applications, 51(3): 936–955, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:1986:ODP

[Wan86] Yuchung J. Wang. Orderdependent parameterization of multinomial distributions. Scandinavian Journal of Statistics. Theory and Applications, 13(3): 199–205, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616025.

Wang:1987:EIN

[Wan87] Jane-Ling Wang. Estimating IFRA and NBU survival curves based on censored data. Scandinavian Journal of Statistics. Theory and Applications, 14(3): 199–210, December 1987. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616063.

Wand:1990:ERC

[Wan90] Matthew P. Wand. On exact L_{1} rates of convergence in non-parametric kernel regression. Scandinavian Journal of Statistics. Theory and Applications, 17 (3):251–256, March 1990. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616172.

Wang:1995:ECD

[Wan95]  Jane-Ling Wang.  M-

estimators for censored data: Strong consistency. Scandinavian Journal of Statistics. Theory and Applications, 22(2):197–205, June 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616353.

Wang:1999:APE

[Wan99] Jane-Ling Wang. Asymptotic properties of Mestimators based on estimating equations and censored data. Scandinavian Journal of Statistics. Theory and Applications, 26(2): 297–318, June 1999. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616557.

Wang:2000:ARE

[Wan00] Jinde Wang. Approximate representation of estimators in constrained regression problems. Scandinavian Journal of Statistics. Theory and Applications, 27(1):21–33, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616587.

Wang:2006:QAL

[Wan06] Jinfang Wang. Quadratic artificial likelihood functions using estimating functions. Scandinavian Journal of Statistics. Theory and Applications, 33(2): 379–390, June 2006. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616930.

Wang:2008:NPM

[Wan08]  C. Y. Wang.  Non-

parametric maximum likelihood estimation for Cox regression with subjectspecific measurement error. Scandinavian Journal of Statistics. Theory and Applications, 35(4):613–628, December 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000292.

Wallin:2015:GMU

[WB15] Jonas Wallin and David Bolin. Geostatistical modelling using non-Gaussian Matérn fields. Scandinavian Journal of Statistics. Theory and Applications, 42(3): 872–890, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wen:2012:CSA

[WC12] Chi-Chung Wen and Yi-Hau Chen. Conditional score approach to errorsin-variable current status data under the proportional odds model. Scandinavian Journal of Statistics. Theory and Applications, 39(4): 635–644, December 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23357215.

Wang:2020:CBE

[WC20] Shao-Hsuan Wang and Chin-Tsang Chiang. Concordancebased estimation approaches for the optimal sufficient dimension reduction score. Scandinavian Journal of Statistics. Theory and Applications, 47(3):662–689, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Weng:2021:BAL

[WC21] Ruby Chiu-Hsing Weng and D. Stephen Coad. Bias approximations for likelihood-based estimators. Scandinavian Journal of Statistics. Theory and Applications, 48(4): 1474–1497, December 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2018:ALT

[WCJ18] Antai Wang, Krishnendu Chandra, and Xieyang Jia. The analysis of left truncated bivariate data using frailty models. Scandinavian Journal of Statistics. Theory and Applications, 45(4):847–860, December 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2023:DNN

[WCSftADNI23] Shuoyang Wang, Guanqun Cao, Zuofeng Shang, and for the Alzheimer’s Disease Neuroimaging Initiative. Deep neural network classifier for multidimensional functional data. Scandinavian Journal of Statistics. Theory and Applications, 50(4):1667–1686, December 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2015:IDC

[WCXS15] Antai Wang, Krishnendu Chandra, Ruihua Xu, and Junfeng Sun. The identifiability of dependent competing risks models induced by bivariate frailty models. Scandinavian Journal of Statistics. Theory and Applications, 42(2): 427–437, June 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2022:ESE

[WCY22] Jinyang Wang, Piao Chen, and Zhisheng Ye. Efficient semiparametric estimation of time-censored intensityreduction models for repairable systems. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1860–1888, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Walker:1998:FBN

[WD98] Stephen Walker and Paul Damien. A full Bayesian non-parametric analysis involving a neutral to the right process. Scandinavian Journal of Statistics. Theory and Applications, 25(4): 669–680, December 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616531.

Weber:1981:IDS

[Web81] N. C. Weber. Incomplete degenerate U-statistics. Scandinavian Journal of Statistics. Theory and Applications, 8(2):120–123, June 1981. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615821.

Weits:1993:SOO

[Wei93]  Ello Weits.  The sec-

ond order optimality of a smoothed Kaplan–Meier estimator. Scandinavian Journal of Statistics. Theory and Applications, 20 (2):111–132, March 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616268.

Winkler:1979:SEP

[WF79] W. Winkler and J. Franz.

Sequential estimation problems for the exponential class of processes with independent increments. Scandinavian Journal of Statistics. Theory and Applications, 6(3):129–139, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615748.

Wang:2016:FPL

[WFC16] Guochang Wang, Xiang-Nan Feng, and Min Chen. Functional partial linear single-index model. Scandinavian Journal of Statistics. Theory and Applications, 43(1):261–274, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:1996:ORE

[WG96] Kongming Wang and Theo Gasser. Optimal rate for estimating local bandwidth in kernel estimators of regression functions. Scandinavian Journal of Statistics. Theory and Applications, 23(3):303–312, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616406.

Wang:2019:ETE

[WGT19] Jun Wang, Wei Gao, and Man-Lai Tang. Estimation of treatment effects for heterogeneous matchedpairs data with probit models. Scandinavian Journal of Statistics. Theory and Applications, 46(2): 575–594, June 2019. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wei:1998:GLA

[WHF98] Bo-Cheng Wei, Yue-Qing Hu, and Wing-Kam Fung. Generalized leverage and its applications. Scandinavian Journal of Statistics. Theory and Applications, 25(1):25–37, March 1998. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616482.

Whitmore:1986:NGM

[Whi86] G. A. Whitmore. Normalgamma mixtures of inverse Gaussian distributions. Scandinavian Journal of Statistics. Theory and Applications, 13(3): 211–220, September 1986. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616027.

Wen:2022:MRE

[WHR22] Lan Wen, Miguel A.

Hernán, and James M. Robins. Multiply robust estimators of causal effects for survival outcomes. Scandinavian Journal of Statistics. Theory and Applications, 49(3):1304–1328, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2020:SIM

[WHZ20] Wu Wang, Xuming He, and Zhongyi Zhu. Statistical inference for multiple changepoint models. Scandinavian Journal of Statistics. Theory and Applications, 47(4): 1149–1170, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wijers:1995:CNP

[Wij95] Bart J. Wijers. Consistent non-parametric estimation for a one-dimensional line segment process observed in an interval. Scandinavian Journal of Statistics. Theory and Applications, 22(3): 335–360, September 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616364.

Willassen:1977:ISD

[Wil77]  Yngve Willassen.  On

identifiability of stochastic difference equations with errors-in-variables in relation to identifiability of the classical errors-in-variables (EIV) models. Scandinavian Journal of Statistics. Theory and Applications, 4(3):119–124, December 1977. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4615662.

Willassen:1979:ESR

[Wil79]  Yngve Willassen.  Ex-

tension of some results by reiersøl to multivariate models. Scandinavian Journal of Statistics. Theory and Applications, 6 (2):89–91, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615738.

Wintenberger:2013:CIS

[Win13] Olivier Wintenberger. Continuous invertibility and stable QML estimation of the EGARCH(1,1) model. Scandinavian Journal of Statistics. Theory and Applications, 40(4):846–867, December 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586703.

Wu:2004:BRV

[WL04] Hulin Wu and Hua Liang.

Backfitting random varyingcoefficient models with time-dependent smoothing covariates. Scandinavian Journal of Statistics. Theory and Applications, 31(1): 3–19, March 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616808.

Wang:2018:HMM

[WL18] Qihua Wang and Yongjin Li. How to make modelfree feature screening approaches for full data applicable to the case of missing response? Scandinavian Journal of Statistics. Theory and Applications, 45(2):324–346, June 2018.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2015:FMA

[WLS15] Haiying Wang, Yang Li, and Jianguo Sun. Focused and model average estimation for regression analysis of panel count data. Scandinavian Journal of Statistics. Theory and Applications, 42(3):732–745, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wu:2015:ELC

[WLT15] Tong Tong Wu, Gang Li, and Chengyong Tang. Empirical likelihood for censored linear regression and variable selection. Scandinavian Journal of Statistics. Theory and Applications, 42(3):798–812, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2019:REE

[WLX19] Li Wang, Zhi Liu, and Xiaochao Xia. Rate efficient estimation of realized Laplace transform of volatility with microstructure noise. Scandinavian Journal of Statistics. Theory and Applications, 46(3): 920–953, September 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wolfer:2024:EID

[Wol24] Geoffrey Wolfer. Empirical and instance-dependent estimation of Markov chain and mixing time. Scandinavian Journal of Statistics. Theory and Applications, 51(2):557–589, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wermuth:1993:ESM

[WR93] Nanny Wermuth and Helmut Rüssmann. Eigenanalysis of symmetrizable matrix products: a result with statistical applications. Scandinavian Journal of Statistics. Theory and Applications, 20(4): 361–367, September 1993. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616290.

Wang:2002:ELB

[WR02] Qihua Wang and J. N. K.

Rao. Empirical likelihoodbased inference in linear models with missing data. Scandinavian Journal of Statistics. Theory and Applications, 29(3):563–576, September 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616732.

Wretman:1978:SDA

[Wre78] Jan Wretman. A simple derivation of the asymptotic distribution of a sample quantile. Scandinavian Journal of Statistics. Theory and Applications, 5(2): 123–124, September 1978. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4615699.

Wu:2013:MEG

[Wu13] Rongning Wu. M-estimation for general ARMA processes with infinite variance. Scandinavian Journal of Statistics. Theory and Applications, 40(3): 571–591, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/24586687.

Wong:2001:AIF

[WW01] A. C. M. Wong and J. Wu. Approximate inference for the factor loading of a simple factor analysis model. Scandinavian Journal of Statistics. Theory and Applications, 28(2):407–414, June 2001. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616666.

Wu:2011:MEA

[WW11] Ximing Wu and Suojin Wang. Maximum entropy approximations for asymptotic distributions of smooth functions of sample means. Scandinavian Journal of Statistics. Theory and Applications, 38 (1):130–146, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/23015387.

Wade:2014:PSD

[WWP14] Sara Wade, Stephen G.

Walker, and Sonia Petrone. A predictive study of Dirichlet process mixture models for curve fitting. Scandinavian Journal of Statistics. Theory and Applications, 41(3):580–605, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586748.

Wen:2015:LEP

[WWW15] Canhong Wen, Xueqin Wang, and Shaoli Wang. Laplace error penalty-based variable selection in high dimension. Scandinavian Journal of Statistics. Theory and Applications, 42(3): 685–700, September 2015.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2003:SMR

[WY03] Yan Wang and Paul S. F. Yip. A semiparametric model for recapture experiments. Scandinavian Journal of Statistics. Theory and Applications, 30(4): 667–676, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616795.

Wylupek:2016:ATU

[Wył16] Grzegorz Wyłupek. An automatic test for the umbrella alternatives. Scandinavian Journal of Statistics. Theory and Applications, 43(4):1103–1123, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wu:2010:OBE

[WZ10] Ping Wu and Li Xing Zhu. An orthogonalitybased estimation of moments for linear mixed models. Scandinavian Journal of Statistics. Theory and Applications, 37(2): 253–263, June 2010. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000878.

Wang:2022:LSC

[WZ22] Pengfei Wang and Wensheng Zhu. Large-scale covariate-assisted two-sample inference under dependence. Scandinavian Journal of Statistics. Theory and Applications, 49(4): 1421–1447, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Wang:2016:ESI

[WZH16] Qihua Wang, Tao Zhang, and Wolfgang Karl Härdle. An extended single-index model with missing response at random. Scandinavian Journal of Statistics. Theory and Applications, 43(4):1140–1152, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xin-Bing:2015:FDN

[XBQF15] Kong Xin-Bing and Xu Qin-Feng. On false discovery and non-discovery proportions of the dynamic adaptive procedure. Scandinavian Journal of Statistics. Theory and Applications, 42(2):530–544, June 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xie:2009:CDA

[XG09] Xianchao Xie and Zhi Geng. Collapsibility for directed acyclic graphs. Scandinavian Journal of Statistics. Theory and Applications, 36(2):185–203, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000317.

Xiang:1994:BRK

[Xia94]  Xiaojing Xiang.  Ba-

hadur representation of kernel quantile estimators. Scandinavian Journal of Statistics. Theory and Applications, 21(2): 169–178, June 1994. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616308.

Xie:1988:NNM

[Xie88] Min Xie. A note on the natvig measure. Scandinavian Journal of Statistics. Theory and Applications, 15(3):211–214, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616103.

Xie:1989:STT

[Xie89] Min Xie. Some total time on test quantities useful for testing constant against bathtub-shaped failure rate distributions. Scandinavian Journal of Statistics. Theory and Applications, 16(2):137–144, June 1989. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616129.

Xue:2010:PSE

[XL10] Lan Xue and Hua Liang.

Polynomial spline estimation for a generalized additive coefficient model. Scandinavian Journal of Statistics. Theory and Applications, 37(1):26– 46, March 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000914.

Xu:2016:LES

[XLS16] Yuhang Xu, Yehua Li, and Xiao Song. Locally efficient semiparametric estimators for proportional hazards models with measurement error. Scandinavian Journal of Statistics. Theory and Applications, 43(2):558–572, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xu:2020:VSS

[XLY20] Jinfeng Xu, Wai Keung Li, and Zhiliang Ying. Variable screening for survival data in the presence of heterogeneous censoring. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1171–1191, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xu:2015:IAR

[XMW15] Kun Xu, Yanyuan Ma, and Liqun Wang. Instrument assisted regression for errors in variables models with binary response. Scandinavian Journal of Statistics. Theory and Applications, 42 (1):104–117, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xia:2023:SIC

[XNL23] Lu Xia, Bin Nan, and Yi Li. Statistical inference for Cox proportional hazards models with a diverging number of covariates. Scandinavian Journal of Statistics. Theory and Applications, 50(2):550–571, June 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xue:2020:LWL

[XT20] Yujie Xue and Masanobu Taniguchi. Local Whittle likelihood approach for generalized divergence. Scandinavian Journal of Statistics. Theory and Applications, 47(1):182–195, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xue:2009:ELC

[Xue09] Liugen Xue. Empirical likelihood confidence intervals for response mean with data missing at random. Scandinavian Journal of Statistics. Theory and Applications, 36(4):671–685, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000346.

Xue:2010:ELL

[Xue10] Liugen Xue. Empirical likelihood local polynomial regression analysis of clustered data. Scandinavian Journal of Statistics. Theory and Applications, 37(4): 644–663, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41000413.

Xu:2014:FIC

[XWH14] Ganggang Xu, Suojin Wang, and Jianhua Z. Huang. Focused information criterion and model averaging based on weighted composite quantile regression. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 365–381, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        24586658.

Xu:2015:TUE

[XY15] Ke-Li Xu and Jui-Chung Yang. Towards uniformly efficient trend estimation under weak/strong correlation and non-stationary volatility. Scandinavian Journal of Statistics. Theory and Applications, 42(1): 63–86, March 2015. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Xu:2009:KBG

[XZ09] Wangli Xu and Lixing Zhu. Kernel-based generalized cross-validation in non-parametric mixedeffect models. Scandinavian Journal of Statistics. Theory and Applications, 36(2):229–247, June 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41000319.

Yata:2020:GCP

[YA20] Kazuyoshi Yata and Makoto Aoshima. Geometric consistency of principal component scores for highdimensional mixture models and its application. Scandinavian Journal of Statistics. Theory and Applications, 47(3):899–921, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yao:1996:CBC

[Yao96] Qiwei Yao. Conditional boundary crossing probabilities and two-stage tests for a change-point. Scandinavian Journal of Statistics. Theory and Applications, 23(4):511–525, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616423.

Yang:2022:TPV

[YC22] Lu Yang and Claudia Czado. Two-part D-vine copula models for longitudinal insurance claim data. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1534–1561, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yuan:2007:SRK

[YD07] Ao Yuan and Jan G. De Gooijer. Semiparametric regression with kernel error model. Scandinavian Journal of Statistics. Theory and Applications, 34(4): 841–869, December 2007. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/41548584.

Yanagihara:2012:IBC

[YF12] Hirokazu Yanagihara and Hironori Fujisawa. Iterative bias correction of the cross-validation criterion. Scandinavian Journal of Statistics. Theory and Applications, 39(1):116–130, March 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41411160.

Ye:2020:LQE

[YH20] Zi Ye and Giles Hooker. Local quadratic estimation of the curvature in a functional single index model. Scandinavian Journal of Statistics. Theory and Applications, 47(4):1307–1338, December 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Jonsdottir:2013:LBM

[ýJRNMJ13] Kristjana ýr Jónsdóttir, Anders Rønn-Nielsen, Kim Mouridsen, and Eva B. Vedel Jensen. Lévy-based modelling in brain imaging. Scandinavian Journal of Statistics. Theory and Applications, 40(3):511–529, September 2013. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586684.

Yang:2016:LBI

[YK16] Shu Yang and Jae Kwang Kim. Likelihood-based inference with missing data under missing-at-random. Scandinavian Journal of Statistics. Theory and Applications, 43(2):436–454, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yang:2020:ATI

[YK20] Shu Yang and Jae Kwang Kim. Asymptotic theory and inference of predictive mean matching imputation using a superpopulation model framework. Scandinavian Journal of Statistics. Theory and Applications, 47(3):839–861, September 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yu:1996:ADB

[YL96] Philip L. H. Yu and K. Lam. Analysis of duplicate bridge tournament data. Scandinavian Journal of Statistics. Theory and Applications, 23(4): 621–633, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616429.

Yu:2004:LLA

[YL04] Keming Yu and Zudi Lu. Local linear additive quantile regression. Scandinavian Journal of Statistics. Theory and Applications, 31(3):333–346, September 2004. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616834.

Yao:2014:AP

[YL14a] Weixin Yao and Longhai Li. Acknowledgement of priority. Scandinavian Journal of Statistics. Theory and Applications, 41(4):1195, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586828.

Yao:2014:NRM

[YL14b] Weixin Yao and Longhai Li. A new regression model: Modal linear regression. Scandinavian Journal of Statistics. Theory and Applications, 41(3):656–671, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586752.

Yin:2020:NCE

[YLGL20] Yunjian Yin, Lan Liu, Zhi Geng, and Peng Luo. Novel criteria to exclude the surrogate paradox and their optimalities. Scandinavian Journal of Statistics. Theory and Applications, 47 (1):84–103, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yu:2000:CSC

[YLW00] Qiqing Yu, Linxiong Li, and George Y. C. Wong. On consistency of the self-consistent estimator of survival functions with interval-censored data. Scandinavian Journal of Statistics. Theory and Applications, 27(1):35–44, March 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616588.

Yu:2019:AMM

[YLZ^{+}19] Guanglei Yu, Yang Li, Liang Zhu, Hui Zhao, Jianguo Sun, and Leslie L. Robison. An additive– multiplicative mean model for panel count data with dependent observation and dropout processes. Scandinavian Journal of Statistics. Theory and Applications, 46(2):414–431, June 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yang:2022:UDG

[YM22] Kaixu Yang and Tapabrata Maiti. Ultrahigh-dimensional generalized additive model: Unified theory and methods. Scandinavian Journal of Statistics. Theory and Applications, 49(3): 917–942, September 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ytterstad:1991:PLP

[Ytt91] E. Ytterstad. Predictive likelihood predictors for the AR(1) model: a small sample simulation. Scandinavian Journal of Statistics. Theory and Applications, 18(2):97–110, December 1991. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616194.

Yu:2011:BJT

[Yu11] Menggang Yu. Buckley– James type estimator for censored data with covariates missing by design. Scandinavian Journal of Statistics. Theory and Applications, 38(2): 252–267, June 2011. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015493.

Yu:2016:CAI

[Yu16]  Dalei Yu.  Conditional

Akaike information criteria for a class of Poisson mixture models with random effects. Scandinavian Journal of Statistics. Theory and Applications, 43(4): 1214–1235, December 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Yukich:1992:WCS

[Yuk92] J. E. Yukich. Weak convergence of smoothed empirical processes. Scandinavian Journal of Statistics. Theory and Applications, 19(3): 271–279, September 1992. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616244.

Yu:2006:CSP

[YWK06] Qiqing Yu, George Y. C.

Wong, and Fanhui Kong. Consistency of the semiparametric MLE in linear regression models with interval-censored data. Scandinavian Journal of Statistics. Theory and Applications, 33(2):367–378, June 2006. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        4616929.

Yu:2015:MCD

[YY15] Menggang Yu and Constantin T. Yiannoutsos. Marginal and conditional distribution estimation from double-sampled semi-competing risks data. Scandinavian Journal of Statistics. Theory and Applications, 42 (1):87–103, March 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

You:2007:SEM

[YZ07] Jinhong You and Haibo Zhou. On semiparametric EV models with serially correlated errors in both regression models and mismeasured covariates. Scandinavian Journal of Statistics. Theory and Applications, 34(2): 365–383, June 2007. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41548557.

Yang:2012:CST

[YZ12] Song Yang and Yichuan Zhao. Checking the shortterm and long-term hazard ratio model for survival data. Scandinavian Journal of Statistics. Theory and Applications, 39(3): 554–567, September 2012. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/41679812.

Yang:2023:MRM

[YZ23] Shu Yang and Yunshu Zhang. Multiply robust matching estimators of average and quantile treatment effects. Scandinavian Journal of Statistics. Theory and Applications, 50 (1):235–265, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

You:2011:SEP

[YZZ11] Jinhong You, Xian Zhou, and Yong Zhou. Series estimation in partially linear in-slide regression models. Scandinavian Journal of Statistics. Theory and Applications, 38(1):89– 107, March 2011. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
23015385.

Zahl:1996:LNP

[Zah96] Per-Henrik Zahl. A linear non-parametric regression model for the excess intensity. Scandinavian Journal of Statistics. Theory and Applications, 23(3):

353–364, September 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616409.

Zhang:2020:CLF

[ZBS20] X. Zhang, B. Beranger, and S. A. Sisson. Constructing likelihood functions for interval-valued random variables. Scandinavian Journal of Statistics. Theory and Applications, 47(1):1–35, March 2020. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhu:2003:SPR

[ZC03] Lixing Zhu and Hengjian Cui. A semi-parametric regression model with errors in variables. Scandinavian Journal of Statistics. Theory and Applications, 30(2):429–442, June 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

                                        https://www.jstor.org/
                                        stable/4616773.

Zhu:2022:SAI

[ZCL22] Yayuan Zhu, Ziqi Chen, and Jerald F. Lawless. Semiparametric analysis of interval-censored failure time data with outcomedependent observation schemes. Scandinavian Journal of Statistics. Theory and Applications, 49(1):236–264, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:2024:TTP

[ZD24] Junyi Zhang and Angelos Dassios. Truncated two-parameter Poisson– Dirichlet approximation for Pitman–Yor process hierarchical models. Scandinavian Journal of Statistics. Theory and Applications, 51(2):590–611, June 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zetterqvist:1988:ADM

[Zet88] Lena Zetterqvist. Asymptotic distribution of Mann’s test for trend for mdependent seasonal observations. Scandinavian Journal of Statistics. Theory and Applications, 15 (2):81–95, September 1988. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616089.

Zhang:2003:SEL

[ZG03] Jian Zhang and Irène Gijbels. Sieve empirical likelihood and extensions of the generalized least squares. Scandinavian Journal of Statistics. Theory and Applications, 30(1): 1–24, March 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616746.

Zhou:2022:RPS

[ZGZ22] Niwen Zhou, Xu Guo, and Lixing Zhu. The role of propensity score structure in asymptotic efficiency of estimated conditional quantile treatment effect. Scandinavian Journal of Statistics. Theory and Applications, 49(2): 718–743, June 2022. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:1995:APE

[Zha95]  Ping Zhang.  Assess-

ing prediction error in non-parametric regression. Scandinavian Journal of Statistics. Theory and Applications, 22(1):83–94, March 1995. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616342.

Zhang:1996:LIL

[Zha96] Biao Zhang. A law of the iterated logarithm for kernel estimators of hazard functions under random censorship. Scandinavian Journal of Statistics. Theory and Applications, 23(1):37–47, March 1996.

CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616384.

Zhang:1998:LI

[Zha98] Li-Chun Zhang. Likelihood imputation. Scandinavian Journal of Statistics. Theory and Applications, 25(2): 401–414, June 1998. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616510.

Zhang:2000:EUT

[Zha00] Biao Zhang. M-estimation under a two-sample semiparametric model. Scandinavian Journal of Statistics. Theory and Applications, 27(2):263–280, June 2000. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
4616602.

Zhang:2008:PEE

[Zha08] Chunming Zhang. Prediction error estimation under Bregman divergence for non-parametric regression and classification. Scandinavian Journal of Statistics. Theory and Applications, 35(3):496–523, September 2008. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000284.

Zhu:2003:LIA

[ZHF03] Zhong-Yi Zhu, Xuming He, and Wing-Kam Fung. Local influence analysis for penalized Gaussian likelihood estimators in partially linear models. Scandinavian Journal of Statistics. Theory and Applications, 30(4): 767–780, December 2003. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616801.

Zhang:2010:SBS

[ZHH10] Ying Zhang, Lei Hua, and Jian Huang. A splinebased semiparametric maximum likelihood estimation method for the Cox model with interval-censored data. Scandinavian Journal of Statistics. Theory and Applications, 37(2):338–354, June 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000883.

Zhong:2015:TCH

[ZHL15] Ping-Shou Zhong, Tao Hu, and Jun Li. Tests for coefficients in high-dimensional additive hazard models. Scandinavian Journal of Statistics. Theory and Applications, 42(3):649–664, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zeng:2017:GDF

[ZHL17] Peng Zeng, Qinqin Hu, and Xiaoyu Li. Geometry and degrees of freedom of linearly constrained generalized lasso. Scandinavian Journal of Statistics. Theory and Applications, 44(4): 989–1008, December 2017. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:2022:PBC

[ZHS22] Haixiang Zhang, Jian Huang, and Liuquan Sun. Projection-based and crossvalidated estimation in high-dimensional Cox model. Scandinavian Journal of Statistics. Theory and Applications, 49(1):353–372, March 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:2019:UEL

[ZHW19] Shixiao Zhang, Peisong Han, and Changbao Wu. A unified empirical likelihood approach for testing MCAR and subsequent estimation. Scandinavian Journal of Statistics. Theory and Applications, 46 (1):272–288, March 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhu:2009:DMG

[ZIS09] Hongtu Zhu, Joseph G. Ibrahim, and Xiaoyan Shi. Diagnostic measures for generalized linear models with missing covariates. Scandinavian Journal of Statistics. Theory and Applications, 36(4):686–712, December 2009. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

                                        /www.jstor.org/stable/
                                        41000347.

Zhou:2024:ELE

[ZKP^{+}24] Xingcai Zhou, Dehan Kong, Matthew Stephen Pietrosanu, Linglong Kong, and Rohana J. Karunamuni. Empirical likelihood Mestimation for the varyingcoefficient model with functional response. Scandinavian Journal of Statistics. Theory and Applications, 51(3):1357–1387, September 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhou:2010:BRD

[ZL10] Hua Zhou and Kenneth L. Lange. On the bumpy road to the dominant mode. Scandinavian Journal of Statistics. Theory and Applications, 37(4):612–631, December 2010. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
41000411.

Zhang:2014:OWA

[ZL14] Jin-Ting Zhang and Xuehua Liang. One-way ANOVA for functional data via globalizing the pointwise F -test. Scandinavian Journal of Statistics. Theory and Applications, 41(1):51–71, March 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586730.

Zhang:2018:PLA

[ZL18] Jun Zhang and Heng Lian.

Partially linear additive models with unknown link functions. Scandinavian Journal of Statistics. Theory and Applications, 45 (2):255–282, June 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:2022:FEH

[ZL22] Jian Zhang and Jie Li. Factorized estimation of highdimensional nonparametric covariance models. Scandinavian Journal of Statistics. Theory and Applications, 49(2):542–567, June 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhong:2021:MAV

[ZLK21] Ping-Shou Zhong, Jun Li, and Piotr Kokoszka. Multivariate analysis of variance and change points estimation for high-dimensional longitudinal data. Scandinavian Journal of Statistics. Theory and Applications, 48(2):375–405, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:2016:SSC

[ZLL^{+}16] Xinyu Zhang, Hua Liang, Anna Liu, David Ruppert, and Guohua Zou. Selection strategy for covariance structure of random effects in linear mixedeffects models. Scandinavian Journal of Statistics. Theory and Applications, 43(1):275–291, March 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhang:2014:SNS

[ZLS14] Xianyang Zhang, Bo Li, and Xiaofeng Shao. Selfnormalization for spatial data. Scandinavian Journal of Statistics. Theory and Applications, 41(2): 311–324, June 2014. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586655.

Zhou:2014:SLV

[ZLSL14] Ling Zhou, Huazhen Lin, Xinyuan Song, and Yi Li. Selection of latent variables for multiple mixed-outcome models. Scandinavian Journal of Statistics. Theory and Applications, 41(4): 1064–1082, December 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586821.

Zheng:2014:SSP

[ZLY14] Gang Zheng, Qizhai Li, and Ao Yuan. Some statistical properties of efficiency robust tests with applications to genetic association studies. Scandinavian Journal of Statistics. Theory and Applications, 41(3): 762–774, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/24586757.

Zhu:2021:TCI

[ZLZZ21] Xuehu Zhu, Jun Lu, Jun Zhang, and Lixing Zhu. Testing for conditional independence: a groupwise dimension reductionbased adaptive-to-model approach. Scandinavian Journal of Statistics. Theory and Applications, 48 (2):549–576, June 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhao:2024:NSM

[ZMH^{+}24] Ge Zhao, Yanyuan Ma, Jill Schnall Hasler, Scott Damrauer, Michael Levin, and Jinbo Chen. A nested semiparametric method for case-control study with missingness. Scandinavian Journal of Statistics. Theory and Applications, 51 (1):201–219, March 2024. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Ziegel:2015:EPS

[ZNJ15] Johanna F. Ziegel, Jens R. Nyengaard, and Eva B. Vedel Jensen. Estimating particle shape and orientation using volume tensors. Scandinavian Journal of Statistics. Theory and Applications, 42(3):813–831, September 2015. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhao:2024:CCE

[ZSJT24] Zixuan Zhao, Yanglei Song, Wenyu Jiang, and Dongsheng Tu. Consistent covariances estimation for stratum imbalances under minimization method for covariate-adaptive randomization. Scandinavian Journal of Statistics. Theory and Applications, 51(2): 861–890, June 2024. CO-DEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zurbenko:1979:EES

[Žur79] Igor G. Žurbenko. On the efficiency of estimates of a spectral density. Scandinavian Journal of Statistics. Theory and Applications, 6 (2):49–56, December 1979. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4615731.

Zaidi:2021:IIL

[ZV21] Jaffer M. Zaidi and Tyler J.

VanderWeele. On the identification of individual level pleiotropic, pure direct, and principal stratum direct effects without cross world assumptions. Scandinavian Journal of Statistics. Theory and Applications, 48(3): 881–907, September 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhao:2022:ECQ

[ZVD22] Yue Zhao, Ingrid Van Keilegom, and Shanshan Ding. Envelopes for censored quantile regression. Scandinavian Journal of Statistics. Theory and Applications, 49(4):1562–1585, December 2022. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zwanzig:2016:DPA

[Zwa16] Silvelyn Zwanzig. Discussion of the paper “Asymptotic Theory of Outlier Detection Algorithms for Linear Time Series Regression Models” by Søren Johansen and Bent Nielsen. Scandinavian Journal of Statistics. Theory and Applications, 43(2):371–373, June 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). See [JN16a].

Zhao:2019:ELW

[ZWS19] Puying Zhao, Lei Wang, and Jun Shao. Empirical likelihood and Wilks phenomenon for data with nonignorable missing values. Scandinavian Journal of Statistics. Theory and Applications, 46(4):1003–1024, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhao:1996:MLE

[ZX96] M. Zhao and M. Xie. On maximum likelihood estimation for a general nonhomogeneous Poisson process. Scandinavian Journal of Statistics. Theory and Applications, 23(4): 597–607, December 1996. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616427.

Zografi:2019:MSO

[ZX19] Mimoza Zografi and Evdokia Xekalaki. Modeling spatial overdispersion via the generalized Waring point process. Scandinavian Journal of Statistics. Theory and Applications, 46(4): 1098–1116, December 2019. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zheng:2018:EAS

[ZXL^{+}18] Gang Zheng, Jinghong Xiong, Qizhai Li, Jinfeng Xu, Ao Yuan, and Joe L. Gastwirth. Evaluating the accuracy of small P -values in genetic association studies using Edgeworth expansions. Scandinavian Journal of Statistics. Theory and Applications, 45(1):1– 33, March 2018. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhu:2023:MPF

[ZXLL23] Weicheng Zhu, Sheng Xu, Catherine C. Liu, and Yehua Li. Minimax powerful functional analysis of covariance tests with application to longitudinal genome-wide association studies. Scandinavian Journal of Statistics. Theory and Applications, 50 (1):266–295, March 2023. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhu:2002:RMT

[ZYT02] L. X. Zhu, K. C. Yuen, and N. Y. Tang. Resampling methods for testing a semiparametric random censorship model. Scandinavian Journal of Statistics. Theory and Applications, 29 (1):111–123, March 2002. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL

https://www.jstor.org/
stable/4616703.

Zhu:2014:SIS

[ZYX14] Liping Zhu, Jinhong You, and Qunfang Xu. Statistical inference for singleindex panel data models. Scandinavian Journal of Statistics. Theory and Applications, 41(3):830–843, September 2014. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https:/

/www.jstor.org/stable/
24586761.

Zhang:2021:FSC

[ZZLC21] Jing Zhang, Haibo Zhou, Yanyan Liu, and Jianwen Cai. Feature screening for case-cohort studies with failure time outcome. Scandinavian Journal of Statistics. Theory and Applications, 48(1):349–370, March 2021. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).

Zhu:2016:PWL

[ZZLZ16] Rong Zhu, Guohua Zou, Hua Liang, and Lixing Zhu. Penalized weighted least squares to small area estimation. Scandinavian Journal of Statistics. Theory and Applications, 43(3): 736–756, September 2016. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic).